A Complete Bibliography of Publications in Statistics and Probability Letters: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.06

Title word cross-reference

(N, d) [WC97]. (n, f, k) [CCH99]. 0 < r < 1/2 [San91]. 1 [MH96]. 2 [EK96b, Ete99, SN96]. 2^{β} [PP92b]. 2^{k} [Kou96c]. 2m [Arc98a]. 3^{k} [Kou96c]. 4 [Kab99]. 5 [Shi99]. _{a} [PP92a]. A [Brz96, MH90]. α [AM94a, Alb97, CR92, Lie99, Mas96, Sam94, Sap99, Sha93, ZW97a, ZW97b]. L̄ [BS96a]. X̄ [KT93]. β [Hil99]. modn [DH97]. c [Rüs96]. C(α) [Wei92]. χ^{2} [Cha90d]. D [BH90, Brz96, Cha99c, CJ98, Kou96b, Kou96f, MZ97, Sch97c, Won95, MV95b]. D[0, 1] [BP93]. D_{+} [Fie90]. δ [Fal94, Kau95]. E [Brz95, Brz96, MZ97]. E(X|X + W) [Beh99]. E_{R} [Koz96]. ϵ [Bor96, Wu92]. F [Ext91, Run93, Sil96a, Zog94]. F^{α} [Bal94]. G [Imh97, Won95, Kra98]. I′ [Fie90, RT93]. k [BP99b, BCQY97, BGP94, CSC93, Cha90b, CAGM98, HMG98, HT91, NNR97, PA90, PAT90, RC94, SB95a, SS95, Tso96, Vas90]. k(≥ 2) [MvdM97]. K × K × K [EZ99]. L [Abr93, BSW97, Brz96, Don98, Ger93, GR93a, PJ94, RW94, Sha91, Xia95].

L^{P} [Lin95, Web92]. L^{p}(0, 1) [OS98]. L_{1} [AM94b, BW97, Dav93, EH99b, Hub97, KP94b, LM92, ST96]. L_{2} [LH98a]. L_{∞} [CAGM98]. L_{p} [Ahm97, Che96, GKL96, Leb96, Wu90, Hei93]. L_{q} [Joh91]. L_{s} [FPMP97]. λ_{3} [HK91]. M [Arc94b, Arc98b, BlZ93, BKR92, Cro94, Fer98, Gen98, Jon94, LZ91, RZ92a, RC94, WW90b, Zha97a, Bo98, Che97b, GB96, Sco93, Wan99d, Wie96]. R^{2} [Whi99]. R^{d} [FM99]. R^{k} [Jen93, Mee91, Mee93b]. R^{m} [EH99a]. Z [How97]. Z^{2} [Aue90]. C_{∞} [Zha94b]. L [BS96a]. L_{∞} [Kul95]. L_{√} [YM95]. T_{⊔} [AB97].

μ′Σ^{−1}μ [Mar94]. N [KK96a, RS99, SS95]. n^{r} [San91]. N_{p}(μ, Σ) [Mar94]. n: F [BP99b]. n ≡ 2 mod 4 [Kou96b]. P [BD90a, ZR96, ARV97, CF96b, Khr98, Lee98a, San92, Sap95, Shi97b, Sub94, Tho97, WR90]. ϕ [HL92]. ϕ^{o}∗ [Qiy95]. ψ [Che92b]. Q [BD90a]. R [Mik98, Ras95, WW92a, BB93b, DH90, Rao99a, SB91]. ρ [PS96b]. S [Ege92, Hös92]. S^{2} [KT93, Son99]. SαS [Rut95]. t [AVB95, Bal95, CS97, GS93a, Jen94a, RZ94, Sep96, SMP95, Sri98]. τ [ABY94, dE93]. × [GPI99, Man99]. U [AJ95, Arc94a, Arc95b, Arc95c, Arc98a, BW99, Chr90, Chr91, Chr94, Chr97, CS98a, dE97, Eic98, HJ93, JVS90, KRW98, KH91, Qiy95, Rem98, Sch97a, Wan99d, CC98a]. U(p; m, n) [Bil96]. V [LN95, Shi94, KP96b]. W [Beh99]. X [BS94a, Beh99]. |X − y| [BS94a].

-Adic [Khr98]. -ancillarity [ZR96]. -ARCH [Hil99]. -band [Shi99]. -based [CAGM98]. -class [CSC93, Don98]. -classes [BS96a]. -connections [ZW97a]. -Consistency [San91]. -contamination [Bor96, Wu92]. -criterion [Imh97]. -dependent [Bo98, Che97b, Wan99d]. -designs [Sri98]. -dimension [SN96]. -dimensional [MH96]. -dissimilarity [Zog94]. -distribution [AVB95, Ext91]. -divergence [ZW97b]. -efficiency [Brz96]. -estimate [RZ92a]. -estimates [ABY94, Che92b, LZ91]. -estimation [Wie96, EH99b]. -estimator [BSW97, Hub97]. -estimators [Arc94b, Arc98b, BKR92, Cro94, Gen98, Hös92, Ras95, RC94, WW92a, WW90b, Xia95]. -exchangeable [EK96b, Ete99]. -exp [Gel96]. -functionals [Ahm97, GR93a, Zha97a]. -geometry [ZW97b]. -independent [Kab99]. -losses [Leb96]. -manifold [ZW97a]. -matches [BGP94, HMG98]. -max [CF96b, Sub94]. -mean [SB95a]. -median [LM92]. -mixing [CR92, HL92, Lie99, PS96b, Qiy95, Sha93, Ger93]. -models [Hof99]. -monotone [Kra98]. -monotonicity [Sap95, Sap99]. -nearest-neighbor [BCQY97]. -neighborhoods [Fal94, Kau95]. -nets [CAGM98]. -NN [Cha90b]. -norm [YM95, BW97, KP94b]. -norms [GKL96, Web92]. -optimal [Rüs96, Cha99c, CJ98, Kou96b, Kou96f, LN95, MH90, Won95]. -optimality [Brz95, Sch97c]. -optimum [MZ97]. -out-of- [SS95]. -parameter [Tso96]. -processes [Arc94a, Arc95b, Eic98, KH91]. -projections [RT93]. -quantiles [AJ95, Che96, Jon94]. -sample [Vas90]. -scheme [Bal94]. -sense [AM94b, FPMP97]. -shaped [Ege92]. -smoothers [BlZ93]. -smoothing [Abr93]. -space [MV95b]. -space-valued [WRL95]. -stable [Alb97, KP96b, Sam94]. -statistic [CS97, Sep96, Run93]. -statistical [BW99]. -statistics [RZ94, Arc95b, Arc95c, Arc98a, Chr90, Chr91, Chr94, Chr97, CS98a, HJ93, JVS90, KRW98, Qiy95, Rem98, RW94, Sch97a, Sha91, Shi94, Wan99d]. -step [RC94]. -sufficiency [ZR96]. -test [GS93a]. -th [KK96a]. -type [Fer98, Sil96a]. -unimodal [AM94a]. -unimodality [Mas96]. -values [San92, Tho97]. -variant [ST96]. -wise [Arc98a]. -x [Gel96]. -year [Bal95].

22 [BL99b]. 28 [Rad99]. 2SHI [Hoa93].

3 [GS93b]. 31 [HV00]. 35 [Lin00]. 36 [Ham98a]. 38 [HV00]. 3R [BP92].

4253H [JPAS97]. 43 [LGC99a, Pal01].

A.S. [RR90]. ABO [SPH91]. above [LMF97]. absolute [Arc95c, HT90, MN93, Men93, Tab94a, Tab94b]. absolutely [Han93, Vie99]. Accelerated [BN97, LSR90]. accuracy [Huw95]. Accurate [AAK97, BB97, Rob90b]. Adaptive [CK94a, DJ97, Fox97, Le 92, Ber99a, Efr94, WC90]. Adaptivity [GT99]. Addelman [SG96b]. Addendum [Ano99a, HV00, Pap98]. adding [Nyc90]. Addition [DM98]. additive [Bos96, CPPP98, CF95, Kro97, Sch96a]. adequacy [NY90]. Adic [Khr98]. adjacent [YL96]. adjusted [Koz91, Mül97]. Adjusting [Por92]. adjustment [SS99b, Wan94a]. adjustments [CC98a]. Admissibility [Wan98b, Yu93, ZL97, HM96, IM97b, Kus98, MZ97]. admissible [JK99b, Mar98]. admits [Bar92a]. Affine [Sta92, Jey91]. affine-equivariant [Jey91]. after [DF95, GS93a]. against [Bal97, Bal98, Bha95, EK95, EZ99, Sil96b, SW90a, Zam90]. Age [Mi99, Coo93a, Lee98b]. age-dependent [Coo93a, Lee98b]. Age-smooth [Mi99]. ageing [BCR96]. aggregation [FB96]. aging [PS97b]. AICu [MST97]. AIDS [Tan95]. Akaike [Cav97, Kan93a, Res96b]. Alarm [MCWD95]. algebraic [NS96c, Uno96]. algorithm [Cho97, Dor97a, FY93, Hol98, Hug97, LM96, Mkh98, Nyc90, Qia96b, Shi99, Wri97, WPH96]. Algorithms [SA97, dC98, Qia92, Zaj96]. aligned [Vas90]. allocation [Ger96]. allocations [Puk97]. Almost [BCH98, BCC99, BlZ93, ÖH96, FS98, Fer94, HL92, IL98, Kou96f, LP90, Lat98, Mat92, Mór92, PS95a, Pet95, Ros93]. along [AC99]. Alternative [Mad93, Lov98, Neu94, Wan94a]. alternatives [Bal97, Bal98, Che93, CS95b, EK95, Epp99, SW99, Szy91]. Altham [Fro95, Lov98]. Amari [ZW97a]. among [KC97b]. amplitude [Kat97]. analog [SE97]. analyses [PL92]. Analysis [AC96, JW99, PL93, Ras95, SP98, ADD93, AFH^{+}96, Che98a, DB94, End95, Fad96, Fad98, Fun92, Gei93, GRS99, GP96, Hec92, IK92, IK95, Jan94, KPTY97, LPB90, LM96, Leo93, Oka94, Pru92, Ruk96, SK93, Spi91, Tom92a, WL95, YB99, Zha94a, Zho99]. ancillarity [ZR96]. ancillary [GW95b, PRS92]. Andersen [Eis92]. Anderson [McC90, PK96]. Andersson [SW93]. angular [CH96]. annealing [Dor97a, Löw97]. annihilating [Haz98]. anomaly [MH97a]. ANOVA [BB95b, CC98b, MM95, Wan91a, Wan91b]. answer [Als91a]. anticipating [Alb97, HN98]. application [AFH^{+}96, Bal93b, BLB90, BG91, Cas90, Che96, FN95, Gu95, HP94, KPTY97, LGC99a, LGC99b, Lou97, MW90, Shi99, YFL96, Yan96a]. Applications [Bar98a, FLP99, Bur93, CT99, Chu96, CS96b, GK97, GSW93, Gut96, IA97, JdKP95, Kim92, Kou97, KW93, Li94a, LZ94, Olo99, PKM93, Pap95, Rob90a, ZN95]. applications-motivated [PKM93]. approach [CD95, GR93b, GLM94, Hey94, Hsi95, HJ93, Hwa99, Jum99, MP99, ÖH96, Rei92, Roo95, Sta97, Sut97, WW96, WL95, WL93, ZK93, Zho99]. Appropriate [ZK93]. Approximate [Bat90, KS95a, Arc98b, Bar95a, Huw95, PS95b, Rob95a]. approximately [Bas94, MN93, Wie91, Wie96]. Approximating [Res96a, SW99, Kuk99]. Approximation [KR98, OW91, BD90a, Bla91, BX95, BGP94, Ehm91, Elt92, FR92, Ger91, HM90, HS99, Hür95, JR99, KK96b, Mon93, Nik99, Olo99, PL98, Pau96, Ral92, Roo95, VC99, Wic90, Yan91, Yu96, vRR98]. Approximations [CG99, San92, CCHP99, Cek98, CH91, CH98a, DW92, DMY91, DMLP97, KRW98, Kol95, QT92, SS97d, SS98, Wan92c]. approximative [Don91, RMNV99]. Arbitrarily [Ves94, Zha92b]. arbitrary [BB93b, BB94, Ete97, Pri99, Wen94, WZ96]. ARCH [Lu96, Hil99]. area [Neu98]. arg [Arc92]. ARIMA [SS93a, SS94a]. arising [BR98, Hei93, JK95a, Par96, She98, Sta95b]. arithmetic [Abt94]. ARMA [Cha99a, CD94, CK94b, VDO94, Zha94a]. arrangement [Bol98]. array [Chr90]. arrays [ARV97, Dav93, DM96, Gut92, HO95b, Hon96, HSV98, HV00, Sun98, Sun99, Wan96a, ZB94, dJ96]. artificial [End95, Oka94]. aspects [Mel95]. assassination [AK90]. Assessing [Gle96, KPTY97, LZ97, ST91, TW92c]. associable [Brz97]. associate [Brz91, GS93b]. associated [BR91a, BC98, CS98c, FLW99, IR99]. Association [KD96, CR97, CS98c, DK92, HH99, Nel92, Rou91a, Rou95, WMN94]. assumption [AB97, Bra96a, Che93, Liu94b]. assumptions [Bir92, Coo96, Zha96]. Asymmetric [MH97a, RS99, Str96, WK99]. asymmetrical [AS96, DM96]. asymmetry [XM91]. Asymptotic [Arc95a, Bar92b, Cai98, Cha90b, CM99, CR91, Don91, Dro99, ES96, Fal97, GL95a, Gen98, GKS96b, GP96, Har96, HL96, IK92, Kay92, KNN97, Kun93, KV96, Lie99, LY94, Llo91, PG94b, Phe95, Rou95, Rou97, Sha90, Tak94, Tso98, WW90b, YSLW98, Zha97b, Zog94, dC93, Abr93, Abr99, AB92b, Bos95, Cek95, Cha98c, CM91a, CM91b, DL91a, Erm92, FH99, FS93, Gan95, GT99, GCL90, GS99, Gui96, Haw90, HP93a, HB97, Jia97b, KP96a, Lee96b, LZ91, Lia95, MRN93, MP99, Mie97, MR99, NS96c, Ouk92, Par96, Res96b, Run93, SQ90, SQ97, Ser96b, SB97b, Sho97, SH99, Ste93, Wu90, Yao90, YB99, Zha94a, ZC95]. asymptotical [JS97]. Asymptotically [Cha94, Irl96, KPM94, Pen98, Rod93, Wei91, Dat95, HT95, LP91a, Wef91]. Asymptotics [Lee98a, Sep96, SB95a, Wie96, CL98, Koz97, PS96a, Rie97, SX97, Tab94a]. attracted [Sch95]. attraction [CF96b, Gel96, Jak94, Mee91, Mee93b, MS97b, Sep96, Sub94]. attributes [Wri97]. augmentation [Jen96]. Augmented [Dav97]. Author [Ano90b, Ano90a, Ano92c, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano94c, Ano94a, Ano94b, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano99b]. autocorrelated [Lee93]. autocorrelation [CR91, EPB94, Kou96a, Kou96d, Kou96c, Sch94]. autocorrelations [Mar97a, TV97]. autocovariance [Gen99, HK95]. automatic [Jon92]. autoregression [BKR96, GL95a, HK96]. autoregressions [PB97]. autoregressive [AD99, AC97a, BM99, BL95, BL99b, BS93, CC94a, Che98a, FB96, Hug96, HB97, IWLX99, Kak98, Kul92, Lee97c, Lee98a, Lu96, MM98b, Pap96, Pap97, Pap98, PS96a, PJ94, Qia96a, RR94a, SB96, SP93, SS99a, SNN97, TW92c, VNS96]. auxiliary [Zha97a]. Availability [SC99]. Average [LM96, Rog95, AM93, CF95, CF97, LL93, LRW92, Pap96, Pap97, Pap98, PS96a, Rao95, Rig95, Wan92a, Wen94, WS98, Zha96]. Average-case [LM96]. averages [Alb97, BL99a, BH96, Ete99, Lin93, ND93, Rie93]. Axiomatic [BN96].

B [AB92b, KL98b]. B-spline [KL98b]. B-splines [AB92b]. B.I.B. [Kan91a]. Backward [LS97, Mat97]. Bahadur [Baj92, Bar95a, BV92, CS97, Cha92a, Cha99b, DS92, Fal90b, Lah92, iY95]. balanced [Afs90a, Cha96a, Das92, Das98, DGS99, Kag91, Sia94, SJK96, Tal91, Wan91a, Wan91b]. ball [Wan97]. Ballot [Kon95]. Banach [ARV97, Che97b, JK98, Lec90, Nor95b, Rot93b, Shi97b, WR90]. Banach-space [Lec90]. Banach-space-valued [Rot93b]. band [Ruy96, Shi99]. band-limited [Ruy96]. Bandit [Her97]. bands [MX96, SZ98]. Bandwidth [KW98, Sun97b, CM90, ĆM93b, FH99, Jon92, KPM94, LP91a, Mam90, Mie90]. bandwidths [CM90, Haz98, Rob95b]. bar [LL97, Sil96a]. barrier [Chu96]. Bartlett [CC94a, Cor93a, Gle96, JW97, Li96, Muk92b, TG96, Wan94a]. Bartlett-type [Li96, Muk92b, TG96, Wan94a]. base [Sch98a]. based [AM98, Ahm97, ACGM99, AL97b, Bal93b, BH90, BH92, Bar95a, Bar99, BS94c, BHB96, BM93, BS98, Cav99, Cha94, Cha98b, CD95, CL98, CB96, Chr90, Chr94, Coo96, CAGM98, DJ95, EvZ92, ED94, ED96, Fut99, GR91, Gor91, Hua97b, Joe91, JL96, Kab99, KMC98, Kim99b, KA97, Kri90, Kri93, Lee96b, MHL95, MM96, MM98a, MMPZ99, Pap90, Par99, Raq97, RW94, Ryc95, SQ97, Sap95, SA97, SS97a, ST91, Siv99, Son99, VA97, Vel93b, Wan98c, XY97, YC90]. baseline [CC94c]. bases [Mas94]. basis [Lea91]. Baum [Lan98]. Bayes [Ang96, BM96, BM93, Bis99, Coo98, DLS90, Dal90, FR95, Hua97b, Hwa99, Leo93, LB91a, Lia99, Pru92, Res96a, Rou97, Ruk96, SB96, SV99, Tho97, YLB99, ZK97a]. Bayesian [Bas96a, BR91b, Bor96, Bor97a, Bru92, Che98a, De 96a, DB94, Dup97, Gei93, GRS99, JEG99, JK99b, Mac96, MN97a, MZ91, Mee93a, NS96a, NS96b, NS97a, PR97, RCD93, SE97, SHC99, Siv99, Spi91, Tom92a, WW96, Wan98c, ZRB98]. Bayesian-feasible [ZRB98]. BC [PP92a]. be [CC98a, Haw97, NN96, ZZ93]. Behavior [WW95a, Cro98, CM91a, CM91b, Deh91, DS92, DMWZ95, Epp99, Gan95, HL96, KP96b, LW92, MN97a, QW97b, QW97a, RW99, Res96b, Ros93, RS98, RA92, Yao90]. behaviour [DL91a, DM93, GG91, KNN97, Khr98, MR99, Ouk92, Pet95, Vel98]. behaviours [ZC95]. Behrens [Jan97, Lin91]. being [PW92]. Bell [Yua96]. Benford [Sch98a]. Bernoulli [BVB93, Ber97, Chr94, Dzi95, ES90, Khr98, Lin92b, Mus96, TYS95, Yan91]. Bernstein [Arc95b, Che95a]. Bernstein-type [Arc95b]. Berry [Che98d, RZ94, Wan99d]. Besov [KP92, KP93a]. Bessel [Cha90d, Rob90a]. Best [Hit90, Zie99, AB95, BSW97, CAGM98, HS99, HH94a, HH94b, KPM94, MS99, Jia97a]. beta [Kry99, Wal98, YM91]. better [LY99]. Between [BH96, Hüs94, Als91b, BJK98, Bry93, Cha90d, CC98b, CM90, Gil92, GO97, Ham95, HH98b, Hud97, KL99, KK99, Kim99a, KT93, Koc99, Kou90, Koz97, Mon93, NN96, PK96, Poo91, Pou93, RSC94, TG95, Tia98, YL96, Zie90, Zog94]. Bhattacharya [CR98]. Bias [CK94b, CV97, FH92, Haz98, Cha91, Che98c, CH99, Elt92, FBCCN96, GS95, KMC98, LN94, Men93, O’N92, Pit98, RC94, Woo92, Xia95]. bias-corrected [KMC98]. biased [Wu97b]. biases [Hal95]. biasing [Dav97]. BIB [Poh92]. BIBDs [Zha92a]. BIC [Res96b]. Bienaymé [Ber99a]. bifurcating [Hug96]. bilateral [Cho98]. bilinear [CFF98, GW96a, Mar97a]. bimonthly [Fra92b]. binary [ANS93, BC92, BT91, Bow99, CC98b, DK92, DT98, End95, Fad98, Glo99, HB99b, JL96, Kag91, Lau92, SJ98, WZ96]. Binomial [Ehm91, AEE92, Bal93b, DeR94, DeR96, Fer97, Fro95, Ham95, HT91, Huw95, Kru99a, KS95b, Lov98, Nee91, Ong95, Poo91, Sap99, Ser96a, TYS92, Wan96b]. binomial-negative [Bal93b]. bioassay [Mey99, Tsu92]. Birgé [Dev95]. Birth [KS92, AK90, Kyr94, Mud93, Mud94, PT96]. birth-and-assassination [AK90]. Birth-death [KS92]. birthday [Hen98b]. Bivariate [Hür93, AM97, BS97, Bol98, CS95a, Deh91, EK95, Gar98, HK91, HH96, Hüs94, Kru99a, KV96, Lai95, Lee97a, LB91b, LR96, Mar91, Nad99, NOT90, PD95, Pen99, PG99, Pru92, Spi91, SB95b, Wu97a, Zim94]. bivariate-multivariate [KV96]. Block [HRS99, BD91, Brz96, BCK97, GDD91, Koz96, Lah93, MM95, PS95b, Ras95, SG96a, Zha92b, Han93]. blocks [AC96]. blockwise [Bo98]. BLUE [Jia99]. BLUP [Jia97a, Jia99]. BM [Chu96]. Board [Ano90d, Ano91a, Ano91b, Ano92d, Ano92e, Ano93e, Ano93f, Ano94d, Ano94e, Ano95e, Ano95f, Ano95g, Ano95h, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano98f, Ano98g, Ano98h, Ano98i, Ano99c, Ano93d, Ano94f]. Bonferroni [GS96a, GR91, Lee97a, Sib91]. Bonferroni-type [GS96a, Lee97a, Sib91].

Bootstrap [Sha92b, AGMCA96b, AGMCA96a, Cha98d, CLK96, Csö92, Dat95, Fal90a, HJ93, Ind94, Lah91, Lah93, PP92a, Rad96, Rad99, Sep94, Wan92c]. bootstrapped [BG91, HS99]. Bootstrapping [CPS93, GGR98, HK96, Kni98, Lah98, Pap96, Pap97, Pap98, Zar99, AJ95]. both [Shi93]. bound [Bal90, CW97, Chu92c, GW93, Ham95, Hu97, HH94a, HH94b, Irl97, Kag91, Kul95, Pal92, Pog92, SE97, Uno96]. boundaries [HP93b]. Boundary [Kim92, DH90, FM92, LMF97, Nor95a]. bounded [Ber97, Dre99, FG92, Fis91, HLL96, IM97b, MS96b, PARA97, SC91b, Sil96b, Sta92, Stu98]. boundedness [Lat98]. Bounds [AB97, CG92, GC92, KPP96, MGH99, Rao99a, Raq97, Ryc95, Szr92, AM94a, BB93a, Bal93b, CS95a, Chr91, FK95, Gho93, Her97, Hua97a, Jia97b, Kim93, Kor97, Koz96, Lug95, MV95a, PP96a, Pit90, RZ94, RJ91, Rob90b, Rom99, RR91, Sia94, Wan99b, Wri97, Wu90]. Box [ST92b, Bal97, Bal98, KC95, Vel93a, Vel93b]. branching [GG91, GLM94, Gri96, KD92, KD93, Rév96, Sri91]. breakage [Hol90]. Breakdown [He94a, He94b, CMC94, Cro94, Dav94, Gor91, Hub97, NOT90, Rou93, Rou94, Ste91b, Van93]. breast [AFH^{+}96]. bridge [HZ99]. bridges [BD90a]. Brownian [AHZ97, AK95, BD90a, Bél91b, Ber96, BG99b, Che97a, Chu96, EN93, EN99, FN95, GA99, HZ99, KP93b, Kro97, Le 98, LL93, MM92, MS96a, Neu98, NV99, Str96, Stu97, Xia97, Yin99]. Bucy [Rut93]. Buehler [RJ91]. Buffon [WR98]. bulk [ADD93]. Bump [Hec92]. Burger [Kar96]. Burman [Joh96]. BVE [dE93].

calculating [Gug98]. calculation [CH94b]. calculations [GPI99, Wan98c]. calculus [Der95]. calibration [BS91, Wan99a]. Call [Ano90c]. Cam [Cek96]. Can [Haw97, KZ91, CC98a, EG99, NN96]. cancer [AFH^{+}96, Lee98b, ZRB98]. Cantelli [AGMCA96b, BR97, EvZ92, Kri90]. Cantelli-type [AGMCA96b, EvZ92]. Cantor [GP96, Hos94, KP96a, LT92]. capability [SHC99]. capture [YLB99]. capture-recapture [YLB99]. cardinal [AB92b]. Carlen [KL97]. Carlo [HRS99, Kuk99, RQ99]. carriers [Wie96]. case [BLB90, Beh99, CG94, De 96a, Fis91, Hua91, Ioa92, KT98, LM96, NS97a, RG99, WW95b, YSLW98]. case-control [WW95b]. cases [BB94, Ber99b, FR95, Kni98, vE97]. catastrophes [Kyr94]. categorical [Krz94]. categorised [RM90]. Cauchy [FRS93, HMT96, Mor94a]. causal [Cho97]. Causality [Pet96b]. cause [Aic98, Gas94]. censored [Akr92, Cai98, CB96, CC99b, Cro92, Elf97, Gas92, Gas94, Gei93, Ger95, GRS99, Har96, Ind94, KPTY97, Kul95, Lee99, Li95, LT99, Mos94, NNR97, PCK98, PP95, Phe90, PG99, RG95, Sun97b, SZ98, VNS96, Wan99b, WS90, Wil98, XY97, YSLW98, Zho96, Zho99]. censoring [CL98, Coo93a, FZ92, GK99, JV91, MX96, MSS90, Yan96a]. censorship [Cha90a, Gas96, Gu95, NS97b, WGR94, Yan90, dUÁGM98, dUÁGM99]. censorship-truncation [Gu95]. center [Hsi95]. central [BJK98, Bél91a, Cha90d, Rob90b, Szr92, Arc94a, Arc95c, BC96, BCH98, BP93, BP92, Bry93, FZF93, FZL97, HH98a, Hor90, IL98, Joh94a, LP90, Lee97d, Mik97, Mik98, PS95a, Qui99, Rog95, Sai95, Yan94, dN95]. centrality [Szr92]. centre [Hud97]. certain [Als91b, Ame90, BS97, CC94b, Dat95, DS94, FL91, GG91, GM93, GW96a, HH98a, How97, JdKP95, Ole97, OW91, Ros93, SQ90, San91, Sch99a, Wan98b, vEK98]. Cesàro [LH98b, vRR98]. CF [Nag99]. chain [BP99a, CH98b, HA98, HRS99, Lou97, Mat93, Pen96]. chains [Ban90, BR95, BT91, Bra99a, Cse93, Dup97, HLL96, Las97, Led95, LL95, MW97, MMPZ99, PT96, QT90, RCD93, Sen93, Spr98, Tan96, Wen94]. Chambers [Wer96]. Change [KL98a, CK94a, CH96, Fer94, Gom94, Hus96, Lee95, Lee96a, MS96a, MS97c, Ruk96, Wat95, Yao90, dC93]. Change-point [KL98a, CH96, Fer94, Lee96a, MS96a, MS97c, Ruk96, Yao90]. change-points [Gom94]. Changepoint [Szy91]. changes [Ser96a]. chaotic [BG95, Bos95]. characterisation [Tra96, WM97, Wal98]. Characteristic [Ram96, Ram97, Ame90, Li93, Li94c, MHL95, SA93]. Characterization [AC99, Fan91, FPMP97, GK99, HJM98, Mar96, Szr96, ANS93, ACS94a, BCR96, Bel99b, Cab90, CATD93, FB91, Kag90, Lee90, Lee98b, Lo91, MPSG90, Ole97, Sap99, SB95b, Tso96, Wes90, Wes94, Wis92, Wu97a, XY97, YFL96]. Characterizations [BK93, Bar98b, Bar99, Don98, LFR99, LB98, PS97b, RG99, Sap95, Ala93, Ala94, AVB95, BHK^{+}99, Mal96, McK91, Pap90, Pap93, RH91, YM91]. Characterizing [Bis96a]. Charlier [Epp99]. chart [CC98a, Rig95]. charts [AL97b, MCWD95]. Chebyshev [Sto96]. checks [NS97b]. chemotherapy [Tan95]. Chen [BX95]. Chernoff [BS98, Kou90]. Chernoff-type [BS98]. Chi [Lor92, AB92a, Cas90, Kal90, Lee96b, LL97, Rea93, Rob90b, San92, Sil96a, Szr92]. chi-bar [LL97]. chi-bar-squared [Sil96a]. Chi-square [Lor92, AB92a, Kal90, San92, Szr92]. chi-squared [Cas90, Lee96b, Rea93]. Chiang [She93]. Chibisov [Szy94]. Chiu [Dev94]. choice [Bes92, CF98, ĆM93b, DJ97, Jon92, KS99c, Mie90, Sun97b]. choices [Kay92, Wu97b]. Cholesky [ST92a]. choosing [HN97]. Chow [Che92a]. Christmas [Mun99a]. Chung [Deg93]. circle [Aue90, CG99, Gne98b, Hol90, Woo95]. Circular [Afs90a]. class [ANS93, ACH97, BB95a, BLB90, Bor96, CSC93, Cha90a, CPPP98, Chr97, CV97, Das98, Don98, Gen99, Gne95, Gne96, HB97, Khu97, Kra98, KV96, Lec90, Lee98a, LL95, MVA93, Nob95, PD95, PJ94, Rou97, RKD90, SJ97, SMP95, Sri98, Ste90, WZ95, Yao90, Zaj96, Zha97b]. classes [AA92, BLE90, BCR96, BS96a, Brz91, GS93b]. classification [KM98, MN97b, Moj97, PK96]. classifying [Leu96]. clinical [AFH^{+}96, MR97b]. closed [BB95a, Gas96, Mol93]. closed-form [Gas96]. closeness [BDG93, Nay98, SS93b]. Closure [Jen94a]. cloud [Mül99]. CLT [Rac95]. clues [Dal90]. cluster [SR97, Tal91]. clustered [Bow99, YR96]. clustering [BCQY97, CAGM98, EN96, SB95a]. clusters [Mee99]. codes [DGA95]. coefficient [BS96a, dE93, Fuj96, Gar98, GS93a, Haw90, Hof97, HL92, Hür95, HH94a, HB97, Jey92, LS97, MV95a, Mat97, Pen99, Qia96a, RP95, SS96a, Sch94, SA97, SNN97]. coefficients [CSC93, DS91, Far96, Kni91, Lee93, RR94a, Rot93b, SS93b, Str99]. Coexistence [LC92]. coherent [FK95, JDP90, KPP96, NJS98]. column [Koz91, Man99, TLC97]. columns [Man99]. combination [Bho92, Gro98]. combinations [ACS96, Bal93a]. combinatorial [Due98, MVA93, Muk95]. combined [Moj97]. Combining [ADEM94, GS96b, JK95b, San92]. come [Sta91]. comment [RJ91, Whi90]. Comments [FRS93]. commercial [Leo93]. common [AC91, EJ96, Hol95, Kim99a, Kle94, KP94a, MvdM97]. Commutative [Neu96]. compact [HJR93, Neu94]. compared [Bos96]. Comparing [Ber97, Coo96, De 96a, DR96b, Gei98, Mil92, Omo97, TG96]. Comparison [HH98b, Mar98, PKM93, BW93, Car92, Cha99a, Coh90, FP94, FF97, FK94, Gru91, Joe91, Kay92, MM94, Mey99, Pfl91, Sar96, Som93, Ste97, TP98, Tia98, Wu99b]. Comparisons [MS99, SJ98, DT93, LS98]. Compatibility [Liu99a]. compatible [JL96]. competing [NS96a, NS96b, NS97a, SD99]. competition [LC92]. competitive [CH94a]. Complete [Als91a, Kag90, LRW92, LS99, Sha93, Zha96, BSW97, Cha94, HSV98, HV00, Hüs94]. completely [SS96b]. Completeness [CR98, Mos94, BD92a]. Complex [Jum99, Cha92c, Far96, Li96, Ram99]. Complex-valued [Jum99]. component [CF96a, Gas94, KM93, Lu97, PW92, Sut97]. components [AB97, Bal97, Bal98, Cha91, DMG90, Das92, FK95, GS95, Kai99, KZ96, Kor97, Mar99a, MM95, PT98, RW92, Sgi98a, Spi91, Tak94, UU95, ZW99]. composed [Neu98]. composite [RM90, Wei91]. composition [Bar98a, JdKP95]. compound [Cek96, CS93b, DLS90, MGH99, VC99, WJ93]. compounded [PD95]. computation [CHW93, Mof97, Shi99, Sil96a]. Computing [Hug97, CS96a, Cor93a, Rou93]. concave [KW93, Wan94b]. Concavity [KC95, BB93b, SB91, Sze90]. Concentration [CT99, Jen93, MS90]. concerning [Dai98, Mie90, Nob95]. concomitance [BB95a]. concomitants [LV99]. condensed [MMP97]. condition [BES95, GW95b, Lie99, Sch98b, SG96b, Ste93]. Conditional [Che96, CB96, KT98, Tsa98, AW94, Arc98d, ACS94a, ACS94b, ACS96, BS94a, Beh99, BD92a, BS92, CA97, FS93, Gan95, IK92, IK95, KW96, Liu99a, Lu96, MM92, Muk92a, NG95, Sev93, Swa96, SV92]. conditionally [DF95, JL96, Vei90]. conditionals [ACS93, Bis96a, Hür93]. Conditioned [Kro95, KP93b]. Conditions [Elt92, BW97, BCH98, Bra99a, CM95, DeR94, DeR96, DT98, Fan91, FS93, HRG97, HH99, Kau95, Lan98, LVCW93, Liu99a, Mik98, Rad96, Rad99, Sep94, Sgi98a, Str94]. cones [DL91b, LL97]. Confidence [FTV98, Fis91, HP93b, Jey92, Kab99, Ped97, AB92b, BM99, Bax93, Cha94, CC94b, CH94b, Dre99, GH96, GH97, HS91, Hei93, Hor95, HkL92, KM97, Liu99b, MX96, PP92a, Qin97, RJ91, Sta92, SZ98, Tsa98, Wan98b, Wan98c, Wri97, Zha97a]. congruential [PP92b]. conjecture [Li94b, RHO99]. Conjugate [MW91, Coo93b, Spi91, Tsu92]. connected [FL91]. Connectedness [HRG97]. Connection [Koz97, Bry93, Cha90d, Gil92, MPS97]. Connections [KC97b, ZW97a].

Connectivity [BCQY97]. Consecutive [BP99b]. consensus [Sha92a]. conservatism [Som93]. conservative [San92]. Conservativeness [Kru99a]. Consistency [Giv95, Ser96a, AAJ97, BW97, BR92, cIC99, CY99, CZ97, Dev94, GW97, Ho95a, KM90, KC96, Lec90, MGH99, MH96, PCK98, RZ92a, Rou91a, Sri91, Stu92, Zie98, dUÁGM99, San91]. Consistent [Rao95, SS97b, CW94, GH96, GH97, Moj97, Sch96b, Sch96a]. constant [DB97, Hit90, KS92, Ram96, Wan98b]. constant-coverage [Wan98b]. constants [GN94]. constrained [Git97, HRS99, Mey99]. constructed [Kou96c]. Constructing [CS91]. Construction [DM96, Lai95, Man99, BLE90, MRR92, Wan96a]. Constructions [DK93, RR90]. Constructive [LDG97, TN93]. consumer [AAK97]. contain [SQC94]. contaminated [CS97, SBB93]. contamination [Bor96, Wu92]. contiguity [PV90]. contiguous [Che93, Szy91, Szy94]. contingency [AG94, CS98c, EK95, EZ99, Hub97, Par98b]. Continuity [HN98, CmHP98, SA93]. Continuous [VDO94, ADD93, BP99a, BB97, Cor93b, Dor97a, Han93, KD93, Kro97, Krz94, Le 92, LS97, Mat97, Ouk94, SH99]. Continuous-time [VDO94, BB97, SH99]. continuously [ADD93]. Contorted [DvC95]. contoured [Gen99, GV91, LF95b]. Control [AL97b, Liu97, Ber99a, CWM91, Le 92, MCWD95, Rig95, WW95b]. controlled [ADD93, TN93]. converge [Lah98]. Convergence [Arc98d, AB92b, ES91, GN94, Gu95, Gu98, Muk92a, New90, Rot93b, Stu94, TT91, vRR98, AdlC93, ARV97, Arc94b, Arc96, Arc98c, BMT97, BR95, BlZ93, Che95b, Chr90, Chr97, Dai98, Dav93, Dev95, Fal90a, Fal90b, FH99, FB95, Fer94, God90a, GR93b, GLM94, Gui96, Ham98a, Ham98b, HJR93, HRG97, Hol98, HSV98, HV00, IL98, IM97a, JN97, Kau95, KH91, Kor99, LRW92, LS99, Lin93, MR97a, Maj99, MS90, Mat92, Mkh98, Moh92, Neu97, OS98, Rac95, Rav92, RA93, RR94b, Ruk96, Sha93, Ste90, Sun97a, WR90, Wu99a, ZB94, Zha96, dUÁGM98]. converges [Str94]. Converse [BH97, BH99]. convex [Arc98c, Bal93b, Bel99b, BH97, BH99, FP94, GW93, Kla93, Mas99, Mol93, Raq97, Vit96, Zar99]. Convolution [SB99, Ext91, KS95a, Ong95]. convolutions [AM94a, KM99, Pur98]. Cook [Kim96]. cooperative [ASC98]. Corrected [CC99b, FC96, BS96b, Cav97, KMC98]. corrected-score [BS96b]. Correcting [CH99, Kal98]. Correction [Pal01, SS98, CK94b, CV97, FH92, JW97, McQ99, TG96]. corrections [Cor93a]. correlated [Bis96b, Chu96, DBD95, FNARP97, GL95b, HL92, LW92, NG95, Pek99, SC90, Sim92, SJ98, SP99]. correlatedness [Mór92]. Correlation [PK96, AT96, Alt93, BL91, CC98b, dE93, Gar98, Gen99, Gne98b, Haw90, HOGFLC98, HH94a, HH94b, Jey92, KG90a, Kim99a, KT93, KMS96, LP91b, NS96c, RB96, SS96a, SQ90, SQ97, Sch96d, Sch97b, Vit99, WMN94, YC90]. Correlations [NN96, CM90, DBD95, HT96, ME92]. Correspondence [End95, Oka94]. corresponding [NOT90]. cotype [WR90]. count [Fad96, FY93, Omo97]. countable [LL95]. counterexample [Imh97, LM92, Li94b, Nob95, RR90, Wis92]. counterexamples [HS91]. counting [Cse93, GKS97]. counts [SB97a]. couplings [LS98]. covariables [dUÁGM98, dUÁGM99]. Covariance [Rao99b, RSC94, Vit96, YR96, Zha94c, Bra96a, Jin93, JDP90, KL99, LV99, Mar91, PFvR98, PG94a, Pou93, cQ94, ST92a, WL95, Wan99a, Woo95, YSM94, YB99]. covariate [LVC90, RG95]. covariates [JL96, LSR90, Sep92]. coverage [SV92, Tsa98, Wan98b]. covered [Aue90]. covering [Hol90, Pal92]. Cox [Bal97, Bal98, CC94c, JW99, KC95, OM92, Vel93a, Vel93b, Wan99a]. Cramer [Sch98b, ST96, BD94, Gho93, Zim94]. Cramér-Rao [Gho93]. Cramér-von [Zim94]. crash [MR99]. credible [SBB93]. Criteria [Ete97, Moc94, Cav97, GN94, God90a]. criterion [Cav99, CLK96, FN95, Imh97, Kan93a, MST97, McQ99, Moh92, Nay98, PG98, SG91, VA97, WS98, ZK93]. Critical [CF96a, Zim94]. cross [CD95, DDD98, Dro99, GHW92, Pap96, Pap97, Pap98, Wir96]. cross-sectional [Wir96]. cross-validation [CD95, Dro99, GHW92]. Crossing [Mi96, HT95, HB99a, Kim92]. Cryer [Rie97]. Csiszár [CS98c]. cumulant [Rao99a]. cumulants [BJK98]. cumulative [Bor97b, Gu95, Rog95]. cure [Tso98]. curious [HMT96]. current [Abr99, AQ95, Grü94, KS99b]. current-status [Abr99]. curve [AGV97, Leo99, LY99, PFvR98, Pru92]. curves [Cha94, CZ97, Del93, FNARP97, KHW91, KW98]. cutting [Sai99]. Cycle [BP99b]. cylinder [End95, Oka94]. cylindrical [AC97b, Gaw97].

D [Poh92]. D-optimality [Poh92]. d’Agostino [BH90]. Darling [BM99]. Data [ĆM93b, Abr99, AAJ97, Ahm91, Akr92, AT96, AD99, AQ95, AFH^{+}96, Bos95, Bow99, Bru92, Cai98, Cha94, CL98, CC98b, CM90, CC99b, Cro92, CH96, End95, Fad96, Fad98, Fan99, FY93, Fuj96, Gar98, Gas96, GKS96a, Ger95, GRS99, Glo99, GSW93, Har96, Ind94, JS98, Kni91, KS99b, KPTY97, Kru99a, Kru99b, Krz94, Kul95, Lee99, Leo96, Li95, LT99, Lor95, LP91b, MM98a, MMP97, Mie90, Mie97, Mos94, NNR97, Oka94, PCK98, PP95, Pit98, PL93, PG99, RG95, SS96a, SQC94, SJ98, Son99, ST92b, Sun97b, SZ98, SP99, TV98, TG96, VNS96, Wan99b, Wan91a, Wan91b, WS90, Wil98, Wir96, WZ99, Wu97b, YLB99, YSLW98, Yua96, Zho96, Zho99]. Data-dependent [ĆM93b, Mie90]. data-determined [CM90]. data-driven [Fan99]. David [Bal90, BB93a]. Dawson [CS95a]. death [Fad98, GLM94, KS92, Kyr94, PT96]. decaying [ME92]. decide [KZ91]. decision [BMvdM97, McC90, ZK93]. decomposability [BLBL92]. decomposition [Mac96, Par96, ST92a, UU95, Wan91a, Wan91b]. deconvolution [Pat96, Ste90, vEK98]. deconvolving [Wan98a]. decoupling [Due98, Hit90]. decreasing [AL98, Als96, EZ99, Imh97]. decrement [NS96a, NS97a]. deepest [Nol99]. defining [BM97]. definitions [LDG97]. deformation [PS99a]. Degenerate [God90a]. degrees [Zha92a]. Dekkers [Tob96]. deleted [RR94b, Rot97]. Deletion [Jen96, DM98]. Demon [DG97]. densities [BD99, Coo93b, CAF97, FN95, HLL96, HL92, KK96b, MX96, NS98, PP96b, Spi91, SZ98, Vid96, vRR98]. Density [Eng90, Kat99, KP92, KP93a, Mül97, Wal99, Ahm92, Als96, AB92b, Bar92a, Bas96a, BB97, Bos95, Bru92, CTW97, CW94, Che98d, CF98, ĆM93b, DK99, DMY91, Dia90, DJ95, Far90, Fer98, Fut99, GD97, Giv95, Gov97, Hal90, HW95, HMT96, Haz98, HJR93, Ho95a, Jon90, Jon91, Jon92, Kal98, Kan93b, KM90, KPM94, KC96, KT98, KK96b, Kul95, Kut97a, Leb96, MO99, MS90, Mas94, McK96, MS96b, Mie90, Mie97, Moh92, Olk98, PC91, Pat96, PG99, Rao95, Rhe99, She94, SH99, Tra90, Vie96, Wan92a, Wan98a, WW95c, Wu97b, vE97]. density-weighted [Rao95]. denumerable [Led95]. dependence [Bae97, Bir92, BK96, Bra96a, DS94, Hüs94, KC97a, KC96, Kra98, Lah93, Mud93, Mud94, Nel92, OS98, Par98b, Pen99, SB97b, Tak94, Tra90, Wan92a, Zam90, Zha96]. dependences [GC92, Liu94b]. Dependency [Sto95, YSM94, Zie90]. Dependent [Sai95, Bal94, BS97, Bo98, BD90b, BD92b, Cai98, CG92, CGT97, Che97b, CL98, CM99, Coo93a, ĆM93b, DR96a, FM94, GM99a, Gen98, Gha94, HL97, Ho95a, HH98b, JWR92, KL98a, Kou92a, Lee98b, LV99, LSR90, Lin93, Mas94, Mat92, Mie90, Mie97, Muk94, ND93, PS95a, Ryc92, Ryc95, SS96a, SC91b, SK93, WW90a, Wan99d, ZB94]. depending [BN96, SC93]. depth [Düm92]. derivation [Bal93b, Jia97a, RMV96]. Derivations [WJ93, Cav97]. derivative [Rao95, Ryc90, SS96b]. derivatives [HW95, HMT96, KM90, PC91]. derived [Sho97]. design [AGV97, CB99, Cha98c, CW98, Imh97, Ioa92, Kag91, Kor99, LN95, Mül97, Rod93, RHO99, Rou93, Wan91a, Wan91b]. design-unbiased [Rod93]. designed [But99]. Designs [DGA95, Wie91, Afs90a, Afs90b, AS96, AK99, Akr93, BD91, Bis96b, Brz91, Brz95, Brz96, Brz97, BCK97, Cha96a, Cha99c, Che98b, CJ98, Das98, DDD98, DW99, DK93, GDD91, Hic99, Hud97, Jen94b, Joh96, Kag90, Kan91a, Kan91b, Kou95, Kou96a, Kou96b, Kou96d, Kou96c, Kou96e, Kou96f, Koz91, Koz96, Man99, MVA93, MZ97, MM95, MH90, Poh92, PS95b, Ras95, Sch97c, Sia94, SJK96, SG96a, Sri98, SW92, TLC97, Tom92a, Vie99, WJ99, Won93b, Won95, YL99, Zha92b]. Desirable [Dav94]. Detecting [Wat95, GPI99, YSM94]. detection [BCQY97, CT94, Fra91, LX97]. determinant [Pro98, WH91]. determination [Cha99a, PT98, Sha92a]. determine [CL94, KM98]. determined [CM90, CH98b]. deterministic [AC93, CPA95]. detour [PS95b]. deviation [BP99a, BS93, Bry93, FP99, MN93, PGTD92, Wen97]. deviations [AA98a, AA98b, BL99a, Bis99, BD90b, BD92b, Che97a, Che97b, Dar97, Dar98, DZ96, Eic96, Eic98, GRZ99, HT90, Iof93, JWR92, KT98, Löw96, Lug95, Mik97, Tab94a, Tab94b, Tud98, Zaj96]. Devroye [Kal98]. DFR [PS97b]. diagnostic [Fun92, Gei98, HT95, PZ97, YC90]. diagnostics [Shi93, Wei90, WS90, WL93]. diagonal [LR96]. diallel [DDD98]. diameter [HkL92]. diatomic [Hür93]. dichotomy [HS94a]. Dickey [MR99]. difference [Hud97, PR98b, Poo91, Wan96a]. differences [BDS99, Qin97]. different [Yat96]. differentiability [Che92b]. Differential [Der95, Bis99, GA99, Kim99c, LS97, Mat97, RA93, Rut90, She93, Stu98]. Differentiation [RMNV99, Pér93]. diffuse [AV97]. Diffusion [CCHP99, DW92, GCL90, Hof97, Kut94, Kut97b, Kut97a, Sal93, Sho97, Str99, Zha94b]. Diffusions [Gzy90, Cha90d]. dilation [BPRS97]. dimension [AC99, SN96, Xia97]. dimensional [Bal95, CG96, CH98b, Gut97, Kim99c, Kul92, Lef96, MH96, Stu98]. dimensionality [Bes92]. dimensions [Hee91, MPS97, Wir96]. Dirac [PP96b]. Dirac-function [PP96b]. direct [WW96]. directional [Fer96, JS98, Liu97]. Dirichlet [Gug98, Jia91, LM98, Lo91, Maj92, PR97, SW90b, Yam90, Zha97b]. discontinuous [Str99]. discrepancies [Hic99]. Discrete [CS98b, PJ95, Ahm91, BHB96, BSW94, CG96, CG99, Cse93, Cse95, Dev93, HK95, Hua91, KJ91, Kru98, Kru99b, KV96, Leb96, Lia99, Lor95, Mas96, NS97a, Olo99, PG94b, Qi97, RG95, Rei92, Rom99, RG99, Sap95, Sho97, TV98, Wal98, Yu93]. discrete-parameter [Cse93, Cse95]. discrete-time [Leb96]. discriminant [Cha90b, Cha92c, DS91, Fun92, Krz94, Leu96, O’N92, Zha94a]. discrimination [CAF97, Kou92b, O’N94, PG94a]. disease [Yip92]. diseases [Ber99b]. disk [Jia91]. disparity [BS94c, BS94b]. Dispersion [Bur93, BPRS97, BZ99, GM99b, GW95a, GP92, KW92, MS97a, MPSG90, RW94, Zha92b]. Dispersive [BCR96, Koc96, KM99, MM94, MP90, Aly90, BB90, Bar92b, Bar98b, Fan99, MM91, PS97b, RH91]. Dispersivity [AP92]. dissimilarity [Zog94]. distance [BHB96, Ham95, Kan93a, Kan93b, Kim96, Kou92b, Muk94, ÖH96, Pak96, Qia96a, SNN97]. distances [CAF97, CF98, EH99a]. distinct [Cha98a]. distinguished [NS96b]. Distinguishing [How97]. distributed [BK93, CLN95, Mal96, Ros93, SMP95]. Distribution [Jia91, SS96a, Sch91a, Yu93, ADEM94, AL98, AdlC93, AM97, ANS93, Ame90, And92, AEE92, AVB95, ACS94a, BS94a, BSW97, BAC92, Bal93a, BHK^{+}99, Bar92a, BH92, BR90, BA93, BR99, BR91b, Bho92, BS94d, BK96, CG92, Cha94, Cha90c, Cha92b, Cha96c, Cha98c, Cha90d, Che92a, Che97c, CB96, Chu96, Chu92b, Coh90, CS94, Deg93, DP94, Dev90, DY98, Don91, dE93, dE97, Ehm91, EvZ92, Elf97, Ext91, Fer97, Fro95, Gen99, GK99, GS99, GH96, GH97, GS93a, GP96, Gui96, HK91, HW90, HL97, Har91, Haw90, Hen98a, Hor93, Hos94, HT91, Hua97b, Hür95, Hwa99, IU95, Jev90, Jev91, JS99, JV91, Jon90, KP96a, KO96, Kou97, Koz98, Kry99, LT92, LH96, Lia99, Lov98, Luc96, Ma99, MS97a, Mac96, MM90, Maj99, MZM94, McK91]. distribution [MPSG90, NS92, NS96b, NS96c, Ole97, Pak96, PL95, PFvR98, PW97, Par92, PGTD92, RA93, Rob90b, Rob95a, RS91, Rou91b, Rou95, RKD90, RA92, Run93, ST92a, Sal93, SS96b, Sap99, Sco93, SB97b, Sgi96, Sgi98b, SX97, SY95, SW90b, Ste91a, SS97d, SS98, SB95b, Szr96, Tan94, Tan95, TH95, Tsa95, TYS92, WW95a, WJ93, Wan94b, Wan96b, WH91, Won93a, YFL96, Yin99, YSM94, Zim94, dHP99]. distribution-free [Cha94, GH96, GH97]. Distributional [Arc94b, dUÁGM98, El 95, Wes97]. Distributions [Wu92, AA92, AM94a, Abt94, Ahm95, Ala93, Ala94, Alb98, ACS93, ACS96, AFM96, BMT97, BJK98, BB95a, Bar92b, Bar98a, Bar98b, Bél91a, BC96, BD92a, Bil96, Bis96a, Bla91, BDG93, Buc94, CH91, CH98a, CDG93, Coo93b, Deh91, Dev93, DvC95, Don98, DLY96, Dzi95, EJ96, EM99, Fal94, FB91, FPMP97, FLW99, Gne99, God90b, HJM98, Ham95, HA98, Hür93, IZ94, Jen94a, JdKP95, Joe90, JK95a, JEG99, JK99a, Kal90, KK99, Kir96, Kir98, Koc99, KJ91, Kre96, Kru98, KV96, Lai95, Lef97, Leo96, LM98, Li93, Li94c, LF95b, LF95a, LB98, LL97, Lin90, Liu99a, MRR92, MQ91, MM94, Mi96, Mol93, MM98b, Mus96, Nad99, Nag99, Nay98, Nol98, PD95, Pap90, PD99, Per98, PG94c, PA90, PAT90, PJ95, Poo91, Pur98, Rao90]. distributions [Rao99b, RQ99, Rob91, RW94, Ryc92, Ryc95, SB97a, Sap95, SS90a, SJ97, Ses91, Ses92, Szr92, Tan93a, Tar94, TG95, Tsu92, WM97, Wes97, Wir96, Wu97a, YM91, YB99, YM95, Zha94c, ZK97b, Zog94]. disturbance [WK99]. disturbances [Sin94]. divergence [Cav99, CS98c, DB94, TW92a, ZW97b]. divergences [MMPZ99]. divisibility [BKS96, GMS94]. divisible [BLB90, DK93]. divisor [Kim99a]. DNA [Hol90]. DNAs [Sai99]. do [CL94, Lah98]. does [DS92, LL93]. Doléans [DM95]. domain [Gel96, Rév96, Sep96]. domains [CF96b, Jak94, Mee91, Mee93b, MS97b, Sub94]. dominance [ST98b]. dominated [Mac98]. dose [ZRB98]. double [CSC93, Rig95]. double-integral [Rig95]. doubly [Elf97, KPTY97, Lee97b]. drawing [Mil92]. Drawings [Hen98a]. drift [Rut90, Sgi97, Str96, Yin99]. driven [Fan99, Le 98]. drug [Tan95]. dual [Poh92, Ves94]. duality [ED94, ED96, SF94]. Dubins [Isa99, Kha92]. due [CH99]. Dümbgen [Fer94]. duration [Bac97]. dynamic [Pit98, ZW99]. dynamical [BM93, BG95].

E-optimal [DW99]. each [KP94a]. ECME [Mkh98]. EDG [GS93b]. EDG/3 [GS93b]. edge [JS96, PR98a, PRB99, Pal01]. edge-transitive [PRB99, Pal01]. Edgeworth [SS98, Bab91, FJ92, Kol95, Mon93, SS97d]. Editorial [Ano90d, Ano91a, Ano91b, Ano92d, Ano92e, Ano93d, Ano93e, Ano93f, Ano94d, Ano94e, Ano94f, Ano95e, Ano95f, Ano95g, Ano95h, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano98f, Ano98g, Ano98h, Ano98i, Ano99c, Joh94b, Joh98]. education [Dal90]. Effect [CMC94, BW93, Dav97, PP95, SG96b]. effectiveness [Mur92]. effects [BL91, CR91, Das92, MVA93, Omo97, SQC94, SW90a, SP98, Sut97, Wu97b]. efficacy [Kou90]. Efficiencies [Jen94b, Bar95a, CMC94]. Efficiency [ABY94, Jen93, Baj92, Brz96, But99, Cha90c, Das98, Dav94, Dor97a, Gru91, Gui96, He94a, He94b, HP93a, KMC98, Koz96, Lia95, Mor94a, Pak96, SG96a, SS90b, Tso98]. efficiency-balanced [Das98]. Efficient [Cro94, Puk97, Sch99b, Yan90, ZR96, Mof97, ÖH96, Sch96a, Sch97c]. eigenvalue [Düm95]. eigenvalues [Jin93]. Einmahl [Tob96]. element [SC93]. elements [ARV97, Due98]. elicitation [PGTD92]. elimination [Brz95, Cha96a, Sia94]. ellipsoid [CHW93]. elliptical [Chu92b, Kou92b, Nay98, Rao90, Zha94c]. elliptically [Gen99, GV91, LF95b]. embedding [Bél91b]. Empirical [BM96, DLS90, Leo93, NNR97, SB96, WJ99, Zha97a, BP99a, BDG98, BH92, BV92, CMS97, CH98a, Cro98, Dar97, Deg93, DP94, EvZ92, FR92, FR95, GKL96, GS99, Ham98a, Ham98b, HL97, Hor90, HS99, Hua97b, IM97a, Jia99, Jin95, Kni91, Kri90, KZ91, Lah98, Lat98, Lee96b, Lee97c, Lia99, LY99, MHL95, OS98, SV99, Shi97a, Szy94, Wan94b, ZK97a, Yu93]. empiricals [Yan96b]. empty [TLC97]. endpoints [JR99]. energy [Gom96]. engineering [FL91]. ensembles [Jen97]. entropy [EPS94, MMP97, PL95, Par99, Rom99, Stu98, WW95c]. entry [CY99]. environment [BL94, Li97]. epidemic [CS92, Rei92, Tan93a]. equal [dE93, Kre96, O’N92, SC90]. equality [AC91, ABK98, Del93, Gro97, KHW91, KW98, KS95b, NNR97, NS96c, RW94, Sch96d, Sch97b]. equation [BW99, FRS93, GA99, Kar96, LPB90, PL92, PL93, RZ92a, Rig95, RT93, Uno96]. equations [BT91, Bis99, Cho97, Gat93, Hey94, KMC98, Kim99c, LS97, Mat97, Rut90, She93, Stu98, Tud98]. equilibrium [Phe95]. equitable [Isa99]. Equivalence [BB93a, BB94, MS92, Sgi98a, CM95, ES96, GP99, IK92, LZ94, MM91, Mun99a, Won95]. Equivalences [GKS97]. Equivalent [Bra99a, Mik98, Per98]. equivariant [Jey91, KP94a, Nay98, Zie99]. Erdos [BB91a, Wan99c]. Ergodic [LW92, CH98b, Kut97b, Kut97a, Nob95]. ergodicity [ACH97, AC97a, BL95, BL99b, Lu96, Tan96]. Erlang [Luc96, Mil99]. Erratum [Ano96l, Ano99d, Ano99e, BL99b, Ham98a, Lin00, LGC99a, Rad99]. error [Abr93, Alt93, Bal97, Bal98, BKR92, BAV98, Bos95, BS91, CER93, Che93, CH94b, Elt92, Fal90a, Hal95, HJM98, Hua99, Hug96, HS94b, Ioa92, JH90, Kul95, LH98a, Mie97, NG95, PS98, Pog92, Pou93, Rog95, Sar96, SMP95, SS97c, SR97, TW92c, WK99, Wan99a, Wan99b, ZK93, ZL97, vE97]. error-in-variables [ZL97]. errors [Akr96b, BS94b, Cha99b, CD94, CR91, Deo97, FNARP97, GKS96b, Gle93, HkL92, IA97, KL98b, Kou92a, Lee93, Liu97, Muk94, SB96, SS99a, SP98, Tsa98, Zha92a, ZK97a]. errors-in-variables [Akr96b, BS94b, HkL92, Tsa98]. errors-invariables [Gle93]. escalation [ZRB98]. Esseen [Che98d, RZ94, Wan99d]. Esseen-type [Che98d]. Essential [Mór92]. estimability [Nee91]. estimate [BD94, CR97, cIC99, ĆM93b, Dev94, DK99, DM93, Fal90a, Fer98, Fut99, GT99, GW97, Hou97, Ind94, KP94b, Ral92, RZ92a, Rou95, Rut93, She98]. estimated [AD99, Har91, Hud97, NM97, WJT92, Zog94]. estimates [ABY94, BW97, BZ99, CW94, CL98, Che92b, DL91a, FNARP97, Gao95, JH90, Kan91a, Kan91b, KM90, LSZ98, LZ91, Mar99a, MH96, Por92, Rie97, Rob95a, Rou91a, SD99, Sch96c, Sun95, WW95a]. Estimating [Alt93, BS95b, Hua91, Lee93, Lee95, Loh97, MM90, Mar91, Mar99b, Phe90, YF93, dHP99, Ani94, BD99, Fro95, Hey94, HT95, Hsi95, Kag90, KMC98, KS96, Nay93, RZ92a, SZ92, Sut97, Wan98b, WPH96]. Estimation [AD99, BR91a, CPPP98, DJ95, Elf97, Fuj96, GV92, HL92, Hus96, KD92, Kul92, KM93, KL98c, LF95b, Liu96, MO99, MSS90, PC91, Pen99, QR92, RP95, SB97a, Sch94, SQC94, Sin94, SR97, Wan96b, ZW99, AS95, Ahm92, Ahm95, Akr96a, AM93, AV97, AB92b, BG99a, BS94b, BR90, Ber99a, BKR96, BG95, BS96b, CR92, CB99, CTW97, Cha98b, CG94, CS93a, CW94, CD94, CH99, Cro92, CF98, Das92, DK90, DF95, DS91, DGS99, Dia90, DJ97, DMLP97, DT98, DL91b, DLY96, Efr94, EH99b, ED94, ED96, EJ96, Eng90, Fal94, FG92, Far90, FM92, Fer96, FP99, FR95, Fox97, FP98, Gal98, Gar98, Gas94, GD97, Gro98, Gu98, GL95b, HM90, Haz98, Hil99, Ho95a, Hof97, Hof99, HK95, Hu97, Hua97b, HB99a, HP93b, Hwa99, IU95, IZ94, JS96]. estimation [Jin93, JS99, Joh91, Jon90, Jon91, Jon92, JS90, Kak97, KD93, KS95a, Kat97, Kat99, KMC98, KP92, KP93a, KPM94, Kim99c, KA99, KZ96, Koh99, KK96b, KL98b, KP94a, Kuk99, Kut94, Kut97b, Kut97a, Le 98, Leo99, LW92, Li95, Mag90, Mar94, MHL95, Mas94, McK96, Mee93a, Mie90, Mil99, MvdM97, MN97b, Muk94, Nag99, Nay98, NS96a, NS96b, NS97a, O’N94, ÖH96, Pak96, PL95, Pap95, PP95, Par98a, Pat96, Ped97, PS96b, Phe95, QT90, RG95, Rao92, Rao95, RQ99, RCBG96, RCD93, RS91, Rou91b, RMV96, Ruy96, Ryc90, SV91, Sch99a, Sch96b, Sch96a, Sch99b, Sém96, SC90, She94, SHC99, SA97, SG91, SS90b, SS97b, SNN97, TP96, TW92a, Tra90, Uno95, VNS96, Vel93b, Vie99, Vie96, Wal99, WL95, Wan99a, Web92, Wie96, WZ99, Wu97b]. estimation [Yan90, Yan96a, YLB99, ZK97a, dC93]. Estimations [UU95]. estimator [AM98, AT96, Ang96, AEE92, BSW97, Ber94, BR92, Bis99, BG91, Cai98, CSC93, CH94a, Cha96c, CY99, Che95a, CC98b, CC94c, CHW93, DMG90, Deh91, Fer97, FS93, FZ92, GL95a, GS95, GS93a, Giv95, Gla98, Gro97, HKR90, Hub97, IM97b, KS93, Kan93b, KR98, Kul95, Kun93, Leb96, LP91a, MP99, Mar97b, MZ91, Mie97, Oht96, Oht99, PS99b, Pit98, PG99, Pru92, Ser96a, SB97b, SX97, Sho97, SH99, Sri91, SS90b, SS94b, Stu94, Tab94a, Tob96, Tza98, WK99, WC90, Wan98b, Wan99a, Wan99b, Wil98, Yan94, Yan96b, Yua96, Zho96, Zie99, vE97]. Estimators [Cas90, Gle93, HW95, Arc94b, Arc98b, Bas94, Ber90, BKR92, Bis99, BB97, Bly94, Bos95, Cha99b, Che98d, CMC94, Coh90, Cro94, De 96a, ES91, Fer94, FK94, GG91, Gas94, Gen98, GM99b, GS96b, Gho95, GKS96b, Gor91, Gru91, Gu95, Gui96, GP91, Hal90, Hal95, HJR93, Hoa93, Hös92, IK92, Ioa92, JK99b, Kak98, Kal98, KMC98, KC96, KZ96, Kou92a, Kus98, Lec90, Lee96a, LB91a, Lie99, LY99, MS97a, Mad93, Mam90, Mar96, Mar98, Men93, Mey99, MM95, Muk92a, MS97c, PKM93, PKJ97, PS97a, Pen98, Qia96a, Ras95, RC94, Rou94, RKD90, RA92, Ruk96, SV99, San91, Sar96, Sha91, Sha92b, SS97a, SMP95, SS93b, SS94b, Ste90, Ste91b, Sun97b, Tab94b, Van93, Wan92a, Wan98a, WW92a, Wef91, WW90b, Xia95, ZN95, ZL97, vEK98, vRR98]. Euclidean [Gzy90, HJR93, Neu96]. Euler [BA93, Sho97]. Eulerian [FLW99, HP94]. evaluation [Leo93, MZM94]. event [BM96]. events [BB93b, BM97, SB91, Sto95]. everywhere [BR95]. everywhere-positive [BR95]. Exact [Baj92, Cha96c, Cha99c, CM93a, CHW93, DMWZ95, KM97, PS99b, CC98a, FLW99, GS93a, hK99, Mor94a, Szr96, Wri97]. Exactly [PW92, BB93b]. example [BW95]. Examples [BM93, Cha98d, Iof93, Jev90]. exceedances [Wiś96]. exchangeability [BS95a, PR97, vP91]. exchangeable [BB94, BR97, Dar97, Dar98, De 96a, EK96b, Ete99, Gne95, Gne96, Lau92, Leo96, MP93, SV93, Sch91a]. excursion [LMF97]. exist [DS92]. Existence [AEE92, Bas96b, BT91, Che92b, Fie90, HLL96, HkL92, MH96, ST90, ACH97, BES95, JW97, LSZ98, LM92, PRS92]. exp [Gel96]. expansion [BC92, BMS97, Hua99, Kat99, Mon93, RMNV99]. expansions [Bab91, BV92, Cek95, Cha90b, FJ92, Har96, Llo91, Pri99]. expectation [Beh99, BS92, GW93]. expectations [Arc98d, GC92, Gha96, Hua98]. Expected [CF99, GAP99, ASC98, Car92, CS96a, Dev91, Hua97a, Irl97, KP96a, Mol93, PT96, Pro98]. Expectiles [Jon94]. experiment [Ger96, Mac98, YF93]. experimental [Imh97, Kou96f, Tom92a]. experimentation [Sta95b]. experiments [But99, DDD98, GV92, Mac96, MGH99, Ste97, Tom93]. Explicit [KP96a, MV95a, Sgi98b, She93]. Exploiting [Mof97]. explosive [MM98b]. exponent [TT91]. Exponential [EG99, IR99, Sun97a, AB95, AM97, Ala93, Ala94, BLBL92, BR90, Bol98, Bor97a, CH91, Chr91, Coo93b, CC99b, Elf97, FC96, FCUOCN96, GM99a, Gei93, GH92, HK91, Han93, Hwa99, Jev90, Jev91, Joe90, Kak98, KM99, KS96, Lan98, LB91b, LR96, Mag90, MW91, MS90, MV95a, MH96, NS96b, PG94c, Rao99b, SC99, Ses92, Sha92a, Stu97, XY97, Yca92]. exponential-type [Mag90, MW91]. exponentially [CM99, Rig95]. exponentials [GO97]. expression [Neu94, ST98a]. expressions [Mus96]. extended [HL94, hK99]. extendibility [SV93]. Extension [Gaw97, AC97b, BB91a, Cor93b, GS96a, McC90, Pap93, Sib91, WZ96]. Extensions [Zho99, RB96, SX97]. extentions [Lor92]. extinction [Cha96b, GG91, KD92]. extrema [Her97]. Extremal [Mee99, Bin95, GC92, Leo99, Rav92, Ryc92]. Extreme [Fal94, Gom96, Hen97b, Bal90, BD99, Deh91, Dev91, Fal90b, Joe90, Mas99, Nad99, Pen98, Tak94, Wiś96]. extreme-order [Wiś96]. extreme-value [Deh91, Pen98]. Extremes [Alb93, Alb97, Hüs99, DH90, Gan95, Pen99]. Eyraud [Mik97].

F [LH96, Mi96]. Fabian [SE97]. Factions [Sha92a]. factor [DLW96, IK92, IK95, Koz96, Res96a]. Factorial [SW92, Che98b, CJ98, Ger96, Kou96c, Tom93]. factorials [AS96]. factorizable [Wes97]. factors [BX95, Kay92]. fail [EG99]. fails [BW95]. failure [AA92, AL98, Bha95, CL98, CDG93, Gas94, Gas96, Lee99, Ves94]. Families [Joe90, AB95, BLBL92, Bor97a, GH92, Gui96, LR96, MH96, Rot93a, Yca92]. family [BD92a, Coo93b, De 96b, FC96, FCUOCN96, FBCCN96, Lor95, PG94c, Ses91, Sha92a, Spi91]. Farlie [Mik97]. FBSDE [Wu99b]. feasible [SS94b, WK99, ZRB98]. features [Rou94]. Fenchel [ED94, ED96]. field [DP96, GR93b, Rou95]. fields [ARM97, CTW97, Car94, MS96a, Pog92, PR98b, Ste93]. filter [LG93, NG95, RMV96]. Filtering [Le 98]. filtration [Ouk94]. filtrations [Men91]. final [ZK93]. finding [LM96]. Finite [BS95a, Wan92a, Wan98a, vP91, ABY94, BS94d, CMC94, Dai98, HH98a, Hor90, HkL92, KD96, Kul97, Lou97, MMPZ99, NOT90, Pen96, RZ94, Sen93, Spr98, Szr96, Wen94, Wri97]. finite-population [RZ94]. finite-sample [NOT90, Szr96]. finiteness [Haw97]. First [Lef97, Lin98, Als91b, AM93, BES95, CFF98, Cek95, DR96b, GCL90, Gut96, HK96, Kak98, Lef96, POGP96, TW92c, Yam90]. first-order [AM93, CFF98, Kak98, TW92c]. first-passage [DR96b, Lef96, POGP96]. Fisher [CS96a, Ger95, GK99, Jan97, KL97, KL99, hK99, Lin91, Llo91, O’N92, Wu92]. fit [Ahm97, ACGM99, BH92, BS94c, BR99, Che94, CM93a, CS91, DZ99, GH92, Hic99, JW99, JPZ97, Lor95, Mor94a, Par99, RM90, Rea93, ST91, WS90, WJT92]. fitting [AB92b, Geo90, Mül99, NS91]. Fixed [MX96, Dre99, Gov97, Ioa92, KA97, SN96, SP98, WJ99, Wie96]. fixed-length [Dre99]. fixed-width [Gov97]. flat [LL93]. FLIL [EN96]. flows [Bax97]. fluid [PR96]. fold [KG90b]. Ford [RHO99]. forecasting [DD98]. forecasts [GR91]. form [CC94a, Gas96, GV91, KK96a, Lin98, TS93, Zha92b]. formation [NJS98]. forms [GRZ99, Gha94, Gha96, GK98, Jen97, KW96, Zha94c]. formula [BP99a, BW95, CC94a, FLP99, GS96a, Gha96, LM98, LL97, Liu94a, Pér93, Rut93, Sgi98b]. Formulae [HA98, Cor93a, God90b, KP96a]. formulas [Li96]. foundation [Aic98]. four [Cha95]. Fourier [Ang96, HL92, Kni91, Kul97, PG94b, vRR98]. Fourth [MH90, DS92]. Fourth-order [MH90]. fractal [Whi99]. Fractional [BG99b, Che98b, CJ98, EN93, EN99, GA99, Le 98, NV99, VDO94, Xia97]. Fréchet [Che92b]. free [Cha94, DR96b, GH96, GH97]. Freedman [Isa99]. freedom [Zha92a]. Freeman [Rea93]. frequencies [SG96b]. Frequency [CGT97, Lin92b, DS96, ES90, Har91, Kat97]. Frequentist [SE97, De 96a]. Full [LPB90, Dro99, IK95]. Fuller [MR99]. fully [Gas96, ÖH96]. function [AB95, Ani94, BR91a, BP99a, Bar92a, BH92, BCI98, BS94d, BG95, Buc94, Cha92c, Che95a, CW98, CC94c, Chu92c, Cro98, DM95, DeR94, DeR96, DJ95, Elf97, EPB94, FLP99, Fox97, Gen99, Geo90, Gne98b, GV92, GR99, Gu95, Imh97, Ioa92, JS99, Jon90, Kai99, KC95, Lec90, Lin92a, Loh97, MRN93, MHL95, MZM94, Mar97b, MPSG90, MP90, NS96a, NS97a, NY90, NS91, Nor95b, O’N92, PL95, PP96b, cQ94, Rou91b, Rou95, SX97, Sze90, Wan94b, Wan98b, Wan99b, Wil96, Woo95, ZZ93, ZN95, ZW97b, Yu93]. Functional [CFS99, MP99, Als96, BS93, Dor97b, FZL97, FRS93, GRZ99, GT99, GW93, Jia91, KM97, Tob96]. functionals [Ahm97, BMS97, Coh90, CF95, Efr94, Gal98, GGR98, GR93a, Hor93, Kro97, KL98c, LL95, MO99, Ouk92, Pri99, Stu97, Yan94, Zha97a, dN95]. functions [ANS93, Arc94a, Bab91, BLB90, BlZ93, BH97, BH99, CG92, CM95, Chu92a, DS96, Deg93, DP94, DGS99, DMY91, EK96a, Eic96, EvZ92, FH92, HL92, HS94b, Jon90, Kla93, Koh99, KL98b, Kra98, Kru98, Krz94, LH96, Li93, Li94c, LVC90, MS90, Moc94, NN96, Nob95, Par98a, Ram96, Ram97, Rob90a, SA93, SS96b, SS92, Tab94b, Zha97b]. Further [BS98, Kou96a, Oht96, BB93a, Cha98d]. future [JEG99]. fuzzy [Gil92, LDG97].

Galton [Ber99a, Rao92]. gambler [Bos96, RS99]. gambling [WZ96]. game [BCC99, Cha95]. games [CS96b]. gamma [AB92a, Ala93, Ala94, IU95, MZ91, SC99, Sim92, SW90b, Won93a, YM91]. Gauss [BKR96, Gro97]. Gaussian [AM97, AM93, Ani94, Arc92, BCI98, BR91b, BH99, Cha90c, DL91a, Deb99, DMWZ95, Fer99, GRZ99, Gen98, Hor93, HKR90, HCJ97, Hüs99, Kou97, LSZ98, Liu99a, MS97a, MS96a, MS95, Nik99, Rec91, Ses92, TT91, Vit99, Wan97, Wiś96, Zha94a, dN95]. Gaussian-inverted [LSZ98]. GCV [Shi99, WW95a]. GEM [SY95]. General [Cor93a, Ses92, Wan92c, Bar98a, CS92, FS99, GKS96a, Gro97, HK96, JdKP95, Khu97, Mar96, Por92, Rei92, Rou97, Sil96b]. generalised [Ext91, Lai95, Sém96]. generalization [AM97, Cha98a, Chu96, Fro95, HP94, Lin94, Lin92b, MW93, PP96a, SB94, Sre95, Tsa95, Wef91]. Generalized [BR98, FL98, HMG98, HJ93, Rob91, TV98, Web91, ARM97, ACS93, BA93, Bha96, Cha96a, Cha94, CWM91, cIC99, Cha98d, Che98a, CS96b, Fal94, FM92, FPMP97, GHW92, Gug98, HW90, Hen98b, HB97, IL98, Jan97, KA99, Kri90, LZ97, Lia95, Lin98, Luc96, MQ91, Mee93b, MS97b, Pak98, PA90, PL92, Sep92, Sep96, Sha91, SPH91, SMP95, Tom90, TV97, WC90, WL95, Wie96]. generally [Gas92]. generated [Adk93, CDG93, FZL97, Lee97d]. generating [AS96, BH92, Kru98]. generation [LSR90]. generator [MZT90]. generators [PP92b]. genetic [LP97]. Geodesic [De 96b]. Geometric [Hol98, ACH97, BL95, BL99b, Hen98a, Lin90, Lu96, SB95b, WL95, Wen94]. geometrically [CLN95]. geometrics [SB99]. geometry [ZW97b]. Ghosh [Lah92]. Gibbs [HRG97, Mat93, QT90]. given [Arc98d, BS94a, Bas96b, Gro98, MRR92, Ryc95, Str94]. Glivenko [AGMCA96b, BR97, EvZ92, Kri90]. Global [BZ99, Jin96, BH96, GS95, Kru99a]. GLRT [Wei92]. GM [Cha99b]. GM-estimators [Cha99b]. GMLE [YSLW98]. Goodness [Ahm97, ACGM99, BR99, DZ99, Hic99, BH92, BS94c, Che94, CS91, GH92, JW99, JPZ97, LZ97, Lor95, Par99, RM90, Rea93, ST91]. Goodness-of-fit [Ahm97, ACGM99, DZ99, Hic99, BS94c, GH92, JW99, Lor95, Par99, Rea93, ST91]. goodness-of-link [LZ97]. got [WR98]. grade [ĆM93b]. Gradient [NS91]. Gram [Epp99]. graph [BCQY97]. graphical [TV98]. graphs [GAP99, Pal92, PR98a, PRB99, Pal01]. greatest [Bal93b, Raq97]. Green [GS99, Stu92]. Group [OB95, Dai98, DK93]. grouped [Akr96b, WZ99]. grouping [BW93]. groups [CCHP99, MV95b, Neu96]. growth [BG97, FNARP97, PFvR98, QR92]. Guess [Gas94]. guided [Gla98]. Gumbel [Mik97, Bal90, BAC92, BB93a, LB91b]. Györfi [Kal98].

Haan [Tob96]. Haar [Eng90]. Haenszel [BW93]. Hájeck [LGC99b]. Hájek [LGC99a, Shi97b]. half [NOT90]. half-samples [NOT90]. Handling [LG93]. Hanson [EN93]. Hardy [SPH91]. harmonic [Als91b]. harmonizable [Moc94]. Hartley [Bal90, BB93a]. Hastings [dC98, Hol98]. Hausdorff [CF98, vRR98]. having [GS93b, Omo97, Szr92]. Hazard [SS95, CC94c, Gu95, Jan94, LVC90, PG94c, SV91, Sun97b, Wan99b, Yan96b].

Hazards [O’N91, Cha96c, LSR90, MM98a, Tso98, Wei90]. heads [Mór94b]. Hellinger [BHB96, Kan93a, Kan93b, Pak96]. heredity [PS94]. Hessian [NS91]. heterogeneity [Brz95, Cha96a, Sia94]. heterogeneous [Dav93, EH99b, Oht99]. heteroscedastic [DZ99, PW92, Sha90, SP98, TW92c, Won95]. heteroscedasticity [Lu96, YC90]. heteroskedastic [Akr96a]. heteroskedasticity [CA97]. Hidden [AK99, DT98, Hug97, LX97, Liu99b, MW97, QT90, RCD93]. high [CMC94, Cro94, Hee91, KP93b, MPS97, Rou93, Ste91b]. high-breakdown [Cro94, Rou93, Ste91b]. Higher [Zha94a, FH92, Irl97]. Higher-order [Zha94a]. Hilbert [DK99, Gat93, Pér93, ZB94]. Hill [KR98]. histogram [Kan93a]. hit [Hen98a]. hitting [GCL90, GAP99, Lef97, PT96, Yin99]. HIV [Tan93a, Tan94, Tan95]. Hjort [Wal98]. Hodges [Bar95a, KK92]. Hodgson [Gas94]. hold [PK96]. hold-out [PK96]. Hölder [Ham98a, Arc94a, Ham98b]. homogeneity [Cha90a, CS91, Dup97]. homogeneous [AHR91, BS97, MM96, Nor95a, Pog92]. homogeneously [Aue90]. horizon [Deb99]. Horvitz [Cor93b]. host [Yip92]. host-vector [Yip92]. hull [Mas99]. hulls [Mol93, Zar99]. hunting [Hec92]. hypercube [Loh97]. hypercubes [Hud97]. hyperfinite [Men91]. hypergeometric [BM96, SF94]. hypotheses [Bas96a, BHB96, BS95b, Hu95, Jan97, KS99a, PS94, RM90, Sil96b, Sil96a, Wei91]. Hypothesis [Kle94, TP98, Kak97, MR99, NS96c, RW92]. hysteresis [FP98].

i.i.d [Ete99, Jan97, SB95a, Woo92]. I.I.D. [CS98a]. ICF [Epp99]. identical [Liu96, UU95]. identically [BK93, Mal96, Ros93]. Identifiability [IK95, WGR94, Glo99, Hug96]. identifiable [KS95a]. identification [Cha99a]. identifies [PW97]. Identifying [BCI98]. Identities [BB91b, BM97, Lan96, Liu94a, Rao99b, Vit96]. identity [Cas90, El 95, Mof97, PKJ97]. if [Gan98]. IFR [CH91, PS97b]. IFRA [NS92]. Ignatov [Pek99]. ignorable [SQC94]. II [PRB99, Pal01, Afs90a, Akr93, CB96, JV91, KK99, MSS90, NS97a, PAT90, XY97]. III [Ram97]. Illustration [Lan96]. image [GR92, Kor99, Xia97]. imbedding [Lou97]. immigration [Kyr94, Sri91]. immigration-birth-death [Kyr94]. impact [HT96]. implementation [RCD93]. Implementing [Far90]. implications [Hu95, Mat93]. impossibility [BD99]. Imprecise [Coo93b, Coo98]. Improved [Das92, FCUOCN96, Gas94, KZ96, Lee97a, Lug95, PL95, Rie97, Sri95, TYS92, Kag91]. Improvement [MM95]. Improvements [KM90, CW97]. Improving [Bal90, Ber94, Kan91a, Kan91b, Oht96, SV99]. IMRL [Kir96, Kir98]. Inadmissibility [CS95b, DMG90, Ber90, HKR90]. including [GB96]. inclusion [Mil92]. incomplete [AT96, AC96, CP97, Gas96, Glo99, Ras95, SP93, SG96a]. Inconsistent [BKR92, Gho95, BM93, Pru92]. incorrect [Gu98]. increasing [AC99, Bel99b, Bha95, Bol98, DS96, FP94, FM99, Gan98, POGP96]. increments [EN99, Fro98, Gu95, HCJ97]. incubation [Tan93a, Tan94, Tan95]. Independence [YSM94, AL97a, AL91, AL92, BM97, Bow99, Fal97, Hee91, Hüs94, Kan92, Liu94b, Mór92, RW92, Tak94, VA97, WH93]. Independence-distribution-preserving [YSM94]. independent [ADEM94, Arc98a, BK93, Beh99, Ber97, Bho92, CLN95, CM91a, CM91b, FK95, GS96b, GO97, HH98b, JK95b, Kab99, Kor97, Lac95, Lee95, LH96, LT96, Mal96, Mar95, Mar97a, Pin94, RW92, Rot93b, SB91, Sco93, TYS95, dlPG92]. Index [Ano99f, Ano99g, Ano99h, Ano99i, AD99, Ano90b, Ano90a, Ano91c, Ano92c, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano94c, Ano94a, Ano94b, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano99b, Ber99b, Kou90, Mar99b, Pen98, Tob96, dHP99]. indexed [Arc94a, Chr90, DM95, Hon96, IM97a, Zha97b]. Indicator [BB92]. indicators [Mos94]. Indices [CMS97, BHZ99, EM99, FZF93, SHC99]. individual [BL91, DZ96, ST91]. induced [Bal93a, Hor95, KPTY97]. Inequalities [Kal90, KMS96, Bis99, BH97, BS98, Bos96, Chr91, DS94, Ete99, GS96a, GR91, Jen93, KS93, Kla93, Lee97a, Li94a, Lin92a, LP91b, NV99, PR96, Pit90, SC91b, TG95, TYS92]. inequality [Arc95b, BL99a, BH99, Bra96a, Chr94, Düm95, Gad97, Hit90, IR99, JDP90, Joh91, Kha92, LGC99a, LGC99b, Lou99, Pap93, Pin94, Qiy96, Shi97b, Sib91, TYS95]. inexactly [MZ91]. Inference [HB99b, KS99b, LB91b, RHO99, BC98, Ben97, BR91b, Bor97a, Che97c, CB96, Coo98, FM92, Han93, JS98, MN97a, MMP97, MMPZ99, PR97, Sev93, Sil97, Won93a, Yip92]. inferences [Sev98, SS99b]. Infinite [BKS96, Shi94, Deb99, GMS94, Kim99c, Kni91, KH91, Mee99, Run93, SS99a]. infinite-dimensional [Kim99c]. infinitely [BLB90]. infinitesimal [Neu96, Sze90]. Influence [Wei90, WS90, CF99, Cro98, DT98, Kyr94, LZ97, Sil96b, ST91, Tab94b, TW92c]. influential [Fra91]. information [AV97, BN96, CS96a, Cav97, Fer96, Ger95, GK99, Hug97, Kab98, KL97, KL99, Kan93a, Llo91, Pap93, Phe95, PV90, SG91, WW95b, WW95c, Wu92, WS98, Zha97a]. initial [DT98]. innovations [AD99, Rut95]. input [GA99]. inquiry [GS91]. Integer [Yca92, BC92, BKS96, HN97, Jes93, Mil99, SY95, WZ95, Wen97]. integer-valued [Jes93, WZ95, Wen97]. integers [DH97]. integral [Gaw97, GR99, HN98, Lee98b, Loh97, Pri99, Rie97, Rig95, Sav94, She98]. integrals [Ger91, Ger93, GM93, HW95, HS94a, Nik99, OW91, Pér93, SW90b]. Integrated [Ioa92, FRS93, FB96, GT99, Mie97, PC91, vE97]. Integration [ET91, Sav94]. integrative [San92]. integrators [MZ93]. intensity [Bal95, GW96b]. interacting [LS98]. interaction [GPI99]. interactions [Man99, Zel90]. interarrival [Sim92]. interblock [Zha92a]. intercept [Cha99c, Wat95]. interim [Gei93]. Intermediate [Pap95, Cha92a]. interpolated [Nyb92]. interpolation [CP97]. interpretation [SS94b]. interquartiles [Bin96]. intersection [AHZ97]. Interval [Cha98b, Als91a, BJZ90, Huw95, HP93b, JK95a, Lee99, PCK98, Ped97, PL93, WGR94, WC97, Wan98b, YSLW98]. interval-censored [PCK98, YSLW98]. intervals [AB92b, Bax93, Cha94, Dre99, FTV98, Fis91, Gil92, GH96, GH97, GR91, Gov97, Hei93, Jey92, Kab99, PP92a, Qin97, Ram96, Rhe99, Sta92, Tsa98, Wan98c, Zha97a]. Intervention [SK93]. intraclass [CC98b]. Intrinsic [FR95, PP92b]. invariables [Gle93]. Invariance [PR98b, DMY91, HZ99, KW97, Rob95b, RZ92b]. Invariant [Che92a, Las97, Wat93, BG95, GS95, HLL96, Löw97, Tak91, Yu93, Zim94]. inverse [AM97, BV92, BR91b, Cha90c, GO99, HKR90, Kou97, Lin98, MS97a, MZM94, MW90, Olk98, Rob91, Ruy96, San92, Ses92, O’N91]. inversion [Shi99]. inverted [LSZ98]. involving [FRS93]. irreducibility [CmHP98]. irregular [Hof97, vRR98]. irregularly [SS94a]. ISE [Jon91]. Isotonic [HM90, WZ99, CC94c, DL91b, Qia92, Qia96b]. isotonicity [Par98a]. Itô [Tud98]. Iterated [Ber96, Löw96, TV97, AK95, CF97, Dor97b, EN93, GKL96, Gao95, Hen97a, Hou97, JK95a, LT96, Neu98, Nor95b, Shi95a, Tob96, Tom90, Tom92b, VD91]. Iterative [Sun95, CHW93, WK99]. IV [Nag99].

jack [Gru91]. jack-knifed [Gru91]. Jackknife [Sha91, Cha98d, Hou97]. jackknifing [She98]. James [Ber94, GP91]. Jensen [Pit90]. Jensen-type [Pit90]. Joint [Chu96, BL91, Cha94, DY98, HA98, LV99, NS98, Rao99a, ST90, Yin99]. jump [ET91, Pri99]. jumps [Bor97b].

Kalman [LG93, NG95, Rut93]. Kanazawa [Jon95]. Kantorovich [Bel99a, CATD93]. Kaplan [Cai98, Cha96c, SD99, She98, Stu94, Yan94]. kappa [CC98b]. kappa-type [CC98b]. Karhunen [RMNV99]. Katz [Lan98]. Kendall [dE93, Shi98]. Kernel [CTW97, FNARP97, LY99, PS97a, Rou91a, She94, Tra90, Ahm95, AL97a, Bly94, ĆM93b, DR96a, DK99, FH92, Giv95, Haz98, JS99, JH90, Jon90, Jon92, Kan93b, KM90, KMC98, KP93a, KPM94, KC96, Mam90, Mie97, Mül97, Rie97, SX97, SH99, Wu97b, vE97, vEK98]. kernels [Chr94, Sch97a]. Kiefer [BV92, Cha92a, Cha99b, DS92, Kep97, Lah92]. Kiefer-type [BV92]. kills [KV99]. kind [Wen97, Yam90]. kinds [Hen97a]. knifed [Gru91]. known [GG91, Gle93, KM90, Mar94]. Kolmogorov [Cek95, CA97, Chr94, JPZ97, Kim99b, Kri93, Par92, ST98b, Ser96b, She93, Tom92b, TYS92, TYS95, Vas90, Wan99c, ZC95]. Kolmogorov-type [ST98b]. Koul [FZ92, Zho99]. Koziol [GS99, Stu92]. Kramer [Som93]. Kronecker [Brz97]. Kullback [Cav99, Kan93b]. kurtosis [AM94b, TG95, ZK97b].

L [GR99]. L-superadditive [GR99]. lack [Sha92a]. Lagrange [GW96a]. Laplace [Bar98b, Bar99, BV92, BOR99, CB96, JN97]. Laplacian [BG99b]. Large [AGMCA96a, BB91a, Bis99, BD90b, BD92b, Che97a, Chr90, Chr97, Csö92, CS96b, Dar97, Dar98, DZ96, Eic98, EK96b, Ete97, FP99, Gut92, HK91, HO95b, Hon96, Hor90, JK98, Mar95, Mik97, Pet96a, Rem98, SM92, Sun98, Sun99, Tud98, Zaj96, dJ96, Arc98a, AGMCA96b, BMT97, BP99a, Bir92, BB95b, BS93, Bry93, Cav99, Eic96, EN99, Fro98, FL91, GRZ99, Iof93, JV91, KT98, Löw96, TW92b, Tom99]. large-sample [Cav99]. largest [dE97, FB91]. Last [BDG98, cIC99]. late [Ber99b]. late-onset [Ber99b]. later [BVB93, ES90, Lin92b]. Latin [Loh97]. lattice [Afs90b, Bab91, CS98a, HB99b, QT92]. Lattices [IMSK95]. Law [AK95, AGMCA96a, Chr90, Chr97, CS96b, EK96b, Ete97, GP99, Gut92, HO95b, Hon96, JK98, Mar95, Neu98, Pet96a, Rem98, SM92, Sun98, Sun99, VD91, dJ96, ARV97, Arc98a, BP99a, Bo98, Dev93, Dor97b, EN93, GKL96, GT99, Hal90, Hen98b, Mar99b, MM92, Nor95b, Qiy95, Sch98a, Sep96, Tob96, Tom90, Tom92b, Tom96, Tom97, TV97, dC93, BB91a, Csö92, Hor90]. Laws [Bir92, LT96, AGMCA96b, Bin95, CF96b, CH98b, DP94, Fro98, Gao95, GO97, Hen97a, Hen97c, KP93b, LL95, MW90, ND93, PL98, Pak98, Pfl91, Sch95, Sch99a, Sub94, WZ95]. LB [WRL95]. LDP [Deb99]. Leadbetter [Rie97]. leading [Ong95]. Least [MN93, Mül99, Par98a, Vie99, BG99a, BM96, But99, DJ97, Fro95, GS93a, GK97, HT90, Kun93, MP99, PL92, SB91, Sch96c, Tab94a, Tab94b, Van93, WTC90, WCT94, WL95, Yat96]. Least-square [Vie99]. Least-squares [Mül99, BG99a, DJ97, Fro95, WCT94, WL95]. Lebesgue [Mac96]. Leffler [PJ95]. Left [Rut95, PCK98, PG99, Sun97b]. Lehmann [Bar95a, KK92]. Leibler [Kan93b]. length [Dre99, FM99, KA97, LMF97, Rig95, Sai95]. lengths [HA98]. Lett [BL99b, Ham98a, HV00, Lin00, LGC99a, Pal01, Rad99]. level [BvHS99, DGA95, HN97, HB99a, KP93b, Pau96, PL92, Tom93, YL99]. levels [Bal95, HW97, Jia97b]. leverage [CMC94]. Lévy [BD94, Bra99b, Gut96, Gut97, Li94a, Neu98]. Lie [CCHP99]. life [AA92, BN97, Don98, FP94, Ger96, LSR90, Lu97, SC99]. life-testing [Lu97]. Lifelength [BL94]. lifetime [AL98, FPMP97, Grü94]. lifetimes [NS98]. light [Bra99b]. lightly [DB96]. Likelihood [EK95, EZ99, LP97, WZ95, YL96, AM98, AL91, AL92, AM93, AEE92, BT91, Bis99, BG91, cIC99, CD95, CL98, DeR94, DeR96, DT98, ED94, ED96, FM92, FS93, Gar98, Gas94, Gho95, Gu98, Hey94, HKR90, Hu97, IM97b, IZ94, Jin95, KD93, KC95, Kuk99, LSZ98, LPB90, Li95, MGH99, MS92, MH96, Mil99, NNR97, Par96, PL92, Qin97, RS91, SB97b, SA97, SS92, Siv99, Tza98, Wan94a, WJ99, YSM94, Zha97a, ZR96]. likelihood-based [CD95, SA97]. likelihoods [Gan98]. Limit [Arc94a, Arc95c, BC96, BCH98, BP93, BP92, Bry93, Cro98, Düm92, EH99a, FZF93, FZL97, GD97, GO97, Gut90, HH98a, HV97, Hen97c, Hor90, Hus97, IL98, Joh94a, KP93b, LP90, Lee97b, Lee97d, MP93, Mik97, Mik98, PL98, PS95a, Qui99, Rog95, Sai95, Sai99, SH95, Tan93b, Yan94, ZK97b, dN95, AHR91, BCC99, Che97a, CY99, DP94, Dzi95, FS98, Gam97, GKS97, Hen98b, JPAS97, JK95a, Liu94b, MR97a, MM92, MW97, PS99b, QW97b, QW97a, Str94, TT91, WW90a, Wan92b, Wan94b, WRL95, Wen94, WW95c, Wes97, Yan96b, Zho96]. Limitations [Akr93]. Limited [Car92, Ruy96]. Limiting [Bar92a, Epp99, RA93, RW99, Ame90, BES95, Deh91, DS92, MM98b, Ros93, RS98, RR91, TH95, Wir96, Zim94]. limits [AAK97, BH90, ES96, HS91, Hor90, HCJ97, Shi95a]. Lindley [Sta97]. line [Ram96, Woo95]. Linear [Cha92c, GK97, HM96, Jia97a, Kou96b, McK91, Mun99b, NS91, Yat96, AC91, ACGM99, ACH97, AC97a, BW97, BG99a, Bal93a, Bal97, Bal98, BHZ99, BS95a, Bas94, CF99, CFS99, CER93, Cha99b, cIC99, Che93, DL91a, Dav94, DB97, Dro99, EH99b, El 95, Elf97, FZF93, FZL97, FK94, FZ92, GL95a, Gao95, GKS96a, GD97, Gle93, GP99, Gro98, Hoa93, HP93a, HT96, Hus96, HP93b, Iso95, Jen97, JK95b, KW96, KS98a, KM97, KHW91, Kou96a, Kou96d, Kou96e, Kou92a, KL98c, Le 92, Le 98, Leb96, Lee93, LH96, LZ97, LMF97, Lee97d, Leu96, Lia95, Liu96, MP99, Man99, Mar98, MN93, MZ97, Mud93, Mud94, Muk94, Neu94, O’N92, Oht96, Pak96, PP92b, Por92, PW92, QR92, RZ92a, Rut95, SB96, SV99, Sar96, Sat92, Sch96b, Sch96c, Sém96, Sep92]. linear [Sev98, Sil97, SMP95, SP98, Son99, SS94b, Ste93, TW92c, Vas90, WK99, WW92a, WJ99, Wef91, Wie91, Wie96, Wu90, ZW99, dN95]. linear-quadratic [El 95, Le 92]. linearity [CM99, CPPP98]. linearization [Str99]. linearly [FL91, Mar98]. lines [Pau96]. link [LZ97]. linkage [LP97, PS94]. Linking [CG94]. links [DH90]. Linnik [And92, Dev90, KO96, Koz98, Pak98]. lives [BPRS97, BOR99]. Ljung [ST92b]. loan [Leo93]. Local [MS96b, AHZ97, ACGM99, AK95, BH96, CF99, Che97a, CS94, DM93, FN95, FR93, Gam97, Giv95, HT95, HB97, Jan94, KW97, KMC98, MW97, Moh92, PZ97, Sal93, Shi95b, SW99, Str99, TW92c, WW90a, Wan92b, WL93]. Locally [AC91, DBD95, SC91a]. Locating [Shi97a]. location [Ang96, BSW97, BZ99, CH94a, Cha98b, CS93a, FS99, GR93a, Gor91, HL96, JS97, JS90, Krz94, Leu96, LZ91, Ma99, MN97a, MM92, Nay98, ÖH96, Rot93a, Sah97, Str96, YL96, Zam90]. location-scale [BSW97, FS99]. Loève [RMNV99]. Log [SB91, BB93b, FS93, KK96b, KC95, SW99]. Log-concavity [SB91, BB93b]. log-densities [KK96b]. log-likelihood [KC95]. logarithm [AK95, Dor97b, EN93, GKL96, Gao95, Hal90, Hen97a, LT96, Neu98, Nor95b, SH95, Tob96, Tom90, Tom92b, TV97, VD91]. logarithmic [BH96, CF95, CF97, HW90]. logical [Aic98, BM97]. logistic [CH99, JL96, O’N94, Ste91a]. loglinear [Bal97, Bal98]. logrank [LT99]. Long [GM99a, DR96a, Deo97, GKS96b, Kou92a, Lah93, LL93, Mór94b, Muk94, PG94b, SB97b]. Long- [GM99a]. long-memory [Deo97]. long-range [DR96a, Muk94]. longer [LVCW93]. longest [FM99]. look [Cha98d, Mon93]. loop [IWLX99, SS92]. Lorenz [Cha94, CZ97, Wil96]. loss [DGS99, Hür93, Kab98, Kan93b, Mar97b, PL95, WK99]. losses [FR95, Leb96, Rou97]. lotteries [Hen98a]. lotto [Joe93]. Low [Coo98, Coo96]. Lower [Shi95a, GW93, HH94a, HH94b, JPAS97, Jia97b, Qia92, Uno96]. LPQD [Lou99]. LS [GL95a]. LSE [Akr96b]. LTAD [Tab94a]. Lugannani [BW95]. lumpability [Led95, Pen96]. Lundberg [CW97]. Lynden [Yua96].

M [Kou92a, ZN95]. M-estimators [Kou92a]. M-type [ZN95]. MA [Fra92a, SS93a]. MAD [Fal97]. main [SG96b]. main-effect [SG96b]. majorant [Wan94b]. majority [Nol99]. Majorization [Vit99, AP92, KS93, Pin94]. Mallows [Wer96]. manifold [ZW97a]. manifolds [Bax97, HJR93, HL96]. Mann [dE97, Zie90]. Mantel [BW93]. mappings [Git97]. Marcinkiewicz [Bo98]. marginal [AG94, ACS96, CH98b, DMY91, dE97, IK92, IK95, Jev90, Par98b, PP96b]. marginals [MRR92]. marked [BMS97, FP99]. marker [HT95]. Markov [Adk93, BP99a, Ban90, BR95, BT91, BR98, Bra99a, Car94, Cha99b, CH98b, Cse93, Cse95, DK92, DR96b, DT98, Dup97, Fad98, Fer99, FF97, Gro97, HA98, HLL96, Hug97, HRS99, HB99b, KD92, KD96, Las97, Led95, Lim93, LL95, Lou97, MRN93, MM96, Mat93, MW97, MMPZ99, Pen96, POGP96, Pfl91, Phe90, RCD93, Rou91b, Sen93, Spr98, Tan96, Tan93b, Wen94, WW95c, Yca92]. Markovian [AB92a, BVB93, Pet96b, Tan94]. Marshall [Bol98, Wu97a]. Martin [Nor95a]. martingale [ARV97, Cab90, Fie90, PR96, PR98b, Rac95, Sch99c, Wes90, Wes94, Wis92]. martingale-difference [PR98b]. martingales [CM95, DL91a, Eth98, FZL97, Gaw97, Iva96, Joh94a, MZ93, Shi97b]. Mason [BB91a]. mass [CATD93]. Master [Ano91c]. matches [BGP94, HMG98]. Matching [SN96]. Mathisen [SB94]. matrices [BMT97, KL99, Kou95, PG94a, Sch96d, Sch97b, WW95b, Wan96a, Wan91a, Wan91b, YSM94, Zha94c]. matrix [CS96a, Cor93a, Gle93, GP92, GV91, Jin93, KK96a, Li93, Li94c, LF95b, Mar91, Mof97, NS96c, Olk98, PFvR98, ST92a, Sar96, Sat92, Shi99, Str94, TS93, WH91]. matters [He94a, He94b]. max [Arc92, Baj92, CF96b, FS98, LZ91, MR97a, Nol99, Sub94, Won93b]. max-limit [FS98]. max-stable [MR97a]. max-type [Baj92]. maxbias [Bas94]. Maxima [Hüs94, BSW94, GN94, Gut90, HV97, HH96, Kau95, Olo99, PL98, QW97b, QW97a, Qi97, RR91, Rot97, Sgi96, Tom96, Tom97]. Maximal [Ete99, Bar92a, Mac98, NV99, Qiy96, SC93]. Maximin [Sch97c, Bis96b, Lee96b]. maximization [ASC98]. maximizers [Fer99]. Maximizing [Cha96b]. Maximum [Gar98, Hu97, IZ94, KD93, MMP97, Mil99, Pem96, PL92, Sre95, AM93, AEE92, Bis99, BG91, cIC99, CL98, CW98, Chu96, Coh90, DT98, ED94, ED96, Fal90a, FM92, FS93, Gas94, Gho95, GO97, HKR90, Hua98, IM97b, Kuk99, LSZ98, LPB90, LH98b, Li99, MGH99, MM92, MH96, Moh92, Poo91, RS91, SB97b, Shi97a, Sta95a, Sta99, Str96]. Maximum-likelihood [Hu97, CL98]. maximum-order [Hua98]. McCullagh [Ses91]. Mean [BS91, FP94, JH90, Sar96, Abr93, ARV97, Ani94, AGMCA96a, Bal90, BDG98, BDS99, CPS93, CR95, Cha90c, Cro92, Csö92, DB96, Dre99, Elt92, FG92, Fis91, Fox97, Gal98, Gug98, Ham95, Haw97, HL96, HKR90, Hu95, Hua99, Ind94, IU95, IM97b, KL98a, KP94a, KZ91, LB91a, Liu99b, Llo91, Mar94, MZ91, MZ97, Mie97, MvdM97, Oos94, PKM93, PW97, PGTD92, Pit90, Rad96, Rad99, Ryc90, Sai95, Sep94, SB95a, SQC94, SZ92, SS99b, Sri91, SS90b, Sta92, SS97d, SS98, Uno95, WW96, WK99, Won93a]. mean-variance [Pit90]. meaning [KL97]. means [ABK98, Arc95a, BL99a, Omo97, Ped97, PS97b, SS92, TYS95]. Measure [PV90, BDG93, BG95, BN96, GLM94, Jum99, Kni91, Koz97, KZ91, Lee90, Löw97, RW92, Sch99a, Sch99c, Sto95, Swa96, XM91]. measure-valued [GLM94]. measurement [BAV98, CER93, DLW96, Hug96, HS94b, Nav99, SS97c, Wan99a]. measurements [Afs90a, Jen92]. measures [Akr93, Als91b, BN96, Cek96, dC98, Dar97, DB94, EPS94, Fun92, GR93a, Mik98, MH90, Nel92, Ras95, RG99, SQ90, Sch95, Szy94, WTC90, WMN94]. measuring [XM91]. Median [CF90, Tak91, CMS94, Che98c, Cro98, Fal97, Ham95, Hua91, Jey91, LM92, NOT90, Oos94, Qia96b, SB94, Van93]. medians [CC99a, Hua99, Mar91, NNR97]. Meier [Cai98, Cha96c, SD99, She98, Stu94, Yan94]. memory [Deo97, GKS96b, PG94b, SB97b]. ménage [Hol91]. Menten [DW99]. method [AL97a, BB92, CAGM98, FL98, GS96a, HB99a, Jin95, Jin96, Lim93, LN94, Lou97, Mar99b, PK96, PP96b, Rao95, Sho97, TG96, Wer96]. methodology [MS99]. methods [Bru92, Cha99a, GT99, GD97, HN97, Kal98, KP93a, SV99, WCT94]. metric [CR98, CT99, Las97, Maj99, Roo95, vRR98]. Metropolis [dC98, Hol98]. Michaelis [DW99]. mid [Bin95, Bin96]. mid-range [Bin95, Bin96]. Midzuno [Sah97]. MIFR [AB97]. Mills [Wic90]. Min [LZ91, Bas94, Joe90, Nol99]. Min-max [LZ91, Nol99]. min-maxbias [Bas94]. min-stable [Joe90]. mini [Won93b]. mini-max [Won93b]. Minimal [KG90b, Khu97, Won93b, Aic98, BD91, BCH98, Rad96, Rad99, Sch99c]. minimality [Sch99c]. Minimax [FG92, Hof97, Hud97, Mag90, BMvdM97, Fer96, HT95, Kor99, LB91a, PL95, RKD90, Sta92, Wei91, WW90b]. minimaxity [Erm92, Ser96b]. minimizing [Wu92]. Minimum [BS94b, Kou92b, Muk94, Pak96, Qia92, Qia96a, SNN97, BW97, Bar92a, Cha91, CLN95, CHW93, GS95, KP94b, ÖH96, SS90b, SS94b, WK99]. minorants [Bal93b, Raq97]. MINQE [Cha91]. MINQUE [Gro97]. misclassification [Cha90b, CH99]. MISE [Jon91]. Mises [Chr97, Kau95, ST96, Zim94]. Mises-type [Zim94]. missing [CP97, Gar98, LP91b, SS96a, SQC94, ST92b]. misspecified [Oht96, Sém96, Sev98]. Mittag [PJ95]. Mixed [Per98, Wan96a, Cha99a, Gu95, Har91, Jev90, Jev91, Khu97, Krz94, Li96, SB96, Sil97, SW90a]. mixed-effects [SW90a]. mixing [AdlC93, BDG98, BT95, Bos95, Bra99a, CR92, Dzi95, Ger93, HL92, Kim93, Kim99b, Kri90, Kri93, Lee97d, Lie99, MM90, Mat93, PS96b, Per98, Qiy95, Rad96, Rad99, Rao99a, Rie93, Sha93, Vei90, iY95, Yu96]. mixingale [Dav93, dJ96]. Mixture [Koz98, O’N94, Pak98, AS96, AM97, Bas96b, DK90, FY93, Hoa93, KO96, MGH99, Mi99, PD99, PT98, SC91a, Ste91a]. mixtures [AdlC93, BC96, EM99, Jen97, Kou92b, McK96]. Miyakawa [Gas94]. ML [Fer97, Nag99]. ML-estimator [Fer97]. MLE [AB92b, Gan98, PCK98, SS96a, YB99]. MLEs [Gas96]. MM [ABY94].

MM- [ABY94]. mode [AFM96, CH98a, Fut99, Vie96]. Model [HT90, MT99, NS97b, AS96, Aic98, Akr96b, ACH97, AC97a, AM93, AFH^{+}96, BS94b, BKR96, CF99, CFS99, CS96a, Cav99, Cha96b, Cha96c, Cha91, CR91, CT94, CS92, CC94c, CLK96, CK94a, Das92, Dav94, DW99, DB97, DT98, ES96, FS99, FY93, Fra91, Fra92a, Fra93, Fuj96, Gar98, Gas96, GKS96a, GS99, Gho95, Gom96, Gro97, Gu95, Gu98, Han93, HP93a, Hug96, Hug97, HT96, JL96, JM90, JW99, KS99a, KG90b, Kul92, KM93, Kun93, Le 92, Lee93, Lee98b, Lee99, LN95, LB91b, Lu96, LC92, MRN93, ME92, Mar97a, MM96, MM98a, MV95a, MST97, McQ99, MPS97, Mof97, Nad99, Oht96, PKM93, PFvR98, Pit98, Por92, QR92, RZ92a, RG99, SB96, Sém96, Sep92, SB97b, Shi93, SQC94, SA97, SPH91, Sin94, Son99, SS94b, Sri95, SS97b, SS97c, Sta99]. model [Stu92, Sut97, Tan95, Tso98, WW92a, Wei90, WS90, Wie96, WMN94, Yan90, Yan96a, Yan96b, Zam90, Zha92b, ZL97, O’N91]. Modeling [Fra92b, KC97a, AC97b, Lea91, PD99]. modelisation [GR92]. modelled [PW92]. Modelling [Fad96, Fad98]. models [AC91, ACGM99, BW97, BG99a, BS97, Bas94, BS94c, BHB96, BL95, BL99b, BAV98, BR98, CER93, Cha99a, Cha99b, cIC99, Che93, CD94, Che94, Che98a, CM99, CH99, CPPP98, CPA95, Cho97, CK94b, CV97, De 96b, DZ99, Dro99, EH99b, FP94, FC96, FCUOCN96, FBCCN96, FZ92, GL95a, Gao95, Gle93, Glo99, Gne98b, Gne99, GW96a, HJM98, HM96, Hil99, Hoa93, HOGFLC98, Hus96, HkL92, HS94b, IK92, IK95, IWLX99, Jan94, Joe91, JK95b, KS98a, Kan92, Khu97, Kor99, Kou96a, Kou96b, Kou96d, Kou96e, Kou92a, Kuk99, LSZ98, LPB90, LZ97, Lee97c, LSR90, Lia95, Liu96, Mar98, MN93, MZ97, MSS90, Mi99, Mor94a, Mos94, Muk94, MS97c, Mur92, NY90, NG95, Omo97, Ong95, Pak96, PR96, PW97, Pap96, Pap97, Pap98, PL92, PL93, Qia96a, RP95, SC93, Sch96b]. models [Sch96a, Sch96c, Ses92, Sev98, Sha90, SS94a, Sil96b, Sil97, SW90a, SMP95, SP98, Str99, SR97, SNN97, Tan94, TV98, TP96, TP98, TW92a, TW92c, Tsa98, WGR94, WJ99, Won95, WW94, WPH96, WS98, YR96, Yip92, ZW99, Hof99]. Moderate [Che97b, JWR92, Dar97, Dar98]. modes [Nol98, Ves94]. modification [Muk92b]. Modified [Rob90a, ST98a, CF90, CS92, Dat95, HT91, LT99, MG96, RMNV99, YSM94]. Modifying [FZ92]. modulo [PP92b, QW97b, QW97a, Sco93]. Moment [Lin92a, DS92, Fan91, Joh91, KK96a, Kim93, Lan96, Lan98, Lin97, Lin00, Lin98, RB96, TS93, dlPG92]. Moments [Bac97, CL94, Chu92a, Gha94, GK98, Hos94, Sch95, SBD92, Tsu92, Als96, ACH97, BAC92, Bal93b, BJK98, Bél91a, DW92, DB96, Dav93, GM93, GP96, IL98, Irl97, JK95a, KP96b, LT92, LH98b, Neu94, PS99b, Raq97, ST90, ST98a, Sat92, Sgi97, SW90b, SD93, Stu97, TS96, ZK97b]. Monge [Bel99a, CATD93]. Monotone [Lia99, BlZ93, DMLP97, Kra98, Muk92a, PCK98, SS96b]. Monotonicity [HS94b, Koc94, Bra96b, GW96b, Pre93, Sap95, Sap99, YLB99]. Monte [HRS99, Kuk99, RQ99]. Morgenstern [Mik97]. Móri [Hen97c]. Morphological [Car94]. Morris [Sha92a]. most [AC91, BM96, SC91a, Tra96]. motion [AHZ97, Alb93, Alb97, AK95, Ber96, BG99b, Chu96, EN93, EN99, GA99, KP93b, Le 98, MM92, Neu96, Neu98, Str96, Xia97, Yin99]. motions [NV99]. motivated [PKM93]. motivation [GP91]. moving [Alb97, AM93, HV97, Lah93, LRW92, Pap96, Pap97, Pap98, PS96a, Rig95, RR91, Wan92a, Zha96]. MSE [Oht99]. much [KV99]. multi [DLW96, DY98, Stu98]. multi-dimensional [Stu98]. multi-factor [DLW96]. multi-state [DY98]. multicollinearity [Sar96, WTC90]. Multidimensional [Li94a, BHZ99, DL91a, Jak94, Neu97]. multidimensionally [Chr90]. multifractional [BCI98]. multimodality [EM99]. Multinomial [DL91b, DLY96, AAJ97, BS94c, Bha95, Bla91, DLS90, ED94, ED96, hK99, ST92a, SB97a, TW92a]. multiparameter [Iva96]. Multiple [Ger93, Pri99, DT93, Geo90, GR91, HW97, KG90a, Lee96a, LM96, Liu97, MM90, McC90, NS96a, NS97a, Pre93, SS96a, SF94, Som93, Wat95, WL93]. multiple-player [SF94]. multiplicative [BKR92, Fro95, LN94, Lov98, Yan96b]. multiplicative-binomial [Lov98]. multiplier [GW96a]. multiresponse [NY90]. multisample [MM94]. multistage [Kor97]. Multivariate [AW94, ACS93, ACS94b, AM94b, Bal93a, GW95a, PA90, PAT90, YM95, Ahm95, And92, Arc95a, ACS94a, Bae97, BM96, BJK98, BB95a, Bis96a, BS91, Cek96, CC99a, CHW93, CV97, Gne98a, Gor91, GK97, Han93, Hen97b, Hen97c, Hen98a, Iso95, JS97, Jen94a, JN97, Jin93, JS99, Joe90, JL96, JS90, JPZ97, Kru98, KV96, LM98, LB91a, Li96, Lor95, MR97a, MRR92, NS96b, Neu94, OW91, PL95, Par92, PD99, Pou93, RQ99, Rig95, RSC94, RA92, Rut95, Sat92, Sep94, Sep96, Ses92, SMP95, SF94, Ste97, Sut97, Tak94, Vel93a, Wiś96, YFL96, ZK97b]. multivariate- [SMP95]. mutual [BCQY97, WH93]. mutually [Brz97].

n [Sch96b, Sch96a]. Nadaraya [Gla98, LP91a]. Natural [BLBL92, AB95, LR96, Sha92a]. nb [PW97]. Near [Hol95, JR99, Li99, PL98, WW95a]. nearest [BCQY97, DH90]. nearness [Lee90]. necessarily [Imh97, KS98a, Sin94]. Necessary [Sep94, BW97, BES95, Str94]. need [JPAS97]. negative [AEE92, Bal93b, Cha91, CM90, Fer97, Hit90, HH99, Jev91, NN96, ST92a, Sap99, Sgi97, Wan96b]. negatively [Mat92, SC91b]. neighbor [BCQY97, DH90]. neighborhoods [Fal94, Kau95]. neighbourhoods [Wu92]. nested [DK93, KG90b, SR97, Szr96]. Nesting [Joh96]. nets [CAGM98, YM99]. Newcomb [FLW99]. Newton [BKR96]. Neyman [Wei92]. nilpotent [CCHP99]. NN [Cha90b]. no [Liu94b]. nodes [TLC97]. noise [ES96, EH99b, GR92, KS99a, SK93, SS93a, Wal99]. Non [BHZ99, BK96, Dia90, KA97, Abt94, AHZ97, Alb97, ACH97, AC97a, Bab91, BlZ93, Cha90d, Cha91, Che96, Dev94, GG91, HN97, Hit90, HP93b, Jan97, JW97, Kag91, KS95a, Kni98, LVCW93, Löw97, MP99, Mar97a, MM96, PRS92, PS99a, Rob90b, SB95a, SQC94, SJK96, SS97b, Szr92, Szr96, Tan94, Woo92, dN95, vE97]. non-anticipating [Alb97]. non-arithmetic [Abt94]. non-binary [Kag91]. non-central [Cha90d, Rob90b, Szr92]. non-centrality [Szr92]. non-consistency [Dev94]. Non-dependence [BK96]. non-existence [JW97, PRS92]. non-extinction [GG91]. non-homogeneous [MM96]. non-i.i.d [Jan97, SB95a, Woo92]. non-identifiable [KS95a]. non-ignorable [SQC94]. non-independent [Mar97a]. non-integer [HN97]. non-lattice [Bab91]. Non-linear [BHZ99, ACH97, AC97a, HP93b, MP99, dN95]. non-Markovian [Tan94]. non-monotone [BlZ93]. non-negative [Cha91, Hit90]. non-nested [Szr96]. non-normal [SS97b]. Non-parametric [Dia90, KA97, Che96]. non-proper [SJK96]. non-regular [Kni98]. non-responders [LVCW93]. non-reversible [Löw97]. non-smooth [vE97]. non-smoothness [AHZ97]. non-stationary [PS99a]. nonadapted [Eth98]. noncentral [Kal90]. Nonconsistent [Kut94]. Nonconstant [Bal95]. noncontinuity [MW90]. nondecreasing [Cas90]. nondifferentiable [Kim92]. nondifferential [BAV98]. Nonexistence [CW94, CC94b, GS95, GW95b, HkL92]. nonhomogeneous [CS93b, CDG93, Fan91, GW96b, LL95, Tan96, Wen94, WW95c]. nonidentical [Ryc95]. Noninformative [Mee93a]. nonlinear [BL95, BL99b, BKR92, BR92, Bis99, CF96a, CPPP98, CmHP98, CV97, DZ99, Efr94, FC96, Gat93, HH98a, Hof99, Jev90, Kun93, Lee98a, NY90, PP96b, RP95, SK93, Sha90, Str99, TP96, WW94]. nonnegative [GS95, Wen97]. Nonnegativity [Als96]. nonnormal [Dat95, Sch99a, YB99]. nonoptimality [SG96b]. nonorthogonality [SG96b]. nonoverlapping [MRR92]. Nonparametric [Akr96a, Ben97, BG95, Cro92, Deo97, Geo90, IA97, JEG99, Kat97, Koh99, Lee96a, RG95, Rao92, TP96, Ahm92, AL97a, Alt93, Bru92, Bur93, CD95, Che95b, Del93, ES96, FP99, Gal98, GS96b, Hof99, Hol95, Ioa92, JPAS97, KMC98, Kul95, Kut97b, Leo99, Li95, Lie99, LN94, MS96a, Mee93a, Mey99, Mül97, Pap95, PKJ97, PR97, Rob95a, RS91, RW92, Ryc90, Sch99b, ST91, Yao90, Yu93, Zie99, vEK98]. nonresponse [Elt92]. nonstandard [Wei91]. nonstationary [FNARP97, HOGFLC98, Pit98]. nonuniform [Wu90]. norm [BW97, KP94b, Sto96, YM95]. Normal [QT92, ABK98, ACS94a, Bal93a, BC98, BHK^{+}99, Bas96b, Bis96a, CF99, CS97, CG94, Cha92c, Che93, DL91a, DK90, Dre99, EM99, FR95, Gar98, Gne98a, GP92, Hu95, IM97b, Jin93, Kou92b, KP94a, Lee95, LB91a, Lia99, Liu99b, Mad93, MZM94, MvdM97, NS96c, Ole97, PL95, Ral92, RSC94, Rob91, RA92, Sep96, SS92, Sin94, SBD92, SD93, SG91, SS90b, SS97b, Sta92, Ste91a, Ste97, Vit96, WW96, YFL96, YB99]. Normality [KW96, Abr99, AW94, Arc95a, ACS94b, Epp99, FS93, GL95a, GKS96b, GMS94, Hen97b, HB97, KV96, Lie99, MP99, Par99, PG94b, Rou91b, Rou95, SQ90, SQ97, VA97, Vel93a, Wu90]. normalization [TT91]. normalized [AdlC93, ES97, Qui99, Wan99c]. normalizing [GN94]. normed [Mee91, Ros93]. norms [GKL96, Lat98, Stu94, Web92]. Note [AT92, Gas96, MN97b, Nyb92, RT93, Alb98, Ame90, AC97a, AM93, BS94a, BMT97, BJK98, Bas94, BR90, BV92, Ber95, BG97, BMS97, BS96b, CFF98, Cha96a, Cha98a, CC99a, Cha92b, Cha91, CS93a, Cha90e, CS95a, Che98c, CY99, CH91, Che95b, CF96b, CLK96, CH96, Deb99, DF95, Dev90, Ext91, FS93, FD93, FR95, Gan95, Gat93, GGR96, Gil92, Gla98, Gov97, Gri92, Gro97, GHW92, GP92, HKP99, HL97, Haw90, HH96, HS94a, HK95, HT95, Hu95, HSV98, HV00, Huw95, Iso95, IM97b, Jev91, Jin93, Kab98, KW97, Kep97, Kim93, KJ91, Koz96, Kru99b, KS96, Kus98, Lac96, LP90, LSZ98, Lee97c, LH98b, Li99, LP91a, Lu96, MRN93, Mam90, Mar97a, Mat92, McK96, Mun99a, NF97, ND93, PT96, PR96, Pat96, PS95a, PS96b, PR97, Pfl91, Pre93]. note [Qi97, Ram99, Rav92, Rhe99, Run93, Ruy96, Sah97, Sev93, SB94, SHC99, SS93a, Sho97, SPH91, SD93, Son99, Sta95a, Ste91b, ST92b, Str94, Tal91, TS96, TLC97, TN93, Tra96, Tza98, Uno95, Van93, Vel93a, Vel98, Vid96, Vie96, WC90, WR90, Web92, Wic90, WH91, Won93a, Xia95, Yu96, Zho96, dlPG92, vE97]. notes [ZW97b]. notions [PS97b]. nuisance [LVC90, Muk92b]. null [RM90]. number [BD91, BB95b, BSW94, BGP94, Cse95, Dev91, FTV98, Git97, KNN97, Lee95, Li99, MZT90, Mas99, Olo99, PL98, PT98, QW97b, QW97a, Qi97, RCBG96, Ser96a, Sta99, Tra96, Uno96, Won93b]. numbers [Arc98a, AGMCA96b, Bir92, FLW99, Gil92, HP94, Joe93, Yam90, AGMCA96a, BB91a, Chr90, Chr97, Csö92, CS96b, EK96b, Ete97, Gut92, HO95b, Hon96, Hor90, JK98, Mar95, Pet96a, Rem98, SM92, Sun98, Sun99, dJ96].

obliquity [MS92]. observation [Che92b, KP94a, KS99c, MM90]. observational [NG95]. observations [BD91, BM93, Bis96b, FM94, Fra91, Gei93, GCL90, GL95b, KL98a, Kut94, Kut97b, KS99c, LV99, Leu96, Mas94, Ser96a, ST91, Vie99]. observed [Hug97, SB97a, SC93, SS93a, SS94a]. obtained [CC98a, RZ92a]. occupancy [Sta99]. occupation [GLM94, HZ99, Kro95]. occupied [Rév96]. occurrence [BB93b, SB91]. Odd [Bél91a, Als96, CMS94]. odds [FS93, Jan94]. off [SW92]. offspring [Sri91]. Oja [NOT90]. Olkin [Bol98, Wu97a]. Olkin-type [Wu97a]. OLS [Pit98]. omega [Ser96b]. omega-square [Ser96b]. One [Yan96a, AM98, BM99, Coo93b, Das92, Dor97b, FCUOCN96, FBCCN96, Fro98, KP94a, KZ91, Liu99b, PS96a, QW97b, QW97a, Sil96b, Sil96a, Sut97, Tom92a, Tom96, Tom97, Zie90]. one-parameter [Coo93b, FCUOCN96, FBCCN96]. one-sided [Fro98, Liu99b, Sil96b, Sil96a, Zie90]. One-step [Yan96a, AM98]. one-way [Das92, Sut97, Tom92a]. only [BN96, Liu96]. onset [Ber99b]. open [IWLX99]. open-loop [IWLX99]. operating [ADD93]. operations [ANS93, Mel95, Roj93]. operator [Sch95, Sch99a]. operators [Lat98]. Optimal [ASC98, BD91, Bos95, DDD98, DR96a, Gas92, Ger96, Gui96, Kou95, KS99c, PS95b, SC93, SBB93, Tom93, WW92a, ZRB98, AK99, Cha98c, Cha99c, CC94b, CM90, CJ98, DW99, Hol95, Irl96, hK99, Kou96b, Kou96f, LP91a, LN95, Mam90, MH90, Pre93, RJ91, Rob95b, Rüs96, San91, TLC97, Won93b, Won95]. Optimality [Das98, KP93a, Brz95, Dro99, GDD91, Hös92, KK92, Koz96, KW98, Lah91, Poh92, Sch97c, Sch91b, Ste93, Wie91]. optimization [Ano90c]. Optimized [GKS96a]. optimum [Kag90, MZ97]. optional [AC99, Eth98]. options [Chu96]. oracle [HKP99]. order [AM93, AL97b, BSW97, Bal90, Bal93a, Bal93b, BB91b, BB92, BB94, BB95a, Ban90, BK93, Bar95b, Bar98a, Bar98b, BPRS97, Bel99b, BOR99, CFF98, CG92, Cha99a, Cha92a, CT99, CC94c, Chu92a, Chu92b, Chu92c, CS95b, Dav95, DB96, Dav97, FH92, FBCCN96, Gad97, GC92, HK96, Hos94, Hu97, HH98b, HP93a, HT91, Hua97a, Hua98, Hua99, Hud97, Jen94b, JEG99, Kak98, KP96a, KW92, Koc96, KH91, Lah91, Lan96, Lee98a, LH98a, LV99, LM96, Li94b, Lia95, Lin90, LT99, Mal96, MvdM97, MH90, Neu94, Nyb92, Pap95, Pap90, Par96, Par98a, Par99, Ped97, PS97b, PA90, PAT90, cQ94, Rao99a, Rem98, RSC94, Ryc92, Ryc95, ST98a, SD90, Sat92, ST98b, Shi94, Sil97, SW90a, SS92, SW99, SW90b, SBD92, SD93, TS96, TW92b, Tom99]. order [TW92c, Wil96, Wiś96, Woo92, WL93, WS98, Zha94a]. order-restricted [AL97b, CS95b, SW99]. Ordered [SY95, CC98b, DLY96, EK95, FF97, HH94a, HH94b, Koc94, LV99, LZ94, MMP97, Mos94, Par98b, PS94, Pau96, Qia96b, Wan94a]. Ordering [NS98, Aly90, Bal93a, BB90, Bar92b, Bax97, BOR99, Bha95, EK96a, Fan99, Koc96, KM99, MP90, RH91, SS90a, SS95, Wil96]. orderings [BCR96, BPRS97, Bur93, GW95a, JdKP95, KK99, KC97b, Koc99, NJS98, Roj93]. orders [Bar99, BDS99]. ordinal [GKS96a, OB95]. ordinary [But99, WCT94]. O’Reilly [Szy94]. origin [Zim94]. origin-invariant [Zim94]. Ornstein [WC97, KP94b, Lef97]. Orthogonal [Kou96d, Kou96c, ARM97, DM96, Kou96a, Kou96e, MT99, Rao95, Tia98, Wan96a]. orthogonality [DLW96, Koz91, Muk92b]. orthonormal [Mas94]. oscillation [Dev91]. other [GMS94]. Outlier [CT94, BCQY97]. outliers [Dav97, SSL96]. overdispersion [Luc96, Mof97].

p [SS93a]. P.B.I.B. [Kan91b]. Packing [Xia97, Rhe99]. paired [Joe91]. pairs [Kre96]. pairwise [CM91a, CM91b, Kre96, Mar95, Rot93b]. Pál [GW97]. panel [Pit98]. Pao [FL98]. Papathanasiou [BB93a]. papers [Ano90c, FRS93, Jon95]. parabolic [GA99]. paradox [BG99a]. parallel [Lu97, Pau96]. parallelepiped [GB96]. Parameter [CD94, Kak97, Kim99c, KA99, QT90, ADD93, AD99, Ang96, AEE92, BSW97, Cab90, Che97c, Coo93b, Cse93, Cse95, EJ96, FM92, FCUOCN96, FBCCN96, Fer97, Gov97, Huw95, Kak98, Kay92, Kus98, KP94b, Le 98, Liu96, Ma99, MS97a, Mil99, Mud93, Mud94, Muk92b, Rot93a, RKD90, SV91, Sun95, TW92a, Tso96, WW95a, Wan96b, Won93a, Yan90, ZZ93]. Parameterizations [Nol98]. parameters [AC91, BM99, Bha95, BKR96, CS96a, Cha98b, CL98, Che92b, DT98, Elf97, Fro95, HP93b, Irl96, KS95a, KS99b, Kul92, Liu96, Nag99, ÖH96, RG95, Szr92, WJT92, WPH96, YL96]. Parametric [Gne99, HOGFLC98, Muk92b, AFH^{+}96, Bow99, Che96, Dia90, Gas96, Ind94, KS98a, KA97, Phe95, Qin97, Sil96b, Yan96a]. parametrically [Gla98]. Pareto [ACS93, DvC95, EJ96, Fal94]. part [Bra96a, PAT90, Akr93]. Partial [EPB94, DS92, Eic98, EN96, Hon96, HP93a, JK98, JdKP95, Kla93, KJ91, Kul97, NJS98, PR97, RS98, Sgi96, TV97, WJ99, vP91]. partially [FF97, SB97a, SV93]. particle [LS98, Phe95]. particular [Gen99, Kun93]. partitions [SY95]. partly [GL95a, Gao95, Sch96b, Sch96c]. parts [BKS96, ET91]. passage [Als91b, BDG98, BvHS99, DR96b, Gut96, Lef96, POGP96]. past [CP97]. path [GG91, LL93]. paths [AC99, Sam94, ZZ93]. pattern [Cha99a]. payoff [SC93]. PBB [Brz91, Brz97]. PCA [Bes92]. PDF [Nag99]. Peakedness [Jen97]. Peaks [Lea91]. Pearson [Lor92, Nag99]. Pearsons [RB96]. peculiar [DH90]. Penalized [Gu98, BHB96]. penalties [ZK93]. percentile [Bas96b]. percolation [Mee99, Whi99]. perfect [SC99, Tak94]. performance [Jon90, Oht99, Wan92a, Wan98a]. periodic [Bal95, CDG93]. periodically [FB96, HL92, LW92]. periodicities [LX97]. periodogram [Pap96, Pap97, Pap98]. periods [Bal95]. Perlman [SW93]. permanents [RW99]. permissible [DBD95]. permutation [Jan97, SS90a]. permutational [SQ90]. permutations [FLW99]. Persistent [RR94b]. perspective [He94a, He94b]. perturbation [Sen93, WW94]. perturbations [MS95]. perturbed [Deg93, DP94, Gut97, Ral92, Wan92b]. pest [Cha96b]. pest-predator [Cha96b]. Petersburg [BCC99, CS96b]. phase [Kim92, Lee93, ZRB98]. phenomena [DH90]. phenomenon [CF96a]. Pickands [Deh91]. piecewise [Koh99, KS92]. Pierce [ST92b]. Pillers [Muk95]. Piterbarg [Hüs99]. Pitman [BDG93, FK94, Kou90, Lee90, Nay98, SS93b]. place [BES95, Lef97, Yin99]. Plackett [Joh96]. plan [Lu97]. planar [AHZ97]. plane [FD93, Jum99]. plans [Mag90, SG96b]. play [Bac97, MR97b, RS99]. play-the-winner [MR97b]. player [SF94]. players [Cha95, RS99]. plot [BD90a, Gle96, NS98]. plots [Vel98]. plus [KS99a]. PMLE [Lia95]. Poincaré [BH97, BH99]. Poincaré-type [BH97, BH99]. Point [Sim92, Ben97, BMS97, CK94a, Cor93b, CH96, Der95, Dia90, Fer94, FP99, Gor91, Gui96, Haz98, HT95, Hud97, IU95, Joh94a, KL98a, Lee96a, MS96a, MS97c, NOT90, Roo95, Ruk96, Van93, Wiś96, Yao90, ZW97a, ZW97b, dC93]. points [Che97a, CM93a, Dev91, Gom94, GB96, Gut90, HV97, Lee95, LF95a, Mas99, Mee99, Nol99, Tar94, Won93b]. pointwise [BC96]. Poisson [AdlC93, AC99, AA98a, AA98b, AHR91, Bal95, BS97, BH92, BR99, Bla91, BR98, BX95, BGP94, Cek96, Cek98, CS93b, CDG93, DLS90, Ehm91, ED96, FR92, Fan91, FR95, God90a, GW96b, Ham95, Har91, Hen98b, IM97a, KC97a, Kru99a, KL98c, KS99c, McK96, McK91, MZ91, Mof97, Olo99, PL98, PD95, Per98, Poo91, Pri99, Roo95, VC99, WJ93, Yan91]. Poisson-multinomial [DLS90]. Poissonity [GMS94]. Poissonity-normality [GMS94]. policy [SC99]. polish [CF90]. Pólya [DS96, SJ97, WM97]. polygon [Wat93]. polygons [CGT97]. polynomial [BLB90, Cha98c, Cha99c, Che94, Che95a, ES91, LN95, Nad99, Ouk94, Sch91b, Tia98]. polynomial-trigonometric [ES91]. polynomial-type [Cha98c]. polynomials [ACGM99, Far96, GS96a, Ram99]. polytomous [LPB90, PL93]. polytopes [Vit96]. pool [Ahm91]. pooled [LY94]. population [BS94d, BK96, CPS93, HH98a, KP94a, Mar94, Ped97, RZ94, SS90b, Sri95, TG96, YF93, YLB99]. populations [Coo96, Hor90, KS96, Mad93, MM91, MvdM97, Qin97, SZ92, Zog94]. portmanteau [Bat90, ST92b]. positions [Ger96]. positive [Als96, Bar92a, BR95, DeR94, DeR96, Fer97, KC97a, Koz97, Mar97b, Nel92, Ros93, Rou94]. positive-breakdown [Rou94]. Positivity [BSW97, LR96]. possession [GS91]. Possible [Sam94]. possibly [Fra91, Sém96, SS99a]. Posterior [PW97, SV92, Bas96a, Gho93, Ma99, Mac96, MGH99, MS90, MZ97, PFvR98]. pot [Mar99b]. pot-method [Mar99b]. Potential [Jon92, WL93]. Power [DT93, GPI99, GW96a, BM99, Bat90, BW93, CW98, KW98, KV96, PG94c, Rot93b, SS92, TP98, TW92a, Wil96]. power-divergence [TW92a]. power-function [Wil96]. power-quadratic [PG94c]. powerful [AC91, SC91a]. powers [dE97, HW95, PL95, Ram97, SW99]. PP [Hen97a]. pre [ABK98, FH99, Oht99]. pre-asymptotic [FH99]. pre-test [ABK98, Oht99]. predator [Cha96b]. predecessors [Git97]. Predicting [Jes93, BS94d]. Prediction [CP97, DLS90, GR91, Hal95, PS98, Pou93, SB96, Tak91, ZK93]. predictions [Ste93]. Predictive [JS98, BK96, Coo98, JEG99]. Predictor [Jia97a]. predictors [CK94a, Gle93, Jen96, Rod93, Sri95]. Preferred [ZW97a, ZW97b]. preliminary [GS93a]. prepivoted [FJ92]. prescribed [AAK97]. presence [HK95, Irl96, LVC90, Muk92b, Sch94, Wal99, Zha97a]. present [Mun99b, dUÁGM98, dUÁGM99]. Preservation [BS97, NJS98, Roj93]. preserves [GR99]. Preserving [BT95, YSM94]. prespecified [KS95b]. principal [Jen96, Kai99, LF95a, Mar99a, Tar94]. principle [BS93, Bry93, FP99, GRZ99, KW97, MMP97, PR98b, RZ92b]. prior [AV97, Coo93b, Mac96, Maj92, MZ91, PW97, PZ97, PR97, PGTD92]. priors [Bor96, FS99, MW91, MS96b, SBB93]. Probab [BL99b, Ham98a, HV00, Lin00, LGC99a, Pal01, Rad99]. Probabilistic [Yam90, HP94, Hol91, Lin94, MW93]. probabilities [Bho92, CR98, Cha90b, CS92, Khr98, Kim92, Kim99b, KW92, Kri93, Kyr94, Las97, Li95, Lug95, Mil92, Par98b, Pau96, Phe90, Wen94]. Probability [Chr91, CAF97, DS94, Mas94, BB93b, BES95, BN96, CW97, Cha96b, DMY91, HLL96, KD92, KRW98, Kru98, MW90, Mik98, Rob90a, RW94, Sap95, SB91, Sav94, SV92, Tsa98, Wan98b]. probable [BM96, Tra96]. problem [AC99, Als91a, Baj92, Bel99a, BJZ90, BS92, CC94b, CW94, Coo98, CH96, CATD93, DG97, Fer96, Fie90, FP98, Hen98b, Hol90, HT95, Jan97, KW92, Koc94, KM98, Lee98b, Lin91, MM94, Mun99a, NS96a, NS97a, PP95, RS99, Rot93a, SB99, SW93, Tak91, Wan97, Yao90]. problème [Hol91]. Problems [She98, BVB93, BMvdM97, CG94, ES90, Hal95, Kay92, Kut97b, Lef96, Leo99, Lin97, Lin00, Lin92b, NS96b, Pre93, SF94, Ste90, Szy91, Wan96b, Yat91, Yu93, vEK98]. procedure [Akr93, CWM91, CF90, Fra93, Fro95, Gug98, Ind94, McC90, San92, Som93, SS97c, TN93, Xia95, Yip92]. procedures [BM93, CG94, CB96, DT93, DL91b, DLY96, Gor91, LB91b, MN93, Ped97]. process [ADD93, AK90, AA98a, AA98b, Ani94, Arc92, AHR91, BN97, BM99, BCI98, Ben97, BS93, Cab90, CFF98, Cro92, Cse95, CS94, DLS90, DW92, DK92, Dia90, ET91, Fad96, Fad98, FLP99, FB96, GG91, GLM94, GW96b, Gug98, Ham98a, Ham98b, HZ99, HB99b, Jia91, Joh94a, Kak98, KD92, KD93, KR98, Kep97, KD96, Kul97, KP94b, Kut97b, Kyr94, Lac96, Leb96, Lee96b, Lee97c, Lee97d, Lef97, Leo93, Lo91, MM96, Mud93, Mud94, Neu98, OS98, PR97, Phe95, PJ94, Rao92, Rou91b, RMNV99, Shi97a, SHC99, Sri91, Sta95b, Wal98, WJ93, Wes90, Wes94, Wis92, Yam90]. processes [AB92a, Arc94a, Arc95b, Arc96, Arc98c, Bae97, BR91a, Bal95, BDG98, BD90a, Ber99a, BD94, BB97, BMS97, BR98, Bra99b, BD90b, BD92b, CGT97, CC94a, CS93b, CD94, Che97a, Cho98, CDG93, CF97, DD98, Deb99, DS92, Der95, DR96b, DBD95, Eic98, EN96, EPB94, FZF93, FZL97, FR92, Fan91, Fer99, FP99, FF97, FP98, GKL96, GRZ99, Ger93, GS99, GR93b, GLM94, Gut96, GKS97, Gut97, Hor93, HS94a, HS99, HN98, HCJ97, HL92, Hüs99, HB97, IM97a, JK98, Jev90, JWR92, Kim92, Kim99b, KH91, KS92, Kri90, Kri93, Kro97, KL98c, Lah98, Lee97b, Lee98a, Lef96, LW92, LRW92, Li97, Mag90, MW91, Mik97, Moc94, MM98b, Nav99, Nay93, ND93, PS96a, PARA97, POGP96, PS99a, Pfl91, Phe90, Pou93, Rav92, Rec91, Rog95]. processes [Sam94, Shi95a, Shi95b, SP93, SS99a, Sho97, Sim92, Sta91, Szy94, Vei90, VDO94, WC97, Wan97, Wiś96, Yca92, Yu96, Zha94a, Zha96, Zha97b, ZZ93, dN95]. processing [Bha96]. Product [DB96, Abt94, ASC98, BAC92, Brz97, CY99, ED96, FF97, GR91, PS99b, RB96, ST98a, Swa96, TS96, Yan96b, Zho96]. product-limit [CY99, Yan96b, Zho96]. product-type [GR91]. production [BN97]. products [DW92, Gha96, GAP99, Qui99]. progressivity [IMSK95]. Projection [EM99, AK99, KW93, ZC95]. Projections [LL97, RT93]. projective [Che98b]. proof [CS94, DS96, Dev95, Düm95, Muk95, Neu94, Neu97, RHO99, Rut93, SA93, Sib91]. proofs [Pur98, Yam90]. proper [Kag91, SJK96]. Properties [AM97, AK99, AB92b, Cai98, Cha92a, Che98b, CR91, Dav94, GM99b, GK99, Gla98, Hoa93, JH90, Koc94, Koz91, Kry99, LR96, Mi96, Mi99, NS98, Nyc90, RH91, SE97, Sev98, Sho97, Sto96, TH95, Wes97, WW90b, Wie91, YLB99, YSLW98, Yua96, Zha97b]. property [BB93b, Bol98, CMS94, DS96, Deg93, DMY91, GCL90, GR99, HMT96, hK99, Kou97, KW93, Kul97, Lee96b, MZ97, PJ94, Pro98, Rob95b, Sch91b, Sch99c, SS97c, WC97]. Prophet [Bos96, SC91b, Kla93]. proportion [DJ97, MN97b, SC91a]. Proportional [O’N91, Cha96c, LSR90, MM98a, SG96b, Tso98, Wei90, dUÁGM98, dUÁGM99]. proportions [KS95b, YR96]. proving [Str94]. pseudo [GL95a, Olk98]. pseudo-inverse [Olk98]. pseudo-LS [GL95a]. pure [Roj93]. pure-tail [Roj93]. purposes [Gei98]. pursuit [ZC95]. pursuit-type [ZC95].

q [SS93a, SS94a]. Quadratic [FR93, Krz94, AB95, AS96, Bos95, BS93, El 95, GRZ99, GS95, GT99, GD97, Gha94, Gha96, GV91, GK98, KK96a, Le 92, Lin98, Man99, MO99, PG94c, Sha92a, TS93, Zha94c]. quadrature [YM99]. qualitative [Zel90]. qualites [LY94]. quality [CWM91, MCWD95]. Qualms [PP92a]. quantal [Ger95]. Quantile [PP95, Sém96, Vel93b, AT92, Che95a, Che98d, DJ95, Fal90a, Fal94, Fut99, GSW93, HB99a, KS93, LP91a, Muk92a, Yu96, Zie99]. Quantile-based [Vel93b]. quantiles [AJ95, CR97, Cha90a, Che96, DJ95, Fal90b, FJ92, GH96, GH97, JH90, Jon94, Kni98, KS96, Lah92, Mee93a, MH97b, Ral92, Rob90b, iY95, Zie98]. quasi [cIC99, Hey94, MH97b, SJ97, Tza98]. quasi-likelihood [cIC99, Hey94, Tza98]. quasi-Pólya [SJ97]. quasi-quantiles [MH97b]. quasimartingales [Tso96]. quasimedians [MH97a]. queueing [ADD93, DS96]. Quickselect [LM96]. Quicksort [TH95]. quota [Lin92b]. quotas [ES90]. quotation [CH94b].

radiation [KPTY97, Sai99]. radiation-induced [KPTY97]. Radon [LFR99, MW90]. Random [AGV97, DH97, Far96, Lee97d, PR98a, PRB99, Pal01, AT92, Abt94, ARV97, Als96, ARM97, Arc98a, Arc98d, Aue90, BMT97, BM97, BL91, BK93, BC92, BL94, Beh99, Bél91b, Ber97, BR97, BG97, Bo98, BKS96, Bor97b, Bos96, BD90b, BD92b, CB99, CCHP99, CTW97, Car94, Cek98, Che97b, CR91, Chr90, CS98a, CS98b, CLN95, CM91a, CM91b, Dai98, Das92, DP96, Dev91, Dev97, Due98, Dzi95, EK96b, Ete97, Ete99, FZF93, FZL97, FR93, FD93, FLW99, Fuj96, Gas96, Gen98, Gil92, Git97, GAP99, GB96, GR93b, Gri96, GV91, Gut97, HL97, Hit90, HH96, How97, HH98b, HH99, Hür93, Hüs94, HB97, IR99, IMSK95, JK98, Jes93, Jia91, JWR92, JM90, JK95a, Jum99, Kab99, KM99, KMS96, KP96b, Kre96, Lac96, Lan98]. random [Lec90, Lee95, LMF97, LZ94, LT96, Li97, Lin93, LGC99a, LGC99b, LDG97, Mal96, MP93, MZT90, Mar95, Mas99, Mat92, MS96a, MPS97, Mol93, MM95, Mül97, NS97b, Olk98, Omo97, Pal92, PS95a, Pet95, Pin94, Pog92, PR98b, Qia96a, Qui99, RP95, Ram99, RW99, Rév96, Rhe99, Rie93, Ros93, Rot93b, Rou95, Run93, Rüs96, Rut95, RZ92b, Sai99, ST90, SC91b, SV93, Sch97a, Sch91a, Sco93, Sep96, Sgi97, SQC94, SA97, SY95, Sta95a, Ste93, Sto95, Str94, SS97d, SS98, Sun97a, Sun99, Sut97, SNN97, Szr92, Tan93b, TYS95, Uno96, VA97, VD91, Vel93b, VC99, Vid96, Vie99, WW90a, Wan92b, WRL95, Web91, Wen94, WZ95, WZ96, Wen97, Wer96, WH91, Wu97a, Yan90, Yan96a, Yan91, iY95, Zog94]. randomization [Jen92]. randomized [MR97b]. randomly [GSW93, Har96, Hon96, RS98, RR94b, Rot97, dlPG92]. randomness [Gri92, KA97, RR94a]. range [Bin95, Bin96, CW98, DR96a, GM99a, Gri96, Kou92a, Lah93, Muk94]. Rank [Abr99, Akr92, AQ95, GV91, KW92, LP91b, NM97, RR94a, Akr93, Baj92, BB95b, Cha92b, Erm92, Gom94, Grü94, Hor95, HJ93, Hus97, JM90, Kol95, Mun99b, SQ97, SW99, Wu90, YC90]. ranked [KA99]. ranking [BMvdM97, CG94, Rot93a]. rankings [AC96]. ranks [Koc94, TW92b]. Rao [Muk92b, Gho93]. Rapid [MZM94]. rare [WW92b, Wri97]. Rasch [Gho95]. Rate [Chr90, Chr97, Eic96, MR97a, AA92, AL98, BMT97, BP99a, Bha95, CDG93, CZ97, Fer94, HT95, Jan94, MS90, Mkh98, Nay93, PG94c, Rac95, SS95, Sun97b]. Rate-of-convergence [MR97a]. Rates [BR95, EN96, FH99, Ste90, Arc94b, BW93, BlZ93, BB97, CR92, Che95b, Dai98, ES91, Gas96, Gui96, Kak98, Kau95, KS92, Kor99, Kou92b, Lin93, MCWD95, Muk92a, RA93, Ruk96, Wan99d]. Rating [Joe91]. Ratio [Kla93, AS95, AL91, AL92, DMG90, Das92, EK95, EZ99, FM92, FS93, Gha94, KG90a, KS99b, LP97, Li95, LVC90, MS92, Men93, NNR97, PKM93, Qin97, SS92, SS90b, Wan94a, WZ95, Wic90, YSM94, YL96, ZR96]. ratio-type [AS95]. ratios [Bar99, Bas96a, GK98, Isa99, Kim99a]. reach [KP93b]. real [Ram96, Ram97, Uno96]. recapture [YLB99]. reciprocal [New90, Rec91]. recoding [DGA95]. reconstruction [Gas92, Pog92]. record [BAC92, BC98, Cha98b, Pek99, Raq97]. Records [Adk93, Gne98a, KNN97, Li99, NN96, RCBG96]. Recovery [Zha92a, PS97a]. rectangles [SN96]. Rectangular [Afs90b, CC94b]. Recurrence [Gha96, AFH^{+}96, BB92, Dav95, ST98a, SD90, TS96, WC97]. recursion [Sta97]. recursions [HA98]. Recursive [Rou91b, SS99b, Cho97, GL95b, GW97, Zaj96]. reduces [Yu94]. Reducing [PS99a]. reduction [AC99, Ber95, Brz91, FBCCN96, GS96a, Lai95, LN94, Xia95]. Reference [FS99, dC98]. Refinement [Tom96, Tom97, Düm95]. Refinements [Tom92b, MH97b]. Reflected [Mat97]. reformulation [Lor92, Rie97]. regions [Hor95, HkL92, KM97]. Regression [JS96, JS97, Shi93, Zho99, Abr99, AM98, AC91, Ahm92, Akr96a, Alt93, AQ95, AFH^{+}96, BG99a, BHZ99, BS95a, BKR92, BR92, Bly94, CF99, CER93, CSC93, Cha98c, Cha99c, Cha90e, Che96, Che95b, CH99, CC94c, CM93a, CK94a, CV97, DL91a, Dav94, Del93, Deo97, DZ99, DJ95, DM93, Dro99, ES96, ES91, FNARP97, FS93, FZ92, GS96b, GS93a, GKS96b, Gle93, GW97, HJM98, Hec92, Hoa93, Hol95, Hsi95, HT90, HT96, HS94b, Ioa92, IA97, JH90, JW99, Kay92, Kim92, KHW91, Koh99, KL98b, Kou96e, KW93, KW98, Kul92, Kun93, Lec90, Lee93, LN94, Loh97, MT99, Mof97, Mor94a, Mül97, MS97c, NM97, NY90, O’N94, OB95, Oht96, Pak96, PKJ97, Pau96, PS97a, PP96b, PW92, Qia92, Qia96b, RG95, Rob95a, Rou93, Sch94, Sch96b, Sch96a]. regression [Sch96c, Sch99b, Sch91b, ST91, Sin94, SMP95, Son99, SG91, SS93b, SS94b, SR97, Tia98, TW92c, Vas90, Vel98, Vie99, WK99, WTC90, WC90, Wan99a, Wat95, Wie91, Wie96, Wil98, WS98, Yan96b, Yat96, ZN95]. regressions [Bal97, Bal98, JL96, Sar96, Szr96]. regressors [MS95]. Regular [Mee91, Mee93b, Tom99, Kni98, Mac96, Vie99, Wat93]. regularity [Arc95c, Moc94]. regularly [GN94]. Reinforced [BCK97]. rejective [Mil92]. related [Bel99a, Dev93, FB91, GCL90, Jen94a, Kan92, Mar97a, PW97, Pap93, Rao99b, Sai99, Ses91, Ses92, Shi95b, Tsa98, Yat91]. relating [Buc94]. Relation [KL99, BB92, CC98b, Kou90, Par96, ST98a, SD90]. Relations [BAC92, Als91b, BB94, Bar95b, Dav95, NS91, Neu94, PP92b, Pou93, Sat92, TS96, Yam90]. relationship [Bal93b, Mon93]. relationships [BJK98, Iso95, KM97]. Relative [PS98, Der95, Mur92, Wen94].

Relative-error [PS98]. Reliabilities [CCH99, FL91, Gas94]. Reliability [FK95, Kor97, AB97, Bax93, Gle93, HS91, KPP96, KM93, Lau92, LZ94, Lin92a, PD99, RJ91, Roj93, RG99, SB99, Whi90, XM91]. reliable [Ves94]. remainder [Fal90b, Ral92]. Remark [Mie90, AHZ97, Arc98b, Bra96b, Bry93, Kha92]. Remarks [Bra99b, Leo99, MH97a, Rao90, Arc96, BB93a, Bil96, CS93b]. REML [Hey94]. remote [GR92]. removal [YF93]. Renewal [BvHS99, CDG93, AA92, Als91b, DS92, Fad96, GKS97, KJ91, Li97, New90, Phe90, Sta91, Sze90]. Rényi [LGC99a, BB91a, BB91b, LGC99b, Shi97b]. repair [SC99]. repeated [Afs90a, Akr93, HH98a, Jen92, KNN97, Ras95]. replacement [Don91]. replicated [CER93]. replicates [GHW92]. replication [CM93a]. replications [SJK96]. Representation [Kru98, ARM97, Arc98d, AG94, BP99a, Bel99a, Cha92a, Fal90b, GR93a, KO96, Koz98, Lah92, Lov98, OM92, SD99, Spr98, Ste91a, iY95]. Representations [Kro97, Cha99b, GSW93, Lin95, LY94, Pak98, Pet96b]. representatives [Cha98a]. reproducing [PKJ97]. requiring [TW92a]. resampling [Don91, Ped97, Rou93]. rescue [Mel95]. research [San92]. Residual [WL93, BPRS97, BOR99, FP94, Lee97c, MVA93, Vel98]. Residuals [Ahm92, KW93, PS96a]. resistance [Tan95]. resolution [HN97]. respect [AFM96, Gaw97, Ger93, Mac96, Mar98, Por92, Roo95]. responders [LVCW93]. response [Ger95, JL96, MS99]. responses [Hud97]. Restricted [Bha95, ED94, ED96, AL97b, CS95b, HM96, Mar98, MvdM97, Par98b, Sil97, Sin94, SW99]. restriction [CC94c, Hu97, HH94a, HH94b, Par98a, SW90a]. restrictions [GK97, Hol95, Hu97, HP93b, LT99, Ped97, Por92]. result [Dev95, MM92, Muk95, Neu94, Spr98]. Results [GS91, BS98, Dai98, Dar98, EK96a, ES97, GR93a, Han93, Hüs99, Kou96f, Mas96, Pur98, cQ94, WRL95]. return [Bal95]. reversible [Löw97]. Révész [GW97]. revised [JPAS97]. revisited [Kon95, Koz98, PS95b, Rec91]. reweighting [AM98]. rho [RB96]. Rice [BW95]. ridge [Cha90e, Gru91, Mar96, WC90]. right [Bel99b, Cha90a, CB96, FZ92, PP95, PG99, Rut95, Sun97b, Wil98]. right-censored [Wil98]. right-censorship [Cha90a]. rises [SQ97]. Risk [RA92, AAK97, Cas90, GD97, KPTY97, Pem96, SD99, SS94b]. risks [NS96a, NS96b, NS97a, Rou97]. Robbins [BM99]. Robertson [Muk95]. Robust [CER93, DK90, DB94, GL95b, Kan92, MHL95, Sil96b, Wan99a, WMN94, AM98, CS93a, CM99, CHW93, EJ96, Hic99, Jey91, MGH99, Mar99a, ÖH96, SBB93, Siv99, Tom92a, WW92a, Yan90, Yan96a]. Robustified [JS90, CAGM98]. Robustness [GM99b, Zam90, Zie90, BZ99, Bor96, CMS97, CC99a, Che98c]. ROC [LY99]. Rohatgi [Hen97c]. role [NS96b]. roles [Jon91]. Root [Sch96b, Sch96a, CPA95, JW97, PS96a, SS93a, SP93, SS94a, SSL96, SS97a, Wir96]. Root-n [Sch96b]. Root-n-consistent [Sch96a]. roots [AC93, Ame90, Kle94, Ram99, RA93, SS99a, Uno96]. Rosenblatt [Alb98]. Rosenthal [Lou99]. rotatable [MH90]. rotational [ZK97b]. rounding [Puk97, TN93]. routine [Szr96]. row [Koz91, Man99, TLC97]. row-column [Man99, TLC97]. Rüger [Kru99b]. ruin [CW97, RS99]. rule [KM98, Moj97, Oht96, SS93b, SS94b]. ruled [SM92]. rules [ASC98, BM96, BMvdM97, Cha90b, Irl96, Kou92b]. run [ES90, FM99, God90b, Mus96, Rig95]. runs [Ban90, DY98, God90a, HA98, KA97, Lou97, MG96, Mór94b, Sch91a]. Russo [EN93]. Russo-type [EN93]. Ryzin [FZ92, Zho99].

sacrifice [GV92]. Saddlepoint [JR99, BW95, Mon93, Wan92c]. Sample [Wan98c, ABY94, Baj92, BL99a, Bin95, Bin96, BSW94, CMS94, Cav99, CMC94, EPS94, Elt92, FJ92, Gas92, Gen99, Gne98a, Gut90, HK91, HL96, Hua91, HH94a, HH94b, Irl97, JS97, Jin95, JM90, JV91, KK99, Kir96, Kir98, Kni98, Koc94, Koc99, KS99b, KM98, Lah92, Llo91, McQ99, MSS90, Moc94, Mor94a, Neu94, NOT90, Olo99, Oos94, PP95, Par99, Qi97, Ral92, RCBG96, Sai95, Sam94, Sat92, Sha92a, SH95, SS92, SS97d, SS98, Szr96, TV97, Vas90, Wan92a, Wan98a, WCT94, XY97, iY95, Yu94, ZZ93, Zie98]. sampled [MM98a]. sampler [HRG97]. samples [Ahm95, CB99, CB96, Elf97, HT90, JVS90, LY94, NOT90, Qiy95, Ryc92, SS94a, Vel93b, Woo92, Zie90]. sampling [Cor93b, HH98a, HRS99, Hür95, KA99, KV99, Lac96, Leo96, Loh97, Mat93, Mil92, Nee91, Pog92, Puk97, Rod93, Sah97, SB95a, SP93, Son99, Sri95, SR97, SS97d, SS98, Tal91, WW92b, Zog94]. samplings [Hof97]. Sandwich [Ste97, Wan99a]. Sankoff [CS95a]. Sankoff-type [CS95a]. Sanov [GO99, Sch98b]. satisfying [SG96b]. Satorra [Neu94]. Savage [Kha92]. Scalable [Mur92]. scalar [Le 92]. scale [BSW97, Bas96b, Cha98b, Cro94, EJ96, FS99, KS99b, Kus98, LZ91, LF95b, MHL95, ÖH96, RKD90, Ste91a, WW90b, Wie96, Wu92]. scaled [Bor97b]. scaling [Tom93]. scan [CG96, CG99]. sceneries [How97]. Scheffé [LN95]. scheme [AM98, Bal94, GS93b, Sah97, Str99, WW94, Wu97a]. Schmidt [Bra95, Pér93]. Schrödinger [RT93]. Schur [GR99, NS98]. Schwarz [McQ99]. Score [Sep92, BlZ93, BS96b, CR91, CC99b, FC96, FCUOCN96, Mof97, Mun99b, OM92]. scores [NM97]. scorings [GKS96a]. scrambled [YM99]. screening [Gei98]. search [Gas94, Mel95]. seasonal [AC93, CPA95]. seasonality [Fra92b, Fra93]. Second [FBCCN96, HP93a, Lah91, Lia95, Hud97, Jen94b, Neu94, Sat92, ST98b, WL93, dlPG92]. Second- [FBCCN96]. Second-order [Lia95, Hud97, Jen94b, Neu94, Sat92, ST98b]. sectional [Wir96]. see [LL93]. selected [BSW97, Ind94]. Selecting [CSC93, PG94a, AB95, BM96, Dro99]. selection [BS91, Cav99, CG94, CLK96, DF95, Fra91, Fra92a, Fra93, HT90, HT96, IWLX99, KW98, LH98a, Mam90, MST97, MT99, McQ99, PG98, SV91, WS98]. selections [RG99]. selector [FH99]. selectors [KPM94]. Seleznjev [Hüs99]. self [AHZ97, BLBL92, Qui99, Sam94, Shi95b, Wan99c]. self-decomposability [BLBL92]. self-intersection [AHZ97]. self-normalized [Qui99, Wan99c]. self-similar [Sam94, Shi95b]. Semi [Qin97, Cse95, DK92, DR96b, KD96, Lim93, MRN93, Sch95]. semi-Markov [Cse95, DK92, DR96b, KD96, Lim93, MRN93]. Semi-parametric [Qin97]. semi-stable [Sch95]. semigroup [Roo95]. semimartingales [Ouk92, Ouk94]. seminorms [DMWZ95]. Semiparametric [SS97a, Far90, MM96, Sch96a]. semistable [MS97b]. Sen [Sah97, Wil98]. sense [AM94b, FPMP97, Gam97, Hen97c]. sensing [GR92]. Sensitivity [Sen93, MS96b, PZ97]. Sequence [Hua98, Dar97, DY98, GO97, Lee95, Mik98, PA90, RR91, WZ95]. sequences [Adk93, BM99, Bo98, Bos96, Cek98, Dar98, DZ96, Ete97, GM99a, Gne95, Gne96, HV97, Ho95a, Jak94, Kim93, Kou96a, Kou96d, Kou96c, Kri90, Lac95, LS99, Lou99, PS96b, Rad96, Rad99, Sha93, Sun97a, Wen97, Wiś96, iY95, Yu96, YM99]. Sequential [SZ98, BM99, CG94, Gov97, HkL92, Hwa99, Irl97, IU95, Kor99, Liu99b, Mag90, Nee91, RHO99, Sta95b, Uno95, WW96, dC93]. serial [BL91, HT96]. series [AC93, ACH97, Bat90, Cha92c, CA97, CmHP98, DR96a, Eng90, Fer98, Fra92b, Fra93, Gen99, HJM98, HW90, Kak97, KV96, Li96, LX97, PG94b, Rao95, RJ91, Rot93b, SB97b, SSL96, SS99b, TP98, Wir96, WS98]. server [ADD93]. set [DM95, Dev91, IM97a, KA99, Liu99a, Liu99b]. set-indexed [DM95, IM97a]. SETAR [DD98]. Sets [AS92, CC94b, FF97, IMSK95, JK98, Joe93, Mol93, Qia92, RR94b, Rot97, SBB93, Sto95]. setting [KS99b, MM90]. several [EJ96, HA98, KS95b, Mad93, Sch96d, Sch97b]. shape [Che97c, Mey99, Rév96]. shaped [Ege92]. Shapiro [JV91]. Sharp [Hua97a, Jia97b, Pit90, Pog92, Rom99, TG95, Bal93b]. sheet [FN95, MS96a]. shell [Yin99]. Shift [HZ99, Sah97, Sch99b]. shock [BS97, FP94, KM93]. shocks [Mar97a]. short [EH99a, GM99a, Mam90]. short-range [GM99a]. showing [Imh97, Luc96]. Shrinkage [MS97a]. SIC [McQ99]. sided [Fro98, Liu99b, Sil96b, Sil96a, Str96, Zie90]. sides [Shi93]. Siegel [Liu94a]. sieve [BR92]. sign [Gom94]. signal [Bha96, KS99a, PS97a, RMV96, SS93a]. signals [Ruy96]. Significance [HW97]. Silvey [NF97, RHO99]. Simes [HL94]. similar [Sam94, Shi95b]. Similarity [Lef96]. Simon [FLW99]. Simple [Gne98b, Mus96, Pur98, vEK98, AT96, Aly90, Bal93b, CC94a, CH94a, CC98a, CM93a, Düm95, FD93, Gri96, Gro97, Gug98, GP91, Kyr94, Le 98, Muk95, PKM93, Pak96, PZ97, Rut93, SS92, Sta99, Ste93, SS97d, SS98, Wri97]. simplex [Ses92]. simplicial [Düm92]. simplified [Sib91]. simulated [Dor97a, Löw97]. Simulating [Dev97, Wer96]. simulation [Hol98, Lee99, Rom99]. Simulations [Wil98, Kab99]. Simultaneous [GR91, AB92b, Jin93, KZ96, ÖH96]. since [Hen98a]. single [BAC92, ET91]. Singular [Wan91a, Wan91b, NG95, Rut90, Sgi98a]. Singularity [Zha94b, FH98]. size [ABY94, BK96, Irl97, KM98, RCBG96, Rie93, Sha92a, Wan98c, YF93, YLB99]. sizes [GDD91, SS92, Zha92b]. skewed [Alb93, Alb97, Wer96]. skewness [Hen97c, TG95, XM91]. skip [DR96b]. skip-free [DR96b]. Skorohod [Arc98d, FB95]. slope [Bly94, CS97]. slopes [Gle93]. slow [Dev95]. slowly [ME92, Mat93]. slowly-decaying [ME92]. Small [AA98a, AA98b, MS95, Mor94a, Wan97, HT90, Jen94b, McQ99, Zha92a]. small-sample [McQ99]. smaller [RS91]. smallest [GB96, Ham95]. Smirnov [CA97, Deg93, JPZ97, Kim99b, Kri93, Par92, Vas90]. Smooth [CR97, RM90, RB96, Che95a, DMY91, Koh99, Mi99, MS97c, Rou95, vE97]. smoothed [Ham98a, Ham98b, Rut93]. smoother [BP92, JPAS97]. smoothers [BlZ93, KHW91, Mül97]. Smoothing [FM94, SV91, Abr93, AAJ97, AGV97, CD95, Che94, Fer98, Hen97b, Kay92, Kim96, Nyc90, Shi99, Sun95, WW95a, Wan98c]. smoothness [AHZ97, FN95, Nav99, Shi97b]. solution [BT91, CATD93, Lim93]. solutions [Kar96, Lef96, Mat97, She93, Stu98]. solving [Cho97, Lin91, SF94]. Some [AC93, Als91b, Arc96, AHR91, BHK^{+}99, Bil96, CW97, CS93b, Che93, Che98b, Dai98, DW92, DH90, EK96a, ES97, Fun92, GR93a, Han93, Hei93, KS98a, Kou96e, Kou96f, Koz91, Kut97b, Mar99a, Mas96, MH97b, Nyc90, Ong95, Pap90, cQ94, Sev98, SX97, Shi95b, Sia94, SJK96, TH95, WRL95, Wan99b, WW92b, Wen94, Yua96, ZW97b, AK99, AVB95, BB93a, BDG98, Bis96a, Bla91, Brz95, BD90b, BD92b, Cha99a, CM99, CT99, CS95b, Dal90, DP94, Der95, DM96, DT93, FRS93, Gao95, God90b, Gzy90, Hol95, HN98, HP93b, Jen94b, JWR92, Jia97b, KM90, KM97, Kou95, Kry99, Lan96, Lat98, Lee97b, LB98, Liu96, Lor92, NJS98, NV99, Ouk92, Ouk94, Ram97, Rob90b, Roj93, Sar96, Sil97, Ste90, TP96, TP98, Wei91, Wes97, YB99]. Sooner [BVB93, ES90, Lin92b]. source [WW95c]. space [CS96a, CR98, Che97b, CT99, CT94, Cse95, FM92, Gne99, JK98, KD93, Las97, Lec90, MV95b, NG95, Rot93b, Shi97b, WRL95, WPH96, ZB94]. space-valued [JK98]. spaces [ARV97, Gat93, Gzy90, Ham98b, Hei93, KP92, KP93a, LFR99, Nor95a, Nor95b, Shi97b, WR90]. spacing [Bar92a, YL96]. spacings [Bar95a, ES97, EvZ92, KK99, Kir96, Kir98, Koc99, Sai99, SH95, WJT92]. Sparre [Eis92]. Sparse [AAJ97, PB97, TW92a]. spatial [AT92, AC99, BKR96, BMS97, But99, Cho97, EPB94, KC97a, Kul92, PB97]. spatially [SJ98]. Spearman [RB96, YC90]. Specialised [PJ94]. species [Yat91]. Specific [God90b, Che93]. Specification [ACS96, ACS94b]. specified [JL96, MZ91, NS96c]. spectra [HK95, Li96]. spectral [BMT97, DR96a, Fer98, HL92, Kan97, PC91, Rob95b, Sch99a, Zha92b]. spectrum [Nor95a]. sphere [Yin99]. spherical [FB91, Li93, Li94c, LB98, Yin99]. spherically [Bél91b]. spline [Che94, Kim96, KL98b, Shi99]. splines [Abr93, AB92b, Fer98, Wan98c]. splitting [Als91a, BJZ90]. spread [Bel99b, FPMP97, GR93a, MP90]. spuriosity [LG93]. Spurious [Kim99a]. Square [JK99a, AB92a, EK95, Ioa92, Kal90, Lor92, San92, Sar96, Ser96b, Szr92, Vie99]. squared [Abr93, BS91, Cas90, Elt92, FG92, Hua99, JH90, Lee96b, LL97, Loh97, Mie97, PC91, Rea93, Rob90b, Sai95, Sil96a, WK99, vE97]. squares [BG99a, But99, DJ97, Fro95, GS93a, GK97, Kun93, MP99, MN93, Mül99, Par98a, PL92, Sch96c, Tab94b, Van93, WTC90, WCT94, WL95, WL93, Yat96]. St. [BCC99, CS96b]. Stability [Bor97a, MZ91, MZ97, Rot97, TW92b, BR90, Bes92, BS92, Haw90, Koz97, Sre95, Tom99, UU95]. stabilizing [BLE90]. stable [Alb93, Alb97, BC96, Ber96, BDG93, Buc94, CF96b, CS98b, Dev93, HS94a, Joe90, KMS96, KP96b, Kro97, MR97a, Mar99b, Nol98, PARA97, RQ99, ST90, Sam94, Sch95, Sch99a, Sub94, Tud98, Wer96, dHP99]. stage [Gov97, Hwa99, MS97c, Rod93, Son99, SR97]. staggered [CY99]. standard [CH94b, KS95b, RS91, SV99, TG96]. standardized [GS99]. standards [BCK97]. star [Bar98a]. state [CS96a, CT94, Cse95, DY98, HA98, KD93, MGH99, Mud93, Mud94, NG95, WPH96]. state-space [CS96a, NG95, WPH96]. statements [CH94b]. Stationarity [Lac95, CFF98, Liu94b, PS99a]. Stationary [Kyr94, BR91a, Che92b, DBD95, FZF93, FZL97, Fer98, GRZ99, Hor93, Jak94, Kak97, Lac96, Lah91, Mee99, Nav99, PS99a, Pfl91, PJ94, Wiś96, dN95]. Statist [BL99b, Ham98a, HV00, Lin00, LGC99a, Pal01, Rad99]. statistic [CS97, Cha92b, CC98b, Coo93a, DS92, Don91, Haw90, HJ93, JV91, Kri93, Lor92, Muk92b, OM92, PK96, Run93, SQ97, Sep96, She94, Shi98, Sil96a, ST92b, Szr96, TW92a, Wan94a, Zim94]. Statistical [Che97c, FM92, Gro98, KL97, Kar96, MMPZ99, BS95b, BW99, CT99, De 96b, GGR98, Jia97b, Mac98, MMP97, Rom99, Sto96]. statistics [Ahm97, Akr92, Ano90c, Arc95b, Arc95c, Arc98a, Bab91, BSW97, Bal90, Bal93a, Bal93b, BB91b, BB92, BB94, BB95a, Ban90, BK93, CG92, Cha92a, Che92a, CG96, CG99, CM99, Chr90, Chr91, Chr94, Chr97, CS98a, Chu92a, Chu92c, Dat95, Dav95, DB96, Dav97, Gad97, GC92, Gom94, Gro98, GW95b, HH98a, Hen97a, Hic99, Hos94, HH98b, Hua97a, Hua98, HJ93, Hus97, JVS90, JEG99, Kan92, KRW98, KG90b, Khu97, Kim99b, KP96a, Koc96, Kol95, Lan96, LV99, LM96, Li94b, Lor95, MP93, Mun99b, Nyb92, Pap95, Pap96, Pap97, Pap98, Pap90, Par96, Par99, PP96b, PRS92, cQ94, Qiy95, RZ94, Rea93, Rem98, RSC94, Rob90a, RW94, Ryc92, Ryc95, ST98a, SD90, Sch97a, Sha91, Shi94, SW90b, SBD92, SD93, Tak94, TS96, TW92b]. statistics [Tom99, WW92a, Wan99d, WJT92, Wil96, Wiś96, Woo92, Wu90, ZC95]. status [Abr99, AQ95, KS99b]. Stein [Ber94, BX95, GP91, JS90, Liu94a, Oht96, SS93b, SS94b, SS97c]. Stein-rule [Oht96, SS93b, SS94b]. step [AM98, DT93, GK97, Liu97, Nyc90, RC94, Yan96a, ZZ93]. step-up [DT93, Liu97]. sticky [AK95]. Stieltjes [Gug98]. Stirling [Yam90]. Stochastic [ADD93, Bar95b, Bax97, FF97, KK99, LZ94, LS98, Rut90, SS90a, YLB99, AP92, Arc96, Arc98c, BM97, Bar99, Bel99a, Bis99, Bra96b, Cha96b, CF96a, CS92, Coo96, FLP99, Gat93, Gaw97, Ger91, GM93, GP99, GA99, HM90, Her97, Kar96, Kim99c, Koc99, Leb96, Lee97b, LS97, Lin95, Mat97, Ong95, Pér93, PS99a, Pri99, Rei92, RCD93, RMNV99, ST98b, Sta95b, Stu98, Tan95, Whi90, ZZ93]. Stochastically [Ryc92, DLY96, Gan98, RS91]. stop [Hür93]. stop-loss [Hür93]. Stopped [Gut96, Gut97, Men91, ASC98, dlPG92]. Stopping [Iva96, ASC98, Eth98, Irl96, KM98, Pre93, SC93]. storage [Gom96]. straight [WR98]. Strassen [Che97a, Nor95b, RZ92b]. stratified [GPI99, Zog94]. strictly [Imh97]. Strong [BR92, cIC99, Ger91, HJR93, RR91, Stu92, dUÁGM99, Bos95, CR92, Che95b, CZ97, EG99, Fro98, Gam97, GSW93, GKS97, GW97, Ho95a, KM90, Kim93, KC96, Kri90, Kri93, Lee97d, Liu94b, LL95, MH96, Qiy95, Rad96, Rad99, Rao99a, Rou91a, Sri91, WRL95, Wen94, WZ95, Wen97, Wu99a, Yu96, Zie98, AGMCA96a, Chr90, Chr97, CS96b, EK96b, Ete97, JK98, Mar95, Pet96a, Rem98, SM92, dJ96]. stronger [Eic98]. Strongly [GH96, GH97, CM99, Ho95a, Ma99, Rie93, Tar94, iY95]. Strongly-consistent [GH96, GH97]. Structural [SG96a, HS94b, Kan92, LPB90, NS91, PL92, PL93]. structure [Bae97, Coo96, Coo98, Gen99, LV99, MS97b, PP92b, SR97, TW92c, YR96, YB99]. structures [BS97, DS94, FK95, JDP90, Kor97, KPP96, Mal96, WL95, YSM94]. Stuck [Wer96]. Studentized [Jan97, CW98, FJ92, HJ93]. studies [JK95a, WW95b]. study [Akr93, Ber99b, Buc94, GG91]. subclasses [LB98]. subdiagonal [GW96a]. subdivisions [JK95a]. subject [But99, Ped97, TG96]. subject-years [TG96]. subjects [WW92b]. submanifolds [De 96b, Gzy90]. submartingales [DM95]. Submultiplicative [Sgi97]. suboptimal [RMV96]. subordinated [CM99]. subordination [GM99a]. subordinator [Ber96]. subordinators [BvHS99]. subsamples [AS92]. Subsampling [KS93, DZ96]. subsequences [RZ92b, VD91]. subset [Cse95]. subsets [Gne99, Mac98]. substitution [FH99, Sav94]. subsurvival [GV92]. success [God90a, God90b, HA98, Kou92b, KA97, Mus96]. successive [Mil92, PK96]. sufficiency [GCL90, HM96, Kab98, Sev93, ZR96]. Sufficient [DeR94, DeR96, HH99, LVCW93, Aic98, BES95, KG90b, Khu97, Mar98, Sep94, Str94]. sum [Abt94, Cha92b, dE97, EN96, GO97, Hor95, JK98, Kol95, New90, Sco93, TYS95, WL93, Wu99a, dlPG92]. summary [Siv99]. summation [LH98b, Wu97a]. summing [SBD92, SD93]. sums [Ber97, Bor97b, BW99, CS96b, CM91a, CM91b, DS92, Dzi95, Eic98, Fro98, Hon96, Hür93, Kla93, KJ91, Kre96, Kul97, LT96, LS99, Mar95, Mat92, Pet95, Pin94, Ros93, RS98, Sgi96, SM92, VC99, WW90a, Wan99c, Yan91, ZB94]. sup [Stu94]. sup-norms [Stu94]. super [Kro97, Sri95]. super-Brownian [Kro97]. super-stable [Kro97]. superadditive [GR99]. superadditivity [KL97]. supercritical [Rév96]. superior [HCJ97]. superpopulation [ZL97]. Superposition [Li97]. superprocesses [Kro95]. supersaturated [YL99]. support [JR99, Maj92, Won93b]. suprema [Bra99b, GS99]. supremum [Deb99, FN95, Kep97, Sal93, Sgi97, Sgi98b]. sure [BCH98, BCC99, BlZ93, FS98, Fer94, IL98, LP90, Lat98, Mat92, PS95a, Pet95]. surface [MS99]. surrogate [Sep92]. survey [Sri95, WW92b]. surveys [WCT94]. survival [BR91a, BW93, EK96a, GV92, Har96, Jan94, Li95, LVCW93, LT99, MRN93, PCK98, Pru92, TG96]. survival/sacrifice [GV92]. Susarla [Zho99, FZ92]. Symbolic [ST92a]. symmetric [Bél91b, BDG93, CPS93, Cav99, Cho98, KMS96, MP93, Pak98, Rot93a, SS90a, SS97a, Str94, Tar94, YM95, ZK97b]. symmetrical [SZ92]. symmetrically [LV99]. Symmetrization [Due98]. Symmetrized [Bly94, Bin95]. Symmetry [MV95b, Che96, EK95, Hsi95, Hu95, Koc92, MG96]. symptotic [Rou91b]. system [ADD93, AB97, Bax93, BG95, CF96a, GLM94, HS91, KM93, Le 98, NG95, SC99]. systematic [Man99, PW92, SP93]. systems [BM93, BP99b, CCH99, FL91, GP99, Joe91, Lau92, Lim93, LS98, Lu97, NJS98, Neu96, QT92, RJ91, SS95, SPH91, Spi91, Ves94, WZ96, Whi90]. Székely [Hen97c].

table [EK95, Lin91, Par98b]. tables [AG94, CS98c, EZ99, GPI99, Hen98a, Hub97]. Tactical [Mel95]. Tail [KS99a, KRW98, BD99, CL94, DMY91, Kim99b, KP96b, Kri93, MN97a, Pen99, Per98, Roj93, Tob96, ZC95]. tailes [Bra99b]. tails [Mór94b]. Tailweight [AFM96]. take [LL93]. taking [Ram96]. Tan [Sib91]. tau [Shi98]. Taylor [Lin94, MW93]. technique [Lai95]. techniques [Ano90c, Che94, WW92b]. Temporal [FB96, But99]. tends [TT91]. term [Fal90b]. Test [Bha96, AC91, AAK97, ABK98, ACGM99, AL91, AL92, Aly90, BL91, BH90, BH92, BB90, Bar95b, BW93, Bow99, CA97, CW98, CPPP98, CR91, Coo93a, DT93, Dup97, ES97, Epp99, Fan99, Fra92a, FH98, Ger96, GS93a, GH92, GW96a, HL94, Hor95, Iso95, JPZ97, Kai99, hK99, Kan92, KV99, Koc92, Kol95, Kru99b, KW98, LH98a, LT99, Liu99b, MQ91, MV95a, MG96, Mof97, Mud93, Mud94, NNR97, NS92, Oht99, Omo97, Par99, RW94, SC91a, ST96, ST98b, SB94, ST91, Szr96, Vas90, WW96, WW92a, Wan99c, Wei92, YSM94, YL96, Zie90]. test-of-independence [Kan92]. tested [Haw97]. testimator [Cha90c]. Testing [AL97a, ABK98, Bal97, Bal98, Che94, Del93, DB97, Gom94, GW96b, GMS94, Hee91, KHW91, KS95b, MM91, MM98a, NY90, Par98b, Sch96d, Sch97b, SC90, SP93, SW90a, SP99, AC91, BN97, Bas96a, Che96, CH98a, Ger96, Haw90, Hen97b, Irl97, Kak97, Kle94, Lu97, MM94, Mun99a, Pau96, RR94a, SS93a, Sil96a, TP98]. Tests [BHB96, Joe93, MM96, PS94, RW94, WJT92, ADEM94, AC93, AL97b, Baj92, Bar92b, Bar95a, BS94c, BM99, Bat90, Bha95, BB95b, Bur93, Car92, Cha90a, CPA95, CS91, CS95b, CC99b, DZ99, EK95, EZ99, Erm92, FC96, FCUOCN96, Gei98, Gne98b, HK91, HW97, Jan97, JW97, Jia97b, JM90, JK95b, KK92, KM97, KA97, Kru99a, Lee96b, LP97, Lia99, Liu97, Lou97, MS96a, MS92, MR99, NS96c, RR94a, RM90, Sep92, Ser96b, SS94a, SSL96, SS97a, SS99a, Sil96b, SS92, SW99, Wat93, Wei91, Wir96, Yao90, Zel90, Zha92a, ZR96]. th [DH90, KK96a, Lee98a, Rao99a]. th-order [Lee98a]. Theil [Wil98]. their [AdlC93, BL99a, BKS96, Che96, HL92, KK99, Koc99, Kro95, LV99, Li94a, Per98, RSC94, Rob90a]. theorem [ARV97, BB91b, Bar98a, Bel99a, BR97, Bra95, Cek95, Cek96, Che98d, Cor93b, Dav93, Ege92, Eis92, Eth98, FB95, Gam97, GO99, JdKP95, KK95, Kri90, Lan98, Lin94, Liu94b, MR97a, MVA93, MW93, MW97, Neu97, Pek99, Rei92, SA93, Sch98b, SF94, Ste97, Swa96, Szy94, WW90a, Wan92b, WW95c, WZ96, Wen97, Wu99b, ZB94, Arc94a, Arc95c, BC96, BP93, BP92, Bry93, FZF93, FZL97, HH98a, Hor90, IL98, Joh94a, LP90, Lee97d, Mik98, NF97, PS95a, Qui99, Rog95, Sai95, Yan94, dN95]. theorems [AGMCA96b, AHR91, BCC99, Düm92, EH99a, EvZ92, FS98, GKS97, HS94a, Hus97, HkL92, Kon95, Lee97b, MP93, Nob95, Sai99, SH95, Tan93b, TW92b, WRL95, Wen94, BCH98, Mik97]. theoretic [ZK93]. theoretical [DS96]. Theory [HS91, Bax93, BvHS99, Cse93, Eic96, Gom96, Iof93, Kun93, LZ94, PS95b, Sch91a, Sha90, SW92, YB99, Zha94a]. thermal [Gom96]. theta [Dev97]. thin [Gne99]. thinned [Ben97]. Third [Hua99, TS93, FBCCN96]. Third-order [Hua99, FBCCN96]. Thomas [Dal90]. Thompson [Cor93b]. Three [YL99, SR97]. Three-level [YL99]. Threshold [IWLX99, Lea91, CFF98, Che98a, FP98, Rei92]. thresholding [LH98a]. thyropathy [KPTY97]. tie [BES95]. ties [LP91b, Mun99b]. tighter [Wri97]. tightness [EG99, GGR96, Iva96]. tilting [Jin96]. time [ADD93, AC93, AHZ97, ACH97, BVB93, BP99a, Ban90, BDG98, Bar95b, Bat90, BB97, cIC99, Cha92c, CA97, Che97a, CL98, CmHP98, Coo93a, CS94, DR96a, DK92, ES90, ES97, Fer98, FR93, Fra92b, Fra93, Gen99, Ger96, GV92, GLM94, HJM98, HZ99, Kai99, Kak97, Kat97, KD96, Le 92, Leb96, Lee99, LSR90, Li96, LX97, Lin92b, Mos94, NS92, Pal92, PS99a, PG94b, Sal93, SC99, SB97b, SSL96, SH99, SS99b, Sta91, TP98, VDO94, Wal98, Wir96, WS98, Yan96b, Yin99]. time-ordered [Mos94]. time-series [HJM98, SS99b, TP98, Wir96]. time-to-tumor [GV92]. time-varying [Kat97]. times [Als91b, DR96b, GCL90, GAP99, Gut96, Iva96, KW97, Kro95, New90, PT96, POGP96, Pri99, Shi95b, Sho97, Sim92]. TJW [Zho96]. Toeplitz [Shi99]. together [Sta91]. Too [KV99]. tool [PZ97]. topological [Maj92]. topologies [Eic98]. topology [EG99]. Total [Kai99, LR96, Bar95b, Coo93a, ES97, Kag90, Kyr94, Nay93, NS92, Pem96]. totally [Alb93, Alb97]. trace [GP92]. tracking [Ber99a]. trade [SW92]. trade-off [SW92]. training [KM98]. Trajectories [Dar98]. transfer [Yan96b]. transference [CATD93]. transform [Akr92, Akr93, Ang96, Bar95b, Bar98b, BV92, DP96, GR99, Gug98, MW90, NS92, PG94b, Sav94, Shi93]. transform-both-sides [Shi93]. transformation [Shi93, Vel93a, Vel93b]. Transformations [MS96a, BLE90, Web91]. transformed [Sah97]. transforms [Bar99, Bos96, Hür93, JN97]. transient [Lim93]. Transition [CS92, Phe90, Rou91b, Wen94]. transitive [PR98a, PRB99, Pal01]. transposition [EZ99]. tray [JK99a]. treatment [LVCW93, MVA93, PP95]. treatments [BB95b, Man99, MM95]. tree [Mun99a, Qia96b, WM97]. tree-ordered [Qia96b]. trees [BG97]. trend [Sch94]. trending [Dav93]. trends [AC93, CPA95, SP99]. trials [BVB93, DY98, ES90, Lin92b, MR97b, Mus96, ZRB98]. triangular [ADEM94, Neu96, Str96, ZB94, dJ96]. trigonometric [ES91]. Trimmed [CAGM98, JK99b, Oos94, Arc95a, CR95, CS96b, DB96, DJ97, JVS90, Tab94a, Tab94b, Van93]. trimming [DJ97, Gor91]. triples [Koc92]. triptych [Dev93]. trivariate [Lai95]. true [FM92]. truncated [Gui96, GSW93, Kat99, MQ91, OB95, PCK98, PG99, Sun97b, SZ98, Woo95, Yua96, Zho96]. Truncating [Yu94]. truncation [Gu95, JM90, Pog92, Sha92b]. truth [GS91]. TTT [NS98]. TTT-plot [NS98]. Tukey [BP92, Rea93, Som93]. tumor [AFH^{+}96, GV92]. Twice [JPAS97]. Two [CG96, EPS94, Hwa99, Iof93, Jev90, Jin95, JM90, MS97c, Tar94, AC91, ABK98, AEE92, Baj92, Bal95, Bra96a, Brz95, Cab90, Cha96a, CM95, CF96a, CW98, Coo96, CH98b, DLY96, Fer97, Gei98, Gha94, Gov97, Gro98, GK97, Gut97, HA98, Hen97a, Jon95, KG90b, Kim92, KHW91, Koc94, KS99b, Kul92, KM93, Lee93, Lef96, Liu96, Liu99b, Lu97, MM91, MSS90, NS96c, Omo97, Ong95, PP95, Par98a, PL92, Pur98, Qin97, Rod93, RW94, Sgi98a, Sia94, Son99, Spi91, Sta91, Str96, SR97, Stu97, Tom93, Ves94, Wan96b, Wie91, YL96, Zha94b, ZZ93]. two-component [Lu97]. two-components [Spi91]. Two-dimensional [CG96, Bal95, CH98b, Gut97, Kul92, Lef96]. two-fold [KG90b]. two-level [PL92, Tom93]. two-parameter [AEE92, Cab90, Fer97, Wan96b, ZZ93]. two-part [Bra96a]. two-phase [Kim92, Lee93]. Two-sample [Jin95, Baj92, Koc94, KS99b, PP95]. two-sided [Liu99b, Str96]. Two-stage [Hwa99, MS97c, Gov97, Rod93, Son99]. two-state [HA98]. two-way [Brz95, Cha96a, CW98, Sia94]. type [ARV97, AS95, Arc95b, AGMCA96b, Baj92, BM99, BV92, BH97, BH99, BS98, Cha98c, CS95a, CA97, CL98, Che98d, CC98b, EvZ92, EN93, Fer98, GS96a, GR91, Gru91, Isa99, KMS96, KZ91, Kut94, Lee96b, Lee97a, Li96, Mag90, MW91, MSS90, Muk92b, Nag99, Pit90, Rao92, ST98b, Sib91, Sil96a, TG96, Vas90, Wan94a, Wu97a, XY97, ZC95, ZN95, Zim94, CB96, Ger95, GK99, JV91]. Type-I [Ger95, GK99]. Type-II [CB96]. types [FB95, FTV98, Neu97]. typical [AS92].

U [LH96, Nee91]. U-estimability [Nee91]. Uhlenbeck [KP94b, Lef97, WC97]. ultrastructural [SS97b]. UMP [KV99]. UMPU [Mud93, Mud94, Omo97]. unbalanced [Fuj96, KG90b]. Unbiased [AS95, Jia97a, KP94a, Ryc90, BSW97, Ber90, CS91, CK94a, DF95, FK94, Kut97a, Mun99a, Pen98, Rod93, SC91a, SS90b, TN93]. unbiasedness [Jia99, Tak91]. uncertainty [EK96a]. Unconditional [AGMCA96b, AGMCA96a, Gla98]. unconfounded [Man99]. uncontrolled [Leo96]. Unconventional [Rou94]. Unequal [SS92, Cha90a, GDD91, Leu96, PG94a, SJK96, SR97, TYS95, YSM94, Zha92b]. unexpected [Zam90]. unified [Iso95, Moh92]. Uniform [CR92, CS98c, KC96, Lec90, Sco93, Zie98, AdlC93, Afs90a, Arc96, AG94, BL99a, BlZ93, CB99, Cek95, CM99, CY99, DP96, DvC95, DMLP97, EvZ92, Gov97, Ham95, HJR93, Ho95a, Hor90, Hu97, Hua97b, Hür95, JK98, KM90, Kim99b, Lug95, Moh92, ND93, Nob95, OS98, Rou91a, RKD90, Sri91, Wie91]. uniformity [BH90, Joe93, Wat93]. Uniformly [SS90b, RS91]. Unifying [Cav97]. Unimodal [Kan97, AM94a, AFM96, Bél91a, Bru92, Gan98, Ma99, Par98b, Pur98, Rot93a, Tar94, TG95]. unimodality [BT95, DeR94, DeR96, Mas96]. unique [IK92]. Uniqueness [LF95a, AEE92, BT91, Che92b, Fer99, Fie90, Tza98]. Unit [CPA95, SS94a, SSL96, AC93, JW97, Jia91, PKM93, PS96a, Ses92, SS93a, SP93, SS97a, SS99a, Wir96]. unit-root [Wir96]. univariate [Chu92b, LF95a, Pou93, Rao90, Sub94]. Universal [GDD91, GW97, Kus98, MZT90]. universally [TLC97]. universe [Cor93b]. universes [Wri97]. unknown [BH90, Irl96, LZ91, RCBG96]. unobservable [Szr92]. unobserved [Nay93, SB97a]. unordered [SY95]. unseen [FTV98]. unstable [BN97]. unsuspected [HT96]. until [Sta91]. unusual [Arc94b]. updating [HRS99]. upon [ED94, ED96]. upper [AM94a, Ham95, Jia97b, Kag91, Lug95, MQ91, Wri97]. urn [Lee99, SC93, Sta99]. Use [AS96, JK95a, Jen92, Kuk99]. used [AB97, DW92, Gei98]. Usher [Gas94]. Using [Fer98, KM98, PGTD92, ABK98, AD99, BK93, Bas96a, Ber99b, Ber94, CC98a, DB94, Fad96, FM92, FNARP97, GD97, HN97, Hug97, Ind94, KHW91, Kou96c, Mas94, Oht96, Ped97, Ras95, Sah97, WJT92, vRR98]. usual [Gru91, ZL97].

validation [CD95, Dro99, GHW92]. validity [Gne98b, Mar98]. value [Car92, Deh91, Gom96, Hua97a, Hub97, Jes93, Joe90, Nad99, Pen98, Pro98, Wan91a, Wan91b]. valued [Ger91, GLM94, JK98, Jes93, Jum99, Lec90, Pér93, Rot93b, Shi97b, WRL95, WZ95, Wen97, Yca92, ZB94]. values [ASC98, BAC92, BC98, Cha98b, Che97b, CP97, KP96a, PK96, Pek99, Ram96, Raq97, SS96a, San92, SQC94, Tho97, Zim94]. Variability [BDS99, Her97, KC97b]. variable [Cse93, GV91, IWLX99, Jes93, KL98b, Lec90, PG98, Rut95, Sah97, SS96a, Sch98a]. variables [Abt94, Akr96b, Als96, AC97b, Arc98a, Arc98d, Bab91, BK93, BS94b, Beh99, Ber97, BR97, Bo98, BKS96, Bor97b, Bos96, Cek98, Cha99b, Che97b, Chr90, CS98a, CS98b, CLN95, CM91a, CM91b, Dzi95, EK96b, Ete99, Gen98, HL97, Hit90, HH98b, HH99, HkL92, IA97, IR99, Kab99, KC97a, KM99, KJ91, KP96b, Kre96, Lan98, LPB90, Lee95, LZ94, LT96, Lin93, LGC99a, LGC99b, LDG97, Mal96, MP93, Mar95, Mat92, Ong95, PG94a, PS95a, Pet95, Qui99, Rie93, RSC94, Ros93, Run93, Rüs96, Ryc95, ST90, SC91b, Sch91a, Sco93, SC90, SBD92, SD93, Sto95, Sun97a, Sun99, Tsa98, TYS95, VC99, Vit96, WW90a, WRL95, Web91, WZ95, WZ96, Wen97, Wer96, Yan91, iY95, ZK97a, ZL97, dlPG92]. Variance [AM94a, AV97, Chu92c, MR97b, PG99, AB95, AS95, ABK98, BLE90, BLB90, Bas96b, Ber95, Ber94, Bha96, BR92, CR95, Cha91, DMG90, Das92, DB96, DB97, dE97, FLP99, GS95, Gen98, Hou97, Hua91, IK92, KS98b, Kag91, KZ96, Kni91, LZ91, Mad93, MRN93, Mas99, MM95, Neu94, OM92, Oht96, PL95, PP96a, PS96b, Pit90, RA92, Run93, ST92a, Sat92, Sha92a, Sha91, Sha92b, SS99a, SJK96, SH99, SS90b, Sut97, Tom92a, WK99, Yu94, YM99, ZW99]. variance-balanced [Kag91, SJK96]. variance-covariance [ST92a]. variances [AC91, Gho93, Leu96, MvdM97, SC90, SR97]. variancies [O’N92]. variant [ST96]. variants [JV91]. Variate [LSR90]. variates [CG92, Dev97, LH96]. variation [BS96a, But99, Eis92, FR93, Hür95, Mee91, Mee93b, Pem96, PARA97, SK93]. variations [AK95, Jon92]. varietal [DGA95]. variogram [DP96, Mül99]. various [KC97b, Liu94b, Qin97]. varying [GN94, Kat97, Rem98, Rhe99, TW92b]. vector [AT92, Fox97, Ger91, Kak97, KP94a, Mee91, Pap96, Pap97, Pap98, SB97b, Tak94, Yip92, dN95]. vector-normed [Mee91]. vector-valued [Ger91]. vectors [Ete97, HH96, Hüs94, KMS96, Liu96, Pin94, SV93, Sep96, VA97]. venereal [Yip92]. Verifying [CmHP98]. version [CAGM98, Dor97a, JS90, NOT90]. versions [Sai95]. versus [Fra92a, He94a, He94b]. vertices [Wat93]. very [Zha92a]. via [AW94, ACS94b, Bél91b, BG99b, cIC99, Che94, Fan91, GC92, How97, Jum99, KW96, Liu94a, Mal96, Mar99b, PL98, Rei92, Rie97, SV99, Sap99, Sch91a, Vid96, Vit96, Wan97, Wu97a]. view [Sch97c, Yat96]. viewpoint [Hol91]. visits [Cse95]. visual [CF98]. volume [Ano90b, Ano90a, Ano92c, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano94c, Ano94a, Ano94b, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano99b, CHW93, GB96]. volumes [Ano91c]. vs [GMS94].

waiting [BVB93, Ban90, ES90, Lin92b]. Wald [BW99]. walk [Aue90, FR93, FD93, Gri96, Jum99, LMF97, Rév96, Sgi97, Sta95a]. Walker [Cho97]. walks [Bél91b, Bor97b, CCHP99, DH97, Dai98, GAP99, Gut97, How97, Pal92, PR98a, PRB99, Pal01, Tan93b, Wan92b]. Watson [Ber99a, Gla98, Rao92, ZK97b]. Wavelet [CB99, GT99, Hua97b, KK96b, Leb96, AGV97, HN97, HKP99, Kat99, LX97]. wavelets [IWLX99, KP93a, Mas94, Vid96, Wan97]. way [Brz95, Cha96a, CW98, Das92, Sia94, Sut97, Tom92a]. Weak [Arc98c, BAV98, Fal90a, Fal90b, GSW93, Ham98b, KH91, MS90, OS98, ARV97, Arc96, AGMCA96b, BW97, GAP99, GR93b, Hei93, KW97, Led95, Pen96, Rav92, Tan96, Gut92, HO95b, Hon96, Sun98, Sun99]. weakly [Bae97, BD90b, BD92b, CGT97, JWR92, Lah98, Lin93, Moc94, Nav99, PS95a]. Weibull [Che97c, GK99, MQ91, PD99, WS90]. weighing [Kag90, Kou95]. weight [CM95, Kag90]. Weighted [AJ95, Ber99a, Sch96c, Yan96b, AM98, Ahm95, Ahm97, Bho92, EvZ92, EN96, Fro95, IZ94, Kro95, Lah98, LT96, LS99, Rao95, Rig95, RS98, SQ90, Shi98, Stu94, WS98, Wu99a]. weights [Hol95, Jev91, Sil96a, Yu94]. Weinberg [SPH91]. Weyl [LL97]. Where [dE97, CL98]. which [Bab91, BW95]. white [ES96, EH99b]. Whitney [Zie90, dE97]. width [Gov97, MX96]. Wieand [KK95]. Wielandt [Düm95]. Wiener [Cab90, CS94, Jum99, Lef96, Shi95a, Wes90, Wes94, Wis92]. Wilcoxon [Cha92b, Hor95, Zie90]. Wilk [JV91]. will [ZZ93]. winding [Dor97b]. Windings [Bél91b]. window [KS99c]. windows [Kan97]. winner [MR97b]. Winsorized [JVS90]. wise [Arc98a]. Wishart [LSZ98, Tsa95]. without [Cha99c, Don91]. Worthy [MZ93]. x [Gel96]. Xu [Sib91].

Y. [Jon95]. Yang [DM93]. year [Bal95]. years [TG96]. Yin [FL98]. Youden [DG97]. Yu [FL98]. Yule [Cho97].

zero [AK95, Har91, Kou96a, Kou96d, Kou96c, Tom96, Tom97, WW95a]. zeroone [Tom96, Tom97]. zeros [Rie97]. Zhuan [FL98]. Zygmund [Bo98].

References

Abouammoh:1992:RFR

[AA92] A. M. Abouammoh and A. N. Ahmed. On renewal failure rate classes of life distributions. Statistics & Probability Letters, 14(3):211–217, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290024Y.

Alvarez-Andrade:1998:SDPa

[AA98a] Sergio Alvarez-Andrade. Small deviations for the Poisson process. Statistics & Probability Letters, 37(2):195–201, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700117X.

Alvarez-Andrade:1998:SDPb

[AA98b] Sergio Alvarez-Andrade. Small deviations for the Poisson process. Statistics & Probability Letters, 37(3):279–285, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001284.

Aerts:1997:SCS

[AAJ97] Marc Aerts, Ilse Augustyns, and Paul Janssen. Sparse consistency and smoothing for multinomial data. Statistics & Probability Letters, 33(1):41–48, April 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001083.

Albers:1997:ATL

[AAK97] Willem Albers, Gerda R. J. Arts, and Wilbert C. M. Kallenberg. Accurate test limits under prescribed consumer risk.

Statistics & Probability Letters, 34(2):141–149, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001757.

Adke:1992:MCS

[AB92a] S. R. Adke and N. Balakrishna. Markovian chi-square and gamma processes. Statistics & Probability Letters, 15(5):349–356, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290152U.

Atilgan:1992:CPM

[AB92b] Taskin Atilgan and Hamparsum Bozdogan. Convergence properties of MLE’s and asymptotic simultaneous confidence intervals in fitting cardinal B-splines for density estimation. Statistics & Probability Letters, 13(2):89–98, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290082G.

Abughalous:1995:SBN

[AB95] Mansour M. Abughalous and Naveen K. Bansal. On selecting the best natural exponential families with quadratic variance function. Statistics & Probability Letters, 25(4):341–349, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002397.

Arizono:1997:BSR

[AB97] H. Arizono and V. C. Bueno. Bounds on system reliability of used components under MIFR / T_{⊔} assumption. Statistics & Probability Letters, 36(1):49–57, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000485.

Albers:1998:TET

[ABK98] Willem Albers, Pieta C. Boon, and Wilbert C. M. Kallenberg. Testing equality of two normal means using a variance pretest. Statistics & Probability Letters, 38(3):221–227, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000182.

Abramovich:1993:AMS

[Abr93] Felix P. Abramovich. The asymptotic mean squared error of L-smoothing splines. Statistics & Probability Letters, 18 (3):179–182, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293902133.

Abrevaya:1999:RRC

[Abr99] Jason Abrevaya. Rank regression for current-status data: asymptotic normality. Statistics & Probability Letters, 43(3):275–287, July 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002673.

Abt:1994:PSR

[Abt94] Markus Abt. On the product and the sum of random variables with arithmetic and non-arithmetic distributions. Statistics & Probability Letters, 19(4):291–297, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901791.

Adrover:1994:EME

[ABY94] Jorge G. Adrover, Ana M. Bianco, and Víctor J. Yohai. Efficiency of MM- and τ-estimates for finite sample size. Statistics & Probability Letters, 19(5):409–415, April 11, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900094.

Ahmad:1991:LMP

[AC91] Manzoor Ahmad and Yogendra P. Chaubey. Locally most powerful test for testing the equality of variances of two linear models with common regression parameters. Statistics & Probability Letters, 11(2):149–153, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190133C.

Ahn:1993:STU

[AC93] Sung K. Ahn and Sinsup Cho. Some tests for unit roots in seasonal time series with deterministic trends. Statistics & Probability Letters, 16(2):85–95, January 27, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293901518.

Alvo:1996:AIB

[AC96] M. Alvo and P. Cabilio. Analysis of incomplete blocks for rankings. Statistics & Probability Letters, 29(2):177–184, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001719.

An:1997:NEN

[AC97a] H. Z. An and S. G. Chen. A note on the ergodicity of non-linear autoregressive model. Statistics & Probability Letters, 34(4):365– 372, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296002040.

Anderson-Cook:1997:EMC

[AC97b] C. M. Anderson-Cook. An extension to modeling cylindrical variables. Statistics & Probability Letters, 35(3):215–223, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000163.

Aletti:1999:CSP

[AC99] G. Aletti and V. Capasso. Characterization of spatial Poisson along optional increasing paths: a problem of dimension’s reduction. Statistics & Probability Letters, 43(4):343–347, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002685.

Alcala:1999:GFT

[ACGM99] J. T. Alcalá, J. A. Cristóbal, and W. González-Manteiga. Goodness-of-fit test for linear models based on local polynomials. Statistics & Probability Letters, 42(1):39–46, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001849.

An:1997:GEE

[ACH97] Hongzhi An, Min Chen, and Fuchun Huang. The geometric ergodicity and existence of moments for a class of non-linear time series model. Statistics & Probability Letters, 31(3):213–224, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000338.

Arnold:1993:MDG

[ACS93] Barry C. Arnold, Enrique Castillo, and Jose María Sarabia. Multivariate distributions with generalized Pareto conditionals. Statistics & Probability Letters, 17(5):361–368, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390256I.

Arnold:1994:CCM

[ACS94a] Barry C. Arnold, Enrique Castillo, and Jose María Sarabia. A conditional characterization of the multivariate normal distribution. Statistics & Probability Letters, 19(4):313–315, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901821.

Arnold:1994:MNC

[ACS94b] Barry C. Arnold, Enrique Castillo, and JoséMaría Sarabia. Multivariate normality via conditional specification. Statistics & Probability Letters, 20(5):353–354, August 9, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901260.

Arnold:1996:SDC

[ACS96] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Specification of distributions by combinations of marginal and conditional distributions. Statistics & Probability Letters, 26 (2):153–157, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000054.

Allen:1999:EIP

[AD99] Michael R. Allen and Somnath Datta. Estimation of the index parameter for autoregressive data using the estimated innovations. Statistics & Probability Letters, 41(3):315–324, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001679.

Abolnikov:1993:SAC

[ADD93] Lev M. Abolnikov, Jewgeni H. Dshalalow, and Alexander M. Dukhovny. Stochastic analysis of a controlled bulk queueing system with continuously operating server: continuous time parameter queueing process. Statistics & Probability Letters, 16(2):121–128, January 27, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390156D.

Abu-Dayyeh:1994:CIT

[ADEM94] Walid A. Abu-Dayyeh and Abed El-Qader El-Masri. Combining independent tests of triangular distribution. Statistics & Probability Letters, 21(3):195–202, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901155.

Adke:1993:RGM

[Adk93] S. R. Adke. Records generated by Markov sequences. Statistics & Probability Letters, 18(4):257–263, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293900139.

Adell:1993:UCN

[AdlC93] JoséA. Adell and Jesús de la Cal. On the uniform convergence of normalized Poisson mixtures to their mixing distribution. Statistics & Probability Letters, 18(3):227–232, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293902214.

Aragon:1992:EUM

[AEE92] Jorge Aragón, David Eberly, and Shelly Eberly. Existence and uniqueness of the maximum likelihood estimator for the two-parameter negative binomial distribution. Statistics & Probability Letters, 15(5):375–379, December 8, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290157Z.

Asselain:1996:PRM

[AFH^{+}96] B. Asselain, A. Fourquet, T. Hoang, A. D. Tsodikov, and A. Yu. Yakovlev. A parametric regression model of tumor recurrence: An application to the analysis of clinical data on breast cancer. Statistics & Probability Letters, 29(3):271–278, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001824.

Averous:1996:TRM

[AFM96] J. Averous, A.-L. Fougères, and M. Meste. Tailweight with respect to the mode for unimodal distributions. Statistics & Probability Letters, 28(4):367–373, August 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001476.

Afsarinejad:1990:CBU

[Afs90a] Kasra Afsarinejad. Circular balanced uniform repeated measurements designs, II. Statistics & Probability Letters, 9(2):141–143, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290008S.

Afsarinejad:1990:RLD

[Afs90b] Kasra Afsarinejad. Rectangular lattice designs. Statistics & Probability Letters, 10(4):279–281, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900426.

Arnold:1994:UMR

[AG94] Barry C. Arnold and D. V. Gokhale. On uniform marginal representation of contingency tables. Statistics & Probability Letters, 21(4):311–316, November 11, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000247.

Arenal-Gutierrez:1996:USL

[AGMCA96a] Eusebio Arenal-Gutiérrez, Carlos Matrán, and Juan A. Cuesta-Albertos. On the unconditional Strong Law of Large Numbers for the bootstrap mean. Statistics & Probability Letters, 27(1):49– 60, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000437.

Arenal-Gutierrez:1996:UGC

[AGMCA96b] Eusebio Arenal-Gutiérrez, Carlos Matrán, and Juan A. Cuesta-Albertos. Unconditional Glivenko–Cantelli-type theorems and weak laws of large numbers for bootstrap. Statistics & Probability Letters, 26(4):365–375, March 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500033X.

Antoniadis:1997:RDW

[AGV97] A. Antoniadis, G. Grégoire, and P. Vial. Random design wavelet curve smoothing. Statistics & Probability Letters, 35(3):225–232, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000175.

Ahmed:1991:PPD

[Ahm91] S. E. Ahmed. To pool or not to pool: The discrete data. Statistics & Probability Letters, 11(3):233–237, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190148K.

Ahmad:1992:RDE

[Ahm92] Ibrahim A. Ahmad. Residuals density estimation in nonparametric regression. Statistics & Probability Letters, 14(2):133– 139, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290077I.

Ahmad:1995:MKE

[Ahm95] Ibrahim A. Ahmad. On multivariate kernel estimation for samples from weighted distributions. Statistics & Probability Letters, 22(2):121–129, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400057F.

Ahmad:1997:GFS

[Ahm97] Ibrahim A. Ahmad. Goodness-of-fit statistics based on weighted L_{p}-functionals. Statistics & Probability Letters, 35(3):261–268, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000217.

Auer:1991:SLT

[AHR91] P. Auer, K. Hornik, and P. Révész. Some limit theorems for the homogeneous Poisson process. Statistics & Probability Letters, 12(2):91–96, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190051R.

Albeverio:1997:RNS

[AHZ97] Sergio Albeverio, Yaozhong Hu, and Xian Yin Zhou. A remark on non-smoothness of the self-intersection local time of planar Brownian motion. Statistics & Probability Letters, 32(1):57–65, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000569.

Aickin:1998:LFM

[Aic98] Mikel Aickin. A logical foundation for the minimal sufficient cause model. Statistics & Probability Letters, 39(3):271–281, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000662.

Aerts:1995:WBQ

[AJ95] Marc Aerts and Paul Janssen. Weighted bootstrapping for Uquantiles. Statistics & Probability Letters, 22(4):317–323, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400083K.

Aldous:1990:BAP

[AK90] David Aldous and William B. Krebs. The ‘birth-and-assassination’ process. Statistics & Probability Letters, 10 (5):427–430, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715290900242.

Amir:1995:LIL

[AK95] Madjid Amir and Frank B. Knight. Law of the iterated logarithm and local variations at zero of the sticky Brownian motion. Statistics & Probability Letters, 23(3):289–295, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400127T.

Aggarwal:1999:HPP

[AK99] M. L. Aggarwal and Renu Kaul. Hidden projection properties of some optimal designs. Statistics & Probability Letters, 43(1):87– 92, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002491.

Akritas:1992:RTS

[Akr92] Michael G. Akritas. Rank transform statistics with censored data. Statistics & Probability Letters, 13(3):209–221, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290050F.

Akritas:1993:LRT

[Akr93] Michael G. Akritas. Limitations of the rank transform procedure: a study of repeated measures designs, Part II. Statistics & Probability Letters, 17(2):149–156, May 26, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293900098.

Akritas:1996:NEH

[Akr96a] Michael G. Akritas. Nonparametric estimation in heteroskedastic regression. Statistics & Probability Letters, 28(1):23–31, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000771.

Akritas:1996:GLU

[Akr96b] Michael G. Akritas. On the grouped LSE under an errors-in-variables model. Statistics & Probability Letters, 28(2):181–189, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001158.

Allaire:1991:LRT

[AL91] Jérôme Allaire and Yves Lepage. On a likelihood ratio test for independence. Statistics & Probability Letters, 11(5):449–452, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190194V.

Allaire:1992:LRT

[AL92] Jérôme Allaire and Yves Lepage. On a likelihood ratio test for independence. Statistics & Probability Letters, 13(2):167, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290092J.

Ahmad:1997:TIN

[AL97a] Ibrahim A. Ahmad and Qi Li. Testing independence by nonparametric kernel method. Statistics & Probability Letters, 34(2):201– 210, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001836.

Arteaga:1997:CCB

[AL97b] Carmen Arteaga and Johannes Ledolter. Control charts based on order-restricted tests. Statistics & Probability Letters, 32(1):1–10, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529600051X.

Adamidis:1998:LDD

[AL98] K. Adamidis and S. Loukas. A lifetime distribution with decreasing failure rate. Statistics & Probability Letters, 39(1):35–42, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000121.

Alamatsaz:1993:CEG

[Ala93] M. H. Alamatsaz. On characterizations of exponential and gamma distributions. Statistics & Probability Letters, 17(4):315– 319, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293902092.

Alamatsaz:1994:CEG

[Ala94] M. H. Alamatsaz. On characterizations of exponential and gamma distributions. Statistics & Probability Letters, 20(3):251, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294000581.

Albin:1993:ETS

[Alb93] J. M. P. Albin. Extremes of totally skewed stable motion. Statistics & Probability Letters, 16(3):219–224, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390146A.

Albin:1997:ENA

[Alb97] J. M. P. Albin. Extremes for non-anticipating moving averages of totally skewed α-stable motion. Statistics & Probability Letters, 36(3):289–297, December 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000758.

Albin:1998:NRD

[Alb98] J. M. P. Albin. A note on Rosenblatt distributions. Statistics & Probability Letters, 40(1):83–91, September 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001096.

Alsmeyer:1991:CAI

[Als91a] Gerold Alsmeyer. Complete answer to an interval splitting problem. Statistics & Probability Letters, 12(4):285–287, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900926.

Alsmeyer:1991:SRB

[Als91b] Gerold Alsmeyer. Some relations between harmonic renewal measures and certain first passage times. Statistics & Probability Letters, 12(1):19–27, July 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190160S.

Alsmeyer:1996:NOF

[Als96] Gerold Alsmeyer. Nonnegativity of odd functional moments of positive random variables with decreasing density. Statistics & Probability Letters, 26(1):75–82, January 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002541.

Altman:1993:EEC

[Alt93] Naomi Simone Altman. Estimating error correlation in nonparametric regression. Statistics & Probability Letters, 18(3):213–218, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293902199.

Aly:1990:STD

[Aly90] Emad-Eldin A. A. Aly. A simple test for dispersive ordering. Statistics & Probability Letters, 9(4):323–325, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290901403.

Anderson:1993:NML

[AM93] T. W. Anderson and R. P. Metz. A note on maximum likelihood estimation in the first-order Gaussian moving average model. Statistics & Probability Letters, 16(3):205–211, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901448.

Abouammoh:1994:VUB

[AM94a] A. M. Abouammoh and A. F. Mashhour. Variance upper bounds and convolutions of α-unimodal distributions. Statistics & Probability Letters, 21(4):281–289, November 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294000174.

Averous:1994:MKS

[AM94b] Jean Averous and Michel Meste. Multivariate kurtosis in L_{1}- sense. Statistics & Probability Letters, 19(4):281–284, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901775.

Al-Mutairi:1997:PIG

[AM97] Dhaifalla K. Al-Mutairi. Properties of an inverse Gaussian mixture of bivariate exponential distribution and its generalization. Statistics & Probability Letters, 33(4):359–365, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001848.

Agostinelli:1998:OSR

[AM98] Claudio Agostinelli and Marianthi Markatou. A one-step robust estimator for regression based on the weighted likelihood reweighting scheme. Statistics & Probability Letters, 37(4):341– 350, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001363.

Amemiya:1990:NLD

[Ame90] Yasuo Amemiya. A note on the limiting distribution of certain characteristic roots. Statistics & Probability Letters, 9(5):465– 470, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090039A.

Anderson:1992:MLD

[And92] Dale N. Anderson. A multivariate Linnik distribution. Statistics & Probability Letters, 14(4):333–336, July 17, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290067F.

Angers:1996:FTB

[Ang96] Jean-François Angers. Fourier transform and Bayes estimator of a location parameter. Statistics & Probability Letters, 29 (4):353–359, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001913.

Anilkumar:1994:EMF

[Ani94] P. Anilkumar. On estimating the mean function of a Gaussian process. Statistics & Probability Letters, 19(1):77–84, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529490071X.

Anonymous:1990:AIVb

[Ano90a] Anonymous. Author index volume 10 (1990). Statistics & Probability Letters, 10(5):449–451, October 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900275.

Anonymous:1990:AIVa

[Ano90b] Anonymous. Author index volume 9 (1990). Statistics & Probability Letters, 9(5):471–473, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090040E.

Anonymous:1990:CPO

[Ano90c] Anonymous. Call for papers on optimization techniques in statistics. Statistics & Probability Letters, 10(3):271, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090085L.

Anonymous:1990:EB

[Ano90d] Anonymous. Editorial board. Statistics & Probability Letters, 9 (1):ii, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090086M.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial board. Statistics & Probability Letters, 11 (1):ii, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190168Q.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. Statistics & Probability Letters, 12(1):ii, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190157M.

Anonymous:1991:MIV

[Ano91c] Anonymous. Master index volumes 1–10. Statistics & Probability Letters, 11(1):77–98, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190183R.

Anonymous:1992:AIVa

[Ano92a] Anonymous. Author index volume 13 (1992). Statistics & Probability Letters, 13(5):421–422, April 7, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290117N.

Anonymous:1992:AIVb

[Ano92b] Anonymous. Author index volume 14 (1992). Statistics & Probability Letters, 14(5):421–422, July 11, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290104D.

Anonymous:1992:AIV

[Ano92c] Anonymous. Author index volume 15 (1992). Statistics & Probability Letters, 15(5):421–422, December 8, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292901788.

Anonymous:1992:EBa

[Ano92d] Anonymous. Editorial board. Statistics & Probability Letters, 13(1):ii, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290228W.

Anonymous:1992:EBb

[Ano92e] Anonymous. Editorial board. Statistics & Probability Letters, 15(1):ii, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290276B.

Anonymous:1993:AIVa

[Ano93a] Anonymous. Author index volume 16(1993). Statistics & Probability Letters, 16(5):421–422, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390077V.

Anonymous:1993:AIVb

[Ano93b] Anonymous. Author index volume 17 (1993). Statistics & Probability Letters, 17(5):421–422, August 5, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390265K.

Anonymous:1993:AIVc

[Ano93c] Anonymous. Author index volume 18 (1993). Statistics & Probability Letters, 18(5):421–422, December 2, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390037J.

Anonymous:1993:EBa

[Ano93d] Anonymous. Editorial board. Statistics & Probability Letters, 16(1):ii, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390113W.

Anonymous:1993:EBb

[Ano93e] Anonymous. Editorial board. Statistics & Probability Letters, 17 (1):ii, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390186M.

Anonymous:1993:EBc

[Ano93f] Anonymous. Editorial board. Statistics & Probability Letters, 18(1):ii, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293900906.

Anonymous:1994:AIVb

[Ano94a] Anonymous. Author index for volume 20 (1994). Statistics & Probability Letters, 20(5):421–422, August 9, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901341.

Anonymous:1994:AIVc

[Ano94b] Anonymous. Author index for volume 21 (1994). Statistics & Probability Letters, 21(5):421–422, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900019.

Anonymous:1994:AIVa

[Ano94c] Anonymous. Author index volume 19 (1994). Statistics & Probability Letters, 19(5):433–434, April 11, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900116.

Anonymous:1994:EBa

[Ano94d] Anonymous. Editorial board. Statistics & Probability Letters, 19(1):ii, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900604.

Anonymous:1994:EBb

[Ano94e] Anonymous. Editorial board. Statistics & Probability Letters, 20 (1):ii, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294902275.

Anonymous:1994:EBc

[Ano94f] Anonymous. Editorial board. Statistics & Probability Letters, 21(1):ii, September 1, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900507.

Anonymous:1995:AIVa

[Ano95a] Anonymous. Author index for volume 22 (1995). Statistics & Probability Letters, 22(4):349–350, March 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295900217.

Anonymous:1995:AIVb

[Ano95b] Anonymous. Author index for volume 23 (1995). Statistics & Probability Letters, 23(4):381–382, June 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529590011X.

Anonymous:1995:AIVc

[Ano95c] Anonymous. Author index for volume 24 (1995). Statistics & Probability Letters, 24(4):381–382, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295900136.

Anonymous:1995:AIVd

[Ano95d] Anonymous. Author index for volume 25 (1995). Statistics & Probability Letters, 25(4):379–380, December 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295900152.

Anonymous:1995:EBa

[Ano95e] Anonymous. Editorial board. Statistics & Probability Letters, 22 (1):ii, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295900233.

Anonymous:1995:EBb

[Ano95f] Anonymous. Editorial board. Statistics & Probability Letters, 23 (1):ii, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295900179.

Anonymous:1995:EBc

[Ano95g] Anonymous. Editorial board. Statistics & Probability Letters, 24(1):ii, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295900039.

Anonymous:1995:EBd

[Ano95h] Anonymous. Editorial board. Statistics & Probability Letters, 25 (1):ii, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295900195.

Anonymous:1996:AIVa

[Ano96a] Anonymous. Author index for volume 26 (1996). Statistics & Probability Letters, 26(4):389–390, March 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715296900231.

Anonymous:1996:AIVb

[Ano96b] Anonymous. Author index for volume 27 (1996). Statistics & Probability Letters, 27(4):393–394, May 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715296900073.

Anonymous:1996:AIVc

[Ano96c] Anonymous. Author index for volume 28 (1996). Statistics & Probability Letters, 28(4):387–388, August 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296900383.

Anonymous:1996:AIVd

[Ano96d] Anonymous. Author index for volume 29 (1996). Statistics & Probability Letters, 29(4):377–378, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296900279.

Anonymous:1996:AIVe

[Ano96e] Anonymous. Author index for volume 30 (1996). Statistics & Probability Letters, 30(4):381–382, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529690005X.

Anonymous:1996:EBa

[Ano96f] Anonymous. Editorial board. Statistics & Probability Letters, 26 (1):ii, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296900334.

Anonymous:1996:EBb

[Ano96g] Anonymous. Editorial board. Statistics & Probability Letters, 27(1):ii, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296900395.

Anonymous:1996:EBc

[Ano96h] Anonymous. Editorial board. Statistics & Probability Letters, 28 (1):ii, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529690036X.

Anonymous:1996:EBd

[Ano96i] Anonymous. Editorial board. Statistics & Probability Letters, 29(1):ii, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715296900097.

Anonymous:1996:EBe

[Ano96j] Anonymous. Editorial board. Statistics & Probability Letters, 30(1):ii, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296900371.

Anonymous:1996:EBf

[Ano96k] Anonymous. Editorial board. Statistics & Probability Letters, 31(1):ii, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296900346.

Anonymous:1996:Ea

[Ano96l] Anonymous. Erratum. Statistics & Probability Letters, 27(1): 99, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715296887025.

Anonymous:1997:AIVa

[Ano97a] Anonymous. Author index for volume 31 (1997). Statistics & Probability Letters, 31(4):365–366, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297841909.

Anonymous:1997:AIVb

[Ano97b] Anonymous. Author index for volume 32 (1997). Statistics & Probability Letters, 32(4):433–435, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297828386.

Anonymous:1997:AIVc

[Ano97c] Anonymous. Author index for volume 33 (1997). Statistics & Probability Letters, 33(4):439–440, May 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297881886.

Anonymous:1997:AIVd

[Ano97d] Anonymous. Author index for volume 34 (1997). Statistics & Probability Letters, 34(4):425–426, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297895041.

Anonymous:1997:AIVe

[Ano97e] Anonymous. Author index for volume 35 (1997). Statistics & Probability Letters, 35(4):433–434, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297866035.

Anonymous:1997:EBa

[Ano97f] Anonymous. Editorial board. Statistics & Probability Letters, 32(1):ii, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297900018.

Anonymous:1997:EBb

[Ano97g] Anonymous. Editorial board. Statistics & Probability Letters, 33 (1):ii, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715297900110.

Anonymous:1997:EBc

[Ano97h] Anonymous. Editorial board. Statistics & Probability Letters, 34 (1):ii, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715297900031.

Anonymous:1997:EBd

[Ano97i] Anonymous. Editorial board. Statistics & Probability Letters, 35(1):ii, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297900079.

Anonymous:1997:EBe

[Ano97j] Anonymous. Editorial board. Statistics & Probability Letters, 36(1):ii, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297900055.

Anonymous:1998:AIVa

[Ano98a] Anonymous. Author index for volume 36 (1998). Statistics & Probability Letters, 36(4):437–438, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297804440.

Anonymous:1998:AIVb

[Ano98b] Anonymous. Author index for volume 37 (1998). Statistics & Probability Letters, 37(4):431–432, March 30, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297859548.

Anonymous:1998:AIVc

[Ano98c] Anonymous. Author index for volume 38 (1998). Statistics & Probability Letters, 38(4):377–378, July 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800114X.

Anonymous:1998:AIVd

[Ano98d] Anonymous. Author index for volume 39 (1998). Statistics & Probability Letters, 39(4):387–388, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001552.

Anonymous:1998:AIVe

[Ano98e] Anonymous. Author index for volume 40 (1998). Statistics & Probability Letters, 40(4):411–412, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001692.

Anonymous:1998:EBa

[Ano98f] Anonymous. Editorial board. Statistics & Probability Letters, 37(1):ii, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298900013.

Anonymous:1998:EBb

[Ano98g] Anonymous. Editorial board. Statistics & Probability Letters, 38 (1):ii, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298900098.

Anonymous:1998:EBc

[Ano98h] Anonymous. Editorial board. Statistics & Probability Letters, 39 (1):iii, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298900074.

Anonymous:1998:EBd

[Ano98i] Anonymous. Editorial board. Statistics & Probability Letters, 40(1):ii, September 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715298900050.

Anonymous:1999:A

[Ano99a] Anonymous. Addendum. Statistics & Probability Letters, 42(2): 217, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298003022.

Anonymous:1999:AIV

[Ano99b] Anonymous. Author index for volume 41 (1999). Statistics & Probability Letters, 41(4):441–442, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002831.

Anonymous:1999:EB

[Ano99c] Anonymous. Editorial board. Statistics & Probability Letters, 41(1):ii, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299900019.

Anonymous:1999:Ea

[Ano99d] Anonymous. Erratum. Statistics & Probability Letters, 42(2):213–215, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298003010.

Anonymous:1999:Ec

[Ano99e] Anonymous. Erratum. Statistics & Probability Letters, 45(3):283, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001674.

Anonymous:1999:Ia

[Ano99f] Anonymous. Index. Statistics & Probability Letters, 42(4):433– 434, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000401.

Anonymous:1999:Ib

[Ano99g] Anonymous. Index. Statistics & Probability Letters, 43(4):433– 434, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000693.

Anonymous:1999:Ic

[Ano99h] Anonymous. Index. Statistics & Probability Letters, 44(4):417– 418, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299001443.

Anonymous:1999:Id

[Ano99i] Anonymous. Index. Statistics & Probability Letters, 45(4): 383–384, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001686.

Alsina:1993:CCB

[ANS93] Claudi Alsina, Roger B. Nelsen, and Berthold Schweizer. On the characterization of a class of binary operations on distribution functions. Statistics & Probability Letters, 17(2):85–89, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390001Y.

Alzaid:1992:DSM

[AP92] Abdulhamid A. Alzaid and Frank Proschan. Dispersivity and stochastic majorization. Statistics & Probability Letters, 13(4):275–278, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292900343.

Aragon:1995:RRC

[AQ95] Jorge Aragón and Adolfo J. Quiróz. Rank regression for current status data. Statistics & Probability Letters, 24(3):251–256, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400180G.

Arcones:1992:AMG

[Arc92] Miguel A. Arcones. On the arg max of a Gaussian process. Statistics & Probability Letters, 15(5):373–374, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290156Y.

Arcones:1994:CLT

[Arc94a] Miguel A. Arcones. The Central Limit Theorem for U-processes indexed by Hölder’s functions. Statistics & Probability Letters, 20(1):57–62, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294902348.

Arcones:1994:DCE

[Arc94b] Miguel A. Arcones. Distributional convergence of M-estimators under unusual rates. Statistics & Probability Letters, 21(4): 271–280, November 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294000131.

Arcones:1995:ANM

[Arc95a] Miguel A. Arcones. Asymptotic normality of multivariate trimmed means. Statistics & Probability Letters, 25(1):43–53, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400204L.

Arcones:1995:BTI

[Arc95b] Miguel A. Arcones. A Bernstein-type inequality for U-statistics and U-processes. Statistics & Probability Letters, 22(3):239–247, February 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400072G.

Arcones:1995:CLT

[Arc95c] Miguel A. Arcones. On the Central Limit Theorem for Ustatistics under absolute regularity. Statistics & Probability Letters, 24(3):245–249, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400179C.

Arcones:1996:SRU

[Arc96] Miguel A. Arcones. Some remarks on the uniform weak convergence of stochastic processes. Statistics & Probability Letters, 28(1):41–49, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000801.

Arcones:1998:LLN

[Arc98a] Miguel A. Arcones. The law of the large numbers for Ustatistics for 2m-wise independent random variables. Statistics & Probability Letters, 40(1):75–81, September 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001084.

Arcones:1998:RAE

[Arc98b] Miguel A. Arcones. A remark on approximate M-estimators. Statistics & Probability Letters, 38(4):311–321, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000388.

Arcones:1998:WCC

[Arc98c] Miguel A. Arcones. Weak convergence of convex stochastic processes. Statistics & Probability Letters, 37(2):171–182, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001156.

Arcudi:1998:CCE

[Arc98d] Ornella Arcudi. Convergence of conditional expectations given the random variables of a Skorohod representation. Statistics & Probability Letters, 40(1):1–8, September 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000601.

Angulo:1997:ORG

[ARM97] J. M. Angulo and M. D. Ruiz-Medina. On the orthogonal representation of generalized random fields. Statistics & Probability Letters, 31(3):145–153, January 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000260.

Adler:1997:MCT

[ARV97] André Adler, Andrew Rosalsky, and Andrej I. Volodin. A mean convergence theorem and weak law for arrays of random elements in martingale type p Banach spaces. Statistics & Probability Letters, 32(2):167–174, March 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297855939.

Atkins:1992:STS

[AS92] Joel E. Atkins and Gary J. Sherman. Sets of typical subsamples. Statistics & Probability Letters, 14(2):115–117, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290074F.

Agrawal:1995:URT

[AS95] M. C. Agrawal and A. B. Sthapit. Unbiased ratio-type variance estimation. Statistics & Probability Letters, 25(4):361–364, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294002427.

Aggarwal:1996:UAF

[AS96] M. L. Aggarwal and V. Sarin. Use of asymmetrical factorials for generating designs for quadratic mixture model. Statistics & Probability Letters, 27(3):275–279, April 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000844.

Assaf:1998:OCS

[ASC98] David Assaf and Ester Samuel-Cahn. Optimal cooperative stopping rules for maximization of the product of the expected stopped values. Statistics & Probability Letters, 38(1):89–99, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001582.

Abdous:1992:NSQ

[AT92] B. Abdous and R. Theodorescu. Note on the spatial quantile of a random vector. Statistics & Probability Letters, 13(4):333–336, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900435.

Albers:1996:SEC

[AT96] W. Albers and M. F. Teulings. A simple estimator for correlation in incomplete data. Statistics & Probability Letters, 31(1):51–57, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000132.

Auer:1990:CHC

[Aue90] Peter Auer. The circle homogeneously covered by random walk on Z^{2}. Statistics & Probability Letters, 9(5):403–407, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900323.

Arnold:1997:VED

[AV97] Barry C. Arnold and Jose A. Villasenor. Variance estimation with diffuse prior information. Statistics & Probability Letters, 33(1):35–39, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001071.

Arellano-Valle:1995:SCD

[AVB95] Reinaldo B. Arellano-Valle and Heleno Bolfarine. On some characterizations of the t-distribution. Statistics & Probability Letters, 25(1):79–85, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400208P.

Ahsanullah:1994:MNC

[AW94] M. Ahsanullah and Jacek Wesołowski. Multivariate normality via conditional normality. Statistics & Probability Letters, 20(3):235– 238, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900477.

Benkherouf:1993:GED

[BA93] Lakdere Benkherouf and Abdulhamid A. Alzaid. On the generalized Euler distribution. Statistics & Probability Letters, 18(4):323–326, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390023C.

Babu:1991:EES

[Bab91] Gutti Jogesh Babu. Edgeworth expansions for statistics which are functions of lattice and non-lattice variables. Statistics & Probability Letters, 12(1):1–7, July 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190158N.

Balakrishnan:1992:RSP

[BAC92] N. Balakrishnan, M. Ahsanullah, and P. S. Chan. Relations for single and product moments of record values from Gumbel distribution. Statistics & Probability Letters, 15(3):223–227, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901939.

Bach:1997:MDP

[Bac97] Eric Bach. Moments in the duration of play. Statistics & Probability Letters, 36(1):1–7, November 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000424.

Baek:1997:WDS

[Bae97] Jong-Il Baek. A weakly dependence structure of multivariate processes. Statistics & Probability Letters, 34(4):355–363, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002039.

Bajorski:1992:EBE

[Baj92] Piotr Bajorski. Exact Bahadur efficiency of max-type rank tests in the two-sample problem. Statistics & Probability Letters, 13(5):357–363, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290108H.

Balakrishnan:1990:IHD

[Bal90]  N. Balakrishnan.  Improving the Hartley–David–Gumbel

bound for the mean of extreme order statistics. Statistics & Probability Letters, 9(4):291–294, April 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090102D.

Balakrishnan:1993:MND

[Bal93a] N. Balakrishnan. Multivariate normal distribution and multivariate order statistics induced by ordering linear combinations. Statistics & Probability Letters, 17(5):343–350, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390253F.

Balakrishnan:1993:SAB

[Bal93b] N. Balakrishnan. A simple application of binomial-negative binomial relationship in the derivation of sharp bounds for moments of order statistics based on greatest convex minorants. Statistics & Probability Letters, 18(4):301–305, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390019F.

Ballerini:1994:DS

[Bal94] Rocco Ballerini. A dependent F^{α}-scheme. Statistics & Probability Letters, 21(1):21–25, September 1, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900531.

Ballerini:1995:NLY

[Bal95] Rocco Ballerini. Nonconstant levels for t-year return periods for two-dimensional Poisson processes with periodic intensity. Statistics & Probability Letters, 22(1):43–47, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400045A.

Baltagi:1997:TLL

[Bal97] Badi H. Baltagi. Testing linear and loglinear error components regressions against Box–Cox alternatives. Statistics & Probability Letters, 33(1):63–68, April 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001101.

Baltagi:1998:TLL

[Bal98] B. H. Baltagi. Testing linear and loglinear error components regressions against Box–Cox alternatives. Statistics & Probability Letters, 37(3):319, March 16, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529700134X.

Banjevic:1990:OSW

[Ban90] Dragan Banjević. On order statistics in waiting time for runs in Markov chains. Statistics & Probability Letters, 9(2):125–127, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290005P.

Barbe:1992:LDM

[Bar92a] Philippe Barbe. Limiting distribution of the maximal spacing when the density function admits a positive minimum. Statistics & Probability Letters, 14(1):53–60, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290210V.

Bartoszewicz:1992:ADT

[Bar92b] Jarosław Bartoszewicz. Asymptotic distributions of tests for dispersive ordering. Statistics & Probability Letters, 15(1):11–20, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290278D.

Bartoszewicz:1995:BHL

[Bar95a] Jarosław Bartoszewicz. Bahadur and Hodges–Lehmann approximate efficiencies of tests based on spacings. Statistics & Probability Letters, 23(3):211–220, May 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400115O.

Bartoszewicz:1995:SOR

[Bar95b] Jarosław Bartoszewicz. Stochastic order relations and the total time on test transform. Statistics & Probability Letters, 22(2):103–110, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400055D.

Bartoszewicz:1998:AGC

[Bar98a] Jarosław Bartoszewicz. Applications of a general composition theorem to the star order of distributions. Statistics & Probability Letters, 38(1):1–9, May 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001478.

Bartoszewicz:1998:CDO

[Bar98b] Jarosław Bartoszewicz. Characterizations of the dispersive order of distributions by the Laplace transform. Statistics & Probability Letters, 40(1):23–29, September 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000807.

Bartoszewicz:1999:CSO

[Bar99] Jarosław Bartoszewicz. Characterizations of stochastic orders based on ratios of Laplace transforms. Statistics & Probability Letters, 42(2):207–212, April 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002168.

Bassett:1994:NMM

[Bas94] Gilbert W. Bassett, Jr. A note on min-maxbias estimators in approximately linear models. Statistics & Probability Letters, 21(1):27–28, September 1, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529490054X.

Basu:1996:BHT

[Bas96a] Sanjib Basu. Bayesian hypotheses testing using posterior density ratios. Statistics & Probability Letters, 30(1):79–86, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002049.

Basu:1996:ENS

[Bas96b] Sanjib Basu. Existence of a normal scale mixture with a given variance and a percentile. Statistics & Probability Letters, 28(2):115–120, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001042.

Battaglia:1990:APP

[Bat90] Francesco Battaglia. Approximate power of portmanteau tests for time series. Statistics & Probability Letters, 9(4):337– 341, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090143U.

Bolfarine:1998:WNM

[BAV98] Heleno Bolfarine and Reinaldo B. Arellano-Valle. Weak nondifferential measurement error models. Statistics & Probability Letters, 40(3):279–287, October 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001047.

Baxter:1993:TTC

[Bax93] Laurence A. Baxter. Towards a theory of confidence intervals for system reliability. Statistics & Probability Letters, 16(1):29–38, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901183.

Baxter:1997:SOF

[Bax97] Laurence A. Baxter. Stochastic ordering of flows on manifolds. Statistics & Probability Letters, 33(3):259–261, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001356.

Bartoszewicz:1990:TDO

[BB90] Jarosław Bartoszewicz and Tadeusz Bednarski. On a test for dispersive ordering. Statistics & Probability Letters, 10(4):355–362, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090055C.

Bacro:1991:MEE

[BB91a] Jean-Noël Bacro and Margarida Brito. On Mason’s extension of the Erdős–Rényi Law of Large Numbers. Statistics & Probability Letters, 11(1):43–47, January 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190175Q.

Balasubramanian:1991:IOS

[BB91b] K. Balasubramanian and R. B. Bapat. Identities for order statistics and a theorem of Rényi. Statistics & Probability Letters, 12(2):141–143, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190058Y.

Balasubramanian:1992:IMR

[BB92] K. Balasubramanian and N. Balakrishnan. Indicator method for a recurrence relation for order statistics. Statistics & Probability Letters, 14(1):67–69, May 4, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290212N.

Balakrishnan:1993:EHD

[BB93a] N. Balakrishnan and K. Balasubramanian. Equivalence of Hartley–David–Gumbel and Papathanasiou bounds and some further remarks. Statistics & Probability Letters, 16(1):39–41, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901194.

Balasubramanian:1993:LCP

[BB93b] K. Balasubramanian and N. Balakrishnan. A log-concavity property of probability of occurrence of exactly r arbitrary events. Statistics & Probability Letters, 16(3):249–251, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390150H.

Balasubramanian:1994:ERO

[BB94] K. Balasubramanian and N. Balakrishnan. Equivalence of relations for order statistics for exchangeable and arbitrary cases.

Statistics & Probability Letters, 21(5):405–407, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294000387.

Balasubramanian:1995:CMD

[BB95a] K. Balasubramanian and N. Balakrishnan. On a class of multivariate distributions closed under concomitance of order statistics. Statistics & Probability Letters, 23(3):239–242, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400119S.

Boos:1995:ART

[BB95b] Dennis D. Boos and Cavell Brownie. ANOVA and rank tests when the number of treatments is large. Statistics & Probability Letters, 23(2):183–191, May 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400112L.

Blanke:1997:ARD

[BB97] D. Blanke and D. Bosq. Accurate rates of density estimators for continuous-time processes. Statistics & Probability Letters, 33(2):185–191, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001265.

Barbour:1992:BER

[BC92] A. D. Barbour and L. H. Y. Chen. On the binary expansion of a random integer. Statistics & Probability Letters, 14(3):235– 241, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900284.

Berkes:1996:PCL

[BC96] I. Berkes and E. Csáki. On the pointwise Central Limit Theorem and mixtures of stable distributions. Statistics & Probability Letters, 29(4):361–368, September 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001921.

Balakrishnan:1998:NRV

[BC98] N. Balakrishnan and P. S. Chan. On the normal record values and associated inference. Statistics & Probability Letters, 39(1): 73–80, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000480.

Berkes:1999:ASL

[BCC99] István Berkes, Endre Csáki, and Sándor Csörgő. Almost sure limit theorems for the St. Petersburg game. Statistics & Probability Letters, 45(1):23–30, October 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000371.

Berkes:1998:ASC

[BCH98] István Berkes, Endre Csáki, and Lajos Horváth. Almost sure Central Limit Theorems under minimal conditions. Statistics & Probability Letters, 37(1):67–76, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001016.

Benassi:1998:IMF

[BCI98] Albert Benassi, Serge Cohen, and Jacques Istas. Identifying the multifractional function of a Gaussian process. Statistics & Probability Letters, 39(4):337–345, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000789.

Brzeskwiniewicz:1997:RBD

[BCK97] Henryk Brzeskwiniewicz, Bronisław Ceranka, and Jolanta Krzyszkowska. Reinforced block designs with standards. Statistics & Probability Letters, 34(2):187–192, June 2, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001812.

Brito:1997:CMN

[BCQY97] M. R. Brito, E. L. Chávez, A. J. Quiroz, and J. E. Yukich. Connectivity of the mutual k-nearest-neighbor graph in clustering and outlier detection. Statistics & Probability Letters, 35(1):33– 42, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296002131.

Belzunce:1996:DOC

[BCR96] F. Belzunce, J. Candel, and J. M. Ruiz. Dispersive orderings and characterization of ageing classes. Statistics & Probability Letters, 28(4):321–327, August 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001417.

Beirlant:1990:APP

[BD90a] Jan Beirlant and Paul Deheuvels. On the approximation of P –P and Q–Q plot processes by Brownian bridges. Statistics & Probability Letters, 9(3):241–251, March 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090063D.

Burton:1990:LDS

[BD90b] Robert M. Burton and Herold Dehling. Large deviations for some weakly dependent random processes. Statistics & Probability Letters, 9(5):397–401, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715290900312.

Bapat:1991:OBD

[BD91] R. B. Bapat and A. Dey. Optimal block designs with minimal number of observations. Statistics & Probability Letters, 11(5): 399–402, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190188W.

Bhat:1992:CFC

[BD92a] B. R. Bhat and Somnath Datta. On the completeness of a family of conditional distributions. Statistics & Probability Letters, 15(1):27–30, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290280I.

Burton:1992:LDS

[BD92b] Robert M. Burton and Herold G. Dehling. Large deviations for some weakly dependent random processes. Statistics & Probability Letters, 13(1):83, January 2, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292902406.

Bertoin:1994:CEL

[BD94] J. Bertoin and R. A. Doney. Cramér’s estimate for Lévy processes. Statistics & Probability Letters, 21(5):363–365, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294000328.

Beirlant:1999:IED

[BD99] Jan Beirlant and Luc Devroye. On the impossibility of estimating densities in the extreme tail. Statistics & Probability Letters, 43(1):57–64, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002466.

Bose:1993:PMC

[BDG93] Sudip Bose, Gauri Sankar Datta, and Malay Ghosh. Pitman’s measure of closeness for symmetric stable distributions. Statistics & Probability Letters, 17(3):245–251, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390174H.

Barbe:1998:LPT

[BDG98] P. Barbe, M. Doisy, and B. Garel. Last passage time for the empirical mean of some mixing processes. Statistics & Probability Letters, 40(3):237–245, October 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000923.

Bassan:1999:VOM

[BDS99] Bruno Bassan, Michel Denuit, and Marco Scarsini. Variability orders and mean differences. Statistics & Probability Letters, 45(2):121–130, November 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000504.

Behrends:1999:CEC

[Beh99] Ehrhard Behrends. On the conditional expectation E(X|X + W) in the case of independent random variables X, W. Statistics & Probability Letters, 41(4):397–400, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001928.

Belisle:1991:OCM

[Bél91a] Claude Bélisle. Odd central moments of unimodal distributions. Statistics & Probability Letters, 12(2):97–107, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190052S.

Belisle:1991:WSS

[Bél91b] Claude Bélisle. Windings of spherically symmetric random walks via Brownian embedding. Statistics & Probability Letters, 12(4):345–349, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190105Z.

Belili:1999:TSR

[Bel99a] Nacereddine Belili. A theorem of stochastic representation related to Monge–Kantorovich problem. Statistics & Probability Letters, 41(2):139–143, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001308.

Belzunce:1999:CRS

[Bel99b] F. Belzunce. On a characterization of right spread order by the increasing convex order. Statistics & Probability Letters, 45(2):103–110, November 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000486.

Bensaid:1997:NIT

[Ben97] Nadia Bensaïd. Nonparametric inference for thinned point process. Statistics & Probability Letters, 33(3):253–258, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001344.

Berger:1990:IUE

[Ber90] James O. Berger. On the inadmissibility of unbiased estimators. Statistics & Probability Letters, 9(5):381–384, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900286.

Berry:1994:IJS

[Ber94] J. Calvin Berry. Improving the James–Stein estimator using the Stein variance estimator. Statistics & Probability Letters, 20(3):241–245, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900493.

Berlinet:1995:NVR

[Ber95] Alain Berlinet. A note on variance reduction. Statistics & Probability Letters, 25(4):357–360, December 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002415.

Bertoin:1996:IBM

[Ber96] Jean Bertoin. Iterated Brownian motion and stable( 14 ) subordinator. Statistics & Probability Letters, 27(2):111–114, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000518.

Berger:1997:CSI

[Ber97] Erich Berger. Comparing sums of independent bounded random variables and sums of Bernoulli random variables. Statistics & Probability Letters, 34(3):251–258, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001885.

Bercu:1999:WET

[Ber99a] B. Bercu. Weighted estimation and tracking for Bienaymé Galton Watson processes with adaptive control. Statistics & Probability Letters, 42(4):415–421, May 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002387.

Berger:1999:UIC

[Ber99b] Agnes Berger. On using index cases in the study of late-onset diseases. Statistics & Probability Letters, 41(1):31–37, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001199.

Besse:1992:PSC

[Bes92] Philippe Besse. PCA stability and choice of dimensionality. Statistics & Probability Letters, 13(5):405–410, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290115L.

Baryshnikov:1995:NSC

[BES95] Yuliy Baryshnikov, Bennett Eisenberg, and Gilbert Stengle. A necessary and sufficient condition for the existence of the limiting probability of a tie for first place. Statistics & Probability Letters, 23(3):203–209, May 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400114N.

Burke:1991:BML

[BG91] Murray D. Burke and Edit Gombay. The bootstrapped maximum likelihood estimator with an application. Statistics & Probability Letters, 12(5):421–427, November 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190031L.

Bosq:1995:NEC

[BG95] D. Bosq and D. Guégan. Nonparametric estimation of the chaotic function and the invariant measure of a dynamical system. Statistics & Probability Letters, 25(3):201–212, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400223U.

Biggins:1997:NGR

[BG97] J. D. Biggins and D. R. Grey. A note on the growth of random trees. Statistics & Probability Letters, 32(4):339–342, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000922.

Bai:1999:PLS

[BG99a] Z. D. Bai and Meihui Guo. A paradox in least-squares estimation of linear regression models. Statistics & Probability Letters, 42(2):167–174, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002053.

Bojdecki:1999:FBM

[BG99b] Tomasz Bojdecki and Luis G. Gorostiza. Fractional Brownian motion via fractional Laplacian. Statistics & Probability Letters, 44(1):107–108, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000140.

Burghardt:1994:PAN

[BGP94] Patrick D. Burghardt, Anant P. Godbole, and Amy B. Prengaman. A Poisson approximation for the number of kmatches. Statistics & Probability Letters, 21(1):1–8, September 1, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900515.

Baringhaus:1990:TUU

[BH90] L. Baringhaus and N. Henze. A test for uniformity with unknown limits based on d’Agostino’s D. Statistics & Probability Letters, 9(4):299–304, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090135T.

Baringhaus:1992:GFT

[BH92] L. Baringhaus and N. Henze. A goodness of fit test for the Poisson distribution based on the empirical generating function. Statistics & Probability Letters, 13(4):269–274, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900332.

Berkes:1996:BLG

[BH96] István Berkes and Lajos Horváth. Between local and global logarithmic averages. Statistics & Probability Letters, 30(4):369–378, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000041.

Bobkov:1997:CPT

[BH97] S. G. Bobkov and C. Houdré. Converse Poincaré-type inequalities for convex functions. Statistics & Probability Letters, 34(1):37– 42, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001630.

Bobkov:1999:CGP

[BH99] S. G. Bobkov and C. Houdré. A converse Gaussian Poincarétype inequality for convex functions. Statistics & Probability Letters, 44(3):281–290, September 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529900019X.

Bhattacharya:1995:RTA

[Bha95] Bhaskar Bhattacharya. Restricted tests for and against the increasing failure rate ordering on multinomial parameters. Statistics & Probability Letters, 25(4):309–316, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400235X.

Bhandary:1996:TGV

[Bha96] Madhusudan Bhandary. Test for generalized variance in signal processing. Statistics & Probability Letters, 27(2):155–162, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000585.

Basu:1996:THD

[BHB96] Ayanendranath Basu, Ian R. Harris, and Srabashi Basu. Tests of hypotheses in discrete models based on the penalized Hellinger distance. Statistics & Probability Letters, 27(4):367–373, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001018.

Bansal:1999:SCN

[BHK^{+}99] N. Bansal, G. G. Hamedani, Eric S. Key, Hans Volkmer, Hao Zhang, and J. Behboodian. Some characterizations of the normal distribution. Statistics & Probability Letters, 42(4):393–400, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002351.

Bhoj:1992:DWC

[Bho92] Dinesh S. Bhoj. On the distribution of the weighted combination of independent probabilities. Statistics & Probability Letters, 15(1):37–40, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290282A.

Bansal:1999:NLR

[BHZ99] Naveen K. Bansal, G. G. Hamedani, and Hao Zhang. Nonlinear regression with multidimensional indices. Statistics & Probability Letters, 45(2):175–186, November 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000577.

Bilodeau:1996:SRD

[Bil96] Martin Bilodeau. Some remarks on U(p; m, n) distributions. Statistics & Probability Letters, 31(1):41–43, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000119.

Bingham:1995:SMR

[Bin95] N. H. Bingham. The sample mid-range and symmetrized extremal laws. Statistics & Probability Letters, 23(3):281–288, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400126S.

Bingham:1996:SMR

[Bin96] N. H. Bingham. The sample mid-range and interquartiles. Statistics & Probability Letters, 27(2):131–136, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000542.

Birkel:1992:LLN

[Bir92] Thomas Birkel. Laws of large numbers under dependence assumptions. Statistics & Probability Letters, 14(5):355–362, July 11, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290096N.

Bischoff:1996:CMN

[Bis96a] Wolfgang Bischoff. Characterizing multivariate normal distributions by some of its conditionals. Statistics & Probability Letters, 26(2):105–111, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294002584.

Bischoff:1996:MDC

[Bis96b] Wolfgang Bischoff. On maximin designs for correlated observations. Statistics & Probability Letters, 26(4):357–363, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000321.

Bishwal:1999:LDI

[Bis99] J. P. N. Bishwal. Large deviations inequalities for the maximum likelihood estimator and the Bayes estimators in nonlinear stochastic differential equations. Statistics & Probability Letters, 43(2):207–215, June 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002612.

Balakrishnan:1998:NRB

[BJK98] N. Balakrishnan, Norman L. Johnson, and Samuel Kotz. A note on relationships between moments, central moments and cumulants from multivariate distributions. Statistics & Probability Letters, 39(1):49–54, July 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000273.

Bruss:1990:ISP

[BJZ90] F. Thomas Bruss, S. Rao Jammalamadaka, and Xian Zhou. On an interval splitting problem. Statistics & Probability Letters, 10(4):321–324, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090049D.

Bapat:1993:CID

[BK93] R. B. Bapat and Subhash C. Kochar. Characterizations of identically distributed independent random variables using order statistics. Statistics & Probability Letters, 17(3):225–230, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390171E.

Bose:1996:NDP

[BK96] Sudip Bose and Benjamin Kedem. Non-dependence of the predictive distribution on the population size. Statistics & Probability Letters, 27(1):43–47, March 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000429.

Bhattacharyya:1992:IEN

[BKR92] B. B. Bhattacharyya, T. M. Khoshgoftaar, and G. D. Richardson. Inconsistent M-estimators: nonlinear regression with multiplicative error. Statistics & Probability Letters, 14(5):407–411, July 11, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290102B.

Bhattacharyya:1996:GNE

[BKR96] B. B. Bhattacharyya, T. M. Khalil, and G. D. Richardson. Gauss–Newton estimation of parameters for a spatial autoregression model. Statistics & Probability Letters, 28(2):173–179, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500114X.

Bondesson:1996:IDR

[BKS96] Lennart Bondesson, Gundorph K. Kristiansen, and Fred W. Steutel. Infinite divisibility of random variables and their integer parts. Statistics & Probability Letters, 28(3):271–278, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001352.

Baltagi:1991:JTS

[BL91] Badi H. Baltagi and Qi Li. A joint test for serial correlation and random individual effects. Statistics & Probability Letters, 11(3):277–280, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190156L.

Baxter:1994:LRE

[BL94] Laurence A. Baxter and Linxiong Li. Lifelength in a random environment. Statistics & Probability Letters, 20(1):27–35, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902313.

Bhattacharya:1995:GEN

[BL95] Rabi Bhattacharya and Chanho Lee. On geometric ergodicity of nonlinear autoregressive models. Statistics & Probability Letters, 22(4):311–315, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400082J.

See erratum [BL99b].

Bartlett:1999:IUD

[BL99a] Peter Bartlett and Gábor Lugosi. An inequality for uniform deviations of sample averages from their means. Statistics & Probability Letters, 44(1):55–62, August 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002910.

Bhattacharya:1999:EGE

[BL99b] Rabi Bhattacharya and Chanho Lee. Erratum: “On geometric ergodicity of nonlinear autoregressive models” [Statist. Probab. Lett. 22 (1995) 311–315]. Statistics & Probability Letters, 41 (4):439–440, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002582.

See [BL95].

Blanc:1991:PAS

[Bla91] Antoni Sintes Blanc. On the Poisson approximation for some multinomial distributions. Statistics & Probability Letters, 11(1): 1–6, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190169R.

Bar-Lev:1990:CID

[BLB90] Shaul K. Bar-Lev and Daoud Bshouty. A class of infinitely divisible variance functions with an application to the polynomial case. Statistics & Probability Letters, 10(5):377–379, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900172.

Bar-Lev:1992:NEF

[BLBL92] Shaul K. Bar-Lev, Daoud Bshouty, and Gérard Letac. Natural exponential families and self-decomposability. Statistics & Probability Letters, 13(2):147–152, January 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290089N.

Bar-Lev:1990:CCV

[BLE90] Shaul K. Bar-Lev and Peter Enis. On the construction of classes of variance stabilizing transformations. Statistics & Probability Letters, 10(2):95–100, July 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090002O.

Blyth:1994:SKE

[Bly94] Stephen Blyth. Symmetrized kernel estimators of the regression slope. Statistics & Probability Letters, 21(2):167–172, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902267.

Bhattacharya:1993:ASU

[BlZ93] P. K. Bhattacharya and Peng liang Zhao. Almost sure uniform convergence rates for M-smoothers with non-monotone score functions. Statistics & Probability Letters, 16(3):189–196, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901426.

Berliner:1993:EIB

[BM93] L. Mark Berliner and Steven N. MacEachern. Examples of inconsistent Bayes procedures based on observations on dynamical systems. Statistics & Probability Letters, 17(5):355–360, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390255H.

Balakrishnan:1996:EBR

[BM96] N. Balakrishnan and Yimin Ma. Empirical Bayes rules for selecting the most and least probable multivariate hypergeometric event. Statistics & Probability Letters, 27(2):181–188, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000623.

Balek:1997:LII

[BM97] Vladimír Balek and Ivan Mizera. On the logical independence of the identities defining the stochastic independence of random events. Statistics & Probability Letters, 31(4):281–284, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000405.

Basu:1999:DRT

[BM99] A. K. Basu and Indranil Mukhopadhyay. On Darling–Robbins type confidence sequences and sequential tests with power one for parameters of an autoregressive process. Statistics & Probability Letters, 45(3):205–214, November 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000607.

Blaszczyszyn:1997:NEF

[BMS97] Bartlomiej Blaszczyszyn, Ely Merzbach, and Volker Schmidt. A note on expansion for functionals of spatial marked point processes. Statistics & Probability Letters, 36(3):299–306, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700076X.

Bai:1997:NCR

[BMT97] Z. D. Bai, Baiqi Miao, and Jhishen Tsay. A note on the convergence rate of the spectral distributions of large random matrices. Statistics & Probability Letters, 34(1):95–101, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001708.

Bansal:1997:MDR

[BMvdM97] Naveen K. Bansal, Neeraj Misra, and Edward C. van der Meulen. On the minimax decision rules in ranking problems. Statistics & Probability Letters, 34(2):179–186, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001800.

Brezmes:1996:AIM

[BN96] T. Brezmes and G. Naval. Axiomatic information measures depending only on a probability measure. Statistics & Probability Letters, 28(4):329–335, August 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001425.

Bagdonavicius:1997:ALT

[BN97] Vilijandas Bagdonavicius and Mikhail Nikulin. Accelerated life testing when a process of production is unstable. Statistics & Probability Letters, 35(3):269–275, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000229.

Bo:1998:MZL

[Bo98] Zhang Bo. Marcinkiewicz–Zygmund law for sequences of blockwise m-dependent random variables. Statistics & Probability Letters, 38(1):83–88, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297001570.

Boland:1998:AIP

[Bol98] Philip J. Boland. An arrangement increasing property of the Marshall–Olkin bivariate exponential. Statistics & Probability Letters, 37(2):167–170, February 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001144.

Boratynska:1996:BRC

[Bor96] Agata Boratyńska. On Bayesian robustness with the ϵcontamination class of priors. Statistics & Probability Letters, 26(4):323–328, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000275.

Boratynska:1997:SBI

[Bor97a] Agata Boratyńska. Stability of Bayesian inference in exponential families. Statistics & Probability Letters, 36(2):173–178, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000606.

Borovkov:1997:RWJ

[Bor97b] Konstantin Borovkov. On random walks with jumps scaled by cumulative sums of random variables. Statistics & Probability Letters, 35(4):409–416, November 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000394.

Belzunce:1999:LOO

[BOR99] Félix Belzunce, Eva Ortega, and José M. Ruiz. The Laplace order and ordering of residual lives. Statistics & Probability Letters, 42(2):145–156, April 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002028.

Bosq:1995:OAQ

[Bos95] Denis Bosq. Optimal asymptotic quadratic error of density estimators for strong mixing or chaotic data. Statistics & Probability Letters, 22(4):339–347, March 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400086N.

Boshuizen:1996:PCG

[Bos96] Frans A. Boshuizen. Prophet compared to gambler: additive inequalities for transforms of sequences of random variables. Statistics & Probability Letters, 29(1):23–32, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001514.

Bowman:1999:PIT

[Bow99] Dale Bowman. A parametric independence test for clustered binary data. Statistics & Probability Letters, 41(1):1–7, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000868.

Bryc:1992:CLT

[BP92] Włodzimierz Bryc and Magda Peligrad. The Central Limit Theorem for Tukey’s 3R smoother. Statistics & Probability Letters, 13(1):29–37, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290233U.

Bloznelis:1993:CLT

[BP93] M. Bloznelis and V. Paulauskas. On the Central Limit Theorem in D[0, 1]. Statistics & Probability Letters, 17(2):105–111, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900043.

Baldi:1999:RFL

[BP99a] Paolo Baldi and Mauro Piccioni. A representation formula for the large deviation rate function for the empirical law of a continuous time Markov chain. Statistics & Probability Letters, 41(2):107–115, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001242.

Boland:1999:CSC

[BP99b]  Philip J. Boland and Stavros Papastavridis.  Consecu-

tive k out of n: F systems with cycle k. Statistics & Probability Letters, 44(2):155–160, August 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000036.

Belzunce:1997:DOD

[BPRS97] F. Belzunce, Franco Pellerey, J. M. Ruiz, and Moshe Shaked. The dilation order, the dispersion order, and orderings of residual lives. Statistics & Probability Letters, 33(3):263–275, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001368.

Baxter:1990:NSE

[BR90] Laurence A. Baxter and Svetlozar T. Rachev. A note on the stability of the estimation of the exponential distribution. Statistics & Probability Letters, 10(1):37–41, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090109K.

Bagai:1991:ESF

[BR91a] Isha Bagai and B. L. S. Prakasa Rao. Estimation of the survival function for stationary associated processes. Statistics & Probability Letters, 12(5):385–391, November 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190027O.

Betro:1991:BII

[BR91b] B. Betrò and R. Rotondi. On Bayesian inference for the inverse Gaussian distribution. Statistics & Probability Letters, 11(3):219–224, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190146I.

Bhattacharyya:1992:SCS

[BR92] B. B. Bhattacharyya and G. D. Richardson. Strong consistency of a sieve estimator for the variance in nonlinear regression. Statistics & Probability Letters, 14(2):165–168, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290080O.

Baxter:1995:RCE

[BR95] J. R. Baxter and Jeffrey S. Rosenthal. Rates of convergence for everywhere-positive Markov chains. Statistics & Probability Letters, 22(4):333–338, March 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400085M.

Berti:1997:GCT

[BR97] Patrizia Berti and Pietro Rigo. A Glivenko–Cantelli theorem for exchangeable random variables. Statistics & Probability Letters, 32(4):385–391, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000983.

Bosch:1998:GPM

[BR98] Ronald J. Bosch and Louise M. Ryan. Generalized Poisson models arising from Markov processes. Statistics & Probability Letters, 39(3):205–212, August 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000285.

Best:1999:GFP

[BR99] D. J. Best and J. C. W. Rayner. Goodness of fit for the Poisson distribution. Statistics & Probability Letters, 44(3):259–265, September 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000176.

Bradley:1995:TKS

[Bra95] Richard C. Bradley. On a theorem of k. Schmidt. Statistics & Probability Letters, 24(1):9–12, July 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400142U.

Bradley:1996:CIU

[Bra96a] Richard C. Bradley. A covariance inequality under a two-part dependence assumption. Statistics & Probability Letters, 30 (4):287–293, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295002316.

Braverman:1996:RSM

[Bra96b] Michael Braverman. A remark on stochastic monotonicity. Statistics & Probability Letters, 26(3):259–262, February 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000186.

Bradley:1999:EMC

[Bra99a] Richard C. Bradley. Equivalent mixing conditions for Markov chains. Statistics & Probability Letters, 41(1):97–99, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001436.

Braverman:1999:RSL

[Bra99b] Michael Braverman. Remarks on suprema of Lévy processes with light tailes. Statistics & Probability Letters, 43(1):41–48, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002442.

Brunner:1992:BNM

[Bru92] Lawrence J. Brunner. Bayesian nonparametric methods for data from a unimodal density. Statistics & Probability Letters, 14(3):195–199, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290021V.

Bryc:1993:RCB

[Bry93] Włodzimierz Bryc. A remark on the connection between the large deviation principle and the Central Limit Theorem. Statistics & Probability Letters, 18(4):253–256, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900128.

Brzeskwiniewicz:1991:RAC

[Brz91] Henryk Brzeskwiniewicz. On the reduction of associate classes for the PBB designs. Statistics & Probability Letters, 11(4):327– 330, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190043Q.

Brzeskwiniewicz:1995:OST

[Brz95] Henryk Brzeskwiniewicz. On the E-optimality of some two-way elimination of heterogeneity designs. Statistics & Probability Letters, 25(1):9–13, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400199I.

Brzeskwiniewicz:1996:EBD

[Brz96] Henryk Brzeskwiniewicz. On the A-, D-, E- and L-efficiency of block designs. Statistics & Probability Letters, 30(3):205–209, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295002200.

Brzeskwiniewicz:1997:KPM

[Brz97] Henryk Brzeskwiniewicz. Kronecker product of mutually associable PBB designs. Statistics & Probability Letters, 33(4):413– 418, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001551.

Brown:1991:MSE

[BS91] P. J. Brown and C. H. Spiegelman. Mean squared error and selection in multivariate calibration. Statistics & Probability Letters, 12(2):157–159, August 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190061U.

Bryc:1992:SPC

[BS92] Włodzimierz Bryc and Włodzimierz Smolenski. On the stability problem for conditional expectation. Statistics & Probability Letters, 15(1):41–46, September 3, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290283B.

Bryc:1993:LDP

[BS93] Włodzimierz Bryc and Włodzimierz Smolenski. On the large deviation principle for a quadratic functional of the autoregressive process. Statistics & Probability Letters, 17(4):281–285, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390203U.

Bai:1994:NCD

[BS94a] Z. D. Bai and Lawrence A. Shepp. A note on the conditional distribution of X when |X − y| is given. Statistics & Probability Letters, 19(3):217–219, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901074.

Basu:1994:MDE

[BS94b] Ayanendranath Basu and Sahadeb Sarkar. Minimum disparity estimation in the errors-in-variables model. Statistics & Probability Letters, 20(1):69–73, May 4, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902364.

Basu:1994:DBG

[BS94c] Ayenendranath Basu and Sahadeb Sarkar. On disparity based goodness-of-fit tests for multinomial models. Statistics & Probability Letters, 19(4):307–312, March 15, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901813.

Bolfarine:1994:PFP

[BS94d] Heleno Bolfarine and Mônica C. Sandoval. On predicting the finite population distribution function. Statistics & Probability Letters, 19(4):339–347, March 15, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901872.

Bassan:1995:FEL

[BS95a] Bruno Bassan and Marco Scarsini. Finite exchangeability and linear regression. Statistics & Probability Letters, 23(2):105– 110, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400100M.

Blyth:1995:ESH

[BS95b] Colin R. Blyth and Robert G. Staudte. Estimating statistical hypotheses. Statistics & Probability Letters, 23(1):45–52, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400093N.

Bhattacharjee:1996:CVC

[BS96a] Arnab Bhattacharjee and Debasis Sengupta. On the coefficient of variation of the L- and L̄-classes. Statistics & Probability Letters, 27(2):177–180, April 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000615.

Buzas:1996:NCS

[BS96b] J. S. Buzas and L. A. Stefanski. A note on corrected-score estimation. Statistics & Probability Letters, 28(1):1–8, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000747.

Balu:1997:PCD

[BS97] Mini N. Balu and S. V. Sabnis. Preservation of certain dependent structures under bivariate homogeneous Poisson shock models. Statistics & Probability Letters, 35(1):91–100, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297812539.

Borzadaran:1998:FRB

[BS98] G. R. Mohtashami Borzadaran and D. N. Shanbhag. Further results based on Chernoff-type inequalities. Statistics & Probability Letters, 39(2):109–117, August 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000364.

Brands:1994:NMD

[BSW94] J. J. A. M. Brands, F. W. Steutel, and R. J. G. Wilms. On the number of maxima in a discrete sample. Statistics & Probability Letters, 20(3):209–217, June 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900442.

Bai:1997:PBU

[BSW97] Zhidong Bai, Sanat K. Sarkar, and Wenjin Wang. Positivity of the best unbiased L-estimator of the scale parameter with complete or selected order statistics from location-scale distribution. Statistics & Probability Letters, 32(2):181–188, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000715.

Bisgaard:1991:EUS

[BT91] Søren Bisgaard and Laurel E. Travis. Existence and uniqueness of the solution of the likelihood equations for binary Markov chains. Statistics & Probability Letters, 12(1):29–35, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190161J.

Bertin:1995:PUM

[BT95] Emile Bertin and Radu Theodorescu. Preserving unimodality by mixing. Statistics & Probability Letters, 25(3):281–288, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400232W.

Buckle:1994:SFR

[Buc94] D. J. Buckle. The study of a function relating to stable distributions. Statistics & Probability Letters, 20(2):85–90, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900221.

Burger:1993:DOA

[Bur93] H. U. Burger. Dispersion orderings with applications to nonparametric tests. Statistics & Probability Letters, 16(1):1–9, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390114X.

Butler:1999:EOL

[But99] Neil A. Butler. The efficiency of ordinary least squares in designed experiments subject to spatial or temporal variation. Statistics & Probability Letters, 41(1):73–81, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001266.

Beirlant:1992:NBK

[BV92] Jan Beirlant and Ann Vanmarcke. A note on Bahadur–Kiefertype expansions for the inverse empirical Laplace transform. Statistics & Probability Letters, 15(4):305–311, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901674.

Balasubramanian:1993:SLW

[BVB93] K. Balasubramanian, R. Viveros, and N. Balakrishnan. Sooner and later waiting time problems for Markovian Bernoulli trials. Statistics & Probability Letters, 18(2):153–161, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390184K.

Bertoin:1999:RTL

[BvHS99] J. Bertoin, K. van Harn, and F. W. Steutel. Renewal theory and level passage by subordinators. Statistics & Probability Letters, 45(1):65–69, October 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000437.

Berger:1993:EGP

[BW93] Agnes Berger and Sylvan Wallenstein. The effect of grouping on the power of the Mantel–Haenszel test for the comparison of survival rates. Statistics & Probability Letters, 16(1):19–25, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390116Z.

Booth:1995:EWL

[BW95] James G. Booth and Andrew T. A. Wood. An example in which the Lugannani–Rice saddlepoint formula fails. Statistics & Probability Letters, 23(1):53–61, April 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400094O.

Bai:1997:NCW

[BW97] Z. D. Bai and Y. Wu. On necessary conditions for the weak consistency of minimum L_{1}-norm estimates in linear models. Statistics & Probability Letters, 34(2):193–199, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001824.

Borovskikh:1999:WES

[BW99] Yuri V. Borovskikh and Neville C. Weber. On Wald’s equation for U-statistical sums. Statistics & Probability Letters, 44(4):375– 382, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000292.

Brown:1995:SCF

[BX95] Timothy C. Brown and Aihua Xia. On Stein–Chen factors for Poisson approximation. Statistics & Probability Letters, 23(4): 327–332, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400131Q.

Berrendero:1999:GRL

[BZ99] José R. Berrendero and Ruben H. Zamar. Global robustness of location and dispersion estimates. Statistics & Probability Letters, 44(1):63–72, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002922.

Chen:1997:KST

[CA97] Min Chen and Hong Zhi An. A Kolmogorov–Smirnov type test for conditional heteroskedasticity in time series. Statistics & Probability Letters, 33(3):321–331, May 5, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001435.

Cabana:1990:MCT

[Cab90] Enrique M. Cabaña. On a martingale characterization of two-parameter Wiener process. Statistics & Probability Letters, 10(3):263–270, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090084K.

Cuadras:1997:PDD

[CAF97] C. M. Cuadras, R. A. Atkinson, and J. Fortiana. Probability densities from distances and discrimination. Statistics & Probability Letters, 33(4):405–411, May 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600154X.

Cuesta-Albertos:1998:TBN

[CAGM98] J. A. Cuesta-Albertos, A. Gordaliza, and C. Matrán. Trimmed best k-nets: a robustified version of an L_{∞}-based clustering method. Statistics & Probability Letters, 36(4):401–413, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000874.

Cai:1998:APK

[Cai98] Zongwu Cai. Asymptotic properties of Kaplan–Meier estimator for censored dependent data. Statistics & Probability Letters, 37(4):381–389, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297001417.

Carriere:1992:LEV

[Car92] Jacques Carriere. Limited expected value comparison tests. Statistics & Probability Letters, 15(4):321–327, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901696.

Carstensen:1994:MMR

[Car94] Jens Michael Carstensen. Morphological Markov random fields. Statistics & Probability Letters, 20(4):321–326, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900205.

Casella:1990:ENR

[Cas90] George Casella. Estimators with nondecreasing risk: application of a chi-squared identity. Statistics & Probability Letters, 10(2): 107–109, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090004Q.

Cuesta-Albertos:1993:CSM

[CATD93] J. A. Cuesta-Albertos and A. Tuero-Díaz. A characterization for the solution of the Monge–Kantorovich mass transference problem. Statistics & Probability Letters, 16(2):147–152, January 27, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390159G.

Cavanaugh:1997:UDA

[Cav97] Joseph E. Cavanaugh. Unifying the derivations for the Akaike and corrected Akaike information criteria. Statistics & Probability Letters, 33(2):201–208, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001289.

Cavanaugh:1999:LSM

[Cav99] Joseph E. Cavanaugh. A large-sample model selection criterion based on Kullback’s symmetric divergence. Statistics & Probability Letters, 42(4):333–343, May 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002004.

Childs:1996:CIP

[CB96] Aaron Childs and N. Balakrishnan. Conditional inference procedures for the Laplace distribution based on Type-II right censored samples. Statistics & Probability Letters, 31(1):31–39, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000107.

Cai:1999:WES

[CB99] T. Tony Cai and Lawrence D. Brown. Wavelet estimation for samples with random uniform design. Statistics & Probability Letters, 42(3):313–321, April 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002235.

Cavazos-Cadena:1994:SFB

[CC94a]  Rolando Cavazos-Cadena.  A simple form of Bartlett’s

formula for autoregressive processes. Statistics & Probability Letters, 19(3):221–231, February 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901082.

Chatterjee:1994:NCO

[CC94b] Shoutir Kishore Chatterjee and Gaurangadeb Chattopadhyay. On the nonexistence of certain optimal confidence sets for the rectangular problem. Statistics & Probability Letters, 21 (4):263–269, November 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294901368.

Chung:1994:IEB

[CC94c] Daehyun Chung and Myron N. Chang. An isotonic estimator of the baseline hazard function in Cox’s regression model under order restriction. Statistics & Probability Letters, 21(3):223–228, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529490118X.

Chen:1998:ECC

[CC98a] Gemai Chen and Smiley W. Cheng. The exact u chart can be obtained using simple adjustments. Statistics & Probability Letters, 37(4):357–365, March 30, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001387.

Cheng:1998:ORB

[CC98b] Song-Show Cheng and Yu-Chun Cheng. An ordered relation between the ANOVA estimator of the intraclass correlation and a kappa-type statistic in binary data. Statistics & Probability Letters, 38(3):275–280, June 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000315.

Chakraborty:1999:NRM

[CC99a] Biman Chakraborty and Probal Chaudhuri. A note on the robustness of multivariate medians. Statistics & Probability Letters, 45(3):269–276, November 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529900067X.

Cordeiro:1999:CST

[CC99b] Gauss M. Cordeiro and Enrico A. Colosimo. Corrected score tests for exponential censored data. Statistics & Probability Letters, 44(4):365–373, October 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000280.

Chang:1999:RS

[CCH99] Gerard J. Chang, Lirong Cui, and Frank K. Hwang. Reliabilities for (n, f, k) systems. Statistics & Probability Letters, 43(3):237– 242, July 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002636.

Caramellino:1999:DAR

[CCHP99] Lucia Caramellino, Adriana Climescu-Haulica, and Barbara Pacchiarotti. Diffusion approximations for random walks on nilpotent Lie groups. Statistics & Probability Letters, 41(4):363–377, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001515.

Chen:1994:PEA

[CD94] Han-Fu Chen and Claude Deniau. Parameter estimation for ARMA processes with errors in models. Statistics & Probability Letters, 20(2):91–99, May 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529490023X.

Chaudhuri:1995:LBA

[CD95] Probal Chaudhuri and Anup Dewanji. On a likelihood-based approach in nonparametric smoothing and cross-validation. Statistics & Probability Letters, 22(1):7–15, January 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400040F.

Chukova:1993:RNP

[CDG93] Stefanka Chukova, Boyan Dimitrov, and José Garrido. Renewal and nonhomogeneous Poisson processes generated by distributions with periodic failure rate. Statistics & Probability Letters, 17(1):19–25, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390189P.

Cekanavicius:1995:AEF

[Cek95] V. Cekanavicius. On asymptotic expansions in the first uniform Kolmogorov theorem. Statistics & Probability Letters, 25(2):145–151, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400216U.

Cekanavicius:1996:MCT

[Cek96] V. Cekanavicius. On multivariate Le Cam theorem and compound Poisson measures. Statistics & Probability Letters, 28(1): 33–39, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000798.

Cekanavicius:1998:PAS

[Cek98] V. Cekanavicius. Poisson approximations for sequences of random variables. Statistics & Probability Letters, 39(2):101–107, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000236.

Carroll:1993:RLR

[CER93] R. J. Carroll, J. L. Eltinge, and D. Ruppert. Robust linear regression in replicated measurement error models. Statistics & Probability Letters, 16(3):169–175, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390139A.

Chen:1990:MPM

[CF90] Shande Chen and David Farnsworth. Median polish and a modified procedure. Statistics & Probability Letters, 9(1):51–57, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090095O.

Csaki:1995:LAA

[CF95] E. Csáki and A. Földes. On the logarithmic average of additive functionals. Statistics & Probability Letters, 22(4):261– 268, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400075J.

Chen:1996:CPT

[CF96a] Xiong Chen and Shui Feng. Critical phenomenon of a two component nonlinear stochastic system. Statistics & Probability Letters, 30(2):147–155, October 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002138.

Christoph:1996:NDA

[CF96b] Gerd Christoph and Michael Falk. A note on domains of attraction of p-max stable laws. Statistics & Probability Letters, 28(3): 279–284, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001360.

Csaki:1997:LAI

[CF97] Endre Csáki and Antónia Földes. On the logarithmic average of iterated processes. Statistics & Probability Letters, 33(4):347– 358, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001472.

Cuevas:1998:VDD

[CF98] Antonio Cuevas and Ricardo Fraiman. On visual distances in density estimation: the Hausdorff choice. Statistics & Probability Letters, 40(4):333–341, November 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001333.

Cadigan:1999:ELI

[CF99] N. G. Cadigan and P. J. Farrell. Expected local influence in the normal linear regression model. Statistics & Probability Letters, 41(1):25–30, January 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001175.

Cappuccio:1998:NST

[CFF98] Nunzio Cappuccio, Marco Ferrante, and Giovanni Fonseca. A note on the stationarity of a threshold first-order bilinear process. Statistics & Probability Letters, 40(4):379–384, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001448.

Cardot:1999:FLM

[CFS99] Hervé Cardot, Frédéric Ferraty, and Pascal Sarda. Functional linear model. Statistics & Probability Letters, 45(1):11–22, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900036X.

Caraux:1992:BDF

[CG92] G. Caraux and O. Gascuel. Bounds on distribution functions of order statistics for dependent variates. Statistics & Probability Letters, 14(2):103–105, May 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290071C.

Chaturvedi:1994:LER

[CG94] Ajit Chaturvedi and Rahul Gupta. Linking the estimation and ranking and selection problems through sequential procedures: The normal case. Statistics & Probability Letters, 20(4):273– 285, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900159.

Chen:1996:TDD

[CG96] Jie Chen and Joseph Glaz. Two-dimensional discrete scan statistics. Statistics & Probability Letters, 31(1):59–68, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000143.

Chen:1999:ADS

[CG99] Jie Chen and Joseph Glaz. Approximations for discrete scan statistics on the circle. Statistics & Probability Letters, 44 (2):167–176, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529900005X.

Carbon:1997:FPW

[CGT97] Michel Carbon, Bernard Garel, and Lanh Tat Tran. Frequency polygons for weakly dependent processes. Statistics & Probability Letters, 33(1):1–13, April 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001046.

Cheng:1991:NEA

[CH91] Kan Cheng and Zongfu He. A note on the exponential approximations of IFR distributions. Statistics & Probability Letters, 12(4):341–344, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190104Y.

Chan:1994:SCE

[CH94a] Y. M. Chan and Xuming He. A simple and competitive estimator of location. Statistics & Probability Letters, 19(2):137–142, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901465.

Chen:1994:CSE

[CH94b] Song Xi Chen and Peter Hall. On the calculation of standard error for quotation in confidence statements. Statistics & Probability Letters, 19(2):147–151, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901481.

Csorgo:1996:NCP

[CH96] Miklós Csörgő and Lajos Horváth. A note on the changepoint problem for angular data. Statistics & Probability Letters, 27(1):61–65, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000445.

Cheng:1998:MTE

[CH98a] Ming-Yen Cheng and Peter Hall. On mode testing and empirical approximations to distributions. Statistics & Probability Letters, 39(3):245–254, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000583.

Courbage:1998:EMC

[CH98b] M. Courbage and D. Hamdan. An ergodic Markov chain is not determined by its two-dimensional marginal laws. Statistics & Probability Letters, 37(1):35–40, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000965.

Cheng:1999:CBD

[CH99] K. F. Cheng and H. M. Hsueh. Correcting bias due to misclassification in the estimation of logistic regression models. Statistics & Probability Letters, 44(3):229–240, September 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000139.

Chakraborti:1990:CTH

[Cha90a] S. Chakraborti. A class of tests for homogeneity of quantiles under unequal right-censorship. Statistics & Probability Letters, 9(2):107–109, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290002M.

Chanda:1990:AEP

[Cha90b] Kamal C. Chanda. Asymptotic expansions of the probabilities of misclassification for k-NN discriminant rules. Statistics & Probability Letters, 10(4):341–349, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090053A.

Chandra:1990:ETM

[Cha90c] Nimai Kumar Chandra. On the efficiency of a testimator for the mean of an inverse Gaussian distribution. Statistics & Probability Letters, 10(5):431–437, October 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900253.

Chari:1990:CBN

[Cha90d] Ravi Chari. On a connection between the non-central χ^{2} distribution and Bessel diffusions. Statistics & Probability Letters, 10(3):217–219, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090076J.

Chawla:1990:NRR

[Cha90e] J. S. Chawla. A note on ridge regression. Statistics & Probability Letters, 9(4):343–345, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090144V.

Chaubey:1991:NNN

[Cha91] Yogendra P. Chaubey. A note on non-negative minimum bias MINQE in variance components model. Statistics & Probability Letters, 11(5):395–397, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190187V.

Chanda:1992:BKR

[Cha92a] Kamal C. Chanda. Bahadur–Kiefer representation properties of intermediate order statistics. Statistics & Probability Letters, 14(3):175–178, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290018Z.

Chang:1992:NDW

[Cha92b] Derek K. Chang. A note on the distribution of the Wilcoxon rank sum statistic. Statistics & Probability Letters, 13(5):343– 349, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290106F.

Chaudhuri:1992:LDF

[Cha92c] G. Chaudhuri. Linear discriminant function for complex normal time series. Statistics & Probability Letters, 15(4):277–279, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290162X.

Chakraborti:1994:ADF

[Cha94] S. Chakraborti. Asymptotically distribution-free joint confidence intervals for generalized Lorenz curves based on complete data. Statistics & Probability Letters, 21(3):229–235, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901198.

Chang:1995:GFP

[Cha95] Derek K. Chang. A game with four players. Statistics & Probability Letters, 23(2):111–115, May 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400101D.

Chai:1996:NBG

[Cha96a] Feng-Shun Chai. A note on balanced generalized two-way elimination of heterogeneity designs. Statistics & Probability Letters, 29(2):131–141, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001662.

Chan:1996:MPP

[Cha96b] Wenyaw Chan. Maximizing the probability of pest extinction on a stochastic pest-predator model. Statistics & Probability Letters, 30(3):265–270, October 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002294.

Chang:1996:EDK

[Cha96c] Myron N. Chang. Exact distribution of the Kaplan–Meier estimator under the proportional hazards model. Statistics & Probability Letters, 28(2):153–157, June 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001115.

Chai:1998:NGD

[Cha98a] Feng-Shun Chai. A note on generalization of distinct representatives. Statistics & Probability Letters, 39(2):173–177, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000637.

Chan:1998:IEL

[Cha98b] Ping Shing Chan. Interval estimation of location and scale parameters based on record values. Statistics & Probability Letters, 37(1):49–58, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000990.

Chang:1998:ADO

[Cha98c]  Fu-Chuen Chang.  On asymptotic distribution of opti-

mal design for polynomial-type regression. Statistics & Probability Letters, 36(4):421–425, January 3, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000898.

Chatterjee:1998:ALJ

[Cha98d] Snigdhansu Chatterjee. Another look at the jackknife: further examples of generalized bootstrap. Statistics & Probability Letters, 40(4):307–319, November 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001163.

Chan:1999:CSP

[Cha99a] Wai-Sum Chan. A comparison of some of pattern identification methods for order determination of mixed ARMA models. Statistics & Probability Letters, 42(1):69–79, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001953.

Chanda:1999:BKR

[Cha99b] Kamal C. Chanda. Bahadur–Kiefer representations for GMestimators in linear Markov models with errors in variables. Statistics & Probability Letters, 42(4):401–408, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002363.

Chang:1999:EOD

[Cha99c] Fu-Chuen Chang. Exact D-optimal designs for polynomial regression without intercept. Statistics & Probability Letters, 44(2):131–136, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002995.

Chelliah:1992:IDC

[Che92a] Thivvianesan Chelliah. Invariant distribution of Chow statistics. Statistics & Probability Letters, 15(5):401–402, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901744.

Cheng:1992:EUF

[Che92b] Bing Cheng. Existence, uniqueness and Fréchet differentiability of ψ-estimates of parameters in stationary observation. Statistics & Probability Letters, 15(5):339–343, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901504.

Chen:1993:SSC

[Che93] Gemai Chen. Some specific contiguous alternatives to the normal error assumption in linear models. Statistics & Probability Letters, 16(3):239–247, February 19, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390149D.

Chen:1994:TGF

[Che94] Juei-Chao Chen. Testing goodness of fit of polynomial models via spline smoothing techniques. Statistics & Probability Letters, 19(1):65–76, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900701.

Cheng:1995:BPE

[Che95a] Cheng Cheng. The Bernstein polynomial estimator of a smooth quantile function. Statistics & Probability Letters, 24(4):321–330, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400190J.

Cheng:1995:NSC

[Che95b] Philip E. Cheng. A note on strong convergence rates in nonparametric regression. Statistics & Probability Letters, 24(4):357–364, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400195E.

Chen:1996:CQT

[Che96] Zehua Chen. Conditional L_{p}-quantiles and their application to the testing of symmetry in non-parametric regression. Statistics & Probability Letters, 29(2):107–115, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001638.

Chen:1997:LDS

[Che97a] Bin Chen. Large deviations and Strassen’s limit points of Brownian local time processes. Statistics & Probability Letters, 34(4):385–395, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296002064.

Chen:1997:MDD

[Che97b] Xia Chen. Moderate deviations for m-dependent random variables with Banach space values. Statistics & Probability Letters, 35(2):123–134, September 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000059.

Chen:1997:SIA

[Che97c] Zhenmin Chen. Statistical inference about the shape parameter of the Weibull distribution. Statistics & Probability Letters, 36(1):85–90, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000515.

Chen:1998:BAG

[Che98a] Cathy W. S. Chen. A Bayesian analysis of generalized threshold autoregressive models. Statistics & Probability Letters, 40(1):15–22, September 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000777.

Chen:1998:SPP

[Che98b] Hegang Chen. Some projective properties of fractional factorial designs. Statistics & Probability Letters, 40(2):185–188, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001102.

Chen:1998:NBR

[Che98c] Zhiqiang Chen. A note on bias robustness of the median. Statistics & Probability Letters, 38(4):363–368, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000492.

Cheng:1998:BET

[Che98d] Cheng Cheng. A Berry–Esseen-type theorem of quantile density estimators. Statistics & Probability Letters, 39(3):255–262, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000649.

Choi:1997:RAS

[Cho97] ByoungSeon Choi. A recursive algorithm for solving the spatial Yule–Walker equations of causal spatial AR models. Statistics & Probability Letters, 33(3):241–251, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001332.

Choi:1998:SBA

[Cho98] ByoungSeon Choi. On the symmetric bilateral AR processes. Statistics & Probability Letters, 40(2):179–183, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001059.

Christofides:1990:RCS

[Chr90] Tasos C. Christofides. Rate of convergence in the Strong Law of Large Numbers for U-statistics based on a multidimensionally indexed array of random variables. Statistics & Probability Letters, 9(3):267–272, March 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090066G.

Christofides:1991:PIE

[Chr91] Tasos C. Christofides. Probability inequalities with exponential bounds for U-statistics. Statistics & Probability Letters, 12(3):257–261, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715291900878.

Christofides:1994:KIS

[Chr94] Tasos C. Christofides. A Kolmogorov inequality for Ustatistics based on Bernoulli kernels. Statistics & Probability Letters, 21(5):357–362, December 7, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400031X.

Christofides:1997:RCS

[Chr97] Tasos C. Christofides. Rate of convergence in the Strong Law of Large Numbers for a class of U-statistics and von Mises statistics. Statistics & Probability Letters, 31(3):225–231, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529600034X.

Chunsheng:1992:MFO

[Chu92a] Ma Chunsheng. Moments of functions of order statistics. Statistics & Probability Letters, 15(1):57–62, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290285D.

Chunsheng:1992:OUE

[Chu92b] Ma Chunsheng. The order of a univariate elliptical distribution. Statistics & Probability Letters, 14(4):307–310, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290062A.

Chunsheng:1992:VBF

[Chu92c] Ma Chunsheng. Variance bound of function of order statistics. Statistics & Probability Letters, 13(1):25–27, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290232T.

Chuang:1996:JDB

[Chu96] Chin-Shan Chuang. Joint distribution of Brownian motion and its maximum, with a generalization to correlated BM and applications to barrier options. Statistics & Probability Letters, 28(1): 81–90, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000925.

Cook:1993:EIC

[CHW93] R. D. Cook, D. M. Hawkins, and S. Weisberg. Exact iterative computation of the robust multivariate minimum volume ellipsoid estimator. Statistics & Probability Letters, 16(3):213–218, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901459.

Chang:1999:SCM

[cIC99] Yuan chin Ivan Chang. Strong consistency of maximum quasilikelihood estimate in generalized linear models via a last time. Statistics & Probability Letters, 45(3):237–246, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000632.

Chiu:1998:OFF

[CJ98] Wan-Yi Chiu and Peter W. M. John. D-optimal fractional factorial designs. Statistics & Probability Letters, 37(4):367–373, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001399.

Cohen:1994:AUP

[CK94a] Arthur Cohen and Debashis Kushary. Adaptive and unbiased predictors in a change point regression model. Statistics & Probability Letters, 20(2):131–138, May 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900280.

Cordeiro:1994:BCA

[CK94b] Gauss M. Cordeiro and Ruben Klein. Bias correction in ARMA models. Statistics & Probability Letters, 19(3):169–176, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901007.

Chao:1994:MDD

[CL94] Min-Te Chao and Wen-Qi Liang. Moments do not determine tail. Statistics & Probability Letters, 19(4):327–328, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901856.

Chen:1998:AML

[CL98] Di Chen and Jye-Chyi Lu. The asymptotics of maximumlikelihood estimates of parameters based on a data type where the failure and the censoring time are dependent. Statistics & Probability Letters, 36(4):379–391, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000850.

Chung:1996:NBM

[CLK96] Han-Yeong Chung, Kee-Won Lee, and Ja-Yong Koo. A note on bootstrap model selection criterion. Statistics & Probability Letters, 26(1):35–41, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294002495.

Ciardo:1995:MIG

[CLN95] Gianfranco Ciardo, Lawrence M. Leemis, and David Nicol. On the minimum of independent geometrically distributed random variables. Statistics & Probability Letters, 23(4):313–326, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400130Z.

Chiu:1990:NCB

[CM90] Shean-Tsong Chiu and J. S. Marron. The negative correlations between data-determined bandwidths and the optimal bandwidth. Statistics & Probability Letters, 10(2):173–180, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090014X.

Cuesta:1991:ABSa

[CM91a] Juan Antonio Cuesta and Carlos Matrán. On the asymptotic behavior of sums of pairwise independent random variables. Statistics & Probability Letters, 11(3):201–210, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190144G.

Cuesta:1991:ABSb

[CM91b] Juan Antonio Cuesta and Carlos Matrán. On the asymptotic behavior of sums of pairwise independent random variables. Statistics & Probability Letters, 12(2):183, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190065Y.

Coakley:1993:EFP

[CM93a] Clint W. Coakley and Lamine Mili. Exact fit points under simple regression with replication. Statistics & Probability Letters, 17(4):265–271, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390201S.

Cwik:1993:DDB

[ĆM93b] Jan Ćwik and Jan Mielniczuk. Data-dependent bandwidth choice for a grade density kernel estimate. Statistics & Probability Letters, 16(5):397–405, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390074S.

Chang:1995:ETC

[CM95] Xiang Qian Chang and Charles N. Moore. The equivalence of two conditions for weight functions for martingales. Statistics & Probability Letters, 23(2):129–133, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400104G.

Chen:1999:AUL

[CM99] Shijie Chen and Kanchan Mukherjee. Asymptotic uniform linearity of some robust statistics under exponentially subordinated strongly dependent models. Statistics & Probability Letters, 44(2):137–146, August 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298003009.

Coakley:1994:ELF

[CMC94] Clint W. Coakley, Lamine Mili, and Michael G. Cheniae. Effect of leverage on the finite sample efficiencies of high breakdown estimators. Statistics & Probability Letters, 19(5):399–408, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900086.

Cline:1998:VIC

[CmHP98] Daren B. H. Cline and Huay min H. Pu. Verifying irreducibility and continuity of a nonlinear time series. Statistics & Probability Letters, 40(2):139–148, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000819.

Cabrera:1994:OPS

[CMS94] J. Cabrera, G. Maguluri, and K. Singh. An odd property of the sample median. Statistics & Probability Letters, 19(4):349–354, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901880.

Cabrera:1997:IER

[CMS97] J. Cabrera, G. Maguluri, and K. Singh. Indices of empirical robustness. Statistics & Probability Letters, 33(1):49–62, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001095.

Cohen:1990:CEF

[Coh90] Jonathan P. Cohen. A comparison of estimators of functionals of the distribution of a maximum. Statistics & Probability Letters, 10(3):257–261, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090083J.

Cooke:1993:TTT

[Coo93a] Roger M. Cooke. The total time on test statistic and agedependent censoring. Statistics & Probability Letters, 18 (4):307–312, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390020J.

Coolen:1993:ICP

[Coo93b] F. P. A. Coolen. Imprecise conjugate prior densities for the one-parameter exponential family of distributions. Statistics & Probability Letters, 16(5):337–342, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390066R.

Coolen:1996:CTP

[Coo96] F. P. A. Coolen. Comparing two populations based on low stochastic structure assumptions. Statistics & Probability Letters, 29(4):297–305, September 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001867.

Coolen:1998:LSI

[Coo98] F. P. A. Coolen. Low structure imprecise predictive inference for Bayes’ problem. Statistics & Probability Letters, 36(4):349–357, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000813.

Cordeiro:1993:GMF

[Cor93a] Gauss M. Cordeiro. General matrix formulae for computing Bartlett corrections. Statistics & Probability Letters, 16(1):11– 18, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390115Y.

Cordy:1993:EHT

[Cor93b] Clifford B. Cordy. An extension of the Horvitz–Thompson theorem to point sampling from a continuous universe. Statistics & Probability Letters, 18(5):353–362, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390028H.

Cheng:1997:PIP

[CP97] R. Cheng and M. Pourahmadi. Prediction with incomplete past and interpolation of missing values. Statistics & Probability Letters, 33(4):341–346, May 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001460.

Cho:1995:URT

[CPA95] Sinsup Cho, Young J. Park, and Sung K. Ahn. Unit root tests for seasonal models with deterministic trends. Statistics & Probability Letters, 25(1):27–35, October 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400202J.

Cheze-Payaud:1998:ETL

[CPPP98] Nathalie Chèze-Payaud, Jean-Michel Poggi, and Bruno Portier. Estimation and test of linearity for a class of additive nonlinear models. Statistics & Probability Letters, 40(2):189–201, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001138.

Cao:1993:BMS

[CPS93] Ricardo Cao and JoséManuel Prada-Sánchez. Bootstrapping the mean of a symmetric population. Statistics & Probability Letters, 17(1):43–48, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390193M.

Qi:1994:SRC

[cQ94] Yong cheng Qi. Some results on covariance of function of order statistics. Statistics & Probability Letters, 19(2):111–114, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901414.

Chi:1991:APS

[CR91] Eric M. Chi and Gregory C. Reinsel. Asymptotic properties of the score test for autocorrelation in a random effects with AR(1) errors model. Statistics & Probability Letters, 11(5):453– 457, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190195W.

Cai:1992:USE

[CR92] Zongwu Cai and George G. Roussas. Uniform strong estimation under α-mixing, with rates. Statistics & Probability Letters, 15(1):47–55, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290284C.

Caperaa:1995:VTM

[CR95] Philippe Capéraà and Louis-Paul Rivest. On the variance of the trimmed mean. Statistics & Probability Letters, 22(1):79– 85, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400052A.

Cai:1997:SEQ

[CR97] Zongwu Cai and George G. Roussas. Smooth estimate of quantiles under association. Statistics & Probability Letters, 36 (3):275–287, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000746.

Chakraborty:1998:CBM

[CR98] Santanu Chakraborty and B. V. Rao. Completeness of Bhattacharya metric on the space of probabilities. Statistics & Probability Letters, 36(4):321–326, January 3, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000783.

Crowell:1992:NEP

[Cro92] John I. Crowell. Nonparametric estimation of a process mean from censored data. Statistics & Probability Letters, 15(4): 253–257, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292901593.

Croux:1994:EHB

[Cro94] Christophe Croux. Efficient high-breakdown M-estimators of scale. Statistics & Probability Letters, 19(5):371–379, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900051.

Croux:1998:LBE

[Cro98] Christophe Croux. Limit behavior of the empirical influence function of the median. Statistics & Probability Letters, 37(4):331– 340, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001351.

Cohen:1991:CUT

[CS91] Arthur Cohen and H. B. Sackrowitz. Constructing unbiased tests for homogeneity and goodness of fit. Statistics & Probability Letters, 12(4):351–355, October 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291901062.

Choi:1992:TPM

[CS92] Youn J. Choi and Norman C. Severo. Transition probabilities for a modified general stochastic epidemic model. Statistics & Probability Letters, 13(5):401–404, April 7, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290114K.

Chaudhuri:1993:NRE

[CS93a] Probal Chaudhuri and Debapriya SenGupta. A note on robust estimation of location. Statistics & Probability Letters, 18(3):241–244, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293902236.

Chen:1993:SRC

[CS93b] C. S. Chen and T. H. Savits. Some remarks on compound nonhomogeneous Poisson processes. Statistics & Probability Letters, 17(3):179–187, June 18, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390165F.

Csorgo:1994:NPD

[CS94] Miklós Csörgő and Qi-Man Shao. A new proof on the distribution of the local time of a Wiener process. Statistics & Probability Letters, 19(4):285–290, March 15, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901783.

Chen:1995:NBD

[CS95a] Tuhao Chen and E. Seneta. A note on bivariate Dawson–Sankofftype bounds. Statistics & Probability Letters, 24(2):99–104, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400153Y.

Cohen:1995:IST

[CS95b] Arthur Cohen and H. B. Sackrowitz. Inadmissibility of some tests for order-restricted alternatives. Statistics & Probability Letters, 24(2):153–156, August 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400161Z.

Cavanaugh:1996:CEF

[CS96a] Joseph E. Cavanaugh and Robert H. Shumway. On computing the expected Fisher information matrix for state-space model parameters. Statistics & Probability Letters, 26(4):347–355, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000313.

Csorgo:1996:SLL

[CS96b] Sándor Csörgő and Gordon Simons. A Strong Law of Large Numbers for trimmed sums, with applications to generalized St. Petersburg games. Statistics & Probability Letters, 26(1):65– 73, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294002533.

Chaganty:1997:BSS

[CS97] N. Rao Chaganty and Jayaram Sethuraman. Bahadur slope of the t-statistic for a contaminated normal. Statistics & Probability Letters, 34(3):245–250, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001873.

Christofides:1998:SLD

[CS98a] Tasos C. Christofides and Robert Serfling. U-statistics on a lattice of I.I.D. random variables. Statistics & Probability Letters, 40(3):293–303, October 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001114.

Christoph:1998:DSR

[CS98b] Gerd Christoph and Karina Schreiber. Discrete stable random variables. Statistics & Probability Letters, 37(3):243–247, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001235.

Conde:1998:UAC

[CS98c] J. Conde and M. Salicrú. Uniform association in contingency tables associated to Csiszár divergence. Statistics & Probability Letters, 37(2):149–154, February 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001120.

Carter:1993:SDC

[CSC93] R. A. L. Carter, V. K. Srivastava, and A. Chaturvedi. Selecting a double k-class estimator for regression coefficients. Statistics & Probability Letters, 18(5):363–371, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390029I.

Csenki:1993:CVT

[Cse93] Attila Csenki. On a counting variable in the theory of discreteparameter Markov chains. Statistics & Probability Letters, 18 (2):105–112, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390177K.

Csenki:1995:NVS

[Cse95] Attila Csenki. The number of visits to a subset of the state space by a discrete-parameter semi-Markov process. Statistics & Probability Letters, 22(1):71–77, January 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000519.

Csorgo:1992:LLN

[Csö92] Sándor Csörgő. On the Law of Large Numbers for the bootstrap mean. Statistics & Probability Letters, 14(1):1–7, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290203H.

Chib:1994:ODS

[CT94] Siddhartha Chib and Ram C. Tiwari. Outlier detection in the state space model. Statistics & Probability Letters, 20(2):143– 148, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900302.

Cheng:1999:COM

[CT99] Cheng Cheng and Y. L. Tong. Concentration order on a metric space, with some statistical applications. Statistics & Probability Letters, 44(4):327–335, October 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000231.

Carbon:1997:KDE

[CTW97] Michel Carbon, Lanh Tat Tran, and Berlin Wu. Kernel density estimation for random fields (density estimation for random fields). Statistics & Probability Letters, 36(2):115–125, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000540.

Cordeiro:1997:BCC

[CV97] Gauss M. Cordeiro and Klaus L. P. Vasconcellos. Bias correction for a class of multivariate nonlinear regression models. Statistics & Probability Letters, 35(2):155–164, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000096.

Chen:1994:NCE

[CW94] X. R. Chen and Y. Wu. Nonexistence of consistent estimates in a density estimation problem. Statistics & Probability Letters, 21(2):141–145, September 23, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902216.

Cai:1997:SIL

[CW97] Jun Cai and Yanhong Wu. Some improvements on the Lundberg bound for the ruin probability. Statistics & Probability Letters, 33(4):395–403, May 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001538.

Cheung:1998:PFM

[CW98] Siu Hung Cheung and Yaohua Wu. The power function of the maximum Studentized range test in a two-way design. Statistics & Probability Letters, 38(4):369–376, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000509.

Champ:1991:GQC

[CWM91] Charles W. Champ, William H. Woodall, and Hassan A. Mohsen. A generalized quality control procedure. Statistics & Probability Letters, 11(3):211–218, March 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190145H.

Chen:1999:NUC

[CY99] Kani Chen and Zhiliang Ying. A note on uniform consistency of the product-limit estimator with staggered entry. Statistics & Probability Letters, 42(3):275–281, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529800217X.

Csorgo:1997:RSC

[CZ97] Miklós Csörgö and Ricardas Zitikis. On the rate of strong consistency of Lorenz curves. Statistics & Probability Letters, 34(2):113–121, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297827897.

Dai:1998:SRC

[Dai98] Jack J. Dai. Some results concerning the rates of convergence of random walks on finite group. Statistics & Probability Letters, 37(1):15–17, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297800016.

Dale:1990:TBS

[Dal90] A. I. Dale. Thomas Bayes: some clues to his education. Statistics & Probability Letters, 9(3):289–290, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090069J.

Daras:1997:LMD

[Dar97] Tryfon Daras. Large and moderate deviations for the empirical measures of an exchangeable sequence. Statistics & Probability Letters, 36(1):91–100, November 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000527.

Daras:1998:TES

[Dar98] Tryfon Daras. Trajectories of exchangeable sequences: Large and moderate deviations results. Statistics & Probability Letters, 39(4):289–304, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800025X.

Das:1992:IER

[Das92] Kalyan Das. Improved estimation of the ratio of variance components for a balanced one-way random effects model. Statistics & Probability Letters, 13(2):99–108, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290083H.

Das:1998:OCE

[Das98] Ashish Das. Optimality of a class of efficiency-balanced designs. Statistics & Probability Letters, 39(4):317–326, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000716.

Datta:1995:MBC

[Dat95] Somnath Datta. On a modified bootstrap for certain asymptotically nonnormal statistics. Statistics & Probability Letters, 24(2):91–98, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400152X.

Davidson:1993:CTH

[Dav93] James Davidson. An L_{1}–convergence theorem for heterogeneous mixingale arrays with trending moments. Statistics & Probability Letters, 16(4):301–304, March 16, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293901345.

Davies:1994:DPB

[Dav94] Laurie Davies. Desirable properties, breakdown and efficiency in the linear regression model. Statistics & Probability Letters, 19(5):361–370, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900043.

David:1995:RRO

[Dav95] H. A. David. On recurrence relations for order statistics. Statistics & Probability Letters, 24(2):133–138, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001574.

David:1997:AOS

[Dav97] H. A. David. Augmented order statistics and the biasing effect of outliers. Statistics & Probability Letters, 36(2):199–204, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000667.

Dey:1994:RBA

[DB94] Dipak K. Dey and Lea R. Birmiwal. Robust Bayesian analysis using divergence measures. Statistics & Probability Letters, 20(4):287–294, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900167.

David:1996:PMO

[DB96] H. A. David and N. Balakrishnan. Product moments of order statistics and the variance of a lightly trimmed mean. Statistics & Probability Letters, 29(1):85–87, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500159X.

Diblasi:1997:TCV

[DB97] Angela Diblasi and Adrian Bowman. Testing for constant variance in a linear model. Statistics & Probability Letters, 33(1):95– 103, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001150.

Donahue:1995:PCL

[DBD95] Rafe M. J. Donahue, Peter J. Brockwell, and Richard A. Davis. On permissible correlations for locally correlated stationary processes. Statistics & Probability Letters, 22(1):49–53, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400046B.

deCambry:1993:ALS

[dC93] O. de Cambry. Asymptotic law in sequential estimation of a change point. Statistics & Probability Letters, 16(5):379–390, April 8, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390072Q.

Chen:1998:HMA

[dC98] Pei de Chen. Hastings–Metropolis algorithms and reference measures. Statistics & Probability Letters, 38(4):323–328, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000406.

DeGooijer:1998:FSP

[DD98] Jan G. De Gooijer and Paul T. De Bruin. On forecasting SETAR processes. Statistics & Probability Letters, 37(1):7–14, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000928.

Das:1998:ODD

[DDD98] Ashish Das, Aloke Dey, and Angela M. Dean. Optimal designs for diallel cross experiments. Statistics & Probability Letters, 36(4):427–436, January 3, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000904.

DelaHorra:1996:CBF

[De 96a] Julián De la Horra. Comparing Bayesian and frequentist estimators in the exchangeable case. Statistics & Probability Letters, 29(1):71–78, August 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001573.

DeSanctis:1996:GSF

[De 96b] Angela De Sanctis. Geodesic submanifolds of a family of statistical models. Statistics & Probability Letters, 30(2):127–132, October 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295002103.

Edwardes:1993:KEC

[dE93] Michael D. deB. Edwardes. Kendall’s τ is equal to the correlation coefficient for the BVE distribution. Statistics & Probability Letters, 17(5):415–419, August 5, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390264J.

Edwardes:1997:WVL

[dE97] Michael D. deB. Edwardes. Where variance is largest for Mann and Whitney’s U under a sum of powers marginal distribution. Statistics & Probability Letters, 36(2):195–198, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000643.

Debicki:1999:NLS

[Deb99] Krzysztof Debicki. A note on LDP for supremum of Gaussian processes over infinite horizon. Statistics & Probability Letters, 44(3):211–219, September 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000115.

Degenhardt:1993:CSP

[Deg93] H. J. A. Degenhardt. Chung–Smirnov property for perturbed empirical distribution functions. Statistics & Probability Letters, 16(2):97–101, January 27, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293901529.

Deheuvels:1991:LBP

[Deh91] Paul Deheuvels. On the limiting behavior of the Pickands estimator for bivariate extreme-value distributions. Statistics & Probability Letters, 12(5):429–439, November 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190032M.

Delgado:1993:TEN

[Del93] Miguel A. Delgado. Testing the equality of nonparametric regression curves. Statistics & Probability Letters, 17(3):199–204, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390167H.

Deo:1997:NRL

[Deo97] R. S. Deo. Nonparametric regression with long-memory errors. Statistics & Probability Letters, 33(1):89–94, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001149.

DeRiggi:1994:SCU

[DeR94] Dennis DeRiggi. Sufficient conditions for unimodality of the positive binomial likelihood function. Statistics & Probability Letters, 19(1):1–4, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900612.

Dermoune:1995:DCR

[Der95] A. Dermoune. Differential calculus relative to some point processes. Statistics & Probability Letters, 24(3):233–242, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400177A.

DeRiggi:1996:SCU

[DeR96] Dennis DeRiggi. Sufficient conditions for unimodality of the positive binomial likelihood function. Statistics & Probability Letters, 30(4):379, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000028.

Devroye:1990:NLD

[Dev90] Luc Devroye. A note on Linnik’s distribution. Statistics & Probability Letters, 9(4):305–306, April 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090136U.

Devroye:1991:OEN

[Dev91] Luc Devroye. On the oscillation of the expected number of extreme points of a random set. Statistics & Probability Letters, 11(4):281–286, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190036Q.

Devroye:1993:TDD

[Dev93] Luc Devroye. A triptych of discrete distributions related to the stable law. Statistics & Probability Letters, 18(5):349–351, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390027G.

Devroye:1994:NCE

[Dev94] Luc Devroye. On the non-consistency of an estimate of Chiu. Statistics & Probability Letters, 20(3):183–188, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529490040X.

Devroye:1995:APS

[Dev95] Luc Devroye. Another proof of a slow convergence result of Birgé. Statistics & Probability Letters, 23(1):63–67, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400095P.

Devroye:1997:STR

[Dev97] Luc Devroye. Simulating theta random variates. Statistics & Probability Letters, 31(4):275–279, February 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000399.

Deshpande:1995:NCU

[DF95] Jayant V. Deshpande and T. P. Muhammad Fareed. A note on conditionally unbiased estimation after selection. Statistics & Probability Letters, 22(1):17–23, January 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000427.

Dmitrienko:1997:DPY

[DG97] Alexei Dmitrienko and Z. Govindarajulu. On the “Demon” problem of Youden. Statistics & Probability Letters, 35(1):65–72, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296002179.

Das:1995:DTR

[DGA95] M. N. Das, N. Giri, and M. Ahmed. Designs through recoding of varietal and level codes. Statistics & Probability Letters, 23(4): 371–380, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294001404.

Dey:1999:EBL

[DGS99] Dipak K. Dey, Malay Ghosh, and William E. Strawderman. On estimation with balanced loss functions. Statistics & Probability Letters, 45(2):97–101, November 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000474.

Dette:1990:SPB

[DH90] H. Dette and N. Henze. Some peculiar boundary phenomena for extremes of r th nearest neighbor links. Statistics & Probability Letters, 10(5):381–390, October 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900183.

Dai:1997:RRW

[DH97] Jack J. Dai and Martin V. Hildebrand. Random random walks on the integers modn. Statistics & Probability Letters, 35(4):371–379, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000357.

deHaan:1999:EIS

[dHP99] L. de Haan and T. Themido Pereira. Estimating the index of a stable distribution. Statistics & Probability Letters, 41(1):39–55, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001205.

Dia:1990:NPE

[Dia90] Galaye Dia. Non-parametric estimation of the density of a point process. Statistics & Probability Letters, 10(5):397–405, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900208.

Dodge:1995:EQD

[DJ95] Yadolah Dodge and Jana Jurecková. Estimation of quantile density function based on regression quantiles. Statistics & Probability Letters, 23(1):73–78, April 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400097R.

deJong:1996:SLL

[dJ96] Robert M. de Jong. A Strong Law of Large Numbers for triangular mixingale arrays. Statistics & Probability Letters, 27(1):1–9, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000364.

Dodge:1997:ACT

[DJ97] Yadolah Dodge and Jana Jurecková. Adaptive choice of trimming proportion in trimmed least-squares estimation. Statistics & Probability Letters, 33(2):167–176, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001459.

DeVeaux:1990:REN

[DK90] Richard D. De Veaux and Abba M. Krieger. Robust estimation of a normal mixture. Statistics & Probability Letters, 10 (1):1–7, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090104F.

Dharmadhikari:1992:ATB

[DK92] A. D. Dharmadhikari and M. M. Kuber. The association in time of a binary semi-Markov process. Statistics & Probability Letters, 14(5):343–348, July 11, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290094L.

Duan:1993:CNG

[DK93] Xiaoping Duan and Sanpei Kageyama. Constructions of nested group divisible designs. Statistics & Probability Letters, 18(1):41– 48, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293900973.

Devroye:1999:HKD

[DK99] Luc Devroye and Adam Krzyzak. On the Hilbert kernel density estimate. Statistics & Probability Letters, 44(3):299–308, September 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000218.

Darwich:1991:AAB

[DL91a] A. R. Darwich and A. Le Breton. About the asymptotic behaviour of multidimensional Gaussian martingales and estimates in normal linear regression. Statistics & Probability Letters, 12(4):317–321, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190099D.

Dykstra:1991:MEP

[DL91b] R. L. Dykstra and Chu-In Charles Lee. Multinomial estimation procedures for isotonic cones. Statistics & Probability Letters, 11(2):155–160, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190134D.

delaPena:1992:NSM

[dlPG92] Victor H. de la Peña and Z. Govindarajulu. A note on second moment of a randomly stopped sum of independent variables. Statistics & Probability Letters, 14(4):275–281, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290057C.

Dalal:1990:EBP

[DLS90] Siddhartha R. Dalal, Jack C. Lee, and Darius J. Sabavala. Empirical Bayes prediction for a compound Poisson-multinomial process. Statistics & Probability Letters, 9(5):385–389, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900297.

Deng:1996:MMF

[DLW96] Lih-Yuan Deng, Dennis K. J. Lin, and Jiannong Wang. A measurement of multi-factor orthogonality. Statistics & Probability Letters, 28(3):203–209, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001255.

Dykstra:1996:MEP

[DLY96] R. L. Dykstra, Chu-In Charles Lee, and Xiaosong Yan. Multinomial estimation procedures for two stochastically ordered distributions. Statistics & Probability Letters, 30(4):353–361, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715295002391.

Dominguez-Menchero:1993:LBY

[DM93] J. Santos Domínguez-Menchero. On the local behaviour of the Yang regression estimate. Statistics & Probability Letters, 16(4):321–329, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293901378.

DeGiosa:1995:DFS

[DM95] Marcello De Giosa and Rosamaria Mininni. On the Doléans function of set-indexed submartingales. Statistics & Probability Letters, 24(1):71–75, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294001507.

Dey:1996:CSA

[DM96] Aloke Dey and Chand K. Midha. Construction of some asymmetrical orthogonal arrays. Statistics & Probability Letters, 28(3): 211–217, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001263.

Dey:1998:ADO

[DM98] Aloke Dey and Chand K. Midha. Addition or deletion? Statistics & Probability Letters, 37(4):409–414, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001442.

Das:1990:IER

[DMG90] K. Das, Q. Meneghini, and N. C. Giri. Inadmissibility of an estimator for the ratio of variance components. Statistics & Probability Letters, 10(2):151–157, July 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090011U.

Dominguez-Menchero:1997:EMU

[DMLP97] J. S. Domínguez-Menchero and M. J. López-Palomo. On the estimation of monotone uniform approximations. Statistics & Probability Letters, 35(4):355–362, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000333.

Dembo:1995:EBG

[DMWZ95] Amir Dembo, Eddy Mayer-Wolf, and Ofer Zeitouni. Exact behavior of Gaussian seminorms. Statistics & Probability Letters, 23(3):275–280, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400125R.

DiCiccio:1991:IPM

[DMY91] Thomas J. DiCiccio, Michael A. Martin, and G. Alastair Young. An invariance property of marginal density and tail probability approximations for smooth functions. Statistics & Probability Letters, 12(3):249–255, September 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291900867.

deNaranjo:1995:CLT

[dN95] M. V. Sánchez de Naranjo. A Central Limit Theorem for non-linear functionals of stationary Gaussian vector processes. Statistics & Probability Letters, 22(3):223–230, February 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400070O.

Donegani:1991:AAD

[Don91] M. Donegani. Asymptotic and approximative distribution of a statistic by resampling with or without replacement. Statistics & Probability Letters, 11(2):181–183, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190140M.

Dong:1998:CCL

[Don98] Lin Gwo Dong. Characterizations of the L-class of life distributions. Statistics & Probability Letters, 40(3):259–266, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001023.

Dorea:1997:ECV

[Dor97a] Chang C. Y. Dorea. On the efficiency of a continuous version of the simulated annealing algorithm. Statistics & Probability Letters, 31(4):247–253, February 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000089.

Dorofeev:1997:LIL

[Dor97b] Eugene A. Dorofeev. The law of iterated logarithm for one winding functional. Statistics & Probability Letters, 36(3):213–218, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529700062X.

Denker:1994:SLL

[DP94] Manfred Denker and Madan L. Puri. On some limit laws for perturbed empirical distribution functions. Statistics & Probability Letters, 21(4):317–321, November 11, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529300026Q.

DeCesare:1996:VUT

[DP96] L. De Cesare and D. Posa. The variogram of the uniform transform of a random field. Statistics & Probability Letters, 27(1):25– 29, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000399.

Delgado:1996:OSK

[DR96a] Miguel A. Delgado and Peter M. Robinson. Optimal spectral kernel for long-range dependent time series. Statistics & Probability Letters, 30(1):37–43, September 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001999.

DiCrescenzo:1996:CFP

[DR96b] Antonio Di Crescenzo and Luigi M. Ricciardi. Comparing first-passage times for semi-Markov skip-free processes. Statistics & Probability Letters, 30(3):247–256, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002278.

Drees:1999:FLC

[Dre99] Holger Drees. On fixed-length confidence intervals for a bounded normal mean. Statistics & Probability Letters, 44(4):399–404, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000322.

Droge:1999:AOF

[Dro99] Bernd Droge. Asymptotic optimality of full cross-validation for selecting linear regression models. Statistics & Probability Letters, 44(4):351–357, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000267.

Dey:1991:EDC

[DS91] Dipak K. Dey and C. Srinivasan. On estimation of discriminant coefficients. Statistics & Probability Letters, 11(3):189– 193, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190142E.

Deheuvels:1992:LBB

[DS92] Paul Deheuvels and Josef Steinebach. On the limiting behavior of the Bahadur–Kiefer statistic for partial sums and renewal processes when the fourth moment does not exist. Statistics & Probability Letters, 13(3):179–188, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900457.

Dobbins:1994:PIC

[DS94] Thomas W. Dobbins and Sanat K. Sarkar. Probability inequalities for certain dependence structures. Statistics & Probability Letters, 21(2):85–94, September 23, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902143.

Daduna:1996:QTP

[DS96] Hans Daduna and Ryszard Szekli. A queueing theoretical proof of increasing property of Pólya frequency functions. Statistics & Probability Letters, 26(3):233–242, February 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000151.

Dunnett:1993:PCS

[DT93] Charles W. Dunnett and Ajit C. Tamhane. Power comparisons of some step-up multiple test procedures. Statistics & Probability Letters, 16(1):55–58, January 4, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390123Z.

Dunmur:1998:IIC

[DT98] A. P. Dunmur and D. M. Titterington. The influence of initial conditions on maximum likelihood estimation of the parameters of a binary hidden Markov model. Statistics & Probability Letters, 40(1):67–73, September 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800100X.

deUna-Alvarez:1998:DCU

[dUÁGM98] Jacobo de Uña-Álvarez and Wenceslao González-Manteiga. Distributional convergence under proportional censorship when covariables are present. Statistics & Probability Letters, 39 (4):305–315, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000698.

deUna-Alvarez:1999:SCU

[dUÁGM99] Jacobo de Uña-Álvarez and Wenceslao González-Manteiga. Strong consistency under proportional censorship when covariables are present. Statistics & Probability Letters, 42(3):283– 292, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002181.

Duembgen:1998:SDC

[Due98] Lutz Duembgen. Symmetrization and decoupling of combinatorial random elements. Statistics & Probability Letters, 39(4):355–361, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000832.

Dumbgen:1992:LTS

[Düm92] Lutz Dümbgen. Limit theorems for the simplicial depth. Statistics & Probability Letters, 14(2):119–128, May 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290075G.

Dumbgen:1995:SPR

[Düm95] Lutz Dümbgen. A simple proof and refinement of Wielandt’s eigenvalue inequality. Statistics & Probability Letters, 25 (2):113–115, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400212Q.

Dupuis:1997:BTH

[Dup97] Jérôme A. Dupuis. Bayesian test of homogeneity for Markov chains. Statistics & Probability Letters, 31(4):333–338, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000478.

Dimitrov:1995:CUP

[DvC95] Boyan Dimitrov and Elart von Collani. Contorted uniform and Pareto distributions. Statistics & Probability Letters, 23(2):157– 164, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400108K.

Dalal:1992:SAM

[DW92] S. R. Dalal and S. Weerahandi. Some approximations for the moments of a process used in diffusion of new products. Statistics & Probability Letters, 15(3):181–189, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290187A.

Dette:1999:ODM

[DW99] Holger Dette and Weng Kee Wong. E-optimal designs for the Michaelis–Menten model. Statistics & Probability Letters, 44(4):405–408, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000334.

Doi:1998:JDR

[DY98] Masayuki Doi and Eiji Yamamoto. On the joint distribution of runs in a sequence of multi-state trials. Statistics & Probability Letters, 39(2):133–141, August 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000510.

Dembo:1996:LDS

[DZ96] Amir Dembo and Ofer Zeitouni. Large deviations for subsampling from individual sequences. Statistics & Probability Letters, 27(3):201–205, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000658.

Diebolt:1999:GFT

[DZ99] Jean Diebolt and Jacques Zuber. Goodness-of-fit tests for nonlinear heteroscedastic regression models. Statistics & Probability Letters, 42(1):53–60, March 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001898.

Dziubdziela:1995:LDS

[Dzi95] Wiesław Dziubdziela. On the limit distributions of sums of mixing Bernoulli random variables. Statistics & Probability Letters, 23(2):179–182, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400111K.

ElBarmi:1994:RMM

[ED94] Hammou El Barmi and Richard L. Dykstra. Restricted multinomial maximum likelihood estimation based upon Fenchel duality. Statistics & Probability Letters, 21(2):121–130, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902194.

ElBarmi:1996:RPM

[ED96] Hammou El Barmi and Richard L. Dykstra. Restricted product multinomial and product Poisson maximum likelihood estimation based upon Fenchel duality. Statistics & Probability Letters, 29(2):117–123, August 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001646.

Efromovich:1994:AEN

[Efr94] Sam Efromovich. On adaptive estimation of nonlinear functionals. Statistics & Probability Letters, 19(1):57–63, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900698.

Eichelsbacher:1999:ETC

[EG99] Peter Eichelsbacher and Malte Grunwald. Exponential tightness can fail in the strong topology. Statistics & Probability Letters, 41(1):83–86, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001278.

Egeland:1992:STO

[Ege92] Thore Egeland. On the S-shaped theorem. Statistics & Probability Letters, 13(1):1–4, January 2, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290229X.

Eastwood:1999:LTS

[EH99a] Vera R. Eastwood and Lajos Horváth. Limit theorems for short distances in R^{m}. Statistics & Probability Letters, 45(3):261–268, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000668.

ElBantli:1999:ELM

[EH99b] Faouzi El Bantli and Marc Hallin. L_{1}-estimation in linear models with heterogeneous white noise. Statistics & Probability Letters, 45(4):305–315, December 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000723.

Ehm:1991:BAP

[Ehm91] Werner Ehm. Binomial approximation to the Poisson binomial distribution. Statistics & Probability Letters, 11(1):7–16, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190170V.

Eichelsbacher:1996:RFT

[Eic96] Peter Eichelsbacher. Rate functions in the theory of large deviations. Statistics & Probability Letters, 26(3):279–283, February 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000216.

Eichelsbacher:1998:LDP

[Eic98] Peter Eichelsbacher. Large deviations in partial sums of Uprocesses in stronger topologies. Statistics & Probability Letters, 38(3):207–214, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000157.

Eisenberg:1992:VTS

[Eis92] Bennett Eisenberg. A variation of a theorem of Sparre– Andersen. Statistics & Probability Letters, 14(3):219–222, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290025Z.

Elfessi:1996:REC

[EJ96] Abdulaziz Elfessi and Chun Jin. On robust estimation of the common scale parameter of several Pareto distributions. Statistics & Probability Letters, 29(4):345–352, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001905.

ElBarmi:1995:LRT

[EK95] Hammou El Barmi and Subhash C. Kochar. Likelihood ratio tests for bivariate symmetry against ordered alternatives in a square contingency table. Statistics & Probability Letters, 22(2):167–173, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400064F.

Ebrahimi:1996:SRO

[EK96a] Nader Ebrahimi and S. N. U. A. Kirmani. Some results on ordering of survival functions through uncertainty. Statistics & Probability Letters, 29(2):167–176, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001700.

Etemadi:1996:SLL

[EK96b] N. Etemadi and M. Kaminski. Strong Law of Large Numbers for 2-exchangeable random variables. Statistics & Probability Letters, 28(3):245–250, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529500131X.

ElBarmi:1995:LQD

[El 95] Hammou El Barmi. A linear-quadratic distributional identity. Statistics & Probability Letters, 24(1):33–37, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400145X.

Elfessi:1997:ELF

[Elf97] Abdulaziz Elfessi. Estimation of a linear function of the parameters of an exponential distribution from doubly censored samples. Statistics & Probability Letters, 36(3):251–259, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297814708.

Eltinge:1992:CAB

[Elt92] John L. Eltinge. Conditions for approximation of the bias and mean squared error of a sample mean under nonresponse. Statistics & Probability Letters, 15(4):267–276, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290161W.

Eslava:1999:PIM

[EM99] G. Eslava and F. H. C. Marriott. Projection indices and multimodality for mixtures of normal distributions. Statistics & Probability Letters, 43(3):289–297, July 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002697.

El-Nouty:1993:HRT

[EN93] Charles El-Nouty. A Hanson–Russo-type law of the iterated logarithm for fractional Brownian motion. Statistics & Probability Letters, 17(1):27–34, May 3, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390190T.

El-Nouty:1996:RCF

[EN96] Charles El-Nouty. Rates of clustering in FLIL for weighted partial sum processes. Statistics & Probability Letters, 30 (2):113–118, October 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295002081.

El-Nouty:1999:LIF

[EN99] Charles El-Nouty. On the large increments of fractional Brownian motion. Statistics & Probability Letters, 41(2):169–178, January 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001400.

Endo:1995:CAA

[End95] Hideki Endo. Correspondence analysis of an artificial binary cylinder data. Statistics & Probability Letters, 25(3):231–240, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400226X.

Engel:1990:DEH

[Eng90] Joachim Engel. Density estimation with Haar series. Statistics & Probability Letters, 9(2):111–117, February 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290003N.

Etchison:1994:PAF

[EPB94] Tonya Etchison, Sastry G. Pantula, and Cavell Brownie. Partial autocorrelation function for spatial processes. Statistics & Probability Letters, 21(1):9–19, September 1, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900523.

Epps:1999:LBI

[Epp99] T. W. Epps. Limiting behavior of the ICF test for normality under Gram–Charlier alternatives. Statistics & Probability Letters, 42(2):175–184, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002077.

Ebrahimi:1994:TMS

[EPS94] Nader Ebrahimi, Kurt Pflughoeft, and Ehsan S. Soofi. Two measures of sample entropy. Statistics & Probability Letters, 20(3):225–234, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900469.

Ermakov:1992:AMR

[Erm92] M. S. Ermakov. On asymptotic minimaxity of rank tests. Statistics & Probability Letters, 15(3):191–196, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290188B.

Ebneshahrashoob:1990:SLW

[ES90] M. Ebneshahrashoob and Milton Sobel. Sooner and later waiting time problems for Bernoulli trials: frequency and run quotas. Statistics & Probability Letters, 9(1):5–11, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090088O.

Eubank:1991:CRT

[ES91] R. L. Eubank and Paul Speckman. Convergence rates for trigonometric and polynomial-trigonometric regression estimators. Statistics & Probability Letters, 11(2):119–124, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190128E.

Efromovich:1996:AEN

[ES96] Sam Efromovich and Alex Samarov. Asymptotic equivalence of nonparametric regression and white noise model has its limits. Statistics & Probability Letters, 28(2):143–145, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001093.

Ebrahimi:1997:SRN

[ES97] Nader Ebrahimi and Fabio Spizzichino. Some results on normalized total time on test and spacings. Statistics & Probability Letters, 36(3):231–243, December 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000680.

Elliott:1991:IPS

[ET91] Robert J. Elliott and Allanus H. Tsoi. Integration by parts for the single jump process. Statistics & Probability Letters, 12(5):363– 370, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190023K.

Etemadi:1997:CSL

[Ete97] N. Etemadi. Criteria for the Strong Law of Large Numbers for sequences of arbitrary random vectors. Statistics & Probability Letters, 33(2):151–157, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600123X.

Etemadi:1999:MIA

[Ete99] N. Etemadi. Maximal inequalities for averages of i.i.d. and 2exchangeable random variables. Statistics & Probability Letters, 44(2):195–200, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000097.

Ethier:1998:OST

[Eth98] S. N. Ethier. An optional stopping theorem for nonadapted martingales. Statistics & Probability Letters, 39(3):283–288, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000686.

Einmahl:1992:GCT

[EvZ92] John H. J. Einmahl and Martien C. A. van Zuijlen. Glivenko– Cantelli-type theorems for weighted empirical distribution functions based on uniform spacings. Statistics & Probability Letters, 13(5):411–419, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290116M.

Exton:1991:NCG

[Ext91] Harold Exton. A note of the convolution of a generalised F - distribution. Statistics & Probability Letters, 12(4):315–316, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190098C.

ElBarmi:1999:LRT

[EZ99] Hammou El Barmi and Dale Zimmerman. Likelihood ratio tests for and against decreasing in transposition in K × K × K contingency tables. Statistics & Probability Letters, 45(1):1–10, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000358.

Faddy:1996:MAC

[Fad96] M. J. Faddy. Modelling and analysis of count data using a renewal process. Statistics & Probability Letters, 31(2):129–132, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000235.

Faddy:1998:MDP

[Fad98] M. J. Faddy. Markov death process modelling and analysis of binary data. Statistics & Probability Letters, 40(1):9–13, September 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000613.

Falk:1990:WCM

[Fal90a] Michael Falk. Weak convergence of the maximum error of the bootstrap quantile estimate. Statistics & Probability Letters, 10(4):301–305, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090046A.

Falk:1990:WCR

[Fal90b] Michael Falk. Weak convergence of the remainder term in the Bahadur representation of extreme quantiles. Statistics & Probability Letters, 9(1):47–50, January 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090094N.

Falk:1994:EQE

[Fal94] Michael Falk. Extreme quantile estimation in δ-neighborhoods of generalized Pareto distributions. Statistics & Probability Letters, 20(1):9–21, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294902291.

Falk:1997:AIM

[Fal97] Michael Falk. Asymptotic independence of median and MAD. Statistics & Probability Letters, 34(4):341–345, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529600199X.

Fang:1991:CNP

[Fan91] Zhaoben Fang. Characterization of nonhomogeneous Poisson processes via moment conditions. Statistics & Probability Letters, 12(1):83–90, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190167P.

Fan:1999:DDT

[Fan99] Yanqin Fan. A data-driven test for dispersive ordering. Statistics & Probability Letters, 41(4):331–336, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000935.

Faraway:1990:ISD

[Far90] Julian Faraway. Implementing semiparametric density estimation. Statistics & Probability Letters, 10(2):141–143, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090009V.

Farahmand:1996:RPC

[Far96] K. Farahmand. Random polynomials with complex coefficients. Statistics & Probability Letters, 27(4):347–355, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000976.

Fang:1991:LCS

[FB91] Kai-Tai Fang and P. M. Bentler. A largest characterization of spherical and related distributions. Statistics & Probability Letters, 11(2):107–110, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190126C.

Fazli:1995:STC

[FB95] K. Fazli and J. Behboodian. Skorohod’s theorem and convergence of types. Statistics & Probability Letters, 24(3):243–244, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400178B.

Franses:1996:TAP

[FB96] Philip Hans Franses and H. Peter Boswijk. Temporal aggregation in a periodically integrated autoregressive process. Statistics & Probability Letters, 30(3):235–240, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002251.

Ferrari:1996:STO

[FBCCN96] Silvia L. P. Ferrari, Denise A. Botter, Gauss M. Cordeiro, and Francisco Cribari-Neto. Second- and third-order bias reduction for one-parameter family models. Statistics & Probability Letters, 30(4):339–345, November 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715295002375.

Ferrari:1996:CST

[FC96] Silvia L. P. Ferrari and Gauss M. Cordeiro. Corrected score tests for exponential family nonlinear models. Statistics & Probability Letters, 26(1):7–12, January 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002460.

Ferrari:1996:IST

[FCUOCN96] Silvia L. P. Ferrari, Gauss M. Cordeiro, Miguel A. Uribe-Opazo, and Francisco Cribari-Neto. Improved score tests for one-parameter exponential family models. Statistics & Probability Letters, 30(1):61–71, September 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002022.

Fotopoulos:1993:NSR

[FD93] S. B. Fotopoulos and D. Y. Downham. A note on the simple random walk in the plane. Statistics & Probability Letters, 17(3):221–224, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390170N.

Ferger:1994:RAS

[Fer94] Dietmar Ferger. On the rate of almost sure convergence of Dümbgen’s change-point estimators. Statistics & Probability Letters, 19(1):27–31, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900647.

Ferguson:1996:PME

[Fer96] Thomas S. Ferguson. A problem of minimax estimation with directional information. Statistics & Probability Letters, 26 (3):205–211, February 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000119.

Ferreri:1997:MEP

[Fer97] Carlo Ferreri. On the ML-estimator of the positive and negative two-parameter binomial distribution. Statistics & Probability Letters, 33(2):129–134, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001204.

Ferreira:1998:UTS

[Fer98] Eva Ferreira. Using M-type smoothing splines to estimate the spectral density of a stationary time series. Statistics & Probability Letters, 38(2):197–205, June 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800011X.

Ferger:1999:UMM

[Fer99] Dietmar Ferger. On the uniqueness of maximizers of Markov– Gaussian processes. Statistics & Probability Letters, 45(1):71–77, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000449.

Forbes:1997:SCM

[FF97] Florence Forbes and Olivier François. Stochastic comparison for Markov processes on a product of partially ordered sets. Statistics & Probability Letters, 33(3):309–320, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001423.

Fan:1992:MEB

[FG92] Jianqing Fan and Irène Gijbels. Minimax estimation of a bounded squared mean. Statistics & Probability Letters, 13(5):383–390, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290111H.

Fan:1992:BCH

[FH92] Jianqing Fan and Tien-Chung Hu. Bias correction and higher order kernel functions. Statistics & Probability Letters, 13 (3):235–243, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292900538.

Frigyesi:1998:TSO

[FH98] Attila Frigyesi and Ola Hössjer. A test for singularity. Statistics & Probability Letters, 40(3):215–226, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529800087X.

Fan:1999:RCP

[FH99] Jianqing Fan and Li-Shan Huang. Rates of convergence for the pre-asymptotic substitution bandwidth selector. Statistics & Probability Letters, 43(3):309–316, July 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002715.

Fierro:1990:EUM

[Fie90] Raul Fierro. Existence and uniqueness of a martingale problem in D_{+}, I′. Statistics & Probability Letters, 9(4):357–359, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090146X.

Fishman:1991:CIM

[Fis91] George S. Fishman. Confidence intervals for the mean in the bounded case. Statistics & Probability Letters, 12(3):223–227, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715291900823.

Falk:1992:EES

[FJ92] Michael Falk and Daniel Janas. Edgeworth expansions for Studentized and prepivoted sample quantiles. Statistics & Probability Letters, 14(1):13–24, May 4, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290205J.

Fountain:1994:PCU

[FK94] Robert L. Fountain and Jerome P. Keating. The Pitman comparison of unbiased linear estimators. Statistics & Probability Letters, 19(2):131–136, January 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901457.

Fu:1995:RBC

[FK95] J. C. Fu and M. V. Koutras. Reliability bounds for coherent structures with independent components. Statistics & Probability Letters, 22(2):137–148, February 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400060L.

Fu:1991:RCL

[FL91] James C. Fu and W. Y. Lou. On reliabilities of certain large linearly connected engineering systems. Statistics & Probability Letters, 12(4):291–296, October 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291900948.

Fu:1998:GMP

[FL98] James C. Fu and Lung-An Li. Generalized method of Pao–Zhuan Yin–Yu. Statistics & Probability Letters, 37(4):321–329, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001272.

Fan:1999:AFV

[FLP99] Ruzong Fan, Kenneth Lange, and Edsel Peña. Applications of a formula for the variance function of a stochastic process. Statistics & Probability Letters, 43(2):123–130, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001989.

Fu:1999:EDE

[FLW99] James C. Fu, W. Y. Wendy Lou, and Yueh-Jir Wang. On the exact distributions of Eulerian and Simon Newcomb numbers associated with random permutations. Statistics & Probability Letters, 42(2):115–125, April 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800176X.

Feng:1992:SIU

[FM92] Ziding Feng and Charles E. McCulloch. Statistical inference using maximum likelihood estimation and the generalized likelihood ratio when the true parameter is on the boundary of the parameter space. Statistics & Probability Letters, 13(4):325–332, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900424.

Fraiman:1994:SDO

[FM94] R. Fraiman and J. Meloche. Smoothing dependent observations. Statistics & Probability Letters, 21(3):203–214, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901163.

Frolov:1999:LLI

[FM99] Andrei N. Frolov and Alexander I. Martikainen. On the length of the longest increasing run in R^{d}. Statistics & Probability Letters, 41(2):153–161, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001369.

Florit:1995:LCS

[FN95] Carme Florit and David Nualart. A local criterion for smoothness of densities and application to the supremum of the Brownian sheet. Statistics & Probability Letters, 22(1):25–31, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294000438.

Ferreira:1997:KRE

[FNARP97] Eva Ferreira, Vicente Núñez-Antón, and Juan Rodríguez-Póo. Kernel regression estimates of growth curves using nonstationary correlated errors. Statistics & Probability Letters, 34(4):413– 423, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529600209X.

Fox:1997:AEF

[Fox97] Martin Fox. Adaptive estimation of a function of a mean vector. Statistics & Probability Letters, 33(2):109–115, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001174.

Fagiuoli:1994:MRL

[FP94] E. Fagiuoli and F. Pellerey. Mean residual life and increasing convex comparison of shock models. Statistics & Probability Letters, 20(5):337–345, August 9, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901244.

Freidlin:1998:TEP

[FP98] Mark Freidlin and Ruth Pfeiffer. A threshold estimation problem for processes with hysteresis. Statistics & Probability Letters, 36(4):337–347, January 3, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000801.

Florens:1999:LDP

[FP99] Danielle Florens and Huyên Pham. Large deviation principle in nonparametric estimation of marked point processes. Statistics & Probability Letters, 41(4):383–388, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001813.

Fernandez-Ponce:1997:CLD

[FPMP97] J. M. Fernandez-Ponce and J. Muñoz-Perez. Characterization of lifetime distributions in the L_{s}-sense by a generalized spread. Statistics & Probability Letters, 31(4):323–331, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000466.

Falk:1992:PAE

[FR92] Michael Falk and Rolf-Dieter Reiss. Poisson approximation of empirical processes. Statistics & Probability Letters, 14(1):39– 48, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290208M.

Foldes:1993:QVL

[FR93] Antónia Földes and Pál Révész. Quadratic variation of the local time of a random walk. Statistics & Probability Letters, 17(1): 1–12, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390187N.

Fourdrinier:1995:ILE

[FR95] Dominique Fourdrinier and Christian P. Robert. Intrinsic losses for empirical Bayes estimation: a note on normal and Poisson cases. Statistics & Probability Letters, 23(1):35–44, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400092M.

Franses:1991:DOP

[Fra91] Philip Hans Franses. The detection of observations possibly influential for model selection. Statistics & Probability Letters, 11(4): 321–325, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190042P.

Franses:1992:MST

[Fra92a] Philip Hans Franses. A model selection test for an AR (1) versus an MA (1) model. Statistics & Probability Letters, 15(4):281–284, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290163Y.

Franses:1992:MSB

[Fra92b] Philip Hans Franses. Modeling seasonality in bimonthly time series. Statistics & Probability Letters, 15(5):407–415, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901766.

Franses:1993:MSP

[Fra93] Philip Hans Franses. A model selection procedure for time series with seasonality. Statistics & Probability Letters, 16(4):253–258, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901286.

From:1995:WLS

[Fro95] Steven G. From. A weighted least-squares procedure for estimating the parameters of Altham’s multiplicative generalization of the binomial distribution. Statistics & Probability Letters, 25(3):193–199, November 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400222T.

Frolov:1998:OSS

[Fro98] Andrei N. Frolov. On one-sided strong laws for large increments of sums. Statistics & Probability Letters, 37(2):155–165, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001132.

Fosam:1993:CSP

[FRS93] E. B. Fosam, C. R. Rao, and D. N. Shanbhag. Comments on some papers involving the integrated Cauchy functional equation. Statistics & Probability Letters, 17(4):299–302, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390206X.

Forbes:1993:NCA

[FS93] Andrew B. Forbes and Thomas J. Santner. A note on conditions for the asymptotic normality of the conditional maximum likelihood estimator in log odds ratio regression. Statistics & Probability Letters, 18(2):137–146, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390182I.

Fahrner:1998:ASM

[FS98] I. Fahrner and U. Stadtmüller. On almost sure max-limit theorems. Statistics & Probability Letters, 37(3):229–236, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001211.

Fernandez:1999:RPG

[FS99] Carmen Fernández and Mark F. J. Steel. Reference priors for the general location-scale model. Statistics & Probability Letters, 43(4):377–384, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002764.

Finkelstein:1998:CIN

[FTV98] Mark Finkelstein, Howard G. Tucker, and Jerry Alan Veeh. Confidence intervals for the number of unseen types. Statistics & Probability Letters, 37(4):423–430, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001466.

Fujisawa:1996:ERC

[Fuj96] Hironori Fujisawa. Estimation on random coefficient model with unbalanced data. Statistics & Probability Letters, 28(3):251–257, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001328.

Fung:1992:SDM

[Fun92] Wing-Kam Fung. Some diagnostic measures in discriminant analysis. Statistics & Probability Letters, 13(4):279–285, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900354.

Futschik:1999:NEM

[Fut99] A. Futschik. A new estimate of the mode based on the quantile density. Statistics & Probability Letters, 43(2):145–152, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002521.

Fong:1993:EAM

[FY93] Daniel Y. T. Fong and Paul Yip. An EM algorithm for a mixture model of count data. Statistics & Probability Letters, 17(1):53– 60, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390195O.

Fygenson:1992:MKS

[FZ92] Mendel Fygenson and Mai Zhou. Modifying the Koul, Susarla and Van Ryzin estimator for linear regression models with right censoring. Statistics & Probability Letters, 13(4):295–299, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900376.

Fakhre-Zakeri:1993:CLT

[FZF93] Issa Fakhre-Zakeri and Jamshid Farshidi. A Central Limit Theorem with random indices for stationary linear processes. Statistics & Probability Letters, 17(2):91–95, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390002Z.

Fakhre-Zakeri:1997:RFC

[FZL97] Issa Fakhre-Zakeri and Sangyeol Lee. A random functional Central Limit Theorem for stationary linear processes generated by martingales. Statistics & Probability Letters, 35(4):417–422, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000400.

Grecksch:1999:PSD

[GA99] W. Grecksch and V. V. Anh. A parabolic stochastic differential equation with fractional Brownian motion input. Statistics & Probability Letters, 41(4):337–346, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001473.

Gadidov:1997:IOS

[Gad97] Anda Gadidov. An inequality for order statistics. Statistics & Probability Letters, 34(1):1–4, May 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001587.

Galindo:1998:NEM

[Gal98] Christian D. Galindo. On nonparametric estimation of mean functionals. Statistics & Probability Letters, 39(2):143–149, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000522.

Gamkrelidze:1997:LLT

[Gam97] N. G. Gamkrelidze. On a local limit theorem in strong sense. Statistics & Probability Letters, 35(1):79–83, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002192.

Gangopadhyay:1995:NAB

[Gan95] Ashis K. Gangopadhyay. A note on the asymptotic behavior of conditional extremes. Statistics & Probability Letters, 25(2):163–170, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400218W.

Gan:1998:MSI

[Gan98] Gaoxiong Gan. MLE are stochastically increasing if likelihoods are unimodal. Statistics & Probability Letters, 40(3):289–292, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001072.

Gao:1995:LIL

[Gao95] Jiti Gao. The laws of the iterated logarithm of some estimates in partly linear models. Statistics & Probability Letters, 25(2):153–162, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400217V.

Gonzalez-Arevalo:1999:EHT

[GAP99] Bárbara González-Arévalo and José Luis Palacios. Expected hitting times for random walks on weak products of graphs. Statistics & Probability Letters, 43(1):33–39, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002429.

Garren:1998:MLE

[Gar98] Steven T. Garren. Maximum likelihood estimation of the correlation coefficient in a bivariate normal model with missing data. Statistics & Probability Letters, 38(3):281–288, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000352.

Gastaldi:1992:ORG

[Gas92] Tommaso Gastaldi. Optimal reconstruction of a generally censored sample. Statistics & Probability Letters, 14(5):393–399, July 11, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290100J.

Gastaldi:1994:IML

[Gas94] Tommaso Gastaldi. Improved maximum likelihood estimation for component reliabilities with Miyakawa–Usher–Hodgson– Guess’ estimators under censored search for the cause of failure. Statistics & Probability Letters, 19(1):5–18, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900620.

Gastaldi:1996:NCF

[Gas96] Tommaso Gastaldi. Note on closed-form MLEs of failure rates in a fully parametric random censorship model with incomplete data. Statistics & Probability Letters, 26(4):309–314, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000259.

Gatarek:1993:NNS

[Gat93] Dariusz Gatarek. A note on nonlinear stochastic equations in Hilbert spaces. Statistics & Probability Letters, 17(5):387–394, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390259L.

Gawarecki:1997:ESI

[Gaw97] Leszek Gawarecki. Extension of a stochastic integral with respect to cylindrical martingales. Statistics & Probability Letters, 34(2):103–111, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529600171X.

Gonzalez-Barrios:1996:VSP

[GB96] JoséM. González-Barrios. The volume of the smallest parallelepiped including m random points. Statistics & Probability Letters, 30(2):139–145, October 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500212X.

Gascuel:1992:BEO

[GC92] O. Gascuel and G. Caraux. Bounds on expectations of order statistics via extremal dependences. Statistics & Probability Letters, 15(2):143–148, September 28, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290127Q.

Genon-Catalot:1990:ASP

[GCL90] V. Genon-Catalot and C. Larédo. An asymptotic sufficiency property of observations related to the first hitting times of a diffusion. Statistics & Probability Letters, 9(3):253–258, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090064E.

Gerard:1997:LQR

[GD97] Kerkyacharian Gérard and Picard Dominique. Limit of the quadratic risk in density estimation using linear methods. Statistics & Probability Letters, 31(4):299–312, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000442.

Gupta:1991:UOB

[GDD91] V. K. Gupta, A. Das, and A. Dey. Universal optimality of block designs with unequal block sizes. Statistics & Probability Letters, 11(2):177–180, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190139I.

Geisser:1993:BIA

[Gei93] Seymour Geisser. Bayesian interim analysis of censored exponential observations. Statistics & Probability Letters, 18(2):163–168, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390185L.

Geisser:1998:CTT

[Gei98] Seymour Geisser. Comparing two tests used for diagnostic or screening purposes. Statistics & Probability Letters, 40 (2):113–119, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000674.

Geluk:1996:DAE

[Gel96] J. L. Geluk. On the domain of attraction of exp(-exp(- x)). Statistics & Probability Letters, 31(2):91–95, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000181.

Genton:1998:AVE

[Gen98] Marc G. Genton. Asymptotic variance of M-estimators for dependent Gaussian random variables. Statistics & Probability Letters, 38(3):255–261, June 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000261.

Genton:1999:CSS

[Gen99] Marc G. Genton. The correlation structure of the sample autocovariance function for a particular class of time series with elliptically contoured distribution. Statistics & Probability Letters, 41(2):131–137, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001291.

Georgiev:1990:NMF

[Geo90] Alexander A. Georgiev. Nonparametric multiple function fitting. Statistics & Probability Letters, 10(3):203–211, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090074H.

Gerencser:1991:SAV

[Ger91] László Gerencsér. Strong approximation of vector-valued stochastic integrals. Statistics & Probability Letters, 12(3): 201–207, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715291900786.

Gerencser:1993:MIR

[Ger93] László Gerencsér. Multiple integrals with respect to L-mixing processes. Statistics & Probability Letters, 17(1):73–83, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390198R.

Gertsbakh:1995:FIT

[Ger95] I. Gertsbakh. On the Fisher information in Type-I censored and quantal response data. Statistics & Probability Letters, 23(4): 297–306, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400128U.

Gertsbakh:1996:OAT

[Ger96] I. Gertsbakh. Optimal allocation of test positions and testing time in a factorial life testing experiment. Statistics & Probability Letters, 26(3):249–258, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000178.

Gadag:1991:SBC

[GG91] V. G. Gadag and R. P. Gupta. A study of the behaviour of certain known estimators on the non-extinction path of a branching process. Statistics & Probability Letters, 12(2):135–139, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190057X.

Genest:1996:NTO

[GGR96] Christian Genest, Kilani Ghoudi, and Bruno Rémillard. A note on tightness. Statistics & Probability Letters, 27(4):331–339, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000941.

Gamero:1998:BSF

[GGR98] M. D. Jiménez Gamero, J. Muñoz García, and A. Muñoz Reyes. Bootstrapping statistical functionals. Statistics & Probability Letters, 39(3):229–236, August 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000558.

Gombay:1992:GFT

[GH92] Edit Gombay and Lajos Horváth. A goodness-of-fit test for exponential families. Statistics & Probability Letters, 15(3):235–239, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290195B.

Gilat:1996:SCD

[GH96] David Gilat and T. P. Hill. Strongly-consistent, distribution-free confidence intervals for quantiles. Statistics & Probability Letters, 29(1):45–53, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001549.

Gilat:1997:SCD

[GH97] D. Gilat and T. P. Hill. Strongly-consistent, distributionfree confidence intervals for quantiles. Statistics & Probability Letters, 31(3):245–246, January 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296002003.

Ghazal:1994:MRT

[Gha94] G. A. Ghazal. Moments of the ratio of two dependent quadratic forms. Statistics & Probability Letters, 20(4):313–319, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900191.

Ghazal:1996:RFE

[Gha96] G. A. Ghazal. Recurrence formula for expectations of products of quadratic forms. Statistics & Probability Letters, 27(2):101– 109, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529500050X.

Ghosh:1993:CRB

[Gho93] Malay Ghosh. Cramér-Rao bounds for posterior variances. Statistics & Probability Letters, 17(3):173–178, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390164E.

Ghosh:1995:IML

[Gho95] Malay Ghosh. Inconsistent maximum likelihood estimators for the Rasch model. Statistics & Probability Letters, 23(2):165– 170, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400109L.

Gu:1992:NGC

[GHW92] Chong Gu, Nancy Heckman, and Grace Wahba. A note on generalized cross-validation with replicates. Statistics & Probability Letters, 14(4):283–287, July 17, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290058D.

Gil:1992:NCB

[Gil92] María Angeles Gil. A note on the connection between fuzzy numbers and random intervals. Statistics & Probability Letters, 13(4):311–319, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290040C.

Gittenberger:1997:NPC

[Git97] Bernhard Gittenberger. On the number of predecessors in constrained random mappings. Statistics & Probability Letters, 36(1):29–34, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529700045X.

Givens:1995:CLK

[Giv95] Geof H. Givens. Consistency of the local kernel density estimator. Statistics & Probability Letters, 25(1):55–61, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400205M.

Gupta:1997:LRT

[GK97] A. K. Gupta and D. G. Kabe. Linear restrictions and two step multivariate least squares with applications. Statistics & Probability Letters, 32(4):413–416, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001010.

Gupta:1998:MRQ

[GK98] A. K. Gupta and D. G. Kabe. Moments of ratios of quadratic forms. Statistics & Probability Letters, 38(1):69–71, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001557.

Gertsbakh:1999:CWD

[GK99] Ilya Gertsbakh and Abram Kagan. Characterization of the Weibull distribution by properties of the Fisher information under Type-I censoring. Statistics & Probability Letters, 42(1):99– 105, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002120.

Gajek:1996:LIL

[GKL96] L. Gajek, M. Kaluszka, and A. Lenic. The law of the iterated logarithm for L_{p}-norms of empirical processes. Statistics & Probability Letters, 28(2):107–110, June 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000992.

Gautam:1996:OSO

[GKS96a] Shiva Gautam, George Kimeldorf, and Allan R. Sampson. Optimized scorings for ordinal data for the general linear model. Statistics & Probability Letters, 27(3):231–239, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000704.

Giraitis:1996:ANR

[GKS96b] Liudas Giraitis, Hira L. Koul, and Donatas Surgailis. Asymptotic normality of regression estimators with long memory errors. Statistics & Probability Letters, 29(4):317–335, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001883.

Gut:1997:ESL

[GKS97] Allan Gut, Oleg Klesov, and Josef Steinebach. Equivalences in strong limit theorems for renewal counting processes. Statistics & Probability Letters, 35(4):381–394, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000369.

Gao:1995:ANP

[GL95a] Jiti Gao and Hua Liang. Asymptotic normality of pseudo-LS estimator for partly linear autoregression models. Statistics & Probability Letters, 23(1):27–34, April 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400091L.

Guttman:1995:RRE

[GL95b] Irwin Guttman and Dennis K. J. Lin. Robust recursive estimation for correlated observations. Statistics & Probability Letters, 23(1):79–92, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400098S.

Glad:1998:NUP

[Gla98] Ingrid K. Glad. A note on unconditional properties of a parametrically guided Nadaraya–Watson estimator. Statistics & Probability Letters, 37(1):101–108, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001065.

Gleser:1993:ESL

[Gle93] Leon Jay Gleser. Estimators of slopes in linear errors-invariables regression models when the predictors have known reliability matrix. Statistics & Probability Letters, 17(2):113–121, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900054.

Glendinning:1996:ABP

[Gle96] Richard H. Glendinning. Assessing a Bartlett plot. Statistics & Probability Letters, 28(1):73–79, June 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000879.

Gorostiza:1994:OTA

[GLM94] L. G. Gorostiza and J. A. Lopez-Mimbela. An occupation time approach for convergence of measure-valued processes, and the death process of a branching system. Statistics & Probability Letters, 21(1):59–67, September 1, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900582.

Glonek:1999:IMI

[Glo99] G. F. V. Glonek. On identifiability in models for incomplete binary data. Statistics & Probability Letters, 41(2):191–197, January 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001424.

Goldstein:1993:MCS

[GM93] Larry Goldstein and Brendan McCabe. On the moments of certain stochastic integrals. Statistics & Probability Letters, 18(1):65–72, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390100W.

Gajek:1999:LSR

[GM99a] Lesław Gajek and Jan Mielniczuk. Long- and short-range dependent sequences under exponential subordination. Statistics & Probability Letters, 43(2):113–121, June 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001886.

Genton:1999:RPD

[GM99b] Marc G. Genton and Yanyuan Ma. Robustness properties of dispersion estimators. Statistics & Probability Letters, 44(4):343– 350, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000255.

Gupta:1994:TPN

[GMS94] A. K. Gupta, T. F. Móri, and G. J. Székely. Testing for Poissonity-normality vs. other infinite divisibility. Statistics & Probability Letters, 19(3):245–248, February 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901112.

Gan:1994:CCM

[GN94] Gaoxiong Gan and James W. Neill. Convergence criteria for maxima with regularly varying normalizing constants. Statistics & Probability Letters, 20(1):23–26, May 4, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902305.

Gnedin:1995:CES

[Gne95] Alexander V. Gnedin. On a class of exchangeable sequences. Statistics & Probability Letters, 25(4):351–355, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002403.

Gnedin:1996:CES

[Gne96] Alexander V. Gnedin. A class of exchangeable sequences. Statistics & Probability Letters, 28(2):159–164, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001123.

Gnedin:1998:RMN

[Gne98a] Alexander V. Gnedin. Records from a multivariate normal sample. Statistics & Probability Letters, 39(1):11–15, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000042.

Gneiting:1998:STV

[Gne98b] Tilmann Gneiting. Simple tests for the validity of correlation function models on the circle. Statistics & Probability Letters, 39(2):119–122, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800042X.

Gneri:1999:PMT

[Gne99] Mario Antonio Gneri. Parametric models as thin subsets of the space of distributions. Statistics & Probability Letters, 43(1):1– 3, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800203X.

Gomes:1997:LLS

[GO97] João Gomes and Orlando Oliveira. Limit laws for a sequence between the maximum and the sum of independent exponentials. Statistics & Probability Letters, 35(1):25–32, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529600212X.

Ganesh:1999:IST

[GO99] Ayalvadi Ganesh and Neil O’Connell. An inverse of Sanov’s theorem. Statistics & Probability Letters, 42(2):201–206, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002156.

Godbole:1990:DPC

[God90a] Anant P. Godbole. Degenerate and Poisson convergence criteria for success runs. Statistics & Probability Letters, 10(3):247– 255, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090082I.

Godbole:1990:SFS

[God90b] Anant P. Godbole. Specific formulae for some success run distributions. Statistics & Probability Letters, 10(2):119–124, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090006S.

Gombay:1994:TCP

[Gom94] Edit Gombay. Testing for change-points with rank and sign statistics. Statistics & Probability Letters, 20(1):49–55, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529490233X.

Gomes:1996:EVT

[Gom96] João Gomes. Extreme value theory for a thermal energy storage model. Statistics & Probability Letters, 30(1):25–31, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001972.

Gordaliza:1991:BPM

[Gor91] A. Gordaliza. On the breakdown point of multivariate location estimators based on trimming procedures. Statistics & Probability Letters, 11(5):387–394, May 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190186U.

Govindarajulu:1997:NTS

[Gov97] Z. Govindarajulu. A note on two-stage and sequential fixedwidth intervals for the parameter in the uniform density. Statistics & Probability Letters, 36(2):179–188, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000618.

Gupta:1991:SMJ

[GP91] Arjun K. Gupta and Edsel A. Peña. A simple motivation for James–Stein estimators. Statistics & Probability Letters, 12(4):337–340, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190103X.

Guo:1992:NTN

[GP92] Ying (Ingrid) Yueh Guo and Nabendu Pal. A note on the trace of a normal dispersion matrix. Statistics & Probability Letters, 14(5):349–353, July 11, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290095M.

Grabner:1996:AAM

[GP96] P. J. Grabner and H. Prodinger. Asymptotic analysis of the moments of the Cantor distribution. Statistics & Probability Letters, 26(3):243–248, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500016X.

Goldys:1999:LES

[GP99] Beniamin Goldys and Szymon Peszat. Law equivalence of stochastic linear systems. Statistics & Probability Letters, 43(3):265–274, July 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002661.

Gardiner:1999:PCD

[GPI99] Joseph Gardiner, Dorothy Pathak, and Alka Indurkhya. Power calculations for detecting interaction in stratified 2×2 tables. Statistics & Probability Letters, 41(3):267–275, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001631.

Glaz:1991:SPI

[GR91] Joseph Glaz and Nalini Ravishanker. Simultaneous prediction intervals for multiple forecasts based on Bonferroni and producttype inequalities. Statistics & Probability Letters, 12(1):57–63, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190164M.

Granville:1992:NMN

[GR92] V. Granville and J. P. Rasson. A new modelisation of noise in image remote sensing. Statistics & Probability Letters, 14(1):61– 65, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290211M.

Giovagnoli:1993:SRR

[GR93a] Alessandra Giovagnoli and Giuliana Regoli. Some results on the representation of measures of location and spread as Lfunctionals. Statistics & Probability Letters, 16(4):269–278, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390130B.

Gorostiza:1993:RFA

[GR93b] Luis G. Gorostiza and Rolando Rebolledo. A random field approach to weak convergence of processes. Statistics & Probability Letters, 18(1):49–55, August 30, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293900984.

Gopal:1999:SFI

[GR99] G. Gopal and S. Rajalakshmi. L-superadditive function and its integral transform that preserves Schur property. Statistics & Probability Letters, 45(3):191–194, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000528.

Grill:1992:NRO

[Gri92] Karl Grill. A note on randomness. Statistics & Probability Letters, 14(3):229–233, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292900273.

Grill:1996:RSB

[Gri96] Karl Grill. The range of simple branching random walk. Statistics & Probability Letters, 26(3):213–218, February 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000127.

Gross:1997:NEM

[Gro97] Jürgen Groß. A note on equality of MINQUE and simple estimator in the general Gauss–Markov model. Statistics & Probability Letters, 35(4):335–339, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000308.

Gross:1998:SEL

[Gro98] Jürgen Groß. Statistical estimation by a linear combination of two given statistics. Statistics & Probability Letters, 39 (4):379–384, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000972.

Ghosh:1999:BAC

[GRS99] J. K. Ghosh, R. V. Ramamoorthi, and K. R. Srikanth. Bayesian analysis of censored data. Statistics & Probability Letters, 41(3):255–265, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800162X.

Gruber:1991:EJK

[Gru91] Marvin H. J. Gruber. The efficiency of jack-knifed and usual ridge type estimators: a comparison. Statistics & Probability Letters, 11(1):49–51, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190176R.

Grubel:1994:RCL

[Grü94] Rudolf Grübel. The rank of the current lifetime. Statistics & Probability Letters, 20(4):269–271, July 14, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900140.

Gamboa:1999:FLD

[GRZ99] F. Gamboa, A. Rouault, and M. Zani. A functional large deviations principle for quadratic forms of Gaussian stationary processes. Statistics & Probability Letters, 43(3):299–308, July 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002703.

Goldman:1991:RIT

[GS91] Alvin I. Goldman and Moshe Shaked. Results on inquiry and truth possession. Statistics & Probability Letters, 12(5):415–420, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190030U.

Giles:1993:EDL

[GS93a] David E. A. Giles and Virendra K. Srivastava. The exact distribution of a least squares regression coefficient estimator after a preliminary t-test. Statistics & Probability Letters, 16(1):59–64, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901242.

Gupta:1993:ESH

[GS93b] Sudhir Gupta and Kishore Sinha. An EDG/3 scheme having 5 associate classes. Statistics & Probability Letters, 17(4):287– 291, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390204V.

Gao:1995:NGN

[GS95] Sujuan Gao and T. M. F. Smith. On the nonexistence of a global nonnegative minimum bias invariant quadratic estimator of variance components. Statistics & Probability Letters, 25(2):117–120, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400213R.

Galambos:1996:EMP

[GS96a] Janos Galambos and Italo Simonelli. An extension of the method of polynomials and a new reduction formula for Bonferroni-type inequalities. Statistics & Probability Letters, 28(2):147–151, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001107.

Gerard:1996:CIN

[GS96b] Patrick D. Gerard and William R. Schucany. On combining independent nonparametric regression estimators. Statistics & Probability Letters, 26(1):25–34, January 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002487.

Ghorai:1999:ADS

[GS99] J. K. Ghorai and J. Schmitter. The asymptotic distribution of the suprema of the standardized empirical processes under the Koziol–Green model. Statistics & Probability Letters, 41(3):303–313, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001667.

Gurler:1993:WSQ

[GSW93] Ülkü Gürler, Winfried Stute, and Jane-Ling Wang. Weak and strong quantile representations for randomly truncated data with applications. Statistics & Probability Letters, 17(2):139–148, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900087.

Gayraud:1999:WME

[GT99] Ghislaine Gayraud and Karine Tribouley. Wavelet methods to estimate an integrated quadratic functional: Adaptivity and asymptotic law. Statistics & Probability Letters, 44(2):109–122, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800296X.

Gu:1995:CIC

[Gu95] Minggao Gu. Convergence of increments for cumulative hazard function in a mixed censorship-truncation model with application to hazard estimators. Statistics & Probability Letters, 23(2):135– 139, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400105H.

Gu:1998:PLE

[Gu98] Chong Gu. Penalized likelihood estimation: Convergence under incorrect model. Statistics & Probability Letters, 36(4):359–364, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000825.

Guglielmi:1998:SPC

[Gug98] Alessandra Guglielmi. A simple procedure calculating the generalized Stieltjes transform of the mean of a Dirichlet process. Statistics & Probability Letters, 38(4):299–303, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000340.

Guijing:1996:OCR

[Gui96] Chen Guijing. Optimal convergence rates and asymptotic efficiency of point estimators under truncated distribution families. Statistics & Probability Letters, 30(4):321–331, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715295002359.

Gut:1990:LPS

[Gut90] Allan Gut. Limit points of sample maxima. Statistics & Probability Letters, 9(4):331–336, April 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090142T.

Gut:1992:WLL

[Gut92] Allan Gut. The Weak Law of Large Numbers for arrays. Statistics & Probability Letters, 14(1):49–52, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290209N.

Gut:1996:SLP

[Gut96] Allan Gut. Stopped Lévy processes with applications to first passage times. Statistics & Probability Letters, 28(4):345–352, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001441.

Gut:1997:STD

[Gut97] Allan Gut. Stopped two-dimensional perturbed random walks and Lévy processes. Statistics & Probability Letters, 35(4): 317–325, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529700028X.

Gupta:1991:RQF

[GV91] A. K. Gupta and T. Varga. Rank of a quadratic form in an elliptically contoured matrix random variable. Statistics & Probability Letters, 12(2):131–134, August 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190056W.

Gomez:1992:ESF

[GV92] Guadalupe Gómez and John Van Ryzin. Estimation of the subsurvival function for time-to-tumor in survival/sacrifice experiments. Statistics & Probability Letters, 13(1):5–13, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292902303.

Guo:1993:LBE

[GW93] Mei-Hui Guo and Ching-Zong Wei. A lower bound for expectation of a convex functional. Statistics & Probability Letters, 18(3):191–194, October 15, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293902155.

Giovagnoli:1995:MDO

[GW95a] Alessandra Giovagnoli and H. P. Wynn. Multivariate dispersion orderings. Statistics & Probability Letters, 22(4):325–332, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400084L.

Gupta:1995:CNA

[GW95b] Arjun K. Gupta and Yining Wang. A condition for the nonexistence of ancillary statistics. Statistics & Probability Letters, 23(4):367–369, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400137W.

Guegan:1996:PLM

[GW96a] Dominique Guegan and Joseph Ngatchou Wandji. Power of the Lagrange multiplier test for certain subdiagonal bilinear models. Statistics & Probability Letters, 29(3):201–212, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001743.

Guffey:1996:TMI

[GW96b] James M. Guffey and F. T. Wright. Testing for monotonicity in the intensity of a nonhomogeneous Poisson process. Statistics & Probability Letters, 28(3):195–202, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001247.

Gyorfi:1997:SUC

[GW97] László Györfi and Harro Walk. On the strong universal consistency of a recursive regression estimate by Pál Révész. Statistics & Probability Letters, 31(3):177–183, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000302.

Gzyl:1990:DSS

[Gzy90] Henryk Gzyl. Diffusions on some submanifolds of Euclidean spaces. Statistics & Probability Letters, 10(4):317–319, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090048C.

Han:1998:FRJ

[HA98] Qing Han and Sigeo Aki. Formulae and recursions for the joint distributions of success runs of several lengths in a two-state Markov chain. Statistics & Probability Letters, 40(3):203–214, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000856.

Hall:1990:LLD

[Hal90] Peter Hall. On the law of the logarithm for density estimators. Statistics & Probability Letters, 9(3):237–240, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090062C.

Hall:1995:BEE

[Hal95] Peter Hall. On the biases of error estimators in prediction problems. Statistics & Probability Letters, 24(3):257–262, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001817.

Hamza:1995:SUU

[Ham95] Kais Hamza. The smallest uniform upper bound on the distance between the mean and the median of the binomial and Poisson distributions. Statistics & Probability Letters, 23(1):21–25, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400090U.

Hamadouche:1998:EHC

[Ham98a] D. Hamadouche. Erratum: “Hölder convergence of smoothed empirical process: [Statist. Probab. Lett. 36 (1998) 393– 400]. Statistics & Probability Letters, 39(4):385–386, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529800073X. See [Ham98b].

Hamadouche:1998:WCS

[Ham98b] D. Hamadouche. Weak convergence of smoothed empirical process in Hölder spaces. Statistics & Probability Letters, 36(4):393– 400, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167715297000862. See erratum

[Ham98a].

Hanagal:1993:SIR

[Han93] David D. Hanagal. Some inference results in an absolutely continuous multivariate exponential model of Block. Statistics & Probability Letters, 16(3):177–180, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390140E.

Harris:1991:EFZ

[Har91] Ian R. Harris. The estimated frequency of zero for a mixed Poisson distribution. Statistics & Probability Letters, 12(5):371–372, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190024L.

Hardwick:1996:AER

[Har96] Janis Hardwick. Asymptotic expansions for randomly censored survival data. Statistics & Probability Letters, 26(2):187–197, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000097.

Hawkins:1990:NAD

[Haw90] D. L. Hawkins. A note on the asymptotic distribution of a statistic for testing stability of a correlation coefficient. Statistics & Probability Letters, 9(2):149–154, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900103.

Hawkins:1997:CFM

[Haw97] D. L. Hawkins. Can the finiteness of a mean be tested? Statistics & Probability Letters, 32(3):273–277, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000843.

Hazelton:1998:BAB

[Haz98] Martin L. Hazelton. Bias annihilating bandwidths for kernel density estimation at a point. Statistics & Probability Letters, 38(4):305–309, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000376.

Hwang:1997:LAN

[HB97] S. Y. Hwang and I. V. Basawa. The local asymptotic normality of a class of generalized random coefficient autoregressive processes. Statistics & Probability Letters, 34(2):165–170, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001782.

Huang:1999:LCQ

[HB99a] Mei Ling Huang and Percy Brill. A level crossing quantile estimation method. Statistics & Probability Letters, 45(2):111–119, November 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000498.

Hwang:1999:IBL

[HB99b] S. Y. Hwang and I. V. Basawa. Inference for a binary lattice Markov process. Statistics & Probability Letters, 43(1):75–85, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800248X.

Huang:1997:SLI

[HCJ97] Kyo Shin Huang, Yong Kab Choi, and Jong Soo Jung. On superior limits for the increments of Gaussian processes. Statistics & Probability Letters, 35(3):289–296, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000254.

He:1994:BVEa

[He94a] Xuming He. Breakdown versus efficiency — your perspective matters. Statistics & Probability Letters, 19(5):357–360, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900035.

He:1994:BVEb

[He94b] Xuming He. Breakdown versus efficiency — your perspective matters. Statistics & Probability Letters, 21(3):253, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901228.

Heckman:1992:BHR

[Hec92] Nancy E. Heckman. Bump hunting in regression analysis. Statistics & Probability Letters, 14(2):141–152, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290078J.

Heer:1991:TIH

[Hee91] Georg R. Heer. Testing independence in high dimensions. Statistics & Probability Letters, 12(1):73–81, July 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190166O.

Heinkel:1993:SCI

[Hei93] Bernard Heinkel. Some confidence intervals arising from weak l_{p} spaces. Statistics & Probability Letters, 18(1):33–40, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900962.

Hengjian:1997:LIL

[Hen97a] Cui Hengjian. The laws of the iterated logarithm for two kinds of PP statistics. Statistics & Probability Letters, 32(3):235–243, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529600079X.

Henze:1997:EST

[Hen97b] Norbert Henze. Extreme smoothing and testing for multivariate normality. Statistics & Probability Letters, 35(3):203–213, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000151.

Henze:1997:LLM

[Hen97c] Norbert Henze. Limit laws for multivariate skewness in the sense of Móri, Rohatgi and Székely. Statistics & Probability Letters, 33(3):299–307, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001411.

Henze:1998:DSH

[Hen98a] Norbert Henze. “Drawings since hit tables in lotteries” and a new multivariate geometric distribution. Statistics & Probability Letters, 40(4):321–327, November 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001187.

Henze:1998:PLL

[Hen98b] Norbert Henze. A Poisson limit law for a generalized birthday problem. Statistics & Probability Letters, 39(4):333–336, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000765.

Herschkorn:1997:BBS

[Her97] Stephen J. Herschkorn. Bandit bounds from stochastic variability extrema. Statistics & Probability Letters, 35(3):283–288, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000242.

Heyde:1994:QLA

[Hey94] C. C. Heyde. A quasi-likelihood approach to the REML estimating equations. Statistics & Probability Letters, 21(5):381–384, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294000352.

Hwang:1994:BLBa

[HH94a] Tea-Yuan Hwang and Chin-Yuan Hu. The best lower bound of sample correlation coefficient with ordered restriction. Statistics & Probability Letters, 19(3):195–198, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529490104X.

Hwang:1994:BLBb

[HH94b] Tea-Yuan Hwang and Chin-Yuan Hu. The best lower bound of sample correlation with ordered restriction. Statistics & Probability Letters, 21(3):254, October 19, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901236.

Hooghiemstra:1996:NMB

[HH96] G. Hooghiemstra and J. Hüsler. A note on maxima of bivariate random vectors. Statistics & Probability Letters, 31(1):1–6, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000053.

Hawkins:1998:CLT

[HH98a] D. L. Hawkins and Chien-Pai Han. A Central Limit Theorem for certain nonlinear statistics in repeated sampling of a finite population. Statistics & Probability Letters, 39(1):25–34, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000078.

Hu:1998:COS

[HH98b] Taizhong Hu and Jinjin Hu. Comparison of order statistics between dependent and independent random variables. Statistics & Probability Letters, 37(1):1–6, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000916.

Hu:1999:SCN

[HH99] Taizhong Hu and Jinjin Hu. Sufficient conditions for negative association of random variables. Statistics & Probability Letters, 45(2):167–173, November 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000565.

Hickernell:1999:GFS

[Hic99] Fred J. Hickernell. Goodness-of-fit statistics, discrepancies and robust designs. Statistics & Probability Letters, 44(1):73–78, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002934.

Hili:1999:EAM

[Hil99] Ouagnina Hili. On the estimation of β-ARCH models. Statistics & Probability Letters, 45(4):285–293, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900070X.

Hitczenko:1990:BCD

[Hit90] Paweł Hitczenko. Best constant in the decoupling inequality for non-negative random variables. Statistics & Probability Letters, 9(4):327–329, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090141S.

Huskova:1993:GBS

[HJ93] Marie Husková and Paul Jansen. Generalized bootstrap for Studentized U-statistics: a rank statistic approach. Statistics & Probability Letters, 16(3):225–233, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390147B.

Hallin:1998:CED

[HJM98] M. Hallin, J. Jurecková, and X. Milhaud. Characterization of error distributions in time-series regression models. Statistics & Probability Letters, 38(4):335–345, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000443.

Hendriks:1993:SUC

[HJR93] H. Hendriks, J. H. M. Janssen, and F. H. Ruymgaart. Strong uniform convergence of density estimators on compact Euclidean manifolds. Statistics & Probability Letters, 16(4):305–311, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901356.

Hanagal:1991:LST

[HK91] David D. Hanagal and B. K. Kale. Large sample tests of λ_{3} in the bivariate exponential distribution. Statistics & Probability Letters, 12(4):311–313, October 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190097B.

Houdre:1995:NAE

[HK95] Christian Houdré and Benjamin Kedem. A note on autocovariance estimation in the presence of discrete spectra. Statistics & Probability Letters, 24(1):1–8, July 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400141T.

Heimann:1996:BGF

[HK96] Günter Heimann and Jens-Peter Kreiss. Bootstrapping general first order autoregression. Statistics & Probability Letters, 30(1):87–98, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295002057.

Kang:1999:OPE

[hK99] Seung ho Kang. An optimal property of the exact multinomial test and the extended Fisher’s exact test. Statistics & Probability Letters, 44(2):201–207, August 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000103.

Hwang:1992:ENT

[HkL92] Jiunn T. Hwang and Hung kung Liu. Existence and nonexistence theorems of finite diameter sequential confidence regions for errors-in-variables models. Statistics & Probability Letters, 13(1):45–55, January 2, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290235W.

Hall:1999:NWO

[HKP99] Peter Hall, Gérard Kerkyacharian, and Dominique Picard. A note on the wavelet oracle. Statistics & Probability Letters, 43(4):415–420, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800282X.

Hsieh:1990:IML

[HKR90] H. K. Hsieh, R. M. Korwar, and A. L. Rukhin. inadmissibility of the maximum likelihood estimator of the inverse Gaussian mean. Statistics & Probability Letters, 9(1):83–90, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090099S.

Hurd:1992:EFC

[HL92] Harry L. Hurd and Jacek Leskow. Estimation of the Fourier coefficient functions and their spectral densities for ϕ-mixing almost periodically correlated processes. Statistics & Probability Letters, 14(4):299–306, July 17, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292900619.

Hochberg:1994:EST

[HL94] Yosef Hochberg and Uri Liberman. An extended Simes’ test. Statistics & Probability Letters, 21(2):101–105, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529490216X.

Hendriks:1996:ABS

[HL96] Harrie Hendriks and Zinoviy Landsman. Asymptotic behavior of sample mean location for manifolds. Statistics & Probability Letters, 26(2):169–178, February 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000070.

Hanson:1997:NED

[HL97] D. L. Hanson and Gang Li. A note on the empirical distribution of dependent random variables. Statistics & Probability Letters, 34(4):337–340, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001988.

Hernandez-Lerma:1996:EBI

[HLL96] Onésimo Hernández-Lerma and Jean B. Lasserre. Existence of bounded invariant probability densities for Markov chains. Statistics & Probability Letters, 28(4):359–366, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001468.

Hanson:1990:IES

[HM90] D. L. Hanson and Hari Mukerjee. Isotonic estimation in stochastic approximation. Statistics & Probability Letters, 9(3):279– 287, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090068I.

Heiligers:1996:LSA

[HM96] Berthold Heiligers and Augustyn Markiewicz. Linear sufficiency and admissibility in restricted linear models. Statistics & Probability Letters, 30(2):105–111, October 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002405.

Herzog:1998:GM

[HMG98] Jonathan Herzog, Christopher McLaren, and Anant P. Godbole. Generalized k-matches. Statistics & Probability Letters, 38(2):167–175, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297001697.

Hall:1996:CPD

[HMT96] Peter Hall, J. Steven Marron, and D. M. Titterington. A curious property of the derivatives of the Cauchy density. Statistics & Probability Letters, 27(3):225–229, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000690.

Hall:1997:CNI

[HN97] Peter Hall and Guy P. Nason. On choosing a non-integer resolution level when using wavelet methods. Statistics & Probability Letters, 34(1):5–11, May 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001599.

Hu:1998:CSA

[HN98] Yaozhong Hu and David Nualart. Continuity of some anticipating integral processes. Statistics & Probability Letters, 37 (2):203–211, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297001181.

Ho:1995:SUC

[Ho95a] Hwai-Chung Ho. On the strong uniform consistency of density estimation for strongly dependent sequences. Statistics & Probability Letters, 22(2):149–156, February 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400061C.

Hong:1995:WLL

[HO95b] Dug Hun Hong and Kwang Sik Oh. On the Weak Law of Large Numbers for arrays. Statistics & Probability Letters, 22(1):55– 57, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400047C.

Hoa:1993:MPE

[Hoa93] Tran Van Hoa. The mixture properties of the 2SHI estimators in linear regression models. Statistics & Probability Letters, 16(2):111–115, January 27, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390154B.

Hoffmann:1997:MED

[Hof97] Marc Hoffmann. Minimax estimation of the diffusion coefficient through irregular samplings. Statistics & Probability Letters, 32(1):11–24, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000521.

Hoffmann:1999:NEN

[Hof99] Marc Hoffmann. On nonparametric estimation in nonlinear AR(1)-models. Statistics & Probability Letters, 44(1):29–45, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002892.

Hughes-Oliver:1998:PNC

[HOGFLC98] Jacqueline M. Hughes-Oliver, Graciela Gonzalez-Farias, Jye-Chyi Lu, and Di Chen. Parametric nonstationary correlation models. Statistics & Probability Letters, 40(3):267–278, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001035.

Holst:1990:CCP

[Hol90] Lars Holst. A circle covering problem and DNA breakage. Statistics & Probability Letters, 9(4):295–298, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090134S.

Holst:1991:PMP

[Hol91] Lars Holst. On the ‘problème des ménages’ from a probabilistic viewpoint. Statistics & Probability Letters, 11(3):225–231, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190147J.

Holiday:1995:NOW

[Hol95] David B. Holiday. Near optimal weights in nonparametric regression under some common restrictions. Statistics & Probability Letters, 22(1):33–42, January 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000449.

Holden:1998:GCM

[Hol98] Lars Holden. Geometric convergence of the Metropolis–Hastings simulation algorithm. Statistics & Probability Letters, 39 (4):371–377, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000960.

Hong:1996:WLL

[Hon96] Dug Hun Hong. On the Weak Law of Large Numbers for randomly indexed partial sums for arrays. Statistics & Probability Letters, 28(2):127–130, June 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001069.

Horowitz:1990:ULL

[Hor90] Joseph Horowitz. A uniform Law of Large Numbers and empirical Central Limit Theorem for limits of finite populations. Statistics & Probability Letters, 10(2):159–166, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090012V.

Hornik:1993:DFS

[Hor93] Kurt Hornik. On the distribution of functionals of stationary Gaussian processes. Statistics & Probability Letters, 18 (5):389–395, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390033F.

Horstmann:1995:CRI

[Hor95] Vibeke Horstmann. On confidence regions induced by the Wilcoxon rank sum test. Statistics & Probability Letters, 23(4): 333–338, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400132R.

Hossjer:1992:OEO

[Hös92] Ola Hössjer. On the optimality of S-estimators. Statistics & Probability Letters, 14(5):413–419, July 11, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290103C.

Hosking:1994:MOS

[Hos94] J. R. M. Hosking. Moments of order statistics of the Cantor distribution. Statistics & Probability Letters, 19(2):161–165, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901503.

Houdre:1997:IJE

[Hou97] Christian Houdré. The iterated jackknife estimate of variance. Statistics & Probability Letters, 35(2):197–201, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700014X.

Howard:1997:DCR

[How97] C. Douglas Howard. Distinguishing certain random sceneries on Z via random walks. Statistics & Probability Letters, 34(2):123– 132, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001733.

Hua:1993:SOA

[HP93a] Liang Hua and Cheng Ping. Second order asymptotic efficiency in a partial linear model. Statistics & Probability Letters, 18(1):73– 84, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390101N.

Hwang:1993:CIE

[HP93b] J. T. Gene Hwang and Shyamal D. Peddada. Confidence interval estimation under some restrictions on the parameters with nonlinear boundaries. Statistics & Probability Letters, 18(5):397–403, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390034G.

Harris:1994:GEN

[HP94] Bernard Harris and C. J. Park. A generalization of the Eulerian numbers with a probabilistic application. Statistics & Probability Letters, 20(1):37–47, May 4, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902321.

Hobert:1997:CCC

[HRG97] J. P. Hobert, C. P. Robert, and C. Goutis. Connectedness conditions for the convergence of the Gibbs sampler. Statistics & Probability Letters, 33(3):235–240, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001320.

Hum:1999:BUC

[HRS99] Merrilee A. Hum, Håvard Rue, and Nuala A. Sheehan. Block updating in constrained Markov chain Monte Carlo sampling. Statistics & Probability Letters, 41(4):353–361, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529890150X.

Harris:1991:TCC

[HS91] Bernard Harris and Andrew P. Soms. Theory and counterexamples for confidence limits on system reliability. Statistics & Probability Letters, 11(5):411–417, May 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291901903.

Horvath:1994:NDT

[HS94a] Lajos Horváth and Qi-Man Shao. A note on dichotomy theorems for integrals of stable processes. Statistics & Probability Letters, 19(1):45–49, January 3, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900671.

Hwang:1994:MRF

[HS94b] Gene T. Hwang and Leonard A. Stefanski. Monotonicity of regression functions in structural measurement error models. Statistics & Probability Letters, 20(2):113–116, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900256.

Horvath:1999:BAB

[HS99] Lajos Horváth and Josef Steinebach. On the best approximation for bootstrapped empirical processes. Statistics & Probability Letters, 41(2):117–122, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001254.

Hsieh:1995:RAE

[Hsi95] Fushing Hsieh. A regression approach for estimating the center of symmetry. Statistics & Probability Letters, 22(2):157–160, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400062D.

Hu:1998:NCC

[HSV98] T.-C. Hu, D. Szynal, and A. I. Volodin. A note on complete convergence for arrays. Statistics & Probability Letters, 38(1):27– 31, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001503. See [HV00].

Hurvich:1990:MSL

[HT90] Clifford M. Hurvich and Chih-Ling Tsai. Model selection for least absolute deviations regression in small samples. Statistics & Probability Letters, 9(3):259–265, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090065F.

Huang:1991:MBD

[HT91] Wen-Tao Huang and Chiou-Shiang Tsai. On a modified binomial distribution of order k. Statistics & Probability Letters, 11(2):125–131, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190129F.

Hsieh:1995:NLA

[HT95] Fushing Hsieh and Bruce W. Turnbull. A note on the local asymptotically minimax rate for estimating a crossing point in a diagnostic marker problem. Statistics & Probability Letters, 24(2):181–185, August 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001677.

Hurvich:1996:IUS

[HT96] Clifford M. Hurvich and Chih-Ling Tsai. The impact of unsuspected serial correlations on model selection in linear regression. Statistics & Probability Letters, 27(2):115–126, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000526.

Hu:1995:NSN

[Hu95] Xiaomi Hu. A note on the symmetry of normal mean hypotheses and its implications. Statistics & Probability Letters, 25(1):15– 20, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400200R.

Hu:1997:MLE

[Hu97] Xiaomi Hu. Maximum-likelihood estimation under bound restriction and order and uniform bound restrictions. Statistics & Probability Letters, 35(2):165–171, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000102.

Huang:1991:EVS

[Hua91] J. S. Huang. Estimating the variance of the sample median, discrete case. Statistics & Probability Letters, 11(4):291–298, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190038S.

Huang:1997:SBE

[Hua97a] J. S. Huang. Sharp bounds for the expected value of order statistics. Statistics & Probability Letters, 33(1):105–107, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001162.

Huang:1997:WBE

[Hua97b] Su-Yun Huang. Wavelet based empirical Bayes estimation for the uniform distribution. Statistics & Probability Letters, 32(2):141– 146, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000661.

Huang:1998:SEM

[Hua98] J. S. Huang. Sequence of expectations of maximum-order statistics. Statistics & Probability Letters, 38(2):117–123, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001612.

Huang:1999:TOE

[Hua99] J. S. Huang. Third-order expansion of mean squared error of medians. Statistics & Probability Letters, 42(2):185–192, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002119.

Hubert:1997:BVE

[Hub97] Mia Hubert. The breakdown value of the L_{1}-estimator in contingency tables. Statistics & Probability Letters, 33(4):419–425, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001563.

Huda:1997:MSO

[Hud97] S. Huda. Minimax second-order designs over hypercubes for the difference between estimated responses at a point and at the centre. Statistics & Probability Letters, 33(2):193–199, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001277.

Huggins:1996:IME

[Hug96] Richard Huggins. On the identifiability of measurement error in the bifurcating autoregressive model. Statistics & Probability Letters, 27(1):17–23, March 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000380.

Hughes:1997:COI

[Hug97] James P. Hughes. Computing the observed information in the hidden Markov model using the EM algorithm. Statistics & Probability Letters, 32(1):107–114, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000624.

Hurlimann:1993:BDD

[Hür93] Werner Hürlimann. Bivariate distributions with diatomic conditionals and stop-loss transforms of random sums. Statistics & Probability Letters, 17(4):329–335, July 13, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390211Z.

Hurlimann:1995:UAS

[Hür95] Werner Hürlimann. A uniform approximation to the sampling distribution of the coefficient of variation. Statistics & Probability Letters, 24(3):263–268, August 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001828.

Husler:1994:MBR

[Hüs94] J. Hüsler. Maxima of bivariate random vectors: Between independence and complete dependence. Statistics & Probability Letters, 21(5):385–394, December 7, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000360.

Huskova:1996:ECL

[Hus96] Marie Husková. Estimation of a change in linear models. Statistics & Probability Letters, 26(1):13–24, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002479.

Huskova:1997:LTR

[Hus97] M. Husková. Limit theorems for rank statistics. Statistics & Probability Letters, 32(1):45–55, February 14, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000557.

Husler:1999:EGP

[Hüs99] J. Hüsler. Extremes of Gaussian processes, on results of Piterbarg and Seleznjev. Statistics & Probability Letters, 44(3):251–258, September 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000164.

Huwang:1995:NAA

[Huw95] Longcheen Huwang. A note on the accuracy of an approximate interval for the binomial parameter. Statistics & Probability Letters, 24(2):177–180, August 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001666.

Hebbar:1997:LPS

[HV97] H. V. Hebbar and N. Vadiraja. Limit points of sequences of moving maxima. Statistics & Probability Letters, 34(1):13–18, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001605.

Hu:2000:ANC

[HV00] Tien-Chung Hu and Andrei Volodin. Addendum to “A note on complete convergence for arrays”, Statist. Probab. Lett. 38 (1) (1998) 27–31. Statistics & Probability Letters, 47(2):209– 211, April 1, 2000. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299002096. See [HSV98].

Hansen:1990:GLS

[HW90] B. G. Hansen and E. Willekens. The generalized logarithmic series distribution. Statistics & Probability Letters, 9(4):311– 316, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090138W.

Hall:1995:EIP

[HW95] Peter Hall and Rodney C. L. Wolff. Estimators of integrals of powers of density derivatives. Statistics & Probability Letters, 24(2):105–110, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400154Z.

Hart:1997:SLM

[HW97] Sergiu Hart and Benjamin Weiss. Significance levels for multiple tests. Statistics & Probability Letters, 35(1):43–48, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002143.

Hwang:1999:TSA

[Hwa99] Leng-Cheng Hwang. Two-stage approach to Bayes sequential estimation in the exponential distribution. Statistics & Probability Letters, 45(3):277–282, November 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000681.

Howard:1999:SIO

[HZ99] Peter Howard and Kevin Zumbrun. Shift invariance of the occupation time of the Brownian bridge process. Statistics & Probability Letters, 45(4):379–382, December 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000802.

Ioannides:1997:NRE

[IA97] D. A. Ioannides and P. D. Alevizos. Nonparametric regression with errors in variables and applications. Statistics & Probability Letters, 32(1):35–43, February 14, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000545.

Ihara:1992:AEU

[IK92] Masamori Ihara and Yutaka Kano. Asymptotic equivalence of unique variance estimators in marginal and conditional factor analysis models. Statistics & Probability Letters, 14(5):337–341, July 11, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290093K.

Ihara:1995:IFM

[IK95] Masamori Ihara and Yutaka Kano. Identifiability of full, marginal, and conditional factor analysis models. Statistics & Probability Letters, 23(4):343–350, June 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400134T.

Ibragimov:1998:CGM

[IL98] Ildar Ibragimov and Mikhail Lifshits. On the convergence of generalized moments in almost sure Central Limit Theorem. Statistics & Probability Letters, 40(4):343–351, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001345.

Ivanoff:1997:PCS

[IM97a] B. Gail Ivanoff and Ely Merzbach. Poisson convergence for set-indexed empirical processes. Statistics & Probability Letters, 32(1):81–86, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000594.

Iwasa:1997:NAM

[IM97b] Manabu Iwasa and Yoshiya Moritani. A note on admissibility of the maximum likelihood estimator for a bounded normal mean. Statistics & Probability Letters, 32(1):99–105, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000612.

Imhof:1997:CED

[Imh97] Lorens Imhof. A counterexample in experimental design showing that the G-criterion function is not necessarily strictly decreasing. Statistics & Probability Letters, 35(4):341–344, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700031X.

Ivanoff:1995:LRS

[IMSK95] B. Gail Ivanoff, Ely Merzbach, and Ioana Schiopu-Kratina. Lattices of random sets and progressivity. Statistics & Probability Letters, 22(2):97–102, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400054C.

Indurkhya:1994:PBP

[Ind94] Alka Indurkhya. A parametric bootstrap procedure to estimate the selected mean using censored data. Statistics & Probability Letters, 21(4):291–298, November 11, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000220.

Ioannides:1992:ISE

[Ioa92] Dimitrios A. Ioannides. Integrated square error of nonparametric estimators of regression function: the fixed design case. Statistics & Probability Letters, 15(2):85–94, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290118O.

Ioffe:1993:TET

[Iof93] Dmitry Ioffe. Two examples in the theory of large deviations. Statistics & Probability Letters, 18(4):297–300, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390018E.

Ioannides:1999:EIA

[IR99] D. A. Ioannides and G. G. Roussas. Exponential inequality for associated random variables. Statistics & Probability Letters, 42(4):423–431, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002405.

Irle:1996:AOS

[Irl96] A. Irle. Asymptotically optimal stopping rules in the presence of unknown parameters. Statistics & Probability Letters, 30(1):9–16, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001956.

Irle:1997:BHM

[Irl97] A. Irle. A bound for higher moments of expected sample size in sequential testing. Statistics & Probability Letters, 35 (2):135–140, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000060.

Isaac:1999:ERD

[Isa99] Richard Isaac. On equitable ratios of Dubins–Freedman type. Statistics & Probability Letters, 42(1):1–6, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001722.

Isogawa:1995:NUT

[Iso95] Yoshiko Isogawa. A note on a unified test for multivariate linear relationships. Statistics & Probability Letters, 24(2):147–151, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294001596.

Isogai:1995:SPE

[IU95] Eiichi Isogai and Chikara Uno. On the sequential point estimation of the mean of a gamma distribution. Statistics & Probability Letters, 22(4):287–293, March 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400079N.

Ivanoff:1996:STT

[Iva96] B. Gail Ivanoff. Stopping times and tightness for multiparameter martingales. Statistics & Probability Letters, 28(2):111–114, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001034.

Ip:1999:TVS

[IWLX99] Wai-Cheung Ip, Heung Wong, Yuan Li, and Zhongjie Xie. Threshold variable selection by wavelets in open-loop threshold autoregressive models. Statistics & Probability Letters, 42(4):375–392, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002326.

Yoshihara:1995:BRS

[iY95] Ken ichi Yoshihara. The Bahadur representation of sample quantiles for sequences of strongly mixing random variables. Statistics & Probability Letters, 24(4):299–304, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400187D.

Iyengar:1994:MLE

[IZ94] Satish Iyengar and Peng-Liang Zhao. Maximum likelihood estimation for weighted distributions. Statistics & Probability Letters, 21(1):37–47, September 1, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900566.

Jakubowski:1994:MDA

[Jak94] Adam Jakubowski. On multidimensional domains of attraction for stationary sequences. Statistics & Probability Letters, 19(4):321–326, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294901848.

Janssen:1994:LOH

[Jan94] Arnold Janssen. On local odds and hazard rate models in survival analysis. Statistics & Probability Letters, 20(5):355–365, August 9, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901279.

Janssen:1997:SPT

[Jan97] Arnold Janssen. Studentized permutation tests for noni.i.d. hypotheses and the generalized Behrens–Fisher problem. Statistics & Probability Letters, 36(1):9–21, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000436.

Joag-dev:1995:GCT

[JdKP95] Kumar Joag-dev, Subhash Kochar, and Frank Proschan. A general composition theorem and its applications to certain partial orderings of distributions. Statistics & Probability Letters, 22(2):111–119, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400056E.

Joag-Dev:1990:CIC

[JDP90] Kumar Joag-Dev and Frank Proschan. A covariance inequality for coherent structures. Statistics & Probability Letters, 9(3):215– 216, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090058F.

Johnson:1999:NBP

[JEG99] Richard A. Johnson, James W. Evans, and David W. Green. Nonparametric Bayesian predictive distributions for future order statistics. Statistics & Probability Letters, 41(3):247–254, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001618.

Jensen:1992:URR

[Jen92] D. R. Jensen. The use of randomization in repeated measurements. Statistics & Probability Letters, 13(3):229–233, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900527.

Jensen:1993:ECI

[Jen93] D. R. Jensen. Efficiency and concentration inequalities on R^{k}. Statistics & Probability Letters, 18(3):209–211, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293902188.

Jensen:1994:CMR

[Jen94a] D. R. Jensen. Closure of multivariate t and related distributions. Statistics & Probability Letters, 20(4):307–312, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900183.

Jensen:1994:ESS

[Jen94b] D. R. Jensen. Efficiencies of some small second-order designs. Statistics & Probability Letters, 21(4):255–261, November 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529490135X.

Jensen:1996:DAP

[Jen96] D. R. Jensen. Deletion, augmentation and principal predictors. Statistics & Probability Letters, 29(3):263–270, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001816.

Jensen:1997:PLF

[Jen97] D. R. Jensen. Peakedness of linear forms in ensembles and mixtures. Statistics & Probability Letters, 35(3):277–282, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000230.

Jeske:1993:PVI

[Jes93] Daniel R. Jeske. Predicting the value of an integer-valued random variable. Statistics & Probability Letters, 16(4):297–300, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901334.

Jevremovic:1990:TEN

[Jev90] Vesna Jevremović. Two examples of nonlinear processes with a mixed exponential marginal distribution. Statistics & Probability Letters, 10(3):221–224, August 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090077K.

Jevremovic:1991:NME

[Jev91] Vesna Jevremovic. A note on mixed exponential distribution with negative weights. Statistics & Probability Letters, 11(3):259– 265, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190153I.

Jeyaratnam:1991:RAE

[Jey91] S. Jeyaratnam. A robust affine-equivariant median. Statistics & Probability Letters, 11(2):167–168, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190136F.

Jeyaratnam:1992:CIC

[Jey92] S. Jeyaratnam. Confidence intervals for the correlation coefficient. Statistics & Probability Letters, 15(5):389–393, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901722.

Jones:1990:MSE

[JH90] M. C. Jones and Peter Hall. Mean squared error properties of kernel estimates or regression quantiles. Statistics & Probability Letters, 10(4):283–289, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900437.

Jiang:1991:DRF

[Jia91] Tom J. Jiang. Distribution of random functional of a Dirichlet process on the unit disk. Statistics & Probability Letters, 12(3):263–265, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715291900889.

Jiang:1997:DBB

[Jia97a] Jiming Jiang. A derivation of BLUP — Best Linear Unbiased Predictor. Statistics & Probability Letters, 32(3):321–324, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000892.

Jiang:1997:SUL

[Jia97b] Jiming Jiang. Sharp upper and lower bounds for asymptotic levels of some statistical tests. Statistics & Probability Letters, 35(4):395–400, November 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000370.

Jiang:1999:UEB

[Jia99] Jiming Jiang. On unbiasedness of the empirical BLUE and BLUP. Statistics & Probability Letters, 41(1):19–24, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001151.

Jin:1993:NSE

[Jin93] Chun Jin. A note on simultaneous estimation of eigenvalues of a multivariate normal covariance matrix. Statistics & Probability Letters, 16(3):197–203, February 19, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293901437.

Jing:1995:TSE

[Jin95] Bing-Yi Jing. Two-sample empirical likelihood method. Statistics & Probability Letters, 24(4):315–319, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400189F.

Jing:1996:GTM

[Jin96] Bing-Yi Jing. Global tilting method. Statistics & Probability Letters, 28(3):219–226, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001271.

Johnson:1995:UMS

[JK95a] Norman L. Johnson and Samuel Kotz. Use of moments in studies of limit distributions arising from iterated random subdivisions of an interval. Statistics & Probability Letters, 24(2):111–119, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294001552.

Jordan:1995:CIT

[JK95b] Scott M. Jordan and K. Krishnamoorthy. On combining independent tests in linear models. Statistics & Probability Letters, 23(2):117–122, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400102E.

Jang:1998:USL

[JK98] Lee-Chae Jang and Joong-Sung Kwon. A uniform Strong Law of Large Numbers for partial sum processes of Banach space-valued random sets. Statistics & Probability Letters, 38(1):21–25, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001491.

Johnson:1999:STD

[JK99a] Norman L. Johnson and Samuel Kotz. Square tray distributions. Statistics & Probability Letters, 42(2):157–165, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002041.

Jureckova:1999:TBA

[JK99b] Jana Jurećková and Lev B. Klebanov. Trimmed, Bayesian and admissible estimators. Statistics & Probability Letters, 42(1):47– 51, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001874.

Joe:1996:MMB

[JL96] Harry Joe and Ying Liu. A model for a multivariate binary response with covariates based on compatible conditionally specified logistic regressions. Statistics & Probability Letters, 31(2):113–120, December 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000211.

Johnson:1990:TSR

[JM90] Richard A. Johnson and Christopher H. Morrell. Two sample rank tests under a random truncation model. Statistics & Probability Letters, 9(5):409–422, May 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900334.

Jensen:1997:CML

[JN97] S. T. Jensen and B. Nielsen. On convergence of multivariate Laplace transforms. Statistics & Probability Letters, 33(2):125– 128, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001198.

Joe:1990:FMS

[Joe90] Harry Joe. Families of min-stable multivariate exponential and multivariate extreme value distributions. Statistics & Probability Letters, 9(1):75–81, January 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090098R.

Joe:1991:RSB

[Joe91] Harry Joe. Rating systems based on paired comparison models. Statistics & Probability Letters, 11(4):343–347, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190046T.

Joe:1993:TUS

[Joe93] Harry Joe. Tests of uniformity for sets of lotto numbers. Statistics & Probability Letters, 16(3):181–188, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901415.

Johnstone:1991:MIE

[Joh91] Iain M. Johnstone. A moment inequality for L_{q} estimation. Statistics & Probability Letters, 12(4):289–290, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900937.

Johansson:1994:CLT

[Joh94a] Björn Johansson. On the Central Limit Theorem for point process martingales. Statistics & Probability Letters, 20(2):125–130, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900272.

Johnson:1994:E

[Joh94b] Richard A. Johnson. Editorial. Statistics & Probability Letters, 19(5):355, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900027.

John:1996:NPB

[Joh96] Peter W. M. John. Nesting Plackett–Burman designs. Statistics & Probability Letters, 27(3):221–223, April 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000682.

Johnson:1998:E

[Joh98] Richard Johnson. Editorial. Statistics & Probability Letters, 40(4):305, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001709.

Jones:1990:PKD

[Jon90] M. C. Jones. The performance of kernel density functions in kernel distribution function estimation. Statistics & Probability Letters, 9(2):129–132, February 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290006Q.

Jones:1991:RIM

[Jon91] M. C. Jones. The roles of ISE and MISE in density estimation. Statistics & Probability Letters, 12(1):51–56, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190163L.

Jones:1992:PAB

[Jon92] M. C. Jones. Potential for automatic bandwidth choice in variations on kernel density estimation. Statistics & Probability Letters, 13(5):351–356, April 7, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290107G.

Jones:1994:EQQ

[Jon94] M. C. Jones. Expectiles and M-quantiles are quantiles. Statistics & Probability Letters, 20(2):149–153, May 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900310.

Jones:1995:TRP

[Jon95] M. C. Jones. On two recent papers of Y. Kanazawa. Statistics & Probability Letters, 24(3):269–271, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001839.

Janosky:1997:NRL

[JPAS97] J. E. Janosky, T. R. Pellitieri, and Q. M. Al-Shboul. The need for a revised lower limit for the 4253H, Twice nonparametric smoother. Statistics & Probability Letters, 32(3):269–272, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000831.

Justel:1997:MKS

[JPZ97] Ana Justel, Daniel Peña, and Rubén Zamar. A multivariate Kolmogorov–Smirnov test of goodness of fit. Statistics & Probability Letters, 35(3):251–259, October 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000205.

Jin:1999:SAN

[JR99] Rungao Jin and John Robinson. Saddlepoint approximation near the endpoints of the support. Statistics & Probability Letters, 45(4):295–303, December 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000711.

Jureckova:1990:RVS

[JS90] Jana Jurecková and A. K. Md. Ehsanes Saleh. Robustified version of Stein’s multivariate location estimation. Statistics & Probability Letters, 9(4):375–380, April 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290901492.

Jacob:1996:REE

[JS96] P. Jacob and Ch. Suquet. Regression and edge estimation. Statistics & Probability Letters, 27(1):11–15, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000372.

Jacob:1997:RAL

[JS97]  P. Jacob and Ch. Suquet.  Regression and asymptoti-

cal location of a multivariate sample. Statistics & Probability Letters, 35(2):173–179, September 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000114.

Jammalamadaka:1998:PID

[JS98] S. Rao Jammalamadaka and Ashis SenGupta. Predictive inference for directional data. Statistics & Probability Letters, 40(3):247–257, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001011.

Jin:1999:KEM

[JS99] Zhezhen Jin and Yongzhao Shao. On kernel estimation of a multivariate distribution function. Statistics & Probability Letters, 41(2):163–168, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001382.

Jumarie:1999:CVW

[Jum99] Guy Jumarie. Complex-valued Wiener measure: An approach via random walk in the complex plane. Statistics & Probability Letters, 42(1):61–67, March 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001941.

Johnson:1991:LSD

[JV91] Richard A. Johnson and Steve Verrill. The large sample distribution of the Shapiro–Wilk statistic and its variants under Type I or Type II censoring. Statistics & Probability Letters, 12(5):405– 413, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190029Q.

Janssen:1990:SWT

[JVS90] Paul Janssen, Noël Veraverbeke, and Robert Serfling. Ustatistics on Winsorized and trimmed samples. Statistics & Probability Letters, 9(5):439–447, May 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900367.

Jensen:1997:NEB

[JW97] J. L. Jensen and Andrew T. A. Wood. On the non-existence of a Bartlett correction for unit root tests. Statistics & Probability Letters, 35(2):181–187, September 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000126.

Jung:1999:AGF

[JW99] Sin-Ho Jung and Sam Wieand. Analysis of goodness-of-fit for Cox regression model. Statistics & Probability Letters, 41 (4):379–382, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001527.

Jiang:1992:MDS

[JWR92] Tiefeng Jiang, Xiangchen Wang, and M. Bhaskara Rao. Moderate deviations for some weakly dependent random processes. Statistics & Probability Letters, 15(1):71–76, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290287F.

Koutras:1997:NPR

[KA97] M. V. Koutras and V. A. Alexandrou. Non-parametric randomness tests based on success runs of fixed length. Statistics & Probability Letters, 32(4):393–404, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000995.

Kim:1999:PEU

[KA99] YongHee Kim and Barry C. Arnold. Parameter estimation under generalized ranked set sampling. Statistics & Probability Letters, 42(4):353–360, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002259.

Kabaila:1998:NSI

[Kab98] Paul Kabaila. A note on sufficiency and information loss. Statistics & Probability Letters, 37(2):111–114, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001077.

Kabaila:1999:CIS

[Kab99] Paul Kabaila. Confidence intervals from simulations based on 4-independent random variables. Statistics & Probability Letters, 45(2):141–147, November 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529900053X.

Kageyama:1990:CCO

[Kag90] Sanpei Kageyama. Complete characterization of optimum weighing designs for estimating the total weight. Statistics & Probability Letters, 9(1):99–100, January 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090101C.

Kageyama:1991:IUB

[Kag91] Sanpei Kageyama. An improved upper bound in a proper non-binary variance-balanced design. Statistics & Probability Letters, 11(4):319–320, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190041O.

Kaigh:1999:TTT

[Kai99] W. D. Kaigh. Total time on test function principal components. Statistics & Probability Letters, 44(4):337–341, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000243.

Kakizawa:1997:PEH

[Kak97] Yoshihide Kakizawa. Parameter estimation and hypothesis testing in stationary vector time series. Statistics & Probability Letters, 33(3):225–234, May 5, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001319.

Kakizawa:1998:ERE

[Kak98] Yoshihide Kakizawa. On exponential rates of estimators of the parameter in the first-order autoregressive process. Statistics & Probability Letters, 38(4):355–362, July 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000467.

Kallenberg:1990:INC

[Kal90] Wilbert C. M. Kallenberg. Inequalities for noncentral chisquare distributions. Statistics & Probability Letters, 9(3):273– 278, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090067H.

Kaluszka:1998:DGM

[Kal98] M. Kaluszka. On the Devroye–Györfi methods of correcting density estimators. Statistics & Probability Letters, 37(3):249–257, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297841545.

Kanjo:1991:IEB

[Kan91a] A. I. Kanjo. Improving estimates in B.I.B. designs. Statistics & Probability Letters, 12(2):167–174, August 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190063W.

Kanjo:1991:IEP

[Kan91b] A. I. Kanjo. Improving estimates in P.B.I.B. designs. Statistics & Probability Letters, 12(2):175–181, August 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190064X.

Kano:1992:RST

[Kan92] Yutaka Kano. Robust statistics for test-of-independence and related structural models. Statistics & Probability Letters, 15(1):21–26, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290279E.

Kanazawa:1993:HDA

[Kan93a] Yuichiro Kanazawa. Hellinger distance and Akaike’s information criterion for the histogram. Statistics & Probability Letters, 17(4):293–298, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390205W.

Kanazawa:1993:HDK

[Kan93b] Yuichiro Kanazawa. Hellinger distance and Kullback–Leibler loss for the kernel density estimator. Statistics & Probability Letters, 18(4):315–321, November 9, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390022B.

Kanter:1997:USW

[Kan97] Marek Kanter. Unimodal spectral windows. Statistics & Probability Letters, 34(4):403–411, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296002088.

Karczewska:1996:SSS

[Kar96] Anna Karczewska. Statistical solutions for a stochastic Burger’s equation. Statistics & Probability Letters, 27(4):305–311, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000909.

Katkovnik:1997:NET

[Kat97] Vladimir Katkovnik. Nonparametric estimation of the timevarying frequency and amplitude. Statistics & Probability Letters, 35(4):307–315, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000278.

Kato:1999:DET

[Kat99] Takeshi Kato. Density estimation by truncated wavelet expansion. Statistics & Probability Letters, 43(2):159–168, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002557.

Kaufmann:1995:MCN

[Kau95] Edgar Kaufmann. Von Mises conditions, δ-neighborhoods and rates of convergence for maxima. Statistics & Probability Letters, 25(1):63–70, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400206N.

Kay:1992:ACF

[Kay92] Jim Kay. Asymptotic comparison factors for smoothing parameter choices in regression problems. Statistics & Probability Letters, 15(4):329–335, November 13, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290170A.

Kouider:1995:CBC

[KC95] Elies Kouider and Hanfeng Chen. Concavity of Box–Cox log-likelihood function. Statistics & Probability Letters, 25 (2):171–175, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400219X.

Kim:1996:USC

[KC96] Tae Yoon Kim and Dennis D. Cox. Uniform strong consistency of kernel density estimators under dependence. Statistics & Probability Letters, 26(2):179–185, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000089.

Kaiser:1997:MPV

[KC97a] Mark S. Kaiser and Noel Cressie. Modeling Poisson variables with positive spatial dependence. Statistics & Probability Letters, 35(4):423–432, November 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000412.

Kochar:1997:CAV

[KC97b]  Subhash C. Kochar and K. C. Carrière.  Connections

among various variability orderings. Statistics & Probability Letters, 35(4):327–333, November 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000291.

Kale:1992:EEP

[KD92] Mohan M. Kale and S. R. Deshmukh. Estimation of extinction probability in Markov branching process. Statistics & Probability Letters, 15(3):173–175, October 21, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292901858.

Kale:1993:MLE

[KD93] Mohan M. Kale and S. R. Deshmukh. Maximum likelihood estimation in branching process with continuous state space. Statistics & Probability Letters, 18(3):245–251, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293902247.

Kuber:1996:ATF

[KD96] Madhuri Kuber and Avinash Dharmadhikari. Association in time of a finite semi-Markov process. Statistics & Probability Letters, 26(2):125–133, February 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002622.

Keprta:1997:NSK

[Kep97] Stanislav Keprta. A note on supremum of a Kiefer process. Statistics & Probability Letters, 35(1):59–63, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002167.

Kabe:1990:MCR

[KG90a] D. G. Kabe and A. K. Gupta. On a multiple correlation ratio. Statistics & Probability Letters, 9(5):449–451, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900378.

Khuri:1990:MSS

[KG90b] AndréI. Khuri and Malay Ghosh. Minimal sufficient statistics for the unbalanced two-fold nested model. Statistics & Probability Letters, 10(4):351–353, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090054B.

Kohatsu-Hia:1991:WCI

[KH91] Arturo Kohatsu-Hia. Weak convergence of infinite order Uprocesses. Statistics & Probability Letters, 12(2):145–150, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190059Z.

Khan:1992:RDS

[Kha92] Rasul A. Khan. A remark on Dubins–Savage inequality. Statistics & Probability Letters, 13(3):189–192, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900468.

Khrennikov:1998:ABB

[Khr98] Andrew Khrennikov. p-Adic behaviour of Bernoulli probabilities. Statistics & Probability Letters, 37(4):375–379, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001405.

Khuri:1997:MSS

[Khu97] AndréI. Khuri. Minimal sufficient statistics for a general class of mixed models. Statistics & Probability Letters, 35(1):1–7, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002106.

King:1991:TET

[KHW91] Eileen King, Jeffrey D. Hart, and Thomas E. Wehrly. Testing the equality of two regression curves using linear smoothers. Statistics & Probability Letters, 12(3):239–247, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900856.

Kim:1992:BCP

[Kim92] Hyune-Ju Kim. Boundary crossing probabilities by nondifferentiable processes and applications to two-phase regression. Statistics & Probability Letters, 14(2):97–102, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290070L.

Kim:1993:NMB

[Kim93] Tae Yoon Kim. A note on moment bounds for strong mixing sequences. Statistics & Probability Letters, 16(2):163–168, January 27, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390162C.

Kim:1996:CDS

[Kim96] Choongrak Kim. Cook’s distance in spline smoothing. Statistics & Probability Letters, 31(2):139–144, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000259.

Kim:1999:SCB

[Kim99a] Ji-Hyun Kim. Spurious correlation between ratios with a common divisor. Statistics & Probability Letters, 44(4):383–386, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000309.

Kim:1999:TPK

[Kim99b] Tae Yoon Kim. On tail probabilities of Kolmogorov–Smirnov statistics based on uniform mixing processes. Statistics & Probability Letters, 43(3):217–223, July 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002545.

Kim:1999:PEI

[Kim99c] Yoon Tae Kim. Parameter estimation in infinite-dimensional stochastic differential equations. Statistics & Probability Letters, 45(3):195–204, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000590.

Kirmani:1996:SSI

[Kir96] S. N. U. A. Kirmani. On sample spacings from IMRL distributions. Statistics & Probability Letters, 29(2):159–166, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001697.

Kirmani:1998:SSI

[Kir98] S. N. U. A. Kirmani. On sample spacings from IMRL distributions. Statistics & Probability Letters, 37(3):315, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001326.

Kotz:1991:NRP

[KJ91] Samuel Kotz and Norman L. Johnson. A note on renewal (partial sums) distributions for discrete variables. Statistics & Probability Letters, 12(3):229–231, September 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291900834.

Kallenberg:1992:HLO

[KK92] Wilbert C. M. Kallenberg and Stavros Kourouklis. Hodges– Lehmann optimality of tests. Statistics & Probability Letters, 14(1):31–38, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290207L.

Kallenberg:1995:WTO

[KK95] Wilbert C. M. Kallenberg and Alex J. Koning. On Wieand’s theorem. Statistics & Probability Letters, 25(2):121–132, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400214S.

Kang:1996:TMM

[KK96a] Chul Kang and Byung-Chun Kim. The N-th moment of matrix quadratic form. Statistics & Probability Letters, 28(4):291– 297, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001387.

Koo:1996:WDE

[KK96b] Ja-Yong Koo and Woo-Chul Kim. Wavelet density estimation by approximation of log-densities. Statistics & Probability Letters, 26(3):271–278, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000208.

Khaledi:1999:SOB

[KK99] Baha-Eldin Khaledi and Subhash Kochar. Stochastic orderings between distributions and their sample spacings — II. Statistics & Probability Letters, 44(2):161–166, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000048.

Kagan:1997:SMC

[KL97] Abram Kagan and Zinoviy Landsman. Statistical meaning of Carlen’s superadditivity of the Fisher information. Statistics & Probability Letters, 32(2):175–179, March 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000703.

Kokoszka:1998:CPM

[KL98a] Piotr Kokoszka and Remigijus Leipus. Change-point in the mean of dependent observations. Statistics & Probability Letters, 40(4):385–393, November 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800145X.

Koo:1998:BSE

[KL98b] Ja-Yong Koo and Kee-Won Lee. B-spline estimation of regression functions with errors in variable. Statistics & Probability Letters, 40(1):57–66, September 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000984.

Kutoyants:1998:ELF

[KL98c] Yu. A. Kutoyants and F. Liese. Estimation of linear functionals of Poisson processes. Statistics & Probability Letters, 40(1):43–55, September 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298800011.

Kagan:1999:RBC

[KL99] Abram Kagan and Zinoviy Landsman. Relation between the covariance and Fisher information matrices. Statistics & Probability Letters, 42(1):7–13, March 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001783.

Klass:1993:RPI

[Kla93] Michael J. Klass. Ratio prophet inequalities for convex functions of partial sums. Statistics & Probability Letters, 17(3):205– 209, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390168I.

Klein:1994:HTC

[Kle94] André Klein. Hypothesis testing of common roots. Statistics & Probability Letters, 20(2):163–167, May 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900337.

Karunamuni:1990:ISU

[KM90] R. J. Karunamuni and K. L. Mehra. Improvements on strong uniform consistency of some known kernel estimates of a density and its derivatives. Statistics & Probability Letters, 9(2):133– 140, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290007R.

Kunchur:1993:ERS

[KM93] S. H. Kunchur and Surekha B. Munoli. Estimation of reliability in a shock model for a two component system. Statistics & Probability Letters, 17(1):35–38, May 3, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390191K.

Kasala:1997:ECR

[KM97] Subramanyam Kasala and Thomas Mathew. Exact confidence regions and tests in some linear functional relationships. Statistics & Probability Letters, 32(3):325–328, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000909.

Kundu:1998:USR

[KM98] Subrata Kundu and Adam T. Martinsek. Using a stopping rule to determine the size of the training sample in a classification problem. Statistics & Probability Letters, 37(1):19–27, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000947.

Kochar:1999:DOC

[KM99] Subhash Kochar and Chunsheng Ma. Dispersive ordering of convolutions of exponential random variables. Statistics & Probability Letters, 43(3):321–324, July 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800279X.

Kauermann:1998:EBC

[KMC98] Göran Kauermann, Marlene Müller, and Raymond J. Carroll. The efficiency of bias-corrected estimators for nonparametric kernel estimation based on local estimating equations. Statistics & Probability Letters, 37(1):41–47, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000989.

Koldobsky:1996:ICT

[KMS96] A. L. Koldobsky and S. J. Montgomery-Smith. Inequalities of correlation type for symmetric stable random vectors. Statistics & Probability Letters, 28(1):91–97, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000968.

Knight:1991:EMF

[Kni91] Keith Knight. On the empirical measure of the Fourier coefficients with infinite variance data. Statistics & Probability Letters, 12(2):109–117, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190053T.

Knight:1998:BSQ

[Kni98] Keith Knight. Bootstrapping sample quantiles in non-regular cases. Statistics & Probability Letters, 37(3):259–267, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001259.

Khmaladze:1997:ABN

[KNN97] E. Khmaladze, M. Nadareishvili, and A. Nikabadze. Asymptotic behaviour of a number of repeated records. Statistics & Probability Letters, 35(1):49–58, August 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296002155.

Kotz:1996:MRL

[KO96] Samuel Kotz and I. V. Ostrovskii. A mixture representation of the Linnik distribution. Statistics & Probability Letters, 26(1): 61–64, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294002525.

Kochar:1992:TTS

[Koc92] Subhash C. Kochar. On the ‘triples test’ for symmetry. Statistics & Probability Letters, 14(4):311–312, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290063B.

Kochar:1994:MPO

[Koc94] Subhash C. Kochar. Monotonicity properties of the ordered ranks in the two-sample problem. Statistics & Probability Letters, 20(1):63–67, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294902356.

Kochar:1996:DOO

[Koc96] Subhash C. Kochar. Dispersive ordering of order statistics. Statistics & Probability Letters, 27(3):271–274, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000836.

Kochar:1999:SOB

[Koc99] Subhash C. Kochar. On stochastic orderings between distributions and their sample spacings. Statistics & Probability Letters, 42(4):345–352, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002247.

Kohler:1999:NEP

[Koh99] Michael Kohler. Nonparametric estimation of piecewise smooth regression functions. Statistics & Probability Letters, 43(1):49– 55, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002454.

Kolassa:1995:EAR

[Kol95] John E. Kolassa. Edgeworth approximations for rank sum test statistics. Statistics & Probability Letters, 24(2):169–171, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500164H.

Konstantopoulos:1995:BTR

[Kon95] Takis Konstantopoulos. Ballot theorems revisited. Statistics & Probability Letters, 24(4):331–338, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400191A.

Kordecki:1997:RBM

[Kor97] Wojciech Kordecki. Reliability bounds for multistage structures with independent components. Statistics & Probability Letters, 34(1):43–51, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001642.

Korostelev:1999:MRC

[Kor99] Alexander Korostelev. On minimax rates of convergence in image models under sequential design. Statistics & Probability Letters, 43(4):369–375, July 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002752.

Kourouklis:1990:RBC

[Kou90] Stavros Kourouklis. A relation between the Chernoff index and the Pitman efficacy. Statistics & Probability Letters, 9(5):391– 395, May 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090030B.

Koul:1992:MEL

[Kou92a] Hira L. Koul. M-estimators in linear models with long range dependent errors. Statistics & Probability Letters, 14(2):153– 164, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290079K.

Koutras:1992:MDD

[Kou92b] Markos Koutras. Minimum distance discrimination rules and success rates for elliptical normal mixtures. Statistics & Probability Letters, 13(4):259–268, March 13, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290032Z.

Koukouvinos:1995:OWD

[Kou95] Christos Koukouvinos. Optimal weighing designs and some new weighing matrices. Statistics & Probability Letters, 25(1):37– 42, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400203K.

Koukouvinos:1996:FOD

[Kou96a] Christos Koukouvinos. Further orthogonal designs in linear models and sequences with zero autocorrelation. Statistics & Probability Letters, 30(3):215–219, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002227.

Koukouvinos:1996:LMO

[Kou96b] Christos Koukouvinos. Linear models and D-optimal designs for n ≡ 2 mod 4. Statistics & Probability Letters, 26(4):329– 332, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000283.

Koukouvinos:1996:OFD

[Kou96c] Christos Koukouvinos. Orthogonal 2^{k} and 3^{k} factorial designs constructed using sequences with zero autocorrelation. Statistics & Probability Letters, 28(1):59–63, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000828.

Koukouvinos:1996:ODL

[Kou96d] Christos Koukouvinos. Orthogonal designs in linear models and sequences with zero autocorrelation. Statistics & Probability Letters, 26(4):333–338, March 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000291.

Koukouvinos:1996:SNO

[Kou96e] Christos Koukouvinos. Some new orthogonal designs in linear regression models. Statistics & Probability Letters, 29(2):125–129, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001654.

Koukouvinos:1996:SRA

[Kou96f] Christos Koukouvinos. Some results for almost D-optimal experimental designs. Statistics & Probability Letters, 30(3):221–226, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295002235.

Kourouklis:1997:NPI

[Kou97] Stavros Kourouklis. A new property of the inverse Gaussian distribution with applications. Statistics & Probability Letters, 32(2):161–166, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000685.

Kozlowska:1991:SPR

[Koz91] Maria Kozłowska. Some properties of the row and column designs with adjusted orthogonality. Statistics & Probability Letters, 11(3):273–276, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190155K.

Kozlowska:1996:NBE

[Koz96] Maria Kozłowska. A note on the bounds of efficiency factor and E_{R} optimality of block designs. Statistics & Probability Letters, 30(3):199–203, October 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002197.

Kozarovitsky:1997:CBA

[Koz97] Eugene Kozarovitsky. Connection between asymptotics and stability of the positive measure. Statistics & Probability Letters, 34(1):59–65, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001666.

Kozubowski:1998:MRL

[Koz98] Tomasz J. Kozubowski. Mixture representation of Linnik distribution revisited. Statistics & Probability Letters, 38(2):157– 160, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001673.

Kerkyacharian:1992:DEB

[KP92] G. Kerkyacharian and D. Picard. Density estimation in Besov spaces. Statistics & Probability Letters, 13(1):15–24, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290231S.

Kerkyacharian:1993:DEK

[KP93a] Gérard Kerkyacharian and Dominique Picard. Density estimation by kernel and wavelets methods: Optimality of Besov spaces. Statistics & Probability Letters, 18(4):327–336, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390024D.

Klass:1993:LLB

[KP93b] Michael Klass and Jim Pitman. Limit laws for Brownian motion conditioned to reach a high level. Statistics & Probability Letters, 17(1):13–17, May 3, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390188O.

Krishnamoorthy:1994:UEE

[KP94a] K. Krishnamoorthy and Nabendu Pal. Unbiased equivariant estimation of a common normal mean vector with one observation from each population. Statistics & Probability Letters, 19(1):33– 38, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900655.

Kutoyants:1994:MNE

[KP94b] Y. Kutoyants and P. Pilibossian. On minimum L_{1}-norm estimate of the parameter of the Ornstein–Uhlenbeck process. Statistics & Probability Letters, 20(2):117–123, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900264.

Knopfmacher:1996:EAF

[KP96a] Arnold Knopfmacher and Helmut Prodinger. Explicit and asymptotic formulae for the expected values of the order statistics of the Cantor distribution. Statistics & Probability Letters, 27(2):189–194, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000631.

Kozubowski:1996:MTB

[KP96b] Tomasz J. Kozubowski and Anna K. Panorska. On moments and tail behavior of v-stable random variables. Statistics & Probability Letters, 29(4):307–315, September 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001875.

Kim:1994:ABB

[KPM94] W. C. Kim, B. U. Park, and J. S. Marron. Asymptotically best bandwidth selectors in kernel density estimation. Statistics & Probability Letters, 19(2):119–127, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901430.

Koutras:1996:BCR

[KPP96] M. V. Koutras, S. G. Papastavridis, and K. I. Petakos. Bounds for coherent reliability structures. Statistics & Probability Letters, 26(3):285–292, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000224.

Kruglikov:1997:ARD

[KPTY97] Ilya L. Kruglikov, Nikolaj I. Pilipenko, Alexander D. Tsodikov, and Andrej Yu. Yakovlev. Assessing risk with doubly censored data: an application to the analysis of radiation-induced thyropathy. Statistics & Probability Letters, 32(3):223–230, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000776.

Kaufmann:1998:AHE

[KR98] E. Kaufmann and R.-D. Reiss. Approximation of the Hill estimator process. Statistics & Probability Letters, 39(4):347–354, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000790.

Krajka:1998:CMD

[Kra98] A. Krajka. On the class of g-monotone dependence functions. Statistics & Probability Letters, 39(3):213–227, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000479.

Krengel:1996:PSR

[Kre96] U. Krengel. On pairs of sums of random variables with pairwise equal distributions. Statistics & Probability Letters, 27(3):217– 219, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000674.

Krishnaiah:1990:GCT

[Kri90] Y. S. Rama Krishnaiah. On the Glivenko–Cantelli theorem for generalized empirical processes based on strong mixing sequences. Statistics & Probability Letters, 10(5):439–447, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900264.

Krishnaiah:1993:TPK

[Kri93] Y. S. Rama Krishnaiah. On tail probabilities of Kolmogorov– Smirnov statistic based on strong mixing processes. Statistics & Probability Letters, 16(5):369–377, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390071P.

Krone:1995:CST

[Kro95] Stephen M. Krone. Conditioned superprocesses and their weighted occupation times. Statistics & Probability Letters, 22(1):59–69, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400050I.

Krone:1997:RCA

[Kro97] Stephen M. Krone. Representations for continuous additive functionals of super-Brownian and super-stable processes. Statistics & Probability Letters, 34(3):211–223, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297894850.

Krummenauer:1998:RMD

[Kru98] Frank Krummenauer. Representation of multivariate discrete distributions by probability generating functions. Statistics & Probability Letters, 39(4):327–331, August 21, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000728.

Krummenauer:1999:CGT

[Kru99a] Frank Krummenauer. Conservativeness of global tests for bivariate binomial and Poisson data. Statistics & Probability Letters, 41(4):407–412, February 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002260.

Krummenauer:1999:NRT

[Kru99b] Frank Krummenauer. A note on Rüger’s test for discrete data. Statistics & Probability Letters, 41(4):347–351, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001497.

Keener:1998:TPA

[KRW98] Robert W. Keener, John Robinson, and Neville C. Weber. Tail probability approximations for U-statistics. Statistics & Probability Letters, 37(1):59–65, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001004.

Krysicki:1999:SNP

[Kry99] Włodzimierz Krysicki. On some new properties of the beta distribution. Statistics & Probability Letters, 42(2):131–137, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001977.

Krzanowski:1994:QLD

[Krz94] W. J. Krzanowski. Quadratic location discriminant functions for mixed categorical and continuous data. Statistics & Probability Letters, 19(2):91–95, January 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901384.

Kopp-Schneider:1992:BDP

[KS92] Annette Kopp-Schneider. Birth-death processes with piecewise constant rates. Statistics & Probability Letters, 13(2):121–127, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290086K.

Kaigh:1993:SQE

[KS93] W. D. Kaigh and Maria Alejandra Sorto. Subsampling quantile estimator majorization inequalities. Statistics & Probability Letters, 18(5):373–379, December 2, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390030M.

Kale:1995:AEN

[KS95a] B. K. Kale and G. Sebastian. Approximate estimation of non-identifiable parameters in a convolution. Statistics & Probability Letters, 25(4):373–378, December 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002440.

Kulkarni:1995:TES

[KS95b] Pandurang M. Kulkarni and Arvind K. Shah. Testing the equality of several binomial proportions to a prespecified standard. Statistics & Probability Letters, 25(3):213–219, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400224V.

Kumar:1996:NEQ

[KS96] Somesh Kumar and Divakar Sharma. A note on estimating quantiles of exponential populations. Statistics & Probability Letters, 26(2):115–118, February 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002606.

Kagan:1998:SLM

[KS98a] Abram Kagan and Lawrence A. Shepp. Some linear models are necessarily parametric. Statistics & Probability Letters, 37(1):77– 80, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001028.

Kagan:1998:WVO

[KS98b] Abram Kagan and Lawrence A. Shepp. Why the variance? Statistics & Probability Letters, 38(4):329–333, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000418.

Kagan:1999:THS

[KS99a] Abram Kagan and Lawrence A. Shepp. Tail hypotheses in the signal plus noise model. Statistics & Probability Letters, 43(3):317– 319, July 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002739.

Koul:1999:IAR

[KS99b] Hira L. Koul and Anton Schick. Inference about the ratio of scale parameters in a two-sample setting with current status data. Statistics & Probability Letters, 45(4):359–369, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000784.

Kutoyants:1999:OCO

[KS99c] Yury A. Kutoyants and Vladimir Spokoiny. Optimal choice of observation window for Poisson observations. Statistics & Probability Letters, 44(3):291–298, September 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000206.

Knautz:1993:CBX

[KT93] Henning Knautz and Götz Trenkler. On the correlation between X̄ and S^{2}. Statistics & Probability Letters, 16(3):235–237, February 19, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390148C.

Kim:1998:CLD

[KT98] Gie-Whan Kim and Donald R. Truax. Conditional large deviations for density case. Statistics & Probability Letters, 38(2):137– 144, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001648.

Kuk:1999:UAM

[Kuk99] Anthony Y. C. Kuk. The use of approximating models in Monte Carlo maximum likelihood estimation. Statistics & Probability Letters, 45(4):325–333, December 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000747.

Kulkarni:1992:EPT

[Kul92] P. M. Kulkarni. Estimation of parameters of a two-dimensional spatial autoregressive model with regression. Statistics & Probability Letters, 15(2):157–162, September 28, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290129S.

Kulasekera:1995:BEN

[Kul95] K. B. Kulasekera. A bound on the L_{∞}–error of a nonparametric density estimator with censored data. Statistics & Probability Letters, 23(3):233–238, May 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400118R.

Kulperger:1997:PFF

[Kul97] R. J. Kulperger. On a property of the finite Fourier partial sums process. Statistics & Probability Letters, 35(2):101–107, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000023.

Kundu:1993:ATL

[Kun93] Debasis Kundu. Asymptotic theory of least squares estimator of a particular nonlinear regression model. Statistics & Probability Letters, 18(1):13–17, August 30, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390093X.

Kushary:1998:NUA

[Kus98] Debashis Kushary. A note on universal admissibility of scale parameter estimators. Statistics & Probability Letters, 38(1):59– 67, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001545.

Kutoyants:1994:NED

[Kut94] Yu. A. Kutoyants. Nonconsistent estimation by diffusion type observations. Statistics & Probability Letters, 20(1):1–7, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902283.

Kutoyants:1997:UDE

[Kut97a] Yu. A. Kutoyants. On unbiased density estimation for ergodic diffusion. Statistics & Probability Letters, 34(2):133–140, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001745.

Kutoyants:1997:SPN

[Kut97b] Yu. A. Kutoyants. Some problems of nonparametric estimation by observations of ergodic diffusion process. Statistics & Probability Letters, 32(3):311–320, March 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000880.

Kyriakoussis:1996:ANC

[KV96] A. G. Kyriakoussis and M. G. Vamvakari. Asymptotic normality of a class of bivariate-multivariate discrete power series distributions. Statistics & Probability Letters, 27(3):207–216, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000666.

Kim:1999:TMS

[KV99] Yongdai Kim and Joseph S. Verducci. Too much sampling kills the UMP test. Statistics & Probability Letters, 41(2):101–105, January 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000959.

Kochar:1992:ROP

[KW92] Subhash C. Kochar and George Woodworth. Rank order probabilities for the dispersion problem. Statistics & Probability Letters, 13(3):203–208, February 19, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290049B.

Kuhlmann:1993:PPR

[KW93] Dietrich W. Kuhlmann and F. T. Wright. A property of projection residuals with applications to concave regression. Statistics & Probability Letters, 16(4):313–319, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901367.

Kagan:1996:NCN

[KW96] Abram Kagan and Jacek Wesołowski. Normality via conditional normality of linear forms. Statistics & Probability Letters, 29(3):229–232, September 2, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001778.

Kang:1997:NWI

[KW97] Ju-Sung Kang and In-Suk Wee. A note on the weak invariance principle for local times. Statistics & Probability Letters, 32(2):147–159, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000673.

Kulasekera:1998:BSP

[KW98] K. B. Kulasekera and J. Wang. Bandwidth selection for power optimality in a test of equality of regression curves. Statistics & Probability Letters, 37(3):287–293, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297841557.

Kyriakidis:1994:SPS

[Kyr94] E. G. Kyriakidis. Stationary probabilities for a simple immigration-birth-death process under the influence of total catastrophes. Statistics & Probability Letters, 20(3):239–240, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900485.

Kulkarni:1991:COD

[KZ91] Sanjeev R. Kulkarni and Ofer Zeitouni. Can one decide the type of the mean from the empirical measure? Statistics & Probability Letters, 12(4):323–327, October 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291901006.

Klonecki:1996:IES

[KZ96] Witold Klonecki and Stefan Zontek. Improved estimators for simultaneous estimation of variance components. Statistics & Probability Letters, 29(1):33–43, August 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001530.

Lacaze:1995:SIS

[Lac95] B. Lacaze. Stationarity of independent sequences. Statistics & Probability Letters, 23(1):9–11, April 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400088P.

Lacaze:1996:NAS

[Lac96] Pr. B. Lacaze. A note about stationary process random sampling. Statistics & Probability Letters, 31(2):133–137, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000247.

Lahiri:1991:SOO

[Lah91] Soumendra Nath Lahiri. Second order optimality of stationary bootstrap. Statistics & Probability Letters, 11(4):335–341, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190045S.

Lahiri:1992:BGK

[Lah92] S. N. Lahiri. On the Bahadur–Ghosh–Kiefer representation of sample quantiles. Statistics & Probability Letters, 15(2):163–168, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290130W.

Lahiri:1993:MBB

[Lah93] S. N. Lahiri. On the moving block bootstrap under long range dependence. Statistics & Probability Letters, 18(5):405–413, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390035H.

Lahiri:1998:BWE

[Lah98] Soumendra Nath Lahiri. Bootstrapping weighted empirical processes that do not converge weakly. Statistics & Probability Letters, 37(3):295–302, March 16, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297841569.

Lai:1995:CBD

[Lai95] C. D. Lai. Construction of bivariate distributions by a generalised trivariate reduction technique. Statistics & Probability Letters, 25(3):265–270, November 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002306.

Lange:1996:ISM

[Lan96] Kenneth Lange. Illustration of some moment identities for order statistics. Statistics & Probability Letters, 27(1):91–98, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000496.

Lanzinger:1998:BKT

[Lan98] Hartmut Lanzinger. A Baum–Katz theorem for random variables under exponential moment conditions. Statistics & Probability Letters, 39(2):89–95, August 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000091.

Lasserre:1997:IPM

[Las97] Jean B. Lasserre. Invariant probabilities for Markov chains on a metric space. Statistics & Probability Letters, 34(3):259–265, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001897.

Latala:1998:ASB

[Lat98] Rafal Latala. On the almost sure boundedness of norms of some empirical operators. Statistics & Probability Letters, 38(2):177– 182, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000029.

Lau:1992:REB

[Lau92] Tai Shing Lau. The reliability of exchangeable binary systems. Statistics & Probability Letters, 13(2):153–158, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290090R.

Li:1991:BME

[LB91a] Tze Fen Li and Dinesh S. Bhoj. Bayes minimax estimators of a multivariate normal mean. Statistics & Probability Letters, 11(5): 373–377, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190184S.

Lu:1991:IPB

[LB91b] Jye-Chyi Lu and Gouri K. Bhattacharyya. Inference procedures for bivariate exponential model of Gumbel. Statistics & Probability Letters, 12(1):37–50, July 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190162K.

Liang:1998:CSS

[LB98] Jia-Juan Liang and Peter M. Bentler. Characterizations of some subclasses of spherical distributions. Statistics & Probability Letters, 40(2):155–164, September 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000844.

Luo:1992:CCM

[LC92] Xiaolong Luo and Haiyan Cai. Coexistence in a competition model. Statistics & Probability Letters, 15(3):241–243, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290196C.

Lopez-Diaz:1997:CDF

[LDG97] Miguel López-Diaz and Maria Angeles Gil. Constructive definitions of fuzzy random variables. Statistics & Probability Letters, 36(2):135–143, December 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000564.

LeBreton:1992:ACS

[Le 92] Alain Le Breton. Adaptive control in the scalar linearquadratic model in continuous time. Statistics & Probability Letters, 13(3):169–177, February 19, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292900446.

LeBreton:1998:FPE

[Le 98] Alain Le Breton. Filtering and parameter estimation in a simple linear system driven by a fractional Brownian motion. Statistics & Probability Letters, 38(3):263–274, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000297.

Leadbetter:1991:BPT

[Lea91] M. R. Leadbetter. On a basis for ‘peaks over threshold’ modeling. Statistics & Probability Letters, 12(4):357–362, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291901073.

Leblanc:1996:WLD

[Leb96] Frédérique Leblanc. Wavelet linear density estimator for a discrete-time stochastic process: L_{p}-losses. Statistics & Probability Letters, 27(1):71–84, March 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000461.

Lecoutre:1990:UCC

[Lec90] Jean-Pierre Lecoutre. Uniform consistency of a class of regression function estimators for Banach-space valued random variable. Statistics & Probability Letters, 10(2):145–149, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900105.

Ledoux:1995:WLD

[Led95] James Ledoux. On weak lumpability of denumerable Markov chains. Statistics & Probability Letters, 25(4):329–339, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002385.

Lee:1990:CPM

[Lee90] Carl M.-S. Lee. On the characterization of Pitman measure of nearness. Statistics & Probability Letters, 9(1):41–46, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090093M.

Lee:1993:ECT

[Lee93] Tze-San Lee. Estimating coefficients of two-phase linear regression model with autocorrelated errors. Statistics & Probability Letters, 18(2):113–120, September 22, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390178L.

Lee:1995:ENC

[Lee95] Chung-Bow Lee. Estimating the number of change points in a sequence of independent normal random variables. Statistics & Probability Letters, 25(3):241–248, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400227Y.

Lee:1996:NMC

[Lee96a] Chung-Bow Lee. Nonparametric multiple change-point estimators. Statistics & Probability Letters, 27(4):295–304, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000895.

Lee:1996:AMP

[Lee96b] Sangyeol Lee. The asymptotic maximin property of chisquared type tests based on the empirical process. Statistics & Probability Letters, 29(4):285–292, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001840.

Lee:1997:IBB

[Lee97a] Min-Young Lee. Improved bivariate Bonferroni-type inequalities. Statistics & Probability Letters, 31(4):359–364, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000508.

Lee:1997:LTS

[Lee97b] Oesook Lee. Limit theorems for some doubly stochastic processes. Statistics & Probability Letters, 32(2):215–221, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000764.

Lee:1997:NRE

[Lee97c] Sangyeol Lee. A note on the residual empirical process in autoregressive models. Statistics & Probability Letters, 32(4):405– 411, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001009.

Lee:1997:RCL

[Lee97d] Sangyeol Lee. Random Central Limit Theorem for the linear process generated by a strong mixing process. Statistics & Probability Letters, 35(2):189–196, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000138.

Lee:1998:ACT

[Lee98a] Chanho Lee. Asymptotics of a class of p th-order nonlinear autoregressive processes. Statistics & Probability Letters, 40 (2):171–177, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000996.

Lee:1998:ICP

[Lee98b] Chinsan Lee. An integral characterization problem in an agedependent cancer model. Statistics & Probability Letters, 40 (2):121–126, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000704.

Lee:1999:UMS

[Lee99] Chinsan Lee. An urn model in the simulation of interval censored failure time data. Statistics & Probability Letters, 45(2):131–139, November 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000516.

Lefebvre:1996:SSF

[Lef96] Mario Lefebvre. Similarity solutions of first-passage problems for two-dimensional Wiener processes. Statistics & Probability Letters, 29(2):185–189, August 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001727.

Lefebvre:1997:FHP

[Lef97] Mario Lefebvre. First hitting place distributions for the Ornstein–Uhlenbeck process. Statistics & Probability Letters, 34(3):309–312, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001952.

Leonard:1993:EBA

[Leo93] Kevin J. Leonard. Empirical Bayes analysis of the commercial loan evaluation process. Statistics & Probability Letters, 18(4):289–296, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390017D.

Leonard:1996:ESD

[Leo96] Tom Leonard. On exchangeable sampling distributions for uncontrolled data. Statistics & Probability Letters, 26(1):1–6, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002452.

Leonov:1999:REP

[Leo99] Sergei L. Leonov. Remarks on extremal problems in nonparametric curve estimation. Statistics & Probability Letters, 43(2):169– 178, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002569.

Leung:1996:LLD

[Leu96] Chi-Ying Leung. The location linear discriminant for classifying observations with unequal variances. Statistics & Probability Letters, 31(1):23–29, December 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000090.

Li:1995:UPP

[LF95a] Luning Li and Bernard Flury. Uniqueness of principal points for univariate distributions. Statistics & Probability Letters, 25(4):323–327, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294002373.

Li:1995:ESM

[LF95b] Run-Ze Li and Kai-Tai Fang. Estimation of scale matrix of elliptically contoured matrix distributions. Statistics & Probability Letters, 24(4):289–297, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400186C.

Leao:1999:CRS

[LFR99] D. Leão, Jr., M. D. Fragoso, and P. R. C. Ruffino. Characterizations of Radon spaces. Statistics & Probability Letters, 42(4):409– 413, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002375.

Lin:1993:HSK

[LG93] Dennis K. J. Lin and Irwin Guttman. Handling spuriosity in the Kalman filter. Statistics & Probability Letters, 16(4):259–268, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901297.

Liu:1999:EHR

[LGC99a] Jingjun Liu, Shixin Gan, and Pingyan Chen. Erratum: “The Hájek–Rényi inequality for the NA random variables and its application” [Statist. Probab. Lett. 43 (1999) 99– 105]. Statistics & Probability Letters, 44(2):210, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299001169. See [LGC99b].

Liu:1999:HRI

[LGC99b] Jingjun Liu, Shixin Gan, and Pingyan Chen. The Hájeck– Rènyi inequality for the NA random variables and its application. Statistics & Probability Letters, 43(1):99–105, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529800251X. See erratum [LGC99a].

Lee:1996:DLF

[LH96] Jack C. Lee and Ling Hu. On the distribution of linear functions of independent f and u variates. Statistics & Probability Letters, 26(4):339–346, March 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000305.

Lee:1998:ETO

[LH98a] GeungHee Lee and Jeffrey D. Hart. An L_{2} error test with order selection and thresholding. Statistics & Probability Letters, 39(1): 61–72, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800039X.

Li:1998:NMM

[LH98b] Deli Li and Mei Ling Huang. A note on moments of the maximum of Cesàro summation. Statistics & Probability Letters, 38(1):73– 81, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001569.

Li:1993:CFS

[Li93] Run-Ze Li. The characteristic functions of spherical matrix distributions. Statistics & Probability Letters, 17(4):273–279, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390202T.

Li:1994:MLI

[Li94a] Gang Li. Multidimensional Lévy inequalities and their applications. Statistics & Probability Letters, 20(4):327–335, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900213.

Li:1994:CCO

[Li94b] Luning Li. A counterexample to a conjecture on order statistics. Statistics & Probability Letters, 19(2):129–130, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901449.

Li:1994:CFS

[Li94c] Run-Ze Li. The characteristic functions of spherical matrix distributions. Statistics & Probability Letters, 20(3):249, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294000301.

Li:1995:NLR

[Li95] Gang Li. On nonparametric likelihood ratio estimation of survival probabilities for censored data. Statistics & Probability Letters, 25(2):95–104, November 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400210Y.

Li:1996:BTF

[Li96] Ta-Hsin Li. Bartlett-type formulas for complex multivariate time series of mixed spectra. Statistics & Probability Letters, 28(3): 259–268, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001336.

Li:1997:SRP

[Li97] Linxiong Li. Superposition of renewal processes in a random environment. Statistics & Probability Letters, 35(3):297–305, October 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000266.

Li:1999:NNR

[Li99] Yun Li. A note on the number of records near the maximum. Statistics & Probability Letters, 43(2):153–158, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002533.

Liang:1995:SOA

[Lia95] Hua Liang. Second-order asymptotic efficiency of PMLE in generalized linear models. Statistics & Probability Letters, 24(3):273–279, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400184A.

Liang:1999:MEB

[Lia99] TaChen Liang. Monotone empirical Bayes tests for a discrete normal distribution. Statistics & Probability Letters, 44(3):241–249, September 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000152.

Liebscher:1999:ANN

[Lie99] Eckhard Liebscher. Asymptotic normality of nonparametric estimators under α-mixing condition. Statistics & Probability Letters, 43(3):243–250, July 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002648.

Limnios:1993:TSM

[Lim93] N. Limnios. A transient solution method for semi-Markov systems. Statistics & Probability Letters, 17(3):211–220, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390169J.

Ling:1990:GDO

[Lin90] K. D. Ling. On geometric distributions of order (k_{1}, …, k_{m}). Statistics & Probability Letters, 9(2):163–171, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290012T.

Linssen:1991:TSB

[Lin91] H. N. Linssen. A table for solving the Behrens–Fisher problem. Statistics & Probability Letters, 11(4):359–363, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190049W.

Lindqvist:1992:MIR

[Lin92a] Bo Henry Lindqvist. Moment inequalities for the reliability function. Statistics & Probability Letters, 13(2):85–88, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290081F.

Ling:1992:GSL

[Lin92b] K. D. Ling. A generalization of the sooner and later waiting time problems for Bernoulli trials: Frequency quota. Statistics & Probability Letters, 14(5):401–405, July 11, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290101A.

Lin:1993:CRA

[Lin93]  Gwo Dong Lin.  On convergence rates of averages of

weakly dependent random variables. Statistics & Probability Letters, 16(2):159–162, January 27, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390161B.

Lin:1994:PGT

[Lin94] Gwo Dong Lin. On a probabilistic generalization of Taylor’s theorem. Statistics & Probability Letters, 19(3):239–243, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901104.

Lin:1995:SR

[Lin95] S. J. Lin. On L^{P} stochastic representations. Statistics & Probability Letters, 24(3):225–227, August 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001758.

Lin:1997:MPO

[Lin97] Gwo Dong Lin. On the moment problems. Statistics & Probability Letters, 35(1):85–90, August 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296002209. See erratum [Lin00].

Lindoff:1998:FIM

[Lin98] B. Lindoff. First inverse moment of a generalized quadratic form. Statistics & Probability Letters, 40(4):363–370, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001370.

Lin:2000:EMP

[Lin00] Gwo Dong Lin. Erratum: “On the moment problems” [Statist. Probab. Lett. 35 (1997) 85–90]. Statistics & Probability Letters, 50(2):205, November 1, 2000. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771520000153X. See [Lin97].

Liu:1994:SFS

[Liu94a] Jun S. Liu. Siegel’s formula via Stein’s identities. Statistics & Probability Letters, 21(3):247–251, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529490121X.

Liu:1994:SLT

[Liu94b] Wen Liu. A strong limit theorem under no assumption of independence, stationarity or various dependences. Statistics & Probability Letters, 21(2):157–161, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902240.

Liu:1996:EPT

[Liu96] Aiyi Liu. Estimation of the parameters in two linear models with only some of the parameter vectors identical. Statistics & Probability Letters, 29(4):369–375, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500193X.

Liu:1997:CDE

[Liu97] Wei Liu. Control of directional errors with step-up multiple tests. Statistics & Probability Letters, 31(3):239–242, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000363.

Liu:1999:CCS

[Liu99a] Chuanhai Liu. Compatibility conditions for a set of conditional Gaussian distributions. Statistics & Probability Letters, 42(2):127–130, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001825.

Liu:1999:OSC

[Liu99b] Wei Liu. A one-sided confidence set hidden under a twosided sequential test of a normal mean. Statistics & Probability Letters, 41(4):435–438, February 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800234X.

Lewis:1993:HLD

[LL93] Thomas M. Lewis and Wenbo V. Li. How long does it take to see a flat Brownian path on the average? Statistics & Probability Letters, 16(5):347–354, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390068T.

Liu:1995:CSL

[LL95] Wen Liu and Guoxin Liu. A class of strong laws for functionals of countable nonhomogeneous Markov chains. Statistics & Probability Letters, 22(2):87–96, February 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400053B.

Lin:1997:PCC

[LL97] Yong Lin and Bruce G. Lindsay. Projections on cones, chibar squared distributions, and Weyl’s formula. Statistics & Probability Letters, 32(4):367–376, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600096X.

Lloyd:1991:AEF

[Llo91] Chris J. Lloyd. Asymptotic expansions of the Fisher information in a sample mean. Statistics & Probability Letters, 11(2):133– 137, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190130J.

Leon:1992:CEM

[LM92] Carlos Antonio León and Jean-Claude Massé. A counterexample on the existence of the L_{1}-median. Statistics & Probability Letters, 13(2):117–120, January 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290085J.

Lent:1996:ACA

[LM96] Janice Lent and Hosam M. Mahmoud. Average-case analysis of multiple Quickselect: An algorithm for finding order statistics. Statistics & Probability Letters, 28(4):299–310, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001395.

Letac:1998:FMD

[LM98] Gérard Letac and Hélène Massam. A formula on multivariate Dirichlet distributions. Statistics & Probability Letters, 38(3):247–253, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000248.

Lee:1997:LEA

[LMF97] Travis Lee, Max Minzner, and Evan Fisher. The length of an excursion above a linear boundary by a random walk. Statistics & Probability Letters, 34(4):397–402, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002076.

Linton:1994:MBR

[LN94] Oliver Linton and Jens Perch Nielsen. A multiplicative bias reduction method for nonparametric regression. Statistics & Probability Letters, 19(3):181–187, February 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901023.

Liu:1995:ODS

[LN95] Shuangzhe Liu and Heinz Neudecker. A V -optimal design for Scheffé’s polynomial model. Statistics & Probability Letters, 23(3):253–258, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400122O.

Lo:1991:CDP

[Lo91] Albert Y. Lo. A characterization of the Dirichlet process. Statistics & Probability Letters, 12(3):185–187, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900753.

Loh:1997:EIS

[Loh97] Wei-Liem Loh. Estimating the integral of a squared regression function with Latin hypercube sampling. Statistics & Probability Letters, 31(4):339–349, February 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600048X.

Lorenzen:1992:RPC

[Lor92] Gunter Lorenzen. A reformulation of Pearson’s chi-square statistic and some extentions. Statistics & Probability Letters, 14(4):327–331, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290066E.

Lorenzen:1995:NFG

[Lor95] Gunter Lorenzen. A new family of goodness-of-fit statistics for discrete multivariate data. Statistics & Probability Letters, 25(4):301–307, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294002348.

Lou:1997:AMF

[Lou97] W. Y. Wendy Lou. An application of the method of finite Markov chain imbedding to runs tests. Statistics & Probability Letters, 31(3):155–161, January 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000272.

Louhichi:1999:RIL

[Lou99] Sana Louhichi. Rosenthal’s inequality for LPQD sequences. Statistics & Probability Letters, 42(2):139–144, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001990.

Lovison:1998:ARA

[Lov98]  G. Lovison.  An alternative representation of Altham’s

multiplicative-binomial distribution. Statistics & Probability Letters, 36(4):415–420, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000886.

Lowe:1996:ILD

[Löw96] Matthias Löwe. Iterated large deviations. Statistics & Probability Letters, 26(3):219–223, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000135.

Lowe:1997:IMN

[Löw97] Matthias Löwe. On the invariant measure of non-reversible simulated annealing. Statistics & Probability Letters, 36(2):189–193, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000631.

Lacey:1990:NAS

[LP90] Michael T. Lacey and Walter Philipp. A note on the almost sure Central Limit Theorem. Statistics & Probability Letters, 9(3):201–205, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090056D.

Lio:1991:NAO

[LP91a] Y. L. Lio and W. J. Padgett. A note on the asymptotically optimal bandwidth for Nadaraya’s quantile estimator. Statistics & Probability Letters, 11(3):243–249, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190150P.

Loukas:1991:RCI

[LP91b] Sotirios Loukas and Takis Papaioannou. Rank correlation inequalities with ties and missing data. Statistics & Probability Letters, 11(1):53–56, January 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190177S.

Lemdani:1997:LRT

[LP97] Mohamed Lemdani and Odile Pons. Likelihood ratio tests for genetic linkage. Statistics & Probability Letters, 33(1):15–22, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001058.

Lee:1990:FML

[LPB90] Sik-Yum Lee, Wai-Yin Poon, and P. M. Bentler. Full maximum likelihood analysis of structural equation models with polytomous variables. Statistics & Probability Letters, 9(1):91–97, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090100L.

Lu:1996:TPP

[LR96] I-Li Lu and Donald Richards. Total positivity properties of the bivariate diagonal natural exponential families. Statistics & Probability Letters, 26(2):119–124, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002614.

Li:1992:CCM

[LRW92] Deli Li, M. Bhaskara Rao, and Xiangchen Wang. Complete convergence of moving average processes. Statistics & Probability Letters, 14(2):111–114, May 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290073E.

Lepeltier:1997:BSD

[LS97] J. P. Lepeltier and J. San Martin. Backward stochastic differential equations with continuous coefficient. Statistics & Probability Letters, 32(4):425–430, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001034.

Lopez:1998:SCC

[LS98] F. Javier López and Gerardo Sanz. Stochastic comparisons and couplings for interacting particle systems. Statistics & Probability Letters, 40(1):93–102, September 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001126.

Liang:1999:CCW

[LS99] Han-Ying Liang and Chun Su. Complete convergence for weighted sums of NA sequences. Statistics & Probability Letters, 45(1):85–95, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000462.

Leemis:1990:VGA

[LSR90] Lawrence M. Leemis, Li-Hsing Shih, and Kurt Reynertson. Variate generation for accelerated life and proportional hazards models with time dependent covariates. Statistics & Probability Letters, 10(4):335–339, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900529.

Le:1998:NEM

[LSZ98] N. Le, L. Sun, and J. V. Zidek. A note on the existence of maximum likelihood estimates for Gaussian-inverted Wishart models. Statistics & Probability Letters, 40(2):133–137, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000753.

Lad:1992:MCD

[LT92] F. R. Lad and W. F. C. Taylor. The moments of the Cantor distribution. Statistics & Probability Letters, 13(4):307–310, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900398.

Li:1996:LIL

[LT96] Deli Li and R. J. Tomkins. Laws of the iterated logarithm for weighted sums of independent random variables. Statistics & Probability Letters, 27(3):247–254, April 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000720.

Liu:1999:MLT

[LT99] P. Y. Liu and Wei Yann Tsai. A modified logrank test for censored survival data under order restrictions. Statistics & Probability Letters, 41(1):57–63, January 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001217.

Lu:1996:NGE

[Lu96] Zudi Lu. A note on geometric ergodicity of autoregressive conditional heteroscedasticity (ARCH) model. Statistics & Probability Letters, 30(4):305–311, November 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715295002332.

Lu:1997:NPL

[Lu97] Jye-Chyi Lu. A new plan for life-testing two-component parallel systems. Statistics & Probability Letters, 34(1):19–32, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001617.

Luceno:1996:GED

[Luc96] Alberto Luceño. A generalized Erlang distribution showing overdispersion. Statistics & Probability Letters, 28(4):375–386, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001484.

Lugosi:1995:IUB

[Lug95] Gábor Lugosi. Improved upper bounds for probabilities of uniform deviations. Statistics & Probability Letters, 25(1):71–77, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400207O.

Lee:1999:JCS

[LV99]  Hak-Myung Lee and Marlos Viana.  The joint covari-

ance structure of ordered symmetrically dependent observations and their concomitants of order statistics. Statistics & Probability Letters, 43(4):411–414, July 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002818.

Liu:1990:RHF

[LVC90] P. Y. Liu, J. Voelkel, and J. Crowley. On the ratio of hazard functions in the presence of a nuisance covariate. Statistics & Probability Letters, 10(2):135–140, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090008U.

Liu:1993:SCT

[LVCW93] P. Y. Liu, Joseph O. Voelkel, John Crowley, and Michael Wolf. Sufficient conditions for treatment responders to have longer survival than non-responders. Statistics & Probability Letters, 18(3):205–208, October 15, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293902177.

Leskow:1992:EBE

[LW92] Jacek Leskow and Aleksander Weron. Ergodic behavior and estimation for periodically correlated processes. Statistics & Probability Letters, 15(4):299–304, November 13, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292901663.

Li:1997:WDH

[LX97] Yuan Li and Zhongjie Xie. The wavelet detection of hidden periodicities in time series. Statistics & Probability Letters, 35(1):9– 23, August 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296002118.

Liu:1994:ARQ

[LY94] Z. J. Liu and Y. Q. Yin. Asymptotic representations for qualites of pooled samples. Statistics & Probability Letters, 19(4):299– 305, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901805.

Lloyd:1999:KER

[LY99] Chris J. Lloyd and Zhou Yong. Kernel estimators of the ROC curve are better than empirical. Statistics & Probability Letters, 44(3):221–228, September 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000127.

Li:1991:MMA

[LZ91] Bing Li and Ruben H. Zamar. Min-max asymptotic variance of M-estimates of location when scale is unknown. Statistics & Probability Letters, 11(2):139–145, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190131A.

Li:1994:SEO

[LZ94] Haijun Li and Haolong Zhu. Stochastic equivalence of ordered random variables with applications in reliability theory. Statistics & Probability Letters, 20(5):383–393, August 9, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901309.

Lee:1997:AIG

[LZ97] Andy H. Lee and Yuejen Zhao. Assessing influence on the goodness-of-link in generalized linear models. Statistics & Probability Letters, 31(4):351–358, February 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000491.

Ma:1999:PDL

[Ma99] Chunsheng Ma. On the posterior distribution of a location parameter from a strongly unimodal distribution. Statistics & Probability Letters, 42(1):33–37, March 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001837.

Macci:1996:LDP

[Mac96] Claudio Macci. On the Lebesgue decomposition of the posterior distribution with respect to the prior in regular Bayesian experiments. Statistics & Probability Letters, 26(2):147–152, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000046.

Macci:1998:MDS

[Mac98] Claudio Macci. The maximal dominated subsets of a statistical experiment. Statistics & Probability Letters, 39(3):185–193, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000145.

Madi:1993:AEV

[Mad93] Mohamed T. Madi. Alternative estimators for the variance of several normal populations. Statistics & Probability Letters, 17(4):321–328, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390210A.

Magiera:1990:MSE

[Mag90] Ryszard Magiera. Minimax sequential estimation plans for exponential-type processes. Statistics & Probability Letters, 9(2):179–185, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290014V.

Majumdar:1992:TSD

[Maj92] Suman Majumdar. On topological support of Dirichlet prior. Statistics & Probability Letters, 15(5):385–388, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290171Z.

Majumdar:1999:MCD

[Maj99] Suman Majumdar. A metric for convergence in distribution. Statistics & Probability Letters, 42(3):239–247, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001904.

Malov:1996:CII

[Mal96] S. V. Malov. On characterizations of independent identically distributed random variables via order structures. Statistics & Probability Letters, 26(2):97–104, February 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002576.

Mammen:1990:SNO

[Mam90] Enno Mammen. A short note on optimal bandwidth selection for kernel estimators. Statistics & Probability Letters, 9(1):23– 25, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090090T.

Mandeli:1999:CSR

[Man99] John P. Mandeli. Construction of systematic row-column designs with treatments unconfounded with the linear row × columns and quadratic row × columns interactions. Statistics & Probability Letters, 42(3):229–237, April 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001758.

Maritz:1991:ECM

[Mar91] J. S. Maritz. Estimating the covariance matrix of bivariate medians. Statistics & Probability Letters, 12(4):305–309, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190096A.

Marchand:1994:EMP

[Mar94] Eric Marchand. On the estimation of the mean of a N_{p}(μ, Σ) population with μ′Σ^{−1}μ known. Statistics & Probability Letters, 21(1):69–75, September 1, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900590.

Martikainen:1995:SLL

[Mar95] Alexander Martikainen. On the Strong Law of Large Numbers for sums of pairwise independent random variables. Statistics & Probability Letters, 25(1):21–26, October 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400201I.

Markiewicz:1996:CGR

[Mar96] Augustyn Markiewicz. Characterization of general ridge estimators. Statistics & Probability Letters, 27(2):145–148, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000569.

Martins:1997:NAR

[Mar97a] C. M. Martins. A note on the autocorrelations related to a bilinear model with non-independent shocks. Statistics & Probability Letters, 36(3):245–250, December 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000692.

Maruyama:1997:NPE

[Mar97b] Yuzo Maruyama. A new positive estimator of loss function. Statistics & Probability Letters, 36(3):269–274, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000734.

Markiewicz:1998:CLR

[Mar98] Augustyn Markiewicz. Comparison of linear restricted models with respect to the validity of admissible and linearly sufficient estimators. Statistics & Probability Letters, 38(4):347–354, July 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000455.

Marden:1999:SRE

[Mar99a] John I. Marden. Some robust estimates of principal components. Statistics & Probability Letters, 43(4):349–359, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002727.

Marohn:1999:EIS

[Mar99b] Frank Marohn. Estimating the index of a stable law via the potmethod. Statistics & Probability Letters, 41(4):413–423, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002107.

Masry:1994:PDE

[Mas94] Elias Masry. Probability density estimation from dependent observations using wavelets orthonormal bases. Statistics & Probability Letters, 21(3):181–194, October 19, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901147.

Mashhour:1996:SRD

[Mas96] A. F. Mashhour. Some results on discrete α-unimodality. Statistics & Probability Letters, 27(1):31–36, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000402.

Masse:1999:VNE

[Mas99] Bruno Massé. On the variance of the number of extreme points of a random convex hull. Statistics & Probability Letters, 44(2):123–130, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002983.

Matula:1992:NAS

[Mat92] Przemysław Matula. A note on the almost sure convergence of sums of negatively dependent random variables. Statistics & Probability Letters, 15(3):209–213, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901917.

Matthews:1993:SMM

[Mat93] Peter Matthews. A slowly mixing Markov chain with implications for Gibbs sampling. Statistics & Probability Letters, 17(3):231– 236, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390172F.

Matoussi:1997:RSB

[Mat97] Anis Matoussi. Reflected solutions of backward stochastic differential equations with continuous coefficient. Statistics & Probability Letters, 34(4):347–354, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296002027.

McCabe:1990:EAM

[McC90] B. P. M. McCabe. An extension of Anderson’s multiple decision procedure. Statistics & Probability Letters, 9(2):119–124, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290004O.

McKenzie:1991:LCP

[McK91] Ed. McKenzie. Linear characterizations of the Poisson distribution. Statistics & Probability Letters, 11(5):459–461, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190196X.

McKay:1996:NDE

[McK96] Ian McKay. A note on density estimation for Poisson mixtures. Statistics & Probability Letters, 27(3):255–258, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000739.

McQuarrie:1999:SSC

[McQ99] Allan D. McQuarrie. A small-sample correction for the Schwarz SIC model selection criterion. Statistics & Probability Letters, 44(1):79–86, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002946.

Margavio:1995:ARQ

[MCWD95] Thomas M. Margavio, Michael D. Conerly, William H. Woodall, and Laurel G. Drake. Alarm rates for quality control charts. Statistics & Probability Letters, 24(3):219–224, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001747.

Martin:1992:NMS

[ME92] R. J. Martin and J. A. Eccleston. A new model for slowlydecaying correlations. Statistics & Probability Letters, 13 (2):139–145, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290088M.

Meerschaert:1991:RVV

[Mee91] Mark M. Meerschaert. Regular variation in R^{k} and vectornormed domains of attraction. Statistics & Probability Letters, 11(4):287–289, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190037R.

Meeden:1993:NNB

[Mee93a] Glen Meeden. Noninformative nonparametric Bayesian estimation of quantiles. Statistics & Probability Letters, 16(2):103–109, January 27, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390153A.

Meerschaert:1993:RVG

[Mee93b] Mark M. Meerschaert. Regular variation and generalized domains of attraction in R^{k}. Statistics & Probability Letters, 18(3):233–239, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293902225.

Meester:1999:EPI

[Mee99] Ronald Meester. Extremal points of infinite clusters in stationary percolation. Statistics & Probability Letters, 42(4):361–365, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002296.

Meloche:1995:TAS

[Mel95] J. Meloche. Tactical aspects of search and rescue operations. Statistics & Probability Letters, 25(3):271–279, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400231V.

Mendieta:1991:SHF

[Men91] G. R. Mendieta. Stopped hyperfinite filtrations. Statistics & Probability Letters, 11(4):349–351, April 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190047U.

Meng:1993:ABR

[Men93] Xiao-Li Meng. On the absolute bias ratio of ratio estimators. Statistics & Probability Letters, 18(5):345–348, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390026F.

Meyer:1999:CNS

[Mey99] Mary C. Meyer. A comparison of nonparametric shape constrained bioassay estimators. Statistics & Probability Letters, 42(3):267–274, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002132.

Modarres:1996:MRT

[MG96] Reza Modarres and Joseph L. Gastwirth. A modified runs test for symmetry. Statistics & Probability Letters, 31(2):107–112, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529600020X.

Majumdar:1999:BRM

[MGH99] Suman Majumdar, Dennis Gilliland, and James Hannan. Bounds for robust maximum likelihood and posterior consistency in compound mixture state experiments. Statistics & Probability Letters, 41(3):215–227, February 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001588.

Mukerjee:1990:FOR

[MH90] Rahul Mukerjee and S. Huda. Fourth-order rotatable designs: A-optimal measures. Statistics & Probability Letters, 10(2):111– 117, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090005R.

Miao:1996:ESC

[MH96] Weiwen Miao and Marjorie G. Hahn. Existence and strong consistency of maximum likelihood estimates for 1-dimensional exponential families. Statistics & Probability Letters, 28(1):9– 21, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000755.

Mudholkar:1997:AQR

[MH97a] Govind S. Mudholkar and Alan D. Hutson. Asymmetric quasimedians: Remarks on an anomaly. Statistics & Probability Letters, 32(3):261–268, March 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600082X.

Mudholkar:1997:SRQ

[MH97b] Govind S. Mudholkar and Alan D. Hutson. Some refinements of the quasi-quantiles. Statistics & Probability Letters, 32(1):67–74, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000570.

Markatou:1995:RSE

[MHL95] Marianthi Markatou, Joel L. Horowitz, and Russell V. Lenth. Robust scale estimation based on the empirical characteristic function. Statistics & Probability Letters, 25(2):185–192, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400221S.

Mi:1996:CPF

[Mi96] Jie Mi. Crossing properties of f distributions. Statistics & Probability Letters, 27(4):289–294, May 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000887.

Mi:1999:ASP

[Mi99] Jie Mi. Age-smooth properties of mixture models. Statistics & Probability Letters, 43(3):225–236, July 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002624.

Mielniczuk:1990:RCD

[Mie90] Jan Mielniczuk. Remark concerning data-dependent bandwidth choice in density estimation. Statistics & Probability Letters, 9(1):27–33, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090091K.

Mielniczuk:1997:AMI

[Mie97] Jan Mielniczuk. On the asymptotic mean integrated squared error of a kernel density estimator for dependent data. Statistics & Probability Letters, 34(1):53–58, May 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001654.

Mikami:1997:LDC

[Mik97] Toshio Mikami. Large deviations and Central Limit Theorems for Eyraud–Farlie–Gumbel–Morgenstern processes. Statistics & Probability Letters, 35(1):73–78, August 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296002180.

Mikami:1998:ECC

[Mik98] Toshio Mikami. Equivalent conditions on the Central Limit Theorem for a sequence of probability measures on R. Statistics & Probability Letters, 37(3):237–242, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001223.

Milbrodt:1992:CIP

[Mil92] Hartmut Milbrodt. Comparing inclusion probabilities and drawing probabilities for rejective sampling and successive sampling. Statistics & Probability Letters, 14(3):243–246, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900295.

Miller:1999:MLE

[Mil99] Gregory K. Miller. Maximum likelihood estimation for the Erlang integer parameter. Statistics & Probability Letters, 43(4):335– 341, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001862.

Mkhadri:1998:RCE

[Mkh98] Abdallah Mkhadri. On the rate of convergence of the ECME algorithm. Statistics & Probability Letters, 37(1):81–87, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700103X.

Mack:1990:EMD

[MM90] Y. P. Mack and Norman S. Matloff. Estimating a mixing distribution in a multiple observation setting. Statistics & Probability Letters, 10(5):369–376, October 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090016Z.

Marzec:1991:TDE

[MM91] Leszek Marzec and Paweł Marzec. Testing the dispersive equivalence of two populations. Statistics & Probability Letters, 12(3):233–237, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715291900845.

Mathew:1992:CLL

[MM92] George Mathew and William P. McCormick. A conditional limit law result on the location of the maximum of Brownian motion. Statistics & Probability Letters, 13(3):199–202, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290048A.

Marzec:1994:DCD

[MM94] Leszek Marzec and Paweł Marzec. Dispersive comparison of distributions: a multisample testing problem. Statistics & Probability Letters, 21(3):237–245, October 19, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901201.

Molinska:1995:IAE

[MM95] Anna Molińska and Krzysztof Moliński. Improvement of ANOVA estimators of variance components in block designs with random treatments. Statistics & Probability Letters, 22(4):281– 285, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400078M.

Marzec:1996:TSM

[MM96] Leszek Marzec and Paweł Marzec. Tests for semiparametric model based on non-homogeneous Markov process. Statistics & Probability Letters, 27(2):137–143, April 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000550.

Marzec:1998:TBS

[MM98a] Leszek Marzec and Paweł Marzec. Testing based on sampled data for proportional hazards model. Statistics & Probability Letters, 37(3):303–313, March 16, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297841570.

Monsour:1998:LDE

[MM98b] Michael J. Monsour and Piotr W. Mikulski. On limiting distributions in explosive autoregressive processes. Statistics & Probability Letters, 37(2):141–147, February 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001119.

Menendez:1997:MEP

[MMP97] M. Menéndez, D. Morales, and L. Pardo. Maximum entropy principle and statistical inference on condensed ordered data. Statistics & Probability Letters, 34(1):85–93, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001691.

Menendez:1999:SIF

[MMPZ99] M. L. Menéndez, D. Morales, L. Pardo, and K. Zografos. Statistical inference for finite Markov chains based on divergences. Statistics & Probability Letters, 41(1):9–17, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001060.

Mathew:1993:LSL

[MN93] Thomas Mathew and Kenneth Nordström. Least squares and least absolute deviation procedures in approximately linear models. Statistics & Probability Letters, 16(2):153–158, January 27, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390160K.

Main:1997:TBB

[MN97a] P. Maín and H. Navarro. On tail behavior in Bayesian location inference. Statistics & Probability Letters, 35(4):363–370, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000345.

Mkhadri:1997:NCP

[MN97b] Abdallah Mkhadri and Abdelaziz Nasroallah. Note on classification and proportion estimation. Statistics & Probability Letters, 34(2):151–158, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001769.

Martinez:1999:EQF

[MO99] Harú V. Martinez and María M. Olivares. Estimation of quadratic functionals of a density. Statistics & Probability Letters, 42(4):327–332, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800193X.

Moche:1994:CRS

[Moc94] Raymond Moché. Criteria for the regularity of the sample functions of weakly harmonizable processes. Statistics & Probability Letters, 19(1):39–44, January 3, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900663.

Moffatt:1997:EMI

[Mof97] Peter G. Moffatt. Exploiting a matrix identity in the computation of the efficient score test for overdispersion in the Poisson regression model. Statistics & Probability Letters, 32(1):75–79, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000582.

Mohan:1992:UCL

[Moh92] N. R. Mohan. A unified criterion for the local uniform convergence of the density of the maximum. Statistics & Probability Letters, 14(3):189–194, June 24, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292900206.

Mojirsheibani:1997:CCC

[Moj97] M. Mojirsheibani. A consistent combined classification rule. Statistics & Probability Letters, 36(1):43–47, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000473.

Molchanov:1993:DRC

[Mol93] Ilya S. Molchanov. On distributions of random closed sets and expected convex hulls. Statistics & Probability Letters, 17(4):253– 257, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390199S.

Monti:1993:NLR

[Mon93] Anna Clara Monti. A new look at the relationship between Edgeworth expansion and saddlepoint approximation. Statistics & Probability Letters, 17(1):49–52, May 3, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390194N.

Mori:1992:ECA

[Mór92] T. F. Móri. Essential correlatedness and almost independence. Statistics & Probability Letters, 15(3):169–172, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901847.

Morgenthaler:1994:SSE

[Mor94a] Stephan Morgenthaler. Small sample efficiency and exact fit for Cauchy regression models. Statistics & Probability Letters, 19(5):381–385, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529490006X.

Mori:1994:LRH

[Mór94b] Tamás F. Móri. On long runs of heads and tails. Statistics & Probability Letters, 19(2):85–89, January 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901376.

Moser:1994:CTO

[Mos94] Martin Moser. Completeness of time-ordered indicators in censored data models. Statistics & Probability Letters, 21 (2):163–166, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294902259.

Munoz-Perez:1990:DOS

[MP90] J. Muñoz-Perez. Dispersive ordering by the spread function. Statistics & Probability Letters, 10(5):407–410, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090021X.

Mandrekar:1993:LTS

[MP93] V. Mandrekar and R. F. Patterson. Limit theorems for symmetric statistics of exchangeable random variables. Statistics & Probability Letters, 17(2):157–161, May 26, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390010G.

Malyutov:1999:FAA

[MP99] Mikhail B. Malyutov and Rostislav S. Protassov. Functional approach to the asymptotic normality of the non-linear least squares estimator. Statistics & Probability Letters, 44(4):409– 416, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000346.

Meester:1997:RCM

[MPS97] Ronald Meester, Mathew D. Penrose, and Anish Sarkar. The random connection model in high dimensions. Statistics & Probability Letters, 35(2):145–153, September 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000084.

Munoz-Perez:1990:CDF

[MPSG90] J. Muñoz-Perez and A. Sanchez-Gomez. A characterization of the distribution function: the dispersion function. Statistics & Probability Letters, 10(3):235–239, August 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090080Q.

Martinez:1991:TGU

[MQ91] Servet Martínez and Fernando Quintana. On a test for generalized upper truncated Weibull distributions. Statistics & Probability Letters, 12(4):273–279, October 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190090E.

Maejima:1997:RCM

[MR97a] Makoto Maejima and Svetlozar T. Rachev. Rate-of-convergence in the multivariate max-stable limit theorem. Statistics & Probability Letters, 32(2):115–123, March 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000636.

Matthews:1997:VRP

[MR97b] Peter C. Matthews and William F. Rosenberger. Variance in randomized play-the-winner clinical trials. Statistics & Probability Letters, 35(3):233–240, October 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000187.

Montanes:1999:ABD

[MR99] Antonio Montañés and Marcelo Reyes. The asymptotic behaviour of the Dickey–Fuller tests under the crash hypothesis. Statistics & Probability Letters, 42(1):81–89, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002016.

Malani:1993:NAV

[MRN93] Hina M. Malani, William J. Redfearn, and Jens Perch Nielsen. A note on the asymptotic variance of survival function in the semi-Markov model. Statistics & Probability Letters, 18(1):19– 25, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390094Y.

Marco:1992:CMD

[MRR92] J. M. Marco and C. Ruiz-Rivas. On the construction of multivariate distributions with given nonoverlapping multivariate marginals. Statistics & Probability Letters, 15(4):259–265, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292901607.

Maryak:1990:WCE

[MS90] John L. Maryak and James C. Spall. Weak convergence and the exponential rate of concentration for posterior density functions. Statistics & Probability Letters, 10(4):273–278, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900415.

Menendez:1992:ELR

[MS92] J. A. Menéndez and B. Salvador. Equivalence of likelihood ratio tests and obliquity. Statistics & Probability Letters, 14(3):223– 228, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900262.

Millet:1995:SPG

[MS95] Annie Millet and Wlodzmierz Smoleński. Small perturbations of Gaussian regressors. Statistics & Probability Letters, 24(1): 21–31, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400144W.

McKeague:1996:TGR

[MS96a] Ian W. McKeague and Yanqing Sun. Transformations of Gaussian random fields to Brownian sheet and nonparametric changepoint tests. Statistics & Probability Letters, 28(4):311–319, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001409.

Meng:1996:LSD

[MS96b] Q. Meng and S. Sivaganesan. Local sensitivity of density bounded priors. Statistics & Probability Letters, 27(2):163–169, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000593.

MacGibbon:1997:SED

[MS97a] Brenda MacGibbon and Glenn Shorrock. Shrinkage estimators for the dispersion parameter of the inverse Gaussian distribution. Statistics & Probability Letters, 32(2):207–214, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000740.

Meerschaert:1997:SGD

[MS97b] Mark M. Meerschaert and Hans-Peter Scheffler. The structure of generalized domains of semistable attraction. Statistics & Probability Letters, 33(4):367–372, May 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001496.

Muller:1997:TSC

[MS97c] Hans-Georg Müller and Kai-Sheng Song. Two-stage changepoint estimators in smooth regression models. Statistics & Probability Letters, 34(4):323–335, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001976.

Moore:1999:CBR

[MS99] Linda J. Moore and Ping Sa. Comparisons with the best in response surface methodology. Statistics & Probability Letters, 44(2):189–194, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000085.

Mehrotra:1990:ETS

[MSS90] K. G. Mehrotra, A. Schick, and V. Susarla. Estimation in two sample type II censoring models. Statistics & Probability Letters, 9(1):13–22, January 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090089P.

McQuarrie:1997:MSC

[MST97] Allan McQuarrie, Robert Shumway, and Chih-Ling Tsai. The model selection criterion AICu. Statistics & Probability Letters, 34(3):285–292, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001927.

McQuarrie:1999:MSO

[MT99] Allan McQuarrie and Chih-Ling Tsai. Model selection in orthogonal regression. Statistics & Probability Letters, 45(4):341–349, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000760.

Muddapur:1993:UTS

[Mud93] M. V. Muddapur. UMPU test for state dependence of a parameter in a linear birth process. Statistics & Probability Letters, 17(1):39–41, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390192L.

Muddapur:1994:UTS

[Mud94] M. V. Muddapur. UMPU test for state dependence of a parameter in a linear birth process. Statistics & Probability Letters, 20(3):247, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294000492.

Mukerjee:1992:CRM

[Muk92a] Hari Mukerjee. Convergence rates of monotone conditional quantile estimators. Statistics & Probability Letters, 15(5):417–420, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292901777.

Mukerjee:1992:POB

[Muk92b] Rahul Mukerjee. Parametric orthogonality and a Bartlett-type modification for Rao’s statistic in the presence of a nuisance parameter. Statistics & Probability Letters, 13(5):397–400, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290113J.

Mukherjee:1994:MDE

[Muk94] Kanchan Mukherjee. Minimum distance estimation in linear models with long-range dependent errors. Statistics & Probability Letters, 21(5):347–355, December 7, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000298.

Mukerjee:1995:SCP

[Muk95] Hari Mukerjee. A simple combinatorial proof of a result by Robertson and Pillers. Statistics & Probability Letters, 24(2):173–175, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294001655.

Muller:1997:DAK

[Mül97] H.-G. Müller. Density adjusted kernel smoothers for random design nonparametric regression. Statistics & Probability Letters, 36(2):161–172, December 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529700059X.

Muller:1999:LSF

[Mül99] Werner G. Müller. Least-squares fitting from the variogram cloud. Statistics & Probability Letters, 43(1):93–98, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002508.

Munk:1999:NUT

[Mun99a] Axel Munk. A note on unbiased testing for the equivalence problem — another Christmas tree. Statistics & Probability Letters, 41(4):401–406, February 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001965.

Munzel:1999:LRS

[Mun99b] Ullrich Munzel. Linear rank score statistics when ties are present. Statistics & Probability Letters, 41(4):389–395, February 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001850.

Murdoch:1992:SRE

[Mur92] Duncan J. Murdoch. Scalable relative effectiveness models. Statistics & Probability Letters, 13(4):321–324, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900413.

Muselli:1996:SES

[Mus96] Marco Muselli. Simple expressions for success run distributions in Bernoulli trials. Statistics & Probability Letters, 31(2):121–128, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000223.

Masse:1995:EEB

[MV95a] Bruno Massé and Marie-Claude Viano. Explicit and exponential bounds for a test on the coefficient of an AR(1) model. Statistics & Probability Letters, 25(4):365–371, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002439.

Meerschaert:1995:SGS

[MV95b] Mark M. Meerschaert and Jeery Alan Veeh. Symmetry groups in d-space. Statistics & Probability Letters, 22(1):1–6, January 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400039B.

Mashouri:1993:CTC

[MVA93] Gh. Mashouri, M. N. Vartak, and M. L. Aggarwal. A combinatorial theorem for a class of residual treatment effects designs. Statistics & Probability Letters, 17(2):123–125, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900065.

Misra:1997:ECM

[MvdM97] Neeraj Misra and Edward C. van der Meulen. On estimation of the common mean of k(≥ 2) normal populations with order restricted variances. Statistics & Probability Letters, 36(3):261–267, December 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000722.

Mayer-Wolf:1990:APL

[MW90] Eddy Mayer-Wolf. An application to probability laws of the noncontinuity of the inverse Radon transform. Statistics & Probability Letters, 10(1):87–90, June 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090116O.

Magiera:1991:CPE

[MW91] Ryszard Magiera and Maciej Wilczyński. Conjugate priors for exponential-type processes. Statistics & Probability Letters, 12(5):379–384, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190026N.

Massey:1993:PGT

[MW93] William A. Massey and Ward Whitt. A probabilistic generalization of Taylor’s theorem. Statistics & Probability Letters, 16(1):51–54, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390122Y.

Maxwell:1997:LLT

[MW97] Michael Maxwell and Michael Woodroofe. A local limit theorem for hidden Markov chains. Statistics & Probability Letters, 32(2):125–131, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000648.

Martinsek:1996:FWC

[MX96] Adam T. Martinsek and Yi Xu. Fixed width confidence bands for densities under censoring. Statistics & Probability Letters, 30(3):257–264, October 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002286.

Meczarski:1991:SBE

[MZ91] Marek Meczarski and Ryszard Zieliński. Stability of the Bayesian estimator of the Poisson mean under the inexactly specified gamma prior. Statistics & Probability Letters, 12(4):329–333, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190101V.

Merzbach:1993:WMI

[MZ93] Ely Merzbach and Moshe Zakai. Worthy martingales and integrators. Statistics & Probability Letters, 16(5):391–395, April 8, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390073R.

Meczarski:1997:SPM

[MZ97] Marek Meczarski and Ryszard Zieliński. Stability of the posterior mean in linear models: An admissibility property of D-optimum and E-optimum designs. Statistics & Probability Letters, 33(2):117–123, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001186.

Marsaglia:1994:REI

[MZM94] George Marsaglia, Arif Zaman, and John C. W. Marsaglia. Rapid evaluation of the inverse of the normal distribution function. Statistics & Probability Letters, 19(4):259–266, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901740.

Marsaglia:1990:TUR

[MZT90] George Marsaglia, Arif Zaman, and Wai Wan Tsang. Toward a universal random number generator. Statistics & Probability Letters, 9(1):35–39, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090092L.

Nadarajah:1999:PMB

[Nad99] S. Nadarajah. A polynomial model for bivariate extreme value distributions. Statistics & Probability Letters, 42(1):15–25, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001795.

Nagahara:1999:PCP

[Nag99] Yuichi Nagahara. The PDF and CF of Pearson type IV distributions and the ML estimation of the parameters. Statistics & Probability Letters, 43(3):251–264, July 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800265X.

Navarro:1999:SMW

[Nav99] H. Navarro. On smoothness measurement for weakly stationary processes. Statistics & Probability Letters, 44(2):147–154, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000024.

Nayak:1993:ETR

[Nay93] Tapan K. Nayak. On estimating the total rate of the unobserved processes. Statistics & Probability Letters, 17(5):351–354, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390254G.

Nayak:1998:EEL

[Nay98] Tapan K. Nayak. On equivariant estimation of the location of elliptical distributions under Pitman closeness criterion. Statistics & Probability Letters, 36(4):373–378, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000849.

Nobel:1993:NUL

[ND93] Andrew Nobel and Amir Dembo. A note on uniform laws of averages for dependent processes. Statistics & Probability Letters, 17(3):169–172, June 18, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390163D.

Neeman:1991:UEU

[Nee91] Teresa Neeman. U-estimability under sequential binomial sampling. Statistics & Probability Letters, 11(2):185–188, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190141D.

Nelsen:1992:MAM

[Nel92] Roger B. Nelsen. On measures of association as measures of positive dependence. Statistics & Probability Letters, 14(4):269– 274, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290056B.

Neudecker:1994:CEV

[Neu94] Heinz Neudecker. A compact expression for the variance of sample second-order moments in multivariate linear relations an alternative proof of Satorra’s result. Statistics & Probability Letters, 19(2):115–117, January 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901422.

Neuenschwander:1996:CIT

[Neu96] Daniel Neuenschwander. Commutative infinitesimal triangular systems on Euclidean motion groups. Statistics & Probability Letters, 30(1):33–36, September 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001980.

Neuenschwander:1997:NPM

[Neu97] Daniel Neuenschwander. A new proof of the multidimensional convergence of types theorem. Statistics & Probability Letters, 33(1):85–88, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001137.

Neuenschwander:1998:LIL

[Neu98] Daniel Neuenschwander. Law of the iterated logarithm for Lévy’s area process composed with Brownian motion. Statistics & Probability Letters, 40(4):371–377, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001394.

Newman:1990:CSR

[New90] Charles M. Newman. Convergence of the sum of reciprocal renewal times. Statistics & Probability Letters, 10(3):185–187, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090071E.

Neuenschwander:1997:NST

[NF97] Beat E. Neuenschwander and Bernard D. Flury. A note on Silvey’s (1959) Theorem. Statistics & Probability Letters, 36 (3):307–317, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297000771.

Nieto:1995:KFS

[NG95] Fabio H. Nieto and Victor M. Guerrero. Kalman filter for singular and conditional state-space models when the system state and the observational error are correlated. Statistics & Probability Letters, 22(4):303–310, March 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400081I.

Nikouline:1999:AGI

[Nik99] A. M. Nikouline. On approximation of Gaussian integrals. Statistics & Probability Letters, 45(2):159–165, November 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000553.

Nanda:1998:PSP

[NJS98] Asok K. Nanda, Kanchan Jain, and Harshinder Singh. Preservation of some partial orderings under the formation of coherent systems. Statistics & Probability Letters, 39(2):123–131, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000431.

Naranjo:1997:RRE

[NM97] Joshua D. Naranjo and Joseph W. McKean. Rank regression with estimated scores. Statistics & Probability Letters, 33(2):209– 216, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001296.

Nagaraja:1996:CBF

[NN96] H. N. Nagaraja and V. B. Nevzorov. Correlations between functions of records can be negative. Statistics & Probability Letters, 29(2):95–100, August 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001611.

Naik-Nimbalkar:1997:ELR

[NNR97] U. V. Naik-Nimbalkar and M. B. Rajarshi. Empirical likelihood ratio test for equality of k medians in censored data. Statistics & Probability Letters, 34(3):267–273, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001903.

Nobel:1995:CCU

[Nob95] Andrew Nobel. A counterexample concerning uniform ergodic theorems for a class of functions. Statistics & Probability Letters, 24(2):165–168, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294001633.

Nolan:1998:PMS

[Nol98] John P. Nolan. Parameterizations and modes of stable distributions. Statistics & Probability Letters, 38(2):187–195, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000108.

Nolan:1999:MMM

[Nol99] D. Nolan. On min-max majority and deepest points. Statistics & Probability Letters, 43(4):325–333, July 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001734.

Northshield:1995:SMB

[Nor95a] S. Northshield. On the spectrum and Martin boundary of homogeneous spaces. Statistics & Probability Letters, 22(4):275– 279, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400077L.

Norvaisa:1995:SLI

[Nor95b] R. Norvaisa. The Strassen law of the iterated logarithm in Banach function spaces. Statistics & Probability Letters, 25(1):1– 8, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400198H.

Niinimaa:1990:FSB

[NOT90] A. Niinimaa, H. Oja, and Mara Tableman. The finitesample breakdown point of the Oja bivariate median and of the corresponding half-samples version. Statistics & Probability Letters, 10(4):325–328, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090050H.

Neudecker:1991:LSR

[NS91] Heinz Neudecker and Albert Satorra. Linear structural relations: Gradient and Hessian of the fitting function. Statistics & Probability Letters, 11(1):57–61, January 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190178T.

Neath:1992:TTT

[NS92] Andrew A. Neath and Francisco J. Samaniego. On the total time on test transform of an IFRA distribution. Statistics & Probability Letters, 14(4):289–291, July 17, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290059E.

Neath:1996:BEM

[NS96a] Andrew A. Neath and Francisco J. Samaniego. On Bayesian estimation of the multiple decrement function in the competing risks problem. Statistics & Probability Letters, 31(2):75–83, December 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000168.

Neath:1996:DRM

[NS96b] Andrew A. Neath and Francisco J. Samaniego. On the distinguished role of the multivariate exponential distribution in Bayesian estimation in competing risks problems. Statistics & Probability Letters, 31(1):69–74, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000156.

Neudecker:1996:AET

[NS96c] Heinz Neudecker and Albert Satorra. The algebraic equality of two asymptotic tests for the hypothesis that a normal distribution has a specified correlation matrix. Statistics & Probability Letters, 30(2):99–103, October 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002065.

Neath:1997:BEM

[NS97a] Andrew A. Neath and Francisco J. Samaniego. On Bayesian estimation of the multiple decrement function in the competing risks problem, II: The discrete case. Statistics & Probability Letters, 35(4):345–354, November 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000321.

Nikabadze:1997:MCU

[NS97b] A. Nikabadze and W. Stute. Model checks under random censorship. Statistics & Probability Letters, 32(3):249–259, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000818.

Nappo:1998:OPT

[NS98] G. Nappo and F. Spizzichino. Ordering properties of the TTT-plot of lifetimes with Schur joint densities. Statistics & Probability Letters, 39(3):195–203, August 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000212.

Novikov:1999:SMI

[NV99] Alexander Novikov and Esko Valkeila. On some maximal inequalities for fractional Brownian motions. Statistics & Probability Letters, 44(1):47–54, August 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002909.

Neill:1990:TRF

[NY90] James W. Neill and Shie-Shien Yang. Testing regression function adequacy in nonlinear multiresponse models. Statistics & Probability Letters, 9(2):101–105, February 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290001L.

Nyblom:1992:NIO

[Nyb92] Jukka Nyblom. Note on interpolated order statistics. Statistics & Probability Letters, 14(2):129–131, May 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290076H.

Nychka:1990:SPA

[Nyc90] Douglas Nychka. Some properties of adding a smoothing step to the EM algorithm. Statistics & Probability Letters, 9(2):187– 193, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290015W.

ONeill:1995:GTO

[OB95] Terence J. O’Neill and Simon C. Barry. Group truncated ordinal regression. Statistics & Probability Letters, 22(3):195–203, February 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400067I.

Ozturk:1996:AFE

[ÖH96] Ömer Öztürk and Thomas P. Hettmansperger. Almost fully efficient and robust simultaneous estimation of location and scale parameters: a minimum distance approach. Statistics & Probability Letters, 29(3):233–244, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001786.

Ohtani:1996:FIS

[Oht96] Kazuhiro Ohtani. Further improving the Stein-rule estimator using the Stein variance estimator in a misspecified linear regression model. Statistics & Probability Letters, 29(3):191–199, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001735.

Ohtani:1999:MPH

[Oht99] Kazuhiro Ohtani. MSE performance of a heterogeneous pre-test estimator. Statistics & Probability Letters, 41(1):65–71, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001230.

Okamoto:1994:CAA

[Oka94] Masashi Okamoto. Correspondence analysis of an artificial cylinder data. Statistics & Probability Letters, 20(2):101–112, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900248.

Oleszkiewicz:1997:CCN

[Ole97] Krzysztof Oleszkiewicz. On certain characterization of normal distribution. Statistics & Probability Letters, 33(3):277–280, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529600137X.

Olkin:1998:DIP

[Olk98] Ingram Olkin. The density of the inverse and pseudo-inverse of a random matrix. Statistics & Probability Letters, 38(2):131– 135, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001636.

Olofsson:1999:PAA

[Olo99] Peter Olofsson. A Poisson approximation with applications to the number of maxima in a discrete sample. Statistics & Probability Letters, 44(1):23–27, August 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002880.

Oakes:1992:NRC

[OM92] David Oakes and Amita K. Manatunga. A new representation of Cox’s score statistic and its variance. Statistics & Probability Letters, 14(2):107–110, May 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290072D.

Omori:1997:CTM

[Omo97] Yasuhiro Omori. Comparing two means in count models having random effects — a UMPU test. Statistics & Probability Letters, 34(3):225–235, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600185X.

ONeill:1991:IPH

[O’N91] Terence J. O’Neill. The Inverse Proportional Hazards Model. Statistics & Probability Letters, 12(2):125–129, August 1991.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190055V.

ONeill:1992:BFL

[O’N92] Terence J. O’Neill. The bias of Fisher’s linear discriminant function when the variancies are not equal. Statistics & Probability Letters, 14(3):205–210, June 24, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290023X.

ONeill:1994:MLR

[O’N94] Terence J. O’Neill. Mixture or logistic regression estimation for discrimination. Statistics & Probability Letters, 20(2):139– 142, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900299.

Ong:1995:SSM

[Ong95] S. H. Ong. Some stochastic models leading to the convolution of two binomial variables. Statistics & Probability Letters, 22(2):161–166, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400063E.

Oosterhoff:1994:TMS

[Oos94] J. Oosterhoff. Trimmed mean or sample median? Statistics & Probability Letters, 20(5):401–409, August 9, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901325.

Oliveira:1998:WCU

[OS98] P. E. Oliveira and Ch. Suquet. Weak convergence in L^{p}(0, 1) of the uniform empirical process under dependence. Statistics & Probability Letters, 39(4):363–370, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000911.

Ouknine:1992:ABF

[Ouk92] Y. Ouknine. On the asymptotic behaviour of functionals of some semimartingales. Statistics & Probability Letters, 15(5):403–405, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292901755.

Ouknine:1994:PFS

[Ouk94] Y. Ouknine. On polynomial filtration of some continuous semimartingales. Statistics & Probability Letters, 20(3):169–172, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900388.

Olson:1991:ACM

[OW91] Jane M. Olson and Lisa A. Weissfeld. Approximation of certain multivariate integrals. Statistics & Probability Letters, 11(4):309– 317, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190040X.

Philippou:1990:MDOa

[PA90] Andreas N. Philippou and Demetris L. Antzoulakos. Multivariate distributions of order k on a generalized sequence. Statistics & Probability Letters, 9(5):453–463, May 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900389.

Pak:1996:MHD

[Pak96] Ro Jin Pak. Minimum Hellinger distance estimation in simple linear regression models; distribution and efficiency. Statistics & Probability Letters, 26(3):263–269, February 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000194.

Pakes:1998:MRS

[Pak98] Anthony G. Pakes. Mixture representations for symmetric generalized Linnik laws. Statistics & Probability Letters, 37(3):213– 221, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001193.

Palacios:1992:BCT

[Pal92] JoséLuis Palacios. A bound for the covering time of random walks on graphs. Statistics & Probability Letters, 14(1):9–11, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290204I.

Palacios:2001:CRW

[Pal01] José Luis Palacios. Correction on “Random walks on edgetransitive graphs (II)” [Statist. Probab. Lett. 43 (1999) 25– 32]. Statistics & Probability Letters, 53(1):111–112, May 15, 2001. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715200001863. See [PRB99].

Papathanasiou:1990:SCD

[Pap90] V. Papathanasiou. Some characterizations of distributions based on order statistics. Statistics & Probability Letters, 9(2):145– 147, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290009T.

Papathanasiou:1993:EII

[Pap93] V. Papathanasiou. An extension of the information inequality and related characterizations. Statistics & Probability Letters, 18(1):27–32, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390095Z.

Papadatos:1995:IOS

[Pap95] N. Papadatos. Intermediate order statistics with applications to nonparametric estimation. Statistics & Probability Letters, 22(3):231–238, February 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400071F.

Paparoditis:1996:BPC

[Pap96] Efstathios Paparoditis. Bootstrapping periodogram and cross periodogram statistics of vector autoregressive moving average models. Statistics & Probability Letters, 27(4):385–391, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500047X.

Paparoditis:1997:BPC

[Pap97] Efstathios Paparoditis. Bootstrapping periodogram and cross periodogram statistics of vector autoregressive moving average models. Statistics & Probability Letters, 31(3):243, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002015.

Paparoditis:1998:ABP

[Pap98] Efstathios Paparoditis. Addendum to “Bootstrapping periodogram and cross periodogram statistics of vector autoregressive moving average models”. Statistics & Probability Letters, 37(1):109, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297800028.

Paramasamy:1992:MKS

[Par92] S. Paramasamy. On the multivariate Kolmogorov–Smirnov distribution. Statistics & Probability Letters, 15(2):149–155, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290128R.

Park:1996:ARAb

[Par96] Sangun Park. An asymptotic relation arising in the decomposition of the likelihood of order statistics. Statistics & Probability Letters, 29(2):101–106, August 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500162X.

Park:1998:LSE

[Par98a] Chul Gyu Park. Least squares estimation of two functions under order restriction in isotonicity. Statistics & Probability Letters, 37(1):97–100, January 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001053.

Park:1998:TUD

[Par98b] Chul Gyu Park. Testing for unimodal dependence in an ordered contingency table with restricted marginal probabilities. Statistics & Probability Letters, 37(2):121–129, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001090.

Park:1999:GFT

[Par99] Sangun Park. A goodness-of-fit test for normality based on the sample entropy of order statistics. Statistics & Probability Letters, 44(4):359–363, October 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000279.

Perez-Abreu:1997:SPB

[PARA97] Victor Pérez-Abreu and Alfonso Rocha-Arteaga. On stable processes of bounded variation. Statistics & Probability Letters, 33(1):69–77, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001113.

Philippou:1990:MDOb

[PAT90] Andreas N. Philippou, Demetris L. Antzoulakos, and Gregory A. Tripsiannis. Multivariate distributions of order k, part II. Statistics & Probability Letters, 10(1):29–35, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090108J.

Patil:1996:NDD

[Pat96] Prakash Patil. A note on deconvolution density estimation. Statistics & Probability Letters, 29(1):79–84, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001581.

Paula:1996:ALP

[Pau96] Gilberto A. Paula. On approximation of the level probabilities for testing ordered parallel regression lines. Statistics & Probability Letters, 30(4):333–338, November 15, 1996.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715295002367.

Pace:1997:SSA

[PB97] R. Kelley Pace and Ronald Barry. Sparse spatial autoregressions. Statistics & Probability Letters, 33(3):291–297, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529600140X.

Park:1991:EIS

[PC91] Byeong U. Park and Sinsup Cho. Estimation of integrated squared spectral density derivatives. Statistics & Probability Letters, 12(1):65–72, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190165N.

Pan:1998:CMM

[PCK98] Wei Pan, Rick Chappell, and Michael R. Kosorok. On consistency of the monotone MLE of survival for left truncated and interval-censored data. Statistics & Probability Letters, 38(1):49– 57, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001533.

Papageorgiou:1995:CBC

[PD95] H. Papageorgiou and Katerina M. David. On a class of bivariate compounded Poisson distributions. Statistics & Probability Letters, 23(1):93–104, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400099T.

Patra:1999:MMW

[PD99] Kaushik Patra and Dipak K. Dey. A multivariate mixture of Weibull distributions in reliability modeling. Statistics & Probability Letters, 45(3):225–235, November 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000620.

Peddada:1997:CIE

[Ped97] Shyamal Das Peddada. Confidence interval estimation of population means subject to order restrictions using resampling procedures. Statistics & Probability Letters, 31(4):255–265, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000375.

Pekoz:1999:ITC

[Pek99] Erol A. Peköz. Ignatov’s theorem and correlated record values. Statistics & Probability Letters, 43(2):107–111, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001771.

Pemantle:1996:MVT

[Pem96] Robin Pemantle. Maximum variation of total risk. Statistics & Probability Letters, 28(3):285–289, July 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001379.

Peng:1996:WLF

[Pen96] Nan-Fu Peng. On weak lumpability of a finite Markov chain. Statistics & Probability Letters, 27(4):313–318, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000917.

Peng:1998:AUE

[Pen98] L. Peng. Asymptotically unbiased estimators for the extremevalue index. Statistics & Probability Letters, 38(2):107–115, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001600.

Peng:1999:ECT

[Pen99] L. Peng. Estimation of the coefficient of tail dependence in bivariate extremes. Statistics & Probability Letters, 43(4):399– 409, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002806.

Perez:1993:SDF

[Pér93] Josefa Linares Pérez. On a stochastic differentiation formula for Hilbert–Schmidt valued stochastic integrals. Statistics & Probability Letters, 18(2):129–135, September 22, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390181H.

Perline:1998:MPD

[Per98] Richard Perline. Mixed Poisson distributions tail equivalent to their mixing distributions. Statistics & Probability Letters, 38(3):229–233, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000194.

Petrov:1995:ASB

[Pet95] Valentin V. Petrov. On the almost sure behaviour of sums of random variables. Statistics & Probability Letters, 24(3):229–231, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294001769.

Petrov:1996:SLL

[Pet96a] Valentin V. Petrov. On the Strong Law of Large Numbers. Statistics & Probability Letters, 26(4):377–380, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000348.

Petrovic:1996:CMR

[Pet96b] Ljiljana Petrović. Causality and Markovian representations. Statistics & Probability Letters, 29(3):223–227, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500176X.

Pflug:1991:NCS

[Pfl91] Georg Ch. Pflug. A note on the comparison of stationary laws of Markov processes. Statistics & Probability Letters, 11(4):331– 334, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190044R.

Pan:1998:PDC

[PFvR98] Jian-Xin Pan, Kai-Tai Fang, and Dietrich von Rosen. On the posterior distribution of the covariance matrix of the growth curve model. Statistics & Probability Letters, 38(1):33–39, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700151X.

Paranjpe:1994:SVD

[PG94a] S. A. Paranjpe and A. P. Gore. Selecting variables for discrimination when covariance matrices are unequal. Statistics & Probability Letters, 21(5):417–419, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000484.

Pham:1994:AND

[PG94b] Dinh Tuan Pham and Dominique Guégan. Asymptotic normality of the discrete Fourier transform of long memory time series. Statistics & Probability Letters, 21(4):299–309, November 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294000239.

Pham-Gia:1994:HRP

[PG94c] T. Pham-Gia. The hazard rate of the power-quadratic exponential family of distributions. Statistics & Probability Letters, 20(5):375–382, August 9, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294901295.

Philips:1998:NCV

[PG98] R. Philips and I. Guttman. A new criterion for variable selection. Statistics & Probability Letters, 38(1):11–19, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700148X.

Prewitt:1999:VBD

[PG99] Kathryn Prewitt and Ulkü Gürler. Variance of the bivariate density estimator for left truncated right censored data. Statistics & Probability Letters, 45(4):351–358, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000772.

Pham-Gia:1992:UMD

[PGTD92] T. Pham-Gia, N. Turkkan, and Q. P. Duong. Using the mean deviation in the elicitation of the prior distribution. Statistics & Probability Letters, 13(5):373–381, April 7, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290110Q.

Phelan:1990:ETP

[Phe90] Michael J. Phelan. Estimating the transition probabilities from censored Markov renewal processes. Statistics & Probability Letters, 10(1):43–47, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090110S.

Phelan:1995:AIP

[Phe95] Michael J. Phelan. Asymptotic information for parametric estimation from an equilibrium particle process. Statistics & Probability Letters, 24(3):193–198, August 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001699.

Pinelis:1994:MIS

[Pin94] Iosif Pinelis. On a majorization inequality for sums of independent random vectors. Statistics & Probability Letters, 19(2):97– 99, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901392.

Pittenger:1990:SMV

[Pit90] A. O. Pittenger. Sharp mean-variance bounds for Jensen-type inequalities. Statistics & Probability Letters, 10(2):91–94, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090001N.

Pitarakis:1998:BOE

[Pit98] Jean-Yves Pitarakis. On the bias of the OLS estimator in a nonstationary dynamic panel data model. Statistics & Probability Letters, 38(2):145–150, June 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529700165X.

Pillai:1994:SCP

[PJ94] R. N. Pillai and K. Jayakumar. Specialised class L property and stationary autoregressive process. Statistics & Probability Letters, 19(1):51–56, January 3, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529490068X.

Pillai:1995:DML

[PJ95] R. N. Pillai and K. Jayakumar. Discrete Mittag–Leffler distributions. Statistics & Probability Letters, 23(3):271–274, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400124Q.

Park:1996:CBS

[PK96] Pil S. Park and Anant M. Kshirsagar. Correlation between successive values of Anderson’s classification statistic in the hold-out method. Statistics & Probability Letters, 27(3):259–265, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000763.

Park:1997:IRN

[PKJ97] B. U. Park, W. C. Kim, and M. C. Jones. On identity reproducing nonparametric regression estimators. Statistics & Probability Letters, 32(3):279–290, March 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000855.

Page:1993:CMP

[PKM93] Connie Page, David Kreling, and Ella Mae Matsumura. Comparison of the mean per unit and ratio estimators under a simple applications-motivated model. Statistics & Probability Letters, 17(2):97–104, May 26, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293900032.

Poon:1992:MLG

[PL92] Wai-Yin Poon and Sik-Yum Lee. Maximum likelihood and generalized least squares analyses of two-level structural equation models. Statistics & Probability Letters, 14(1):25–30, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290206K.

Poon:1993:ASE

[PL93] Wai-Yin Poon and Yin-Ping Leung. Analysis of structural equation models with interval and polytomous data. Statistics & Probability Letters, 17(2):127–137, May 26, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293900076.

Pal:1995:IME

[PL95] Nabendu Pal and Chiahua Ling. Improved minimax estimation of powers of the variance of a multivariate normal distribution under the entropy loss function. Statistics & Probability Letters, 24(3):205–211, August 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001725.

Pakes:1998:LLN

[PL98] Anthony G. Pakes and Yun Li. Limit laws for the number of near maxima via the Poisson approximation. Statistics & Probability Letters, 40(4):395–401, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001485.

Pogany:1992:STE

[Pog92] Tibor Pogány. Sharp truncation error bound in the sampling reconstruction of homogeneous random fields. Statistics & Probability Letters, 15(5):345–348, December 8, 1992.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290151T.

Perez-Ocon:1996:FPT

[POGP96] Rafael Pérez-Ocón and M. Luz Gámiz-Pérez. On firstpassage times in increasing Markov processes. Statistics & Probability Letters, 26(3):199–203, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000107.

Pohl:1992:DOD

[Poh92] Gerhardt M. Pohl. D-optimality of the dual of BIB designs. Statistics & Probability Letters, 14(3):201–203, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290022W.

Poor:1991:MDB

[Poo91] H. Vincent Poor. The maximum difference between the binomial and Poisson distributions. Statistics & Probability Letters, 11(2):103–106, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190125B.

Pordzik:1992:AEG

[Por92] Paweł R. Pordzik. Adjusting of estimates in general linear model with respect to linear restrictions. Statistics & Probability Letters, 15(2):125–130, September 28, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290124N.

Pourahmadi:1993:RBP

[Pou93] Mohsen Pourahmadi. On relations between prediction error covariance of univariate and multivariate processes. Statistics & Probability Letters, 16(5):355–359, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390069U.

Peddada:1992:QAB

[PP92a] Shyamal Das Peddada and Girish Patwardhan. Qualms about BC_{a} bootstrap confidence intervals. Statistics & Probability Letters, 15(1):77–83, September 3, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290288G.

Percus:1992:IRS

[PP92b] Ora E. Percus and J. K. Percus. Intrinsic relations in the structure of linear congruential generators modulo 2^{β}. Statistics & Probability Letters, 15(5):381–383, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901582.

Park:1995:QET

[PP95] Hyo-Il Park and Sang-Gue Park. Quantile estimation of treatment effect for the two-sample problem with right censored data. Statistics & Probability Letters, 24(2):139–145, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001585.

Papadatos:1996:GVB

[PP96a] N. Papadatos and V. Papathanasiou. A generalization of variance bounds. Statistics & Probability Letters, 28(2):191–194, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001174.

Pazman:1996:DFM

[PP96b] A. Pázman and L. Pronzato. A Dirac-function method for densities of nonlinear statistics and for marginal densities in nonlinear regression. Statistics & Probability Letters, 26(2):159–167, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000062.

Palmowski:1996:NMI

[PR96] Zbigniew Palmowski and Tomasz Rolski. A note on martingale inequalities for fluid models. Statistics & Probability Letters, 31(1):13–21, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000077.

Petrone:1997:NDP

[PR97] Sonia Petrone and Adrian E. Raftery. A note on the Dirichlet process prior in Bayesian nonparametric inference with partial exchangeability. Statistics & Probability Letters, 36(1):69–83, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000503.

Palacios:1998:RWE

[PR98a] JoséLuis Palacios and JoséMiguel Renom. Random walks on edge transitive graphs. Statistics & Probability Letters, 37(1):29–34, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000953.

Poghosyan:1998:IPM

[PR98b] S. Poghosyan and S. Roelly. Invariance principle for martingaledifference random fields. Statistics & Probability Letters, 38(3):235–245, June 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000200.

Palacios:1999:RWE

[PRB99] José Luis Palacios, José Miguel Renom, and Pedro Berrizbeitia. Random walks on edge-transitive graphs (II). Statistics & Probability Letters, 43(1):25–32, May 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002417. See correction [Pal01].

Preater:1993:NMO

[Pre93] J. Preater. A note on monotonicity in optimal multiple stopping problems. Statistics & Probability Letters, 16(5):407–410, April 8, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390075T.

Privault:1999:MSI

[Pri99] Nicolas Privault. Multiple stochastic integral expansions of arbitrary Poisson jump times functionals. Statistics & Probability Letters, 43(2):179–188, June 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002570.

Pronzato:1998:PEV

[Pro98] L. Pronzato. On a property of the expected value of a determinant. Statistics & Probability Letters, 39(2):161–165, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000571.

Pena:1992:NEA

[PRS92] Edsel A. Peña, Vijay K. Rohatgi, and Gábor J. Székely. On the non-existence of ancillary statistics. Statistics & Probability Letters, 15(5):357–360, December 8, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290153V.

Pruitt:1992:IBE

[Pru92] Ronald C. Pruitt. An inconsistent Bayes estimator in bivariate survival curve analysis. Statistics & Probability Letters, 15(3):177–180, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292901869.

Paula:1994:TOH

[PS94] Gilberto A. Paula and Pranab K. Sen. Tests of ordered hypotheses in linkage in heredity. Statistics & Probability Letters, 20(5):395–400, August 9, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294901317.

Peligrad:1995:NAS

[PS95a] Magda Peligrad and Qi-Man Shao. A note on the almost sure Central Limit Theorem for weakly dependent random variables. Statistics & Probability Letters, 22(2):131–136, February 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400059H.

Pukelsheim:1995:OBD

[PS95b] Friedrich Pukelsheim and Bikas Kumar Sinha. Optimal block designs revisited: An approximate theory detour. Statistics & Probability Letters, 24(4):375–380, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400197G.

Park:1996:ARAa

[PS96a] Chul Gyu Park and Dong Wan Shin. On the asymptotics of residuals in autoregressive moving average processes with one autoregressive unit root. Statistics & Probability Letters, 27(4):341– 346, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529500095X.

Peligrad:1996:NEV

[PS96b] Magda Peligrad and Qi-Man Shao. A note on estimation of variance for ρ-mixing sequences. Statistics & Probability Letters, 26(2):141–145, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000038.

Pawlak:1997:KRE

[PS97a] M. Pawlak and U. Stadtmüller. Kernel regression estimators for signal recovery. Statistics & Probability Letters, 31(3):185–198, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000314.

Pellerey:1997:CID

[PS97b] Franco Pellerey and Moshe Shaked. Characterizations of the IFR and DFR aging notions by means of the dispersive order. Statistics & Probability Letters, 33(4):389–393, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001526.

Park:1998:REP

[PS98] Heungsun Park and L. A. Stefanski. Relative-error prediction. Statistics & Probability Letters, 40(3):227–236, October 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000881.

Perrin:1999:RNS

[PS99a] Olivier Perrin and Rachid Senoussi. Reducing non-stationary stochastic processes to stationarity by a time deformation. Statistics & Probability Letters, 43(4):393–397, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002788.

Phadia:1999:EMP

[PS99b] Eswar G. Phadia and Peter Yi-Shi Shao. Exact moments of the product limit estimator. Statistics & Probability Letters, 41(3):277–286, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298001643.

Palacios:1996:NEH

[PT96] JoséLuis Palacios and Prasad Tetali. A note on expected hitting times for birth and death chains. Statistics & Probability Letters, 30(2):119–125, October 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500209X.

Polymenis:1998:DNC

[PT98] A. Polymenis and D. M. Titterington. On the determination of the number of components in a mixture. Statistics & Probability Letters, 38(4):295–298, July 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000303.

Pukelsheim:1997:ERS

[Puk97] Friedrich Pukelsheim. Efficient rounding of sampling allocations. Statistics & Probability Letters, 35(2):141–143, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000072.

Purkayastha:1998:SPT

[Pur98] Sumitra Purkayastha. Simple proofs of two results on convolutions of unimodal distributions. Statistics & Probability Letters, 39(2):97–100, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000133.

Puri:1990:MIC

[PV90] Madan Lal Puri and István Vincze. Measure of information and contiguity. Statistics & Probability Letters, 9(3):223–228, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090060K.

Portnoy:1992:EWB

[PW92] Stephen Portnoy and A. H. Welsh. Exactly what is being modelled by the systematic component in a heteroscedastic linear regression. Statistics & Probability Letters, 13(4):253–258, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290031Y.

Papageorgiou:1997:PMI

[PW97] H. Papageorgiou and Jacek Wesołowski. Posterior mean identifies the prior distribution in nb and related models. Statistics & Probability Letters, 36(2):127–134, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000552.

Pena:1997:SDT

[PZ97] Daniel Peña and Ruben Zamar. A simple diagnostic tool for local prior sensitivity. Statistics & Probability Letters, 36(2):205–212, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000679.

Qi:1997:NNM

[Qi97] Yongcheng Qi. A note on the number of maxima in a discrete sample. Statistics & Probability Letters, 33(4):373–377, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001502.

Qian:1992:MLS

[Qia92] Shixian Qian. Minimum lower sets algorithms for isotonic regression. Statistics & Probability Letters, 15(1):31–35, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292902819.

Qian:1996:MDE

[Qia96a] Lianfen Qian. Minimum distance estimators for random coefficient autoregressive models. Statistics & Probability Letters, 29(3):251–262, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001808.

Qian:1996:ATO

[Qia96b] Shixian Qian. An algorithm for tree-ordered isotonic median regression. Statistics & Probability Letters, 27(3):195–199, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500064X.

Qin:1997:SPL

[Qin97] Yong Song Qin. Semi-parametric likelihood ratio confidence intervals for various differences of two populations. Statistics & Probability Letters, 33(2):135–143, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001216.

Qiying:1995:SLS

[Qiy95] Wang Qiying. The strong law of U-statistics with ϕ^{o}∗-mixing samples. Statistics & Probability Letters, 23(2):151–155, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400107J.

Qiying:1996:MIO

[Qiy96] Wang Qiying. On the maximal inequality. Statistics & Probability Letters, 31(2):85–89, December 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600017X.

Quine:1992:ELG

[QR92] M. P. Quine and J. Robinson. Estimation for a linear growth model. Statistics & Probability Letters, 15(4):293–297, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901652.

Qian:1990:PEH

[QT90] W. Qian and D. M. Titterington. Parameter estimation for hidden Gibbs chains. Statistics & Probability Letters, 10(1):49– 58, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090111J.

Qian:1992:NAL

[QT92] W. Qian and D. M. Titterington. Normal approximations for lattice systems. Statistics & Probability Letters, 13(3):245–252, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900549.

Quine:1999:CLT

[Qui99] M. P. Quine. A Central Limit Theorem for self-normalized products of random variables. Statistics & Probability Letters, 43(2):137–143, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002430.

Qi:1997:LBMb

[QW97a] Y. Qi and R. J. G. Wilms. The limit behavior of maxima modulo one and the number of maxima. Statistics & Probability Letters, 34(1):75–84, May 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600168X.

Qi:1997:LBMa

[QW97b] Yongcheng Qi and R. J. G. Wilms. The limit behavior of maxima modulo one and the number of maxima. Statistics & Probability Letters, 32(4):357–366, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000958.

Rukhin:1992:RBV

[RA92] Andrew L. Rukhin and Malwane M. A. Ananda. Risk behavior of variance estimators in multivariate normal distribution. Statistics & Probability Letters, 13(2):159–166, January 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290091I.

Remadi:1993:LDR

[RA93] Sellem Remadi and Yasuo Amemiya. Limiting distribution of roots with differential rates of convergence. Statistics & Probability Letters, 17(3):237–244, June 18, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390173G.

Rackauskas:1995:RCM

[Rac95] Alfredas Rackauskas. On the rate of convergence in the martingale CLT. Statistics & Probability Letters, 23(3):221–226, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400116P.

Radulovic:1996:BMS

[Rad96] Dragan Radulović. The bootstrap of the mean for strong mixing sequences under minimal conditions. Statistics & Probability Letters, 28(1):65–72, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000852.

See erratum [Rad99].

Radulovic:1999:EBM

[Rad99] Dragan Radulović. Erratum: “The bootstrap of the mean for strong mixing sequences under minimal conditions” [Statist. Probab. Lett. 28 (1996) 65–72]. Statistics & Probability Letters, 44(2):209, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299001157.

See [Rad96].

Ralescu:1992:REN

[Ral92] Stefan S. Ralescu. A remainder estimate for the normal approximation of perturbed sample quantiles. Statistics & Probability Letters, 14(4):293–298, July 17, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290060I.

Ramachandran:1996:CFT

[Ram96] B. Ramachandran. Characteristic functions taking constant values on intervals of the real line. Statistics & Probability Letters, 28(3):269–270, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001344.

Ramachandran:1997:CFS

[Ram97] B. Ramachandran. Characteristic functions with some powers real — III. Statistics & Probability Letters, 34(1):33–36, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001629.

Ramponi:1999:NCR

[Ram99] A. Ramponi. A note on the complex roots of complex random polynomials. Statistics & Probability Letters, 44(2):181–187, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000073.

Rao:1990:RUE

[Rao90] B. L. S. Prakasa Rao. Remarks on univariate elliptical distributions. Statistics & Probability Letters, 10(4):307–315, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090047B.

Rao:1992:NEG

[Rao92] B. L. S. Prakasa Rao. Nonparametric estimation for Galton– Watson type process. Statistics & Probability Letters, 13(4):287– 293, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900365.

Rao:1995:CED

[Rao95] B. L. S. Prakasa Rao. Consistent estimation of densityweighted average derivative by orthogonal series method.

Statistics & Probability Letters, 22(3):205–212, February 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400068J.

Rao:1999:BOJ

[Rao99a] B. L. S. Prakasa Rao. Bounds for r th order joint cumulant under r th order strong mixing. Statistics & Probability Letters, 43(4):427–431, July 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002867.

Rao:1999:CIE

[Rao99b] B. L. S. Prakasa Rao. Covariance identities for exponential and related distributions. Statistics & Probability Letters, 42(3):305– 311, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002223.

Raqab:1997:BBG

[Raq97] Mohammad Z. Raqab. Bounds based on greatest convex minorants for moments of record values. Statistics & Probability Letters, 36(1):35–41, November 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000461.

Rashid:1995:ARM

[Ras95] M. Mushfiqur Rashid. Analysis of repeated measures incomplete block designs using R-estimators. Statistics & Probability Letters, 22(3):213–221, February 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400069K.

Ravi:1992:NWC

[Rav92] S. Ravi. A note on weak convergence to extremal processes. Statistics & Probability Letters, 13(4):301–306, March 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900387.

Rayner:1996:SEP

[RB96] J. C. W. Rayner and D. J. Best. Smooth extensions of Pearsons’s product moment correlation and Spearman’s rho. Statistics & Probability Letters, 30(2):171–177, October 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002162.

Rousseeuw:1994:BSE

[RC94] Peter J. Rousseeuw and Christophe Croux. The bias of k-step M-estimators. Statistics & Probability Letters, 20(5):411–420, August 9, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901333.

Rebollo:1996:EUS

[RCBG96] J. L. Moreno Rebollo, I. Barranco Chamorro, F. López Blázquez, and T. Gómez Gómez. On the estimation of the unknown sample size from the number of records. Statistics & Probability Letters, 31(1):7–12, December 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000065.

Robert:1993:BEH

[RCD93] Christian P. Robert, Gilles Celeux, and Jean Diebolt. Bayesian estimation of hidden Markov chains: a stochastic implementation. Statistics & Probability Letters, 16(1):77–83, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901275.

Read:1993:FTC

[Rea93] Campbell B. Read. Freeman–Tukey chi-squared goodness-of-fit statistics. Statistics & Probability Letters, 18(4):271–278, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390015B.

Recoules:1991:GRP

[Rec91] Raymond Recoules. Gaussian reciprocal processes revisited. Statistics & Probability Letters, 12(4):297–303, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900959.

Reinert:1992:TTG

[Rei92] Gesine Reinert. A threshold theorem for the general stochastic epidemic via a discrete approach. Statistics & Probability Letters, 14(2):85–90, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290068G.

Rempala:1998:SLL

[Rem98] Grzegorz Rempala. Strong Law of Large Numbers for U-statistics of varying order. Statistics & Probability Letters, 39(3):263–270, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000650.

Reschenhofer:1996:ABF

[Res96a] Erhard Reschenhofer. Approximating the Bayes factor. Statistics & Probability Letters, 30(3):241–245, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500226X.

Reschenhofer:1996:ABA

[Res96b] Erhard Reschenhofer. On the asymptotic behavior of Akaike’s BIC. Statistics & Probability Letters, 27(2):171–175, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000607.

Revesz:1996:SDO

[Rév96] P. Révész. On the shape of the domain occupied by a supercritical branching random walk. Statistics & Probability Letters, 30(4):295–303, November 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715295002324.

Rahbar:1995:NER

[RG95] Mohammad H. Rahbar and Joseph C. Gardiner. Nonparametric estimation of regression parameters from censored data with a discrete covariate. Statistics & Probability Letters, 24(1):13–20, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400143V.

Roy:1999:CMS

[RG99] Dilip Roy and R. P. Gupta. Characterizations and model selections through reliability measures in the discrete case. Statistics & Probability Letters, 43(2):197–206, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002600.

Rojo:1991:NPC

[RH91] Javier Rojo and Guo Zhong He. New properties and characterizations of the dispersive ordering. Statistics & Probability Letters, 11(4):365–372, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715291900502.

Rhee:1999:NPR

[Rhe99] WanSoo T. Rhee. A note on packing random intervals with varying density. Statistics & Probability Letters, 41(2):199–208, January 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001461.

Rosenberger:1999:ISD

[RHO99] William F. Rosenberger and Jacqueline M. Hughes-Oliver. Inference from a sequential design: proof of a conjecture by Ford and Silvey. Statistics & Probability Letters, 44(2):177–180, August 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000061.

Rieders:1993:SAS

[Rie93] Eric Rieders. The size of the averages of strongly mixing random variables. Statistics & Probability Letters, 18(1):57–64, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900995.

Riedel:1997:IAZ

[Rie97] Kurt S. Riedel. Improved asymptotics for zeros of kernel estimates via a reformulation of the Leadbetter–Cryer integral. Statistics & Probability Letters, 32(4):351–356, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000946.

Rigdon:1995:DIE

[Rig95] Steven E. Rigdon. A double-integral equation for the average run length of a multivariate exponentially weighted moving average control chart. Statistics & Probability Letters, 24(4):365–373, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400196F.

Reiser:1991:CBO

[RJ91] Benjamin Reiser and Mordechai Jaegar. A comment on Buehler optimal confidence bounds for series systems reliability. Statistics & Probability Letters, 11(1):65–67, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190180Y.

Rukhin:1990:CME

[RKD90] Andrew L. Rukhin, Lynn Kuo, and Dipak K. Dey. A class of minimax estimators of the scale parameter of the uniform distribution. Statistics & Probability Letters, 9(4):317–321, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090139X.

Rayner:1990:SGF

[RM90] J. C. W. Rayner and L. G. McAlevey. Smooth goodness of fit tests for categorised composite null hypotheses. Statistics & Probability Letters, 9(5):423–429, May 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900345.

Ruiz-Molina:1999:DMA

[RMNV99] Juan Carlos Ruiz-Molina, Jesús Navarro, and J. Mariano Valderrama. Differentiation of the modified approximative Karhunen–Loève expansion of a stochastic process. Statistics & Probability Letters, 42(1):91–98, March 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002090.

Ruiz-Molina:1996:DSF

[RMV96] Juan Carlos Ruiz-Molina and Mariano J. Valderrama. On the derivation of a suboptimal filter for signal estimation. Statistics & Probability Letters, 28(3):239–243, July 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001301.

Robert:1990:MBF

[Rob90a] Christian Robert. Modified Bessel functions and their applications in probability and statistics. Statistics & Probability Letters, 9(2):155–161, February 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290011S.

Robert:1990:SAB

[Rob90b] Christian Robert. On some accurate bounds for the quantiles of a non-central chi squared distribution. Statistics & Probability Letters, 10(2):101–106, July 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090003P.

Robert:1991:GIN

[Rob91] Christian Robert. Generalized inverse normal distributions. Statistics & Probability Letters, 11(1):37–41, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190174P.

Robinson:1995:ADN

[Rob95a] P. M. Robinson. The approximate distribution of nonparametric regression estimates. Statistics & Probability Letters, 23(2):193– 201, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400113M.

Robinson:1995:IPO

[Rob95b] P. M. Robinson. An invariance property of optimal spectral bandwidths. Statistics & Probability Letters, 24(4):345–346, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400193C.

Rodrigues:1993:ADU

[Rod93] Josemar Rodrigues. Asymptotically design-unbiased predictors for two-stage sampling. Statistics & Probability Letters, 17(5):411–414, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390263I.

Roginsky:1995:AEC

[Rog95] Allen Roginsky. Average error in the Central Limit Theorem for the cumulative processes. Statistics & Probability Letters, 24(3):199–204, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294001714.

Rojo:1993:PSP

[Roj93] Javier Rojo. On the preservation of some pure-tail orderings by reliability operations. Statistics & Probability Letters, 17(3):189– 198, June 18, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390166G.

Romik:1999:SEB

[Rom99] Dan Romik. Sharp entropy bounds for discrete statistical simulation. Statistics & Probability Letters, 42(3):219–227, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001746.

Roos:1995:SAP

[Roo95] Bero Roos. A semigroup approach to Poisson approximation with respect to the point metric. Statistics & Probability Letters, 24(4):305–314, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400188E.

Rosalsky:1993:ACL

[Ros93] Andrew Rosalsky. On the almost certain limiting behavior of normed sums of identically distributed positive random variables. Statistics & Probability Letters, 16(1):65–70, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901253.

Roters:1993:RPS

[Rot93a] M. Roters. On a ranking problem for symmetric and unimodal location parameter families. Statistics & Probability Letters, 16(1):47–49, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390121X.

Roters:1993:CRP

[Rot93b] Markus Roters. Convergence of random power series with pairwise independent Banach-space-valued coefficients. Statistics & Probability Letters, 18(2):121–123, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390179M.

Rothmann:1997:SMR

[Rot97] Mark D. Rothmann. Stability of maxima over randomly deleted sets. Statistics & Probability Letters, 36(2):145–152, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000576.

Roussas:1991:KEU

[Rou91a] George G. Roussas. Kernel estimates under association: strong uniform consistency. Statistics & Probability Letters, 12(5):393– 403, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190028P.

Roussas:1991:RET

[Rou91b] George G. Roussas. Recursive estimation of the transition distribution function of a Markov process: a symptotic normality. Statistics & Probability Letters, 11(5):435–447, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190193U.

Rousseeuw:1993:RDC

[Rou93] Peter J. Rousseeuw. A resampling design for computing highbreakdown regression. Statistics & Probability Letters, 18(2): 125–128, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390180Q.

Rousseeuw:1994:UFP

[Rou94] Peter J. Rousseeuw. Unconventional features of positivebreakdown estimators. Statistics & Probability Letters, 19 (5):417–431, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900108.

Roussas:1995:ANS

[Rou95] George G. Roussas. Asymptotic normality of a smooth estimate of a random field distribution function under association. Statistics & Probability Letters, 24(1):77–90, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400151W.

Rousseau:1997:ABR

[Rou97] Judith Rousseau. Asymptotic Bayes risks for a general class of losses. Statistics & Probability Letters, 35(2):115–121, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000047.

Ramos:1995:ENR

[RP95] Rogelio Q. Ramos and Sastry G. Pantula. Estimation of nonlinear random coefficient models. Statistics & Probability Letters, 24 (1):49–56, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400147Z.

Ravishanker:1999:MCE

[RQ99] Nalini Ravishanker and Zuqiang Qiou. Monte Carlo EM estimation for multivariate stable distributions. Statistics & Probability Letters, 45(4):335–340, December 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000759.

Rachev:1990:CC

[RR90] S. T. Rachev and L. Rüschendorf. A counterexample to A.S. constructions. Statistics & Probability Letters, 9(4):307–309, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090137V.

Rothmann:1991:SLB

[RR91] Mark. D. Rothmann and Ralph P. Russo. Strong limiting bounds for a sequence of moving maxima. Statistics & Probability Letters, 11(5):403–410, May 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190189X.

Ramanathan:1994:RTT

[RR94a] R. V. Ramanathan and M. B. Rajarshi. Rank tests for testing the randomness of autoregressive coefficients. Statistics & Probability Letters, 21(2):115–120, September 23, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902186.

Rothmann:1994:PCR

[RR94b] Mark D. Rothmann and Ralph P. Russo. Persistent convergence on randomly deleted sets. Statistics & Probability Letters, 20(5):367–373, August 9, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294901287.

Rojo:1991:NML

[RS91] Javier Rojo and Francisco J. Samaniego. On nonparametric maximum likelihood estimation of a distribution uniformly stochastically smaller than a standard. Statistics & Probability Letters, 11(3):267–271, March 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190154J.

Rosalsky:1998:LBR

[RS98] Andrew Rosalsky and M. Sreehari. On the limiting behavior of randomly weighted partial sums. Statistics & Probability Letters, 40(4):403–410, November 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001539.

Rocha:1999:GRP

[RS99] Amy L. Rocha and Frederick Stern. The gambler’s ruin problem with n players and asymmetric play. Statistics & Probability Letters, 44(1):87–95, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002958.

Rinott:1994:CBV

[RSC94] Yosef Rinott and Ester Samuel-Cahn. Covariance between variables and their order statistics for multivariate normal variables. Statistics & Probability Letters, 21(2):153–155, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902232.

Ruschendorf:1993:NSE

[RT93] L. Rüschendorf and W. Thomsen. Note on the Schrödinger equation and I-projections. Statistics & Probability Letters, 17(5):369–375, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390257J.

Rukhin:1996:RCB

[Ruk96] Andrew L. Rukhin. The rates of convergence of Bayes estimators in change-point analysis. Statistics & Probability Letters, 27(4):319–329, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000933.

Runde:1993:NAD

[Run93] Ralf Runde. A note on the asymptotic distribution of the F - statistic for random variables with infinite variance. Statistics & Probability Letters, 18(1):9–12, August 30, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390092W.

Ruschendorf:1996:ORV

[Rüs96] Ludger Rüschendorf. On c-optimal random variables. Statistics & Probability Letters, 27(3):267–270, April 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529500078X.

Rutkowski:1990:SDE

[Rut90] Marek Rutkowski. Stochastic differential equations with singular drift. Statistics & Probability Letters, 10(3):225–229, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090078L.

Rutkowski:1993:SPK

[Rut93] Marek Rutkowski. A simple proof for the Kalman–Bucy smoothed estimate formula. Statistics & Probability Letters, 17(5):377–385, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390258K.

Rutkowski:1995:LRL

[Rut95] Marek Rutkowski. Left and right linear innovations for a multivariate SαS random variable. Statistics & Probability Letters, 22(3):175–184, February 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400065G.

Ruymgaart:1996:NIE

[Ruy96] Frits H. Ruymgaart. A note on the inverse estimation of bandlimited signals. Statistics & Probability Letters, 27(3):241–246, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000712.

Rosenblatt:1992:NMI

[RW92] Murray Rosenblatt and Bruce E. Wahlen. A nonparametric measure of independence under a hypothesis of independent components. Statistics & Probability Letters, 15(3):245–252, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290197D.

Rojo:1994:TBS

[RW94] Javier Rojo and Jinping Wang. Tests based on L-statistics to test the equality in dispersion of two probability distributions. Statistics & Probability Letters, 21(2):107–113, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902178.

Rempala:1999:LBR

[RW99] Grzegorz Rempala and Jacek Wesołowski. Limiting behavior of random permanents. Statistics & Probability Letters, 45(2):149–158, November 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000541.

Rychlik:1990:UNE

[Ryc90] Tomasz Rychlik. Unbiased nonparametric estimation of the derivative of the mean. Statistics & Probability Letters, 10 (4):329–333, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715290900518.

Rychlik:1992:SED

[Ryc92] Tomasz Rychlik. Stochastically extremal distributions of order statistics for dependent samples. Statistics & Probability Letters, 13(5):337–341, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290105E.

Rychlik:1995:BOS

[Ryc95] Tomasz Rychlik. Bounds for order statistics based on dependent variables with given nonidentical distributions. Statistics & Probability Letters, 23(4):351–358, June 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400135U.

Rao:1992:CEL

[RZ92a] C. Radhakrishna Rao and L. C. Zhao. On the consistency of M-estimate in a linear model obtained through an estimating equation. Statistics & Probability Letters, 14(1):79–84, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290214P.

Rychlik:1992:SIP

[RZ92b] Z. Rychlik and J. Zygo. Strassen’s invariance principle for random subsequences. Statistics & Probability Letters, 13 (3):193–197, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292900479.

Rao:1994:BEB

[RZ94] C. R. Rao and L. C. Zhao. Berry–Esseen bounds for finitepopulation t-statistics. Statistics & Probability Letters, 21 (5):409–416, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294000417.

Sadooghi-Alvandi:1993:PCT

[SA93] S. M. Sadooghi-Alvandi. A proof of the continuity theorem for characteristic functions. Statistics & Probability Letters, 16(1):27–28, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715293901172.

Shin:1997:ALB

[SA97] Chungyeol Shin and Yasuo Amemiya. Algorithms for the likelihood-based estimation of the random coefficient model. Statistics & Probability Letters, 32(2):189–199, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000727.

Sahoo:1997:NUL

[Sah97] J. Sahoo. A note on using location shift in the Midzuno– Sen scheme of sampling with a transformed variable. Statistics & Probability Letters, 33(3):285–290, May 5, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001393.

Saikkonen:1995:DVC

[Sai95] Pentti Saikkonen. Dependent versions of a Central Limit Theorem for the squared length of a sample mean. Statistics & Probability Letters, 22(3):185–194, February 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400066H.

Saisho:1999:LTR

[Sai99] Yasumasa Saisho. Limit theorems and random spacings related to cutting of DNAs by radiation. Statistics & Probability Letters, 43(4):361–367, July 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002740.

Salminen:1993:DSD

[Sal93] Paavo Salminen. On the distribution of supremum of diffusion local time. Statistics & Probability Letters, 18(3):219–225, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390220D.

Samorodnitsky:1994:PSP

[Sam94] Gennady Samorodnitsky. Possible sample paths of self-similar αstable processes. Statistics & Probability Letters, 19(3):233–237, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901090.

Sanz:1991:CCO

[San91] Gerardo Sanz. n^{r}-Consistency of certain optimal estimators, 0 < r < 1/2. Statistics & Probability Letters, 11(1):71–76, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190182Q.

Saner:1992:ACI

[San92] Hilary Saner. Approximations to a conservative inverse chisquare procedure for combining p-values in integrative research.

Statistics & Probability Letters, 15(3):215–222, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292901928.

Sapatinas:1995:CPD

[Sap95] Theofanis Sapatinas. Characterizations of probability distributions based on discrete p-monotonicity. Statistics & Probability Letters, 24(4):339–344, September 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400192B.

Sapatinas:1999:CNB

[Sap99] Theofanis Sapatinas. A characterization of the negative binomial distribution via α-monotonicity. Statistics & Probability Letters, 45(1):49–53, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715299000413.

Sarkar:1996:MSE

[Sar96] Nityananda Sarkar. Mean square error matrix comparison of some estimators in linear regressions with multicollinearity. Statistics & Probability Letters, 30(2):133–138, October 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002111.

Satorra:1992:VMS

[Sat92] Albert Satorra. The variance matrix of sample secondorder moments in multivariate linear relations. Statistics & Probability Letters, 15(1):63–69, September 3, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290286E.

Savits:1994:ISP

[Sav94] Thomas H. Savits. On integration, substitution and the probability integral transform. Statistics & Probability Letters, 21 (3):173–179, October 19, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294901139.

Sathe:1991:LCP

[SB91] Y. S. Sathe and S. M. Bendre. Log-concavity of probability of occurrence of at least r independent events. Statistics & Probability Letters, 11(1):63–64, January 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190179U.

Shetty:1994:NGM

[SB94] I. D. Shetty and Sharada V. Bhat. A note on the generalization of Mathisen’s median test. Statistics & Probability Letters, 19(3):199–204, February 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901058.

Serinko:1995:AMC

[SB95a] Regis J. Serinko and Gutti Jogesh Babu. Asymptotics of k-mean clustering under non-i.i.d. sampling. Statistics & Probability Letters, 24(1):57–66, July 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001482.

Sun:1995:CBG

[SB95b] Kai Sun and Asit P. Basu. A characterization of a bivariate geometric distribution. Statistics & Probability Letters, 23(4): 307–311, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400129V.

Sajjan:1996:EBP

[SB96] S. G. Sajjan and I. V. Basawa. Empirical Bayes prediction for a mixed linear model with autoregressive errors. Statistics & Probability Letters, 29(1):1–7, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001492.

Sajjan:1997:EUC

[SB97a] S. G. Sajjan and I. V. Basawa. Estimation of unobserved counts from partially observed multinomial distributions. Statistics & Probability Letters, 34(3):237–243, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001861.

Sethuraman:1997:ADM

[SB97b] S. Sethuraman and I. V. Basawa. The asymptotic distribution of the maximum likelihood estimator for a vector time series model with long memory dependence. Statistics & Probability Letters, 31(4):285–293, February 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000417.

Sen:1999:CGR

[SB99] Ananda Sen and N. Balakrishnan. Convolution of geometrics and a reliability problem. Statistics & Probability Letters, 43(4):421– 426, July 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002843.

Sivaganesan:1993:ORC

[SBB93] Siva Sivaganesan, L. Mark Berliner, and James Berger. Optimal robust credible sets for contaminated priors. Statistics & Probability Letters, 18(5):383–388, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390032E.

Song:1992:MVS

[SBD92] Ruiguang Song, Steven G. Buchberger, and James A. Deddens. Moments of variables summing to normal order statistics. Statistics & Probability Letters, 15(3):203–208, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290190G.

Shapiro:1990:TEE

[SC90] Alexander Shapiro and Ayala Cohen. Testing and estimation of equal variances for correlated variables. Statistics & Probability Letters, 10(3):231–234, August 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090079M.

Salome:1991:LMP

[SC91a] M. Salomé and E. Cabral. A locally most powerful unbiased test for the proportion of mixture. Statistics & Probability Letters, 12(2):119–124, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190054U.

Samuel-Cahn:1991:PIB

[SC91b] Ester Samuel-Cahn. Prophet inequalities for bounded negatively dependent random variables. Statistics & Probability Letters, 12(3):213–216, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190080B.

Samuel-Cahn:1993:OSU

[SC93] Ester Samuel-Cahn. Optimal stopping in urn models with payoff depending on maximal observed element. Statistics & Probability Letters, 16(5):361–368, April 8, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390070Y.

Sarkar:1999:ASG

[SC99] Jyotirmoy Sarkar and Gopal Chaudhuri. Availability of a system with gamma life and exponential repair time under a perfect repair policy. Statistics & Probability Letters, 43(2):189–196, June 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002594.

Schuster:1991:DTR

[Sch91a] Eugene F. Schuster. Distribution theory of runs via exchangeable random variables. Statistics & Probability Letters, 11(5):379–386, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190185T.

Schwarz:1991:OPP

[Sch91b] Gideon Schwarz. An optimality property of polynomial regression. Statistics & Probability Letters, 12(4):335–336, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190102W.

Schick:1994:EAC

[Sch94] Anton Schick. Estimation of the autocorrelation coefficient in the presence of a regression trend. Statistics & Probability Letters, 21(5):371–380, December 7, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000344.

Scheffler:1995:MMA

[Sch95] Hans-Peter Scheffler. Moments of measures attracted to operator semi-stable laws. Statistics & Probability Letters, 24(3):187–192, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294001688.

Schick:1996:RCEb

[Sch96a] Anton Schick. Root-n-consistent and efficient estimation in semiparametric additive regression models. Statistics & Probability Letters, 30(1):45–51, September 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002006.

Schick:1996:RCEa

[Sch96b] Anton Schick. Root-n consistent estimation in partly linear regression models. Statistics & Probability Letters, 28(4):353– 358, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529500145X.

Schick:1996:WLS

[Sch96c] Anton Schick. Weighted least squares estimates in partly linear regression models. Statistics & Probability Letters, 27(3):281– 287, April 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000860.

Schott:1996:TES

[Sch96d] James R. Schott. Testing for the equality of several correlation matrices. Statistics & Probability Letters, 27(1):85–89, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000488.

Schick:1997:SRK

[Sch97a] Anton Schick. On U-statistics with random kernels. Statistics & Probability Letters, 34(3):275–283, June 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001915.

Schott:1997:TES

[Sch97b] James R. Schott. Testing for the equality of several correlation matrices. Statistics & Probability Letters, 36(3):319, December 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529781471X.

Schwabe:1997:MED

[Sch97c] Rainer Schwabe. Maximin efficient designs another view at D-optimality. Statistics & Probability Letters, 35(2):109–114, September 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000035.

Schatte:1998:BLV

[Sch98a] P. Schatte. On Benford’s law to variable base. Statistics & Probability Letters, 37(4):391–397, March 30, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001429.

Schied:1998:CCS

[Sch98b] Alexander Schied. Cramer’s condition and Sanov’s theorem. Statistics & Probability Letters, 39(1):55–60, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000339.

Scheffler:1999:ESM

[Sch99a] Hans-Peter Scheffler. On estimation of the spectral measure of certain nonnormal operator stable laws. Statistics & Probability Letters, 43(4):385–392, July 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002776.

Schick:1999:EES

[Sch99b] Anton Schick. Efficient estimation of a shift in nonparametric regression. Statistics & Probability Letters, 41(3):287–301, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001655.

Schweizer:1999:MPM

[Sch99c] Martin Schweizer. A minimality property of the minimal martingale measure. Statistics & Probability Letters, 42(1):27–31, March 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001801.

Scozzafava:1993:UDS

[Sco93] Paola Scozzafava. Uniform distribution and sum modulo m of independent random variables. Statistics & Probability Letters, 18(4):313–314, November 9, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390021A.

Sathe:1990:RRO

[SD90] Y. S. Sathe and U. J. Dixit. On a recurrence relation for order statistics. Statistics & Probability Letters, 9(1):1–4, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090087N.

Song:1993:NMV

[SD93] Ruiguang Song and James A. Deddens. A note on moments of variables summing to normal order statistics. Statistics & Probability Letters, 17(5):337–341, August 5, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390252E.

Satten:1999:KMR

[SD99] Glen A. Satten and Somnath Datta. Kaplan–Meier representation of competing risk estimates. Statistics & Probability Letters, 42(3):299–304, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800220X.

Sa:1997:FPB

[SE97] Ping Sa and Don Edwards. Frequentist properties of a Bayesian analog to Fabian’s bound. Statistics & Probability Letters, 32(4):417–424, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001022.

Semenou:1996:QEU

[Sém96] M. Séménou. Quantile estimation under possibly misspecified generalised linear model. Statistics & Probability Letters, 27(4):357–365, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529500100X.

Seneta:1993:SFM

[Sen93] E. Seneta. Sensitivity of finite Markov chains under perturbation. Statistics & Probability Letters, 17(2):163–168, May 26, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293900117.

Sepanski:1992:STG

[Sep92] J. H. Sepanski. Score tests in a generalized linear model with surrogate covariates. Statistics & Probability Letters, 15(1):1–10, September 3, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290277C.

Sepanski:1994:NSC

[Sep94] Steven J. Sepanski. Necessary and sufficient conditions for the multivariate bootstrap of the mean. Statistics & Probability Letters, 19(3):205–216, February 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901066.

Sepanski:1996:AMS

[Sep96] Steven J. Sepanski. Asymptotics for multivariate t-statistic for random vectors in the generalized domain of attraction of the multivariate normal law. Statistics & Probability Letters, 30(2):179–188, October 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002170.

Serbinowska:1996:CEN

[Ser96a] Monika Serbinowska. Consistency of an estimator of the number of changes in binomial observations. Statistics & Probability Letters, 29(4):337–344, September 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001891.

Sergeevich:1996:AMK

[Ser96b]  Ermakov Mikhail Sergeevich.  On asymptotic minimax-

ity of Kolmogorov and omega-square tests. Statistics & Probability Letters, 30(3):227–233, October 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002243.

Seshadri:1991:FDR

[Ses91] V. Seshadri. A family of distributions related to the McCullagh family. Statistics & Probability Letters, 12(5):373–378, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190025M.

Seshadri:1992:GEM

[Ses92] V. Seshadri. General exponential models on the unit simplex and related multivariate inverse Gaussian distributions. Statistics & Probability Letters, 14(5):385–391, July 11, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290099Q.

Severini:1993:NSC

[Sev93] Thomas A. Severini. A note on sufficiency and conditional inference. Statistics & Probability Letters, 17(4):303–305, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390207Y.

Severini:1998:SPI

[Sev98] Thomas A. Severini. Some properties of inferences in misspecified linear models. Statistics & Probability Letters, 40(2):149–153, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000820.

Sobel:1994:DTS

[SF94] Milton Sobel and Krzysztof Frankowski. A duality theorem for solving multiple-player multivariate hypergeometric problems. Statistics & Probability Letters, 20(2):155–162, May 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900329.

Soofi:1991:ICN

[SG91] Ehsan S. Soofi and D. V. Gokhale. An information criterion for normal regression estimation. Statistics & Probability Letters, 11(2):111–117, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190127D.

Srinivasan:1996:SEI

[SG96a] M. R. Srinivasan and G. Gopal. Structural efficiency of incomplete block designs. Statistics & Probability Letters, 30 (3):211–214, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295002219.

Srivastava:1996:NNA

[SG96b] J. N. Srivastava and S. Ghosh. On nonorthogonality and nonoptimality of Addelman’s main-effect plans satisfying the condition of proportional frequencies. Statistics & Probability Letters, 26(1): 51–60, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294002517.

Sgibnev:1996:DMP

[Sgi96] M. S. Sgibnev. On the distribution of the maxima of partial sums. Statistics & Probability Letters, 28(3):235–238, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001298.

Sgibnev:1997:SMS

[Sgi97] M. S. Sgibnev. Submultiplicative moments of the supremum of a random walk with negative drift. Statistics & Probability Letters, 32(4):377–383, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000971.

Sgibnev:1998:ETC

[Sgi98a] M. S. Sgibnev. Equivalence of two conditions on singular components. Statistics & Probability Letters, 40(2):127–131, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000741.

Sgibnev:1998:EFD

[Sgi98b] M. S. Sgibnev. On an explicit formula for the distribution of the supremum. Statistics & Probability Letters, 40(4):329–331, November 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298001229.

Shao:1995:LTL

[SH95] Yongzhao Shao and Marjorie G. Hahn. Limit theorems for the logarithm of sample spacings. Statistics & Probability Letters, 24(2):121–132, August 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294001563.

Skold:1999:AVC

[SH99] Martin Sköld and Ola Hössjer. On the asymptotic variance of the continuous-time kernel density estimator. Statistics & Probability Letters, 44(1):97–106, August 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002971.

Shao:1990:ATH

[Sha90] Jun Shao. Asymptotic theory in heteroscedastic nonlinear models. Statistics & Probability Letters, 10(1):77–85, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090115N.

Shao:1991:JVE

[Sha91] Jun Shao. Jackknife variance estimators for generalized Lstatistics. Statistics & Probability Letters, 11(1):27–32, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190172N.

Shanmugam:1992:FMN

[Sha92a] Ramalingam Shanmugam. Factions in Morris’ natural exponential quadratic variance family and a lack of consensus in sample size determination. Statistics & Probability Letters, 15(2):117–120, September 28, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290122L.

Shao:1992:BVE

[Sha92b] Jun Shao. Bootstrap variance estimators with truncation. Statistics & Probability Letters, 15(2):95–101, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290119P.

Shao:1993:CCM

[Sha93] Qi-Man Shao. Complete convergence for α-mixing sequences. Statistics & Probability Letters, 16(4):279–287, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901312.

Shiau:1999:NBE

[SHC99] Jyh-Jen Horng Shiau, Hui-Nien Hung, and Chun-Ta Chiang. A note on Bayesian estimation of process capability indices. Statistics & Probability Letters, 45(3):215–224, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000619.

Shen:1993:CES

[She93] Larry Zaiqian Shen. On Chiang’s explicit solutions for the Kolmogorov differential equations. Statistics & Probability Letters, 17(4):307–313, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390208Z.

Sherman:1994:KED

[She94] Michael Sherman. Kernel estimation of the density of a statistic. Statistics & Probability Letters, 21(1):29–36, September 1, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900558.

Shen:1998:PAJ

[She98] Pao-Sheng Shen. Problems arising from jackknifing the estimate of a Kaplan–Meier integral. Statistics & Probability Letters, 40(4):353–361, November 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001357.

Shih:1993:RTD

[Shi93] Jian-Qing Shih. Regression transformation diagnostics in transform-both-sides model. Statistics & Probability Letters, 16(5):411–420, April 8, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390076U.

Shieh:1994:IOS

[Shi94] Grace S. Shieh. Infinite order V -statistics. Statistics & Probability Letters, 20(1):75–80, May 4, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294902372.

Shi:1995:LLI

[Shi95a] Z. Shi. Lower limits of iterated Wiener processes. Statistics & Probability Letters, 23(3):259–270, May 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400123P.

Shieh:1995:SSS

[Shi95b] Narn-Rueih Shieh. Some self-similar processes related to local times. Statistics & Probability Letters, 24(3):213–218, August 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001736.

Shi:1997:LME

[Shi97a] Z. Shi. Locating the maximum of an empirical process. Statistics & Probability Letters, 31(3):199–211, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000326.

Shixin:1997:HRI

[Shi97b] Gan Shixin. The Hájek–Rènyi inequality for Banach space valued martingales and the p smoothness of Banach spaces. Statistics & Probability Letters, 32(3):245–248, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000806.

Shieh:1998:WKT

[Shi98] Grace S. Shieh. A weighted Kendall’s tau statistic. Statistics & Probability Letters, 39(1):17–24, July 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000066.

Shiau:1999:NAB

[Shi99] Jyh-Jen Horng Shiau. A new algorithm for 5-band Toeplitz matrix inversion with application to GCV smoothing spline computation. Statistics & Probability Letters, 45(4):317–324, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000735.

Shoji:1997:NAP

[Sho97] Isao Shoji. A note on asymptotic properties of the estimator derived from the Euler method for diffusion processes at discrete times. Statistics & Probability Letters, 36(2):153–159, December 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000588.

Siatkowski:1994:SBB

[Sia94] Idzi Siatkowski. Some bounds for balanced two-way elimination of heterogeneity designs. Statistics & Probability Letters, 21(2):95–99, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294902151.

Sibuya:1991:SPB

[Sib91] Masaaki Sibuya. A simplified proof of a Bonferroni-type inequality by Tan and Xu and its extension. Statistics & Probability Letters, 11(3):255–257, March 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190152H.

Silvapulle:1996:TST

[Sil96a] Mervyn J. Silvapulle. On an F -type statistic for testing one-sided hypotheses and computation of chi-bar-squared weights. Statistics & Probability Letters, 28(2):137–141, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001085.

Silvapulle:1996:RBI

[Sil96b] Mervyn J. Silvapulle. Robust bounded influence tests against one-sided hypotheses in general parametric models. Statistics & Probability Letters, 31(1):45–50, December 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000120.

Silvapulle:1997:ORI

[Sil97] Mervyn J. Silvapulle. On order restricted inference in some mixed linear models. Statistics & Probability Letters, 36(1):23–27, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000448.

Sim:1992:PPC

[Sim92] C. H. Sim. Point processes with correlated gamma interarrival times. Statistics & Probability Letters, 15(2):135–141, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290126P.

Singh:1994:ERR

[Sin94] R. Karan Singh. Estimation of restricted regression model when disturbances are not necessarily normal. Statistics & Probability Letters, 19(2):101–109, January 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901406.

Sivaganesan:1999:LBR

[Siv99] Siva Sivaganesan. A likelihood based robust Bayesian summary. Statistics & Probability Letters, 43(1):5–12, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002065.

Sen:1997:CQP

[SJ97] Kanwar Sen and Ritu Jain. A class of quasi-Pólya distributions. Statistics & Probability Letters, 31(4):313–322, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000454.

Sim:1998:CSC

[SJ98] Songyong Sim and Richard A. Johnson. Comparisons of spatially correlated binary data. Statistics & Probability Letters, 39(2):81– 87, August 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800008X.

Sinha:1996:SNN

[SJK96] K. Sinha, B. Jones, and Sanpei Kageyama. Some new non-proper variance-balanced designs with unequal replications. Statistics & Probability Letters, 27(2):149–153, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000577.

Sarkar:1993:IAN

[SK93] A. Sarkar and B. Kartikeyan. Intervention analysis with nonlinear dependent noise variation. Statistics & Probability Letters, 18(2):91–103, September 22, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390176J.

Skovoroda:1992:SLL

[SM92] Boris Skovoroda and Thomas Mikosch. A Strong Law of Large Numbers for ruled sums. Statistics & Probability Letters, 13(2):129–137, January 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290087L.

Singh:1995:GCE

[SMP95] R. Karan Singh, Sheela Misra, and S. K. Pandey. A generalized class of estimators in linear regression models with multivariatet distributed error. Statistics & Probability Letters, 23(2):171– 178, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400110T.

Sheng:1996:MFR

[SN96] Ke-Ning Sheng and Joseph I. Naus. Matching fixed rectangles in 2-dimension. Statistics & Probability Letters, 26(1):83–90, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400255X.

Swaminathan:1997:MDE

[SNN97] V. Swaminathan and U. V. Naik-Nimbalkar. Minimum distance estimation for random coefficient autoregressive models. Statistics & Probability Letters, 34(4):313–322, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001964.

Somerville:1993:CTK

[Som93] Paul N. Somerville. On the conservatism of the Tukey–Kramer multiple comparison procedure. Statistics & Probability Letters, 16(5):343–345, April 8, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390067S.

Song:1999:NLR

[Son99] Seuck Heun Song. A note on S^{2} in a linear regression model based on two-stage sampling data. Statistics & Probability Letters, 43(2):131–135, June 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002272.

Shin:1993:TUR

[SP93] Dongwan Shin and Sastry G. Pantula. Testing for a unit root in autoregressive processes with systematic but incomplete sampling. Statistics & Probability Letters, 18(3):183–190, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293902144.

Smith:1998:AFE

[SP98] Todd Smith and Shyamal D. Peddada. Analysis of fixed effects linear models under heteroscedastic errors. Statistics & Probability Letters, 37(4):399–408, March 30, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001430.

Sun:1999:TTC

[SP99] Hongguang Sun and Sastry G. Pantula. Testing for trends in correlated data. Statistics & Probability Letters, 41(1):87–95, January 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800131X.

Singer:1991:NHW

[SPH91] Julio M. Singer, Clovis A. Peres, and Carlos E. Harle. A note on the Hardy–Weinberg model in generalized ABO systems. Statistics & Probability Letters, 11(2):173–175, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190138H.

Spizzichino:1991:BAT

[Spi91] F. Spizzichino. Bayesian analysis for two-components systems and a conjugate family of bivariate densities. Statistics & Probability Letters, 11(1):17–25, January 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190171M.

Spreij:1998:RRF

[Spr98] Peter Spreij. A representation result for finite Markov chains. Statistics & Probability Letters, 38(2):183–186, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000054.

Salama:1990:APN

[SQ90] Ibrahim A. Salama and Dana Quade. On the asymptotic permutational normality of certain weighted measures of correlation. Statistics & Probability Letters, 10(2):125–133, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090007T.

Salama:1997:ANR

[SQ97] Ibrahim A. Salama and Dana Quade. The asymptotic normality of a rank correlation statistic based on rises. Statistics & Probability Letters, 32(2):201–205, March 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000739.

Shih:1994:EMW

[SQC94] Weichung J. Shih, Hui Quan, and Myron N. Chang. Estimation of the mean when data contain non-ignorable missing values from a random effects model. Statistics & Probability Letters, 19(3):249–257, February 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901120.

Stukel:1997:ERM

[SR97] D. M. Stukel and J. N. K. Rao. Estimation of regression models with nested error structure and unequal error variances under two and three stage cluster sampling. Statistics & Probability Letters, 35(4):401–407, November 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297866023.

Sreehari:1995:MSG

[Sre95] M. Sreehari. Maximum stability and a generalization. Statistics & Probability Letters, 23(4):339–342, June 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400133S.

Sriram:1991:USC

[Sri91] T. N. Sriram. On the uniform strong consistency of an estimator of the offspring mean in a branching process with immigration. Statistics & Probability Letters, 12(2):151–155, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900605.

Srivastava:1995:IPS

[Sri95] A. K. Srivastava. Improved predictors in survey sampling under a super population model. Statistics & Probability Letters, 25(3):249–252, November 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400228Z.

Srivastav:1998:CDO

[Sri98] Sudesh K. Srivastav. On the class of t-designs. Statistics & Probability Letters, 39(2):167–172, August 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000595.

Scarsini:1990:SOP

[SS90a] Marco Scarsini and Moshe Shaked. Stochastic ordering for permutation symmetric distributions. Statistics & Probability Letters, 9(3):217–222, March 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090059G.

Srivastava:1990:UMV

[SS90b] V. K. Srivastava and R. S. Singh. Uniformly minimum variance unbiased estimator of efficiency ratio in estimation of normal population mean. Statistics & Probability Letters, 10(3):241– 245, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090081H.

Singh:1992:PFL

[SS92] Bahadur Singh and Michael J. Schell. On power functions of the likelihood ratio tests for the simple loop order in normal means: Unequal sample sizes. Statistics & Probability Letters, 14(4):253– 267, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290055A.

Shin:1993:NTU

[SS93a] Dongwan Shin and Sahadeb Sarkar. A note on testing for a unit root in an ARIMA(p, 1, 0) signal observed with MA(q) noise. Statistics & Probability Letters, 18(3):195–203, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293902166.

Srivastava:1993:PCS

[SS93b] A. K. Srivastava and V. K. Srivastava. Pitman closeness for Stein-rule estimators of regression coefficients. Statistics & Probability Letters, 18(2):85–89, September 22, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390175I.

Shin:1994:URT

[SS94a] Dong Wan Shin and Sahadeb Sarkar. Unit root tests for ARIMA(0, 1, q) models with irregularly observed samples. Statistics & Probability Letters, 19(3):189–194, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901031.

Srivastava:1994:FMR

[SS94b] V. K. Srivastava and A. K. Srivastava. A feasible minimum risk estimator interpretation to Stein-rule estimators in linear regression model. Statistics & Probability Letters, 19(4):317–319, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529490183X.

Shaked:1995:HRO

[SS95] Moshe Shaked and J. George Shanthikumar. Hazard rate ordering of k-out-of-n systems. Statistics & Probability Letters, 23(1): 1–8, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400087O.

Sakalauskas:1996:DMM

[SS96a] L. Sakalauskas and D. Sukauskaitė. Distribution of MLE of the multiple correlation coefficient for data with missing values in a dependent variable. Statistics & Probability Letters, 30(1):53–59, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295002014.

Sandhya:1996:DFC

[SS96b] E. Sandhya and S. Satheesh. On distribution functions with completely monotone derivative. Statistics & Probability Letters, 27(2):127–129, April 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000534.

Shin:1997:SUR

[SS97a] Dong Wan Shin and Beong-Soo So. Semiparametric unit root tests based on symmetric estimators. Statistics & Probability Letters, 33(2):177–184, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001253.

Srivastava:1997:CEN

[SS97b] Anil K. Srivastava and Shalabh. Consistent estimation for the non-normal ultrastructural model. Statistics & Probability Letters, 34(1):67–73, May 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001678.

Srivastava:1997:NPS

[SS97c] Anil K. Srivastava and Shalabh. A new property of Stein procedure in measurement error model. Statistics & Probability Letters, 32(3):231–234, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000788.

Sugden:1997:EAD

[SS97d] R. A. Sugden and T. M. F. Smith. Edgeworth approximations to the distribution of the sample mean under simple random sampling. Statistics & Probability Letters, 34(3):293–299, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001934. See correction [SS98].

Sugden:1998:CEA

[SS98] R. A. Sugden and T. M. F. Smith. Correction to “Edgeworth approximations to the distribution of the sample mean under simple random sampling”. Statistics & Probability Letters, 37(3): 317, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001338. See [SS97d].

Shin:1999:NTU

[SS99a] Dong Wan Shin and Beong Soo So. New tests for unit roots in autoregressive processes with possibly infinite variance errors. Statistics & Probability Letters, 44(4):387–397, October 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000310.

So:1999:RMA

[SS99b] Beong Soo So and Dong Wan Shin. Recursive mean adjustment in time-series inferences. Statistics & Probability Letters, 43(1): 65–73, May 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298002478.

Shin:1996:URT

[SSL96] Dong Wan Shin, Sahadeb Sarkar, and Jong Hyup Lee. Unit root tests for time series with outliers. Statistics & Probability Letters, 30(3):189–197, October 30, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002189.

Samorodnitsky:1990:EJM

[ST90] Gennady Samorodnitsky and Murad S. Taqqu. Existence of joint moments of stable random variables. Statistics & Probability Letters, 10(2):167–172, July 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090013W.

Simonoff:1991:AII

[ST91] Jeffrey S. Simonoff and Chih-Ling Tsai. Assessing the influence of individual observations on a goodness-of-fit test based on nonparametric regression. Statistics & Probability Letters, 12(1): 9–17, July 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190159O.

Sagae:1992:SCD

[ST92a] Masahiko Sagae and Kunio Tanabe. Symbolic Cholesky decomposition of the variance-covariance matrix of the negative multinomial distribution. Statistics & Probability Letters, 15 (2):103–108, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290120T.

Stoffer:1992:NLB

[ST92b] David S. Stoffer and Clélia M. C. Toloi. A note on the Ljung– Box–Pierce portmanteau statistic with missing data. Statistics & Probability Letters, 13(5):391–396, April 7, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290112I.

Schmid:1996:VCM

[ST96] Friedrich Schmid and Mark Trede. An L_{1}-variant of the Cramer– von Mises test. Statistics & Probability Letters, 26(1):91–96, January 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002561.

Samuel:1998:MER

[ST98a] P. Samuel and P. Y. Thomas. Modified expression for a recurrence relation on the product moments of order statistics. Statistics & Probability Letters, 37(1):89–95, January 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001041.

Schmid:1998:KTT

[ST98b] Friedrich Schmid and Mark Trede. A Kolmogorov-type test for second-order stochastic dominance. Statistics & Probability Letters, 37(2):183–193, February 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001168.

Stadje:1991:TUT

[Sta91] W. Stadje. The time until two renewal processes come together. Statistics & Probability Letters, 12(3):195–200, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900775.

Stark:1992:AMC

[Sta92] Philip B. Stark. Affine minimax confidence intervals for a bounded normal mean. Statistics & Probability Letters, 13(1):39– 44, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290234V.

Stadje:1995:NMR

[Sta95a] W. Stadje. A note on the maximum of a random walk. Statistics & Probability Letters, 23(3):227–231, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400117Q.

Stadje:1995:SPA

[Sta95b] W. Stadje. A stochastic process arising in sequential experimentation. Statistics & Probability Letters, 23(4):359–365, June 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400136V.

Stadje:1997:NAL

[Sta97] Wolfgang Stadje. A new approach to the Lindley recursion. Statistics & Probability Letters, 31(3):169–175, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000296.

Stadje:1999:MON

[Sta99] Wolfgang Stadje. The maximum occupancy number in a simple urn model. Statistics & Probability Letters, 42(2):193–200, April 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002144.

Stefanski:1990:RCS

[Ste90] Leonard A. Stefanski. Rates of convergence of some estimators in a class of deconvolution problems. Statistics & Probability Letters, 9(3):229–235, March 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090061B.

Stefanski:1991:NSM

[Ste91a] Leonard A. Stefanski. A normal scale mixture representation of the logistic distribution. Statistics & Probability Letters, 11(1): 69–70, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190181P.

Stefanski:1991:NHB

[Ste91b] Leonard A. Stefanski. A note on high-breakdown estimators. Statistics & Probability Letters, 11(4):353–358, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190048V.

Stein:1993:SCA

[Ste93] Michael L. Stein. A simple condition for asymptotic optimality of linear predictions of random fields. Statistics & Probability Letters, 17(5):399–404, August 5, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390261G.

Stepniak:1997:STC

[Ste97] C. Stepniak. Sandwich theorem in comparison of multivariate normal experiments. Statistics & Probability Letters, 33(3):281– 284, May 5, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001381.

Stoyanov:1995:DMS

[Sto95] Jordan Stoyanov. Dependency measure for sets of random events or random variables. Statistics & Probability Letters, 23(1):13– 20, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400089Q.

Stoimenova:1996:SPC

[Sto96] E. Stoimenova. On statistical properties of Chebyshev’s norm. Statistics & Probability Letters, 29(1):89–93, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001603.

Strawderman:1994:NNS

[Str94] Robert L. Strawderman. A note on necessary and sufficient conditions for proving that a random symmetric matrix converges to a given limit. Statistics & Probability Letters, 21(5):367–370, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294000336.

Stryhn:1996:LMA

[Str96] Henrik Stryhn. The location of the maximum of asymmetric two-sided Brownian motion with triangular drift. Statistics & Probability Letters, 29(3):279–284, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001832.

Stramer:1999:LLS

[Str99] O. Stramer. The local linearization scheme for nonlinear diffusion models with discontinuous coefficients. Statistics & Probability Letters, 42(3):249–256, April 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001916.

Stute:1992:SCU

[Stu92] Winfried Stute. Strong consistency under the Koziol–Green model. Statistics & Probability Letters, 14(4):313–320, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290064C.

Stute:1994:CKM

[Stu94] Winfried Stute. Convergence of the Kaplan–Meier estimator in weighted sup-norms. Statistics & Probability Letters, 20(3):219– 223, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900450.

Stummer:1997:EMT

[Stu97] Wolfgang Stummer. On exponential moments of two Brownian functionals. Statistics & Probability Letters, 31(3):233–237, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000351.

Stummer:1998:BES

[Stu98] Wolfgang Stummer. On bounded entropy of solutions of multi-dimensional stochastic differential equations. Statistics & Probability Letters, 36(4):327–336, January 3, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000795.

Subramanya:1994:MDA

[Sub94] U. R. Subramanya. On max domains of attraction of univariate pmax stable laws. Statistics & Probability Letters, 19(4):271–279, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294901767.

Sun:1995:IES

[Sun95] Jiayang Sun. Iterative estimates for a smoothing parameter. Statistics & Probability Letters, 24(4):347–356, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400194D.

Sun:1997:ECS

[Sun97a] Jiaming Sun. Exponential convergence for sequences of random variables. Statistics & Probability Letters, 34(2):159–164, June 2, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001770.

Sun:1997:BCH

[Sun97b] Liuquan Sun. Bandwidth choice for hazard rate estimators from left truncated and right censored data. Statistics & Probability Letters, 36(2):101–114, December 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000539.

Sung:1998:WLL

[Sun98] Soo Hak Sung. Weak Law of Large Numbers for arrays. Statistics & Probability Letters, 38(2):101–105, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001594.

Sung:1999:WLL

[Sun99] Soo Hak Sung. Weak Law of Large Numbers for arrays of random variables. Statistics & Probability Letters, 42(3):293–298, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002193.

Sutradhar:1997:MAE

[Sut97] Brajendra C. Sutradhar. A multivariate approach for estimating the random effects variance component in one-way random effects model. Statistics & Probability Letters, 33(4):333–339, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001447.

Sarda:1991:SPS

[SV91] P. Sarda and P. Vieu. Smoothing parameter selection in hazard estimation. Statistics & Probability Letters, 11(5):429–434, May 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190192T.

Swartz:1992:PPC

[SV92] T. Swartz and C. Villegas. Posterior probability and conditional coverage. Statistics & Probability Letters, 14(3):169–173, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290017Y.

Scarsini:1993:EPE

[SV93] Marco Scarsini and Lorenzo Verdicchio. On the extendibility of partially exchangeable random vectors. Statistics & Probability Letters, 16(1):43–46, January 4, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715293901208.

Samaniego:1999:ISE

[SV99] Francisco J. Samaniego and Eric Vestrup. On improving standard estimators via linear empirical Bayes methods. Statistics & Probability Letters, 44(3):309–318, September 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529900022X.

Singh:1990:TAO

[SW90a] Bahadur Singh and F. T. Wright. Testing for and against an order restriction in mixed-effects models. Statistics & Probability Letters, 9(2):195–200, February 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290016X.

Sobel:1990:DIM

[SW90b] Milton Sobel and Martin Wells. Dirichlet integrals and moments of gamma distribution order statistics. Statistics & Probability Letters, 9(5):431–437, May 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900356.

Stufken:1992:FDT

[SW92] John Stufken and Kui-Jang Wang. Factorial designs and the theory of trade-off. Statistics & Probability Letters, 15(5):369–372, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290155X.

Steel:1993:PAP

[SW93] M. A. Steel and G. R. Wood. On a problem of Andersson and Perlman. Statistics & Probability Letters, 18(5):381–382, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390031D.

Singh:1999:APO

[SW99] Bahadur Singh and F. T. Wright. Approximating the powers of order-restricted log rank tests at local alternatives. Statistics & Probability Letters, 44(1):7–17, August 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002855.

Swart:1996:CPM

[Swa96] J. M. Swart. A conditional product measure theorem. Statistics & Probability Letters, 28(2):131–135, June 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001077.

Shao:1997:SEA

[SX97] Yongzhao Shao and Xiaojing Xiang. Some extensions of the asymptotics of a kernel estimator of a distribution function. Statistics & Probability Letters, 34(3):301–308, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001940.

Sibuya:1995:OUR

[SY95] Masaaki Sibuya and Hajime Yamato. Ordered and unordered random partitions of an integer and the GEM distribution. Statistics & Probability Letters, 25(2):177–183, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002203.

Singh:1992:EMS

[SZ92] R. Karan Singh and S. M. H. Zaidi. On estimating the mean of symmetrical populations. Statistics & Probability Letters, 13(1):67–72, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290238Z.

Sun:1998:SCB

[SZ98] Liuquan Sun and Yong Zhou. Sequential confidence bands for densities under truncated and censored data. Statistics & Probability Letters, 40(1):31–41, September 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000893.

Szekli:1990:CIR

[Sze90] R. Szekli. On the concavity of the infinitesimal renewal function. Statistics & Probability Letters, 10(3):181–184, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090070N.

Szroeter:1992:BNC

[Szr92] Jerzy Szroeter. Bounds for non-central chi-square distributions having unobservable random non-centrality parameters. Statistics & Probability Letters, 13(1):73–81, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292902392.

Szroeter:1996:CEF

[Szr96] J. Szroeter. Characterization of the exact finite-sample distribution of a routine test statistic for non-nested regressions. Statistics & Probability Letters, 29(1):9–14, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001506.

Szyszkowicz:1991:CPC

[Szy91] Barbara Szyszkowicz. Changepoint problems and contiguous alternatives. Statistics & Probability Letters, 11(4):299–308, April 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190039T.

Szyszkowicz:1994:COT

[Szy94] Barbara Szyszkowicz. The Chibisov–O’Reilly theorem for empirical processes under contiguous measures. Statistics & Probability Letters, 19(2):153–159, January 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529490149X.

Tableman:1994:ALT

[Tab94a] Mara Tableman. The asymptotics of the least trimmed absolute deviations (LTAD) estimator. Statistics & Probability Letters, 19(5):387–398, April 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900078.

Tableman:1994:IFL

[Tab94b] Mara Tableman. The influence functions for the least trimmed squares and the least trimmed absolute deviations estimators. Statistics & Probability Letters, 19(4):329–337, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901864.

Takada:1991:MUI

[Tak91] Y. Takada. Median unbiasedness in an invariant prediction problem. Statistics & Probability Letters, 12(4):281–283, October 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900915.

Takahashi:1994:AIP

[Tak94] Rinya Takahashi. Asymptotic independence and perfect dependence of vector components of multivariate extreme statistics. Statistics & Probability Letters, 19(1):19–26, January 3, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900639.

Tallis:1991:NBC

[Tal91] G. M. Tallis. A note on balanced cluster sampling. Statistics & Probability Letters, 11(2):169–172, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190137G.

Tan:1993:IDH

[Tan93a] W. Y. Tan. On the incubation distributions of the HIV epidemic. Statistics & Probability Letters, 18(4):279–287, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390016C.

Tang:1993:LTM

[Tan93b] Loon Ching Tang. Limit theorems for Markov random walks. Statistics & Probability Letters, 18(4):265–270, November 9, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390014A.

Tan:1994:HID

[Tan94] W. Y. Tan. On the HIV incubation distribution under non-Markovian models. Statistics & Probability Letters, 21(1):49–57, September 1, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294900574.

Tan:1995:SMD

[Tan95] Wai Y. Tan. A stochastic model for drug resistance in AIDS chemotherapy and the HIV incubation distribution. Statistics & Probability Letters, 25(4):289–299, December 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002336.

Tan:1996:WEN

[Tan96] Choon Peng Tan. On the weak ergodicity of nonhomogeneous Markov chains. Statistics & Probability Letters, 26(4):293–295, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295000232.

Tarpey:1994:TPP

[Tar94] Thaddeus Tarpey. Two principal points of symmetric, strongly unimodal distributions. Statistics & Probability Letters, 20 (4):253–257, July 14, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294900124.

Teuscher:1995:SIB

[TG95] F. Teuscher and V. Guiard. Sharp inequalities between skewness and kurtosis for unimodal distributions. Statistics & Probability Letters, 22(3):257–260, February 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400074I.

Tu:1996:BTC

[TG96] Dongsheng Tu and Alan J. Gross. A Bartlett-type correction for the subject-years method in comparing survival data to a standard population. Statistics & Probability Letters, 29(2):149–157, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001689.

Tan:1995:SPL

[TH95] Kok Hooi Tan and Petros Hadjicostas. Some properties of a limiting distribution in Quicksort. Statistics & Probability Letters, 25(1):87–94, October 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400209Q.

Thompson:1997:BV

[Tho97] Peter Thompson. Bayes p-values. Statistics & Probability Letters, 31(4):267–274, February 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000387.

Tian:1998:CBP

[Tia98] Guo-Liang Tian. The comparison between polynomial regression and orthogonal polynomial regression. Statistics & Probability Letters, 38(4):289–294, July 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000170.

Ting:1997:NUO

[TLC97] Chao-Ping Ting, Bing-Ying L. Lin, and Feng-Shun Chai. A note on universally optimal row-column designs with empty nodes. Statistics & Probability Letters, 33(2):145–149, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001228.

Tiwari:1993:NCP

[TN93] Neeraj Tiwari and A. K. Nigam. A note on constructive procedure for unbiased controlled rounding. Statistics & Probability Letters, 18(5):415–420, December 2, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390036I.

Tobbal:1996:FLI

[Tob96] Khelifa Tobbal. A functional law of the iterated logarithm for the Dekkers–Einmahl–de Haan tail index estimator. Statistics & Probability Letters, 29(1):15–22, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001522.

Tomkins:1990:GLI

[Tom90] R. J. Tomkins. A generalized law of the iterated logarithm. Statistics & Probability Letters, 10(1):9–15, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090105G.

Toman:1992:BRE

[Tom92a] Blaza Toman. Bayesian robust experimental designs for the one-way analysis of variance. Statistics & Probability Letters, 15(5):395–400, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292901733.

Tomkins:1992:RKL

[Tom92b] R. J. Tomkins. Refinements of Kolmogorov’s law of the iterated logarithm. Statistics & Probability Letters, 14(4):321–325, July 17, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290065D.

Toman:1993:OST

[Tom93] Blaza Toman. Optimal scaling of two-level factorial experiments. Statistics & Probability Letters, 16(4):331–336, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901389.

Tomkins:1996:RZO

[Tom96] R. J. Tomkins. Refinement of a zero-one law for maxima. Statistics & Probability Letters, 27(1):67–69, March 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000453.

Tomkins:1997:RZO

[Tom97] R. J. Tomkins. Refinement of a zero-one law for maxima. Statistics & Probability Letters, 32(4):431, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296002222.

Tomkins:1999:RSL

[Tom99] R. J. Tomkins. Regular stability of large order statistics. Statistics & Probability Letters, 41(2):145–151, January 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001321.

Thavaneswaran:1996:NES

[TP96] A. Thavaneswaran and Shelton Peiris. Nonparametric estimation for some nonlinear models. Statistics & Probability Letters, 28(3): 227–233, July 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529500128X.

Thavaneswaran:1998:HTS

[TP98] A. Thavaneswaran and Shelton Peiris. Hypothesis testing for some time-series models: a power comparison. Statistics & Probability Letters, 38(2):151–156, June 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001661.

Tran:1990:KDE

[Tra90] Lanh Tat Tran. Kernel density estimation under dependence. Statistics & Probability Letters, 10(3):193–201, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090073G.

Trajstman:1996:NCM

[Tra96] A. C. Trajstman. A note on the characterisation of the most probable number. Statistics & Probability Letters, 29(4):293–295, September 16, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001859.

Tracy:1993:TMM

[TS93] Derrick S. Tracy and Shagufta A. Sultan. Third moment of matrix quadratic form. Statistics & Probability Letters, 16(1):71–76, January 4, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293901264.

Thomas:1996:NRR

[TS96] P. Yageen Thomas and Philip Samuel. A note on recurrence relations for the product moments of order statistics. Statistics & Probability Letters, 29(3):245–249, September 2, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001794.

Tsai:1995:GWD

[Tsa95] Ming-Tien Tsai. A generalization of the Wishart distribution. Statistics & Probability Letters, 24(1):67–70, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294001493.

Tsao:1998:CCP

[Tsa98] Chen-Hai Andy Tsao. Conditional coverage probability of confidence intervals in errors-in-variables and related models. Statistics & Probability Letters, 40(2):165–170, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529800090X.

Tsoi:1996:CPQ

[Tso96] Allanus H. Tsoi. A characterization of k-parameter quasimartingales. Statistics & Probability Letters, 28(4):337–343, August 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001433.

Tsodikov:1998:AEP

[Tso98] Alexander Tsodikov. Asymptotic efficiency of a proportional hazards model with cure. Statistics & Probability Letters, 39(3):237–244, August 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771529800056X.

Tsutakawa:1992:MUC

[Tsu92] Robert K. Tsutakawa. Moments under conjugate distributions in bioassay. Statistics & Probability Letters, 15(3):229–233, October 21, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290194A.

Terrin:1991:CGL

[TT91] Norma Terrin and Murad S. Taqqu. Convergence to a Gaussian limit as the normalization exponent tends to 12. Statistics & Probability Letters, 11(5):419–427, May 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190191S.

Tudor:1998:LDS

[Tud98] C. Tudor. Large deviations for stable Itô equations. Statistics & Probability Letters, 40(2):103–111, September 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000327.

Truong-Van:1997:ILL

[TV97] B. Truong-Van. Iterated logarithm law for sample generalized partial autocorrelations. Statistics & Probability Letters, 33(2):217–223, April 30, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715297001302.

Teugels:1998:GGM

[TV98] Jozef L. Teugels and Johan Van Horebeek. Generalized graphical models for discrete data. Statistics & Probability Letters, 38(1): 41–47, May 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001521.

Tiwari:1992:PDS

[TW92a] Ram C. Tiwari and Martin T. Wells. On the power-divergence statistic in sparse multinomial models requiring parameter estimation. Statistics & Probability Letters, 13(1):57–60, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290236X.

Tomkins:1992:STL

[TW92b] R. J. Tomkins and Hong Wang. Stability theorems for large order statistics with varying ranks. Statistics & Probability Letters, 14(2):91–95, May 27, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290069H.

Tsai:1992:ALI

[TW92c] Chih-Ling Tsai and Xizhi Wu. Assessing local influence in linear regression models with first-order autoregressive or heteroscedastic error structure. Statistics & Probability Letters, 14(3):247– 252, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715292900309.

Turner:1992:IKI

[TYS92] Danny W. Turner, Dean M. Young, and John W. Seaman, Jr. Improved Kolmogorov inequalities for the binomial distribution. Statistics & Probability Letters, 13(3):223–227, February 19, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715292900516.

Turner:1995:KIS

[TYS95] Danny W. Turner, Dean M. Young, and John W. Seaman, Jr. A Kolmogorov inequality for the sum of independent Bernoulli random variables with unequal means. Statistics & Probability Letters, 23(3):243–245, May 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400120W.

Tzavelas:1998:NUQ

[Tza98] George Tzavelas. A note on the uniqueness of the quasi-likelihood estimator. Statistics & Probability Letters, 38(2):125–130, June 1, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001624.

Uno:1995:NSE

[Uno95] Chikara Uno. A note on sequential estimation of the mean. Statistics & Probability Letters, 23(2):141–149, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400106I.

Uno:1996:LBN

[Uno96] Takashi Uno. On the lower bound of the number of real roots of a random algebraic equation. Statistics & Probability Letters, 30(2):157–163, October 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002146.

Ushakov:1995:EDS

[UU95] N. G. Ushakov and A. P. Ushakova. Estimations of the decomposition stability into identical components. Statistics & Probability Letters, 25(3):221–229, November 15, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400225W.

Valderrama:1997:NCR

[VA97] M. J. Valderrama and A. M. Aguilera. A normality criterion for random vectors based on independence. Statistics & Probability Letters, 33(2):159–165, April 30, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001241.

Vandev:1993:NBP

[Van93] D. Vandev. A note on the breakdown point of the least median of squares and least trimmed squares estimators. Statistics & Probability Letters, 16(2):117–119, January 27, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390155C.

Vasudaven:1990:KST

[Vas90] M. Vasudaven. On Kolmogorov–Smirnov type aligned test in ksample linear regression. Statistics & Probability Letters, 10(1): 23–27, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090107I.

Vellaisamy:1999:CPA

[VC99] P. Vellaisamy and B. Chaudhuri. On compound Poisson approximation for sums of random variables. Statistics & Probability Letters, 41(2):179–189, January 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001412.

Vasudeva:1991:LIL

[VD91] R. Vasudeva and G. Divanji. Law of iterated logarithm for random subsequences. Statistics & Probability Letters, 12(3):189–194, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715291900764.

Viano:1994:CTF

[VDO94] M. C. Viano, C. Deniau, and G. Oppenheim. Continuous-time fractional ARMA processes. Statistics & Probability Letters, 21(4):323–336, November 11, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294000158.

vanEs:1997:NIS

[vE97] Bert van Es. A note on the integrated squared error of a kernel density estimator in non-smooth cases. Statistics & Probability Letters, 35(3):241–250, October 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000199.

Veijanen:1990:CMP

[Vei90] Ari Veijanen. On conditionally mixing processes. Statistics & Probability Letters, 9(1):59–65, January 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090096P.

vanEs:1998:SKE

[vEK98] A. J. van Es and A. R. Kok. Simple kernel estimators for certain nonparametric deconvolution problems. Statistics & Probability Letters, 39(2):151–160, August 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000546.

Velilla:1993:NMB

[Vel93a] Santiago Velilla. A note on the multivariate Box–Cox transformation to normality. Statistics & Probability Letters, 17(4):259– 263, July 13, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715293902003.

Velilla:1993:QBE

[Vel93b] Santiago Velilla. Quantile-based estimation for the Box– Cox transformation in random samples. Statistics & Probability Letters, 16(2):137–145, January 27, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390158F.

Velilla:1998:NBR

[Vel98] Santiago Velilla. A note on the behaviour of residual plots in regression. Statistics & Probability Letters, 37(3):269–278, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001260.

Vesely:1994:ARS

[Ves94] Petr Veselý. Arbitrarily reliable systems with two dual modes of failure. Statistics & Probability Letters, 20(3):203–207, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900434.

Vidakovic:1996:NRD

[Vid96] Brani Vidakovic. A note on random densities via wavelets. Statistics & Probability Letters, 26(4):315–321, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000267.

Vieu:1996:NDM

[Vie96] Philippe Vieu. A note on density mode estimation. Statistics & Probability Letters, 26(4):297–307, March 1, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000240.

Viennet:1999:LSE

[Vie99] Gabrielle Viennet. Least-square estimation for regression on random designs for absolutely regular observations. Statistics & Probability Letters, 43(1):13–23, May 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002211.

Vitale:1996:CIN

[Vit96] Richard A. Vitale. Covariance identities for normal variables via convex polytopes. Statistics & Probability Letters, 30(4):363–368, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529600003X.

Vitale:1999:MGC

[Vit99] Richard A. Vitale. Majorization and Gaussian correlation. Statistics & Probability Letters, 45(3):247–251, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000644.

Vasudaven:1996:ECA

[VNS96] M. Vasudaven, M. G. Nair, and M. M. Sithole. On estimation for censored autoregressive data. Statistics & Probability Letters, 31(2):97–105, December 16, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000193.

vonPlato:1991:FPE

[vP91] Jan von Plato. Finite partial exchangeability. Statistics & Probability Letters, 11(2):99–102, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190124A.

vanRooij:1998:CHM

[vRR98] Arnoud C. M. van Rooij and Frits H. Ruymgaart. Convergence in the Hausdorff metric of estimators of irregular densities, using Fourier–Cesàro approximation. Statistics & Probability Letters, 39(2):179–184, August 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000625.

Walker:1998:CHD

[Wal98] Stephen Walker. A characterisation of Hjort’s discrete time beta process. Statistics & Probability Letters, 37(4):351–355, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001375.

Walter:1999:DEP

[Wal99] Gilbert G. Walter. Density estimation in the presence of noise. Statistics & Probability Letters, 41(3):237–246, February 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298001606.

Wansbeek:1991:SVDa

[Wan91a] Tom Wansbeek. Singular value decomposition of design matrices in ANOVA with balanced data. Statistics & Probability Letters, 11(1):33–36, January 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190173O.

Wansbeek:1991:SVDb

[Wan91b] Tom Wansbeek. Singular value decomposition of design matrices in ANOVA with balanced data. Statistics & Probability Letters, 12(5):441, November 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190033N.

Wand:1992:FSP

[Wan92a] M. P. Wand. Finite sample performance of density estimators under moving average dependence. Statistics & Probability Letters, 13(2):109–115, January 27, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290084I.

Wang:1992:LLT

[Wan92b] Mei Wang. A local limit theorem for perturbed random walks. Statistics & Probability Letters, 15(2):109–116, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290121K.

Wang:1992:GSA

[Wan92c] Suojin Wang. General saddlepoint approximations in the bootstrap. Statistics & Probability Letters, 13(1):61–66, January 2, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290237Y.

Wang:1994:BTA

[Wan94a] Yazhen Wang. A Bartlett-type adjustment for the likelihood ratio statistic with an ordered alternative. Statistics & Probability Letters, 20(5):347–352, August 9, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901252.

Wang:1994:LDC

[Wan94b] Yazhen Wang. The limit distribution of the concave majorant of an empirical distribution function. Statistics & Probability Letters, 20(1):81–84, May 4, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294902380.

Wang:1996:MDM

[Wan96a] J. C. Wang. Mixed difference matrices and the construction of orthogonal arrays. Statistics & Probability Letters, 28(2):121– 126, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001050.

Wang:1996:EPT

[Wan96b] Yining Wang. Estimation problems for the two-parameter negative binomial distribution. Statistics & Probability Letters, 26(2):113–114, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294002592.

Wang:1997:SBP

[Wan97] Yazhen Wang. Small ball problem via wavelets for Gaussian processes. Statistics & Probability Letters, 32(2):133–139, March 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529600065X.

Wand:1998:FSP

[Wan98a] M. P. Wand. Finite sample performance of deconvolving density estimators. Statistics & Probability Letters, 37(2):131–139, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297001107.

Wang:1998:ACC

[Wan98b] Hsiuying Wang. Admissibility of the constant-coverage probability estimator for estimating the coverage function of certain confidence interval. Statistics & Probability Letters, 36(4):365–372, January 3, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715297000837.

Wang:1998:SSC

[Wan98c] Yuedong Wang. Sample size calculations for smoothing splines based on Bayesian confidence intervals. Statistics & Probability Letters, 38(2):161–166, June 1, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297001685.

Wang:1999:RSC

[Wan99a] C. Y. Wang. Robust sandwich covariance estimation for regression calibration estimator in Cox regression with measurement error. Statistics & Probability Letters, 45(4):371–378, December 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000796.

Wang:1999:SBE

[Wan99b] Qi-Hua Wang. Some bounds for the error of an estimator of the hazard function with censored data. Statistics & Probability Letters, 44(4):319–326, October 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002338.

Wang:1999:KET

[Wan99c] Qiying Wang. Kolmogorov and Erdős test for self-normalized sums. Statistics & Probability Letters, 42(3):323–326, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002284.

Wang:1999:BER

[Wan99d] Qiying Wang. On Berry–Esseen rates for m-dependent Ustatistics. Statistics & Probability Letters, 41(2):123–130, January 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771529800128X.

Watson:1993:ITU

[Wat93] Geoffrey S. Watson. Invariant tests for uniformity on the vertices of a regular polygon. Statistics & Probability Letters, 17(1):67– 71, May 3, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390197Q.

Watson:1995:DCI

[Wat95] G. S. Watson. Detecting a change in the intercept in multiple regression. Statistics & Probability Letters, 23(1):69–72, April 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400096Q.

Wang:1990:NAG

[WC90] Song-Gui Wang and Shein-Chung Chow. A note on adaptive generalized ridge regression estimator. Statistics & Probability Letters, 10(1):17–21, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090106H.

Wang:1997:IRP

[WC97] H. Wang and X. Chen. On the interval recurrence property of (N, d)–Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 33(1):79–84, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296001125.

Wang:1994:OLS

[WCT94] Song-Gui Wang, Shein-Chung Chow, and Siu-Keung Tse. On ordinary least-squares methods for sample surveys. Statistics & Probability Letters, 20(3):173–182, June 22, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900396.

Weber:1991:GTR

[Web91] James Weber. Generalized transformations of random variables. Statistics & Probability Letters, 12(2):161–166, August 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190062V.

Weba:1992:NEN

[Web92] Michael Weba. A note on the estimation of L^{p}-norms. Statistics & Probability Letters, 15(2):131–133, September 28, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290125O.

Wefelmeyer:1991:GAL

[Wef91] W. Wefelmeyer. A generalization of asymptotically linear estimators. Statistics & Probability Letters, 11(3):195–199, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529190143F.

Weissfeld:1990:IDP

[Wei90] L. A. Weissfeld. Influence diagnostics for the proportional hazards model. Statistics & Probability Letters, 10(5):411–417, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090022Y.

Weiss:1991:AMT

[Wei91] Lionel Weiss. Asymptotically minimax tests of some nonstandard composite hypotheses. Statistics & Probability Letters, 11(3):251–254, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190151G.

Weiss:1992:NTG

[Wei92] Lionel Weiss. Neyman’s C(α) test as a GLRT. Statistics & Probability Letters, 15(2):121–124, September 28, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290123M.

Wen:1994:SSL

[Wen94] Liu Wen. Some strong limit theorems relative to the geometric average of random transition probabilities of arbitrary finite nonhomogeneous Markov chains. Statistics & Probability Letters, 21(1):77–83, September 1, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294000026.

Wen:1997:KSD

[Wen97] Liu Wen. A kind of strong deviation theorem for the sequences of nonnegative integer-valued random variables. Statistics & Probability Letters, 32(4):343–349, April 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000934.

Weron:1996:CMS

[Wer96] Rafal Weron. On the Chambers–Mallows–Stuck method for simulating skewed stable random variables. Statistics & Probability Letters, 28(2):165–171, June 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001131.

Wesolowski:1990:MCW

[Wes90] Jacek Wesołowski. A martingale characterization of the Wiener process. Statistics & Probability Letters, 10(3):213–215, August 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090075I.

Wesolowski:1994:MCW

[Wes94] Jacek Wesołowski. A martingale characterization of the Wiener process. Statistics & Probability Letters, 19(2):167, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901511.

Wesolowski:1997:SDL

[Wes97] Jacek Wesołowski. On some distributional and limit properties of factorizable distributions. Statistics & Probability Letters, 32(4):329–337, April 1, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000910.

Wang:1994:IIC

[WGR94] Zhiming Wang, J. C. Gardiner, and R. V. Ramamoorthi. Identifiability in interval censorship models. Statistics & Probability Letters, 21(3):215–221, October 19, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901171.

Wise:1991:NDD

[WH91] Gary L. Wise and Eric B. Hall. A note on the distribution of the determinant of a random matrix. Statistics & Probability Letters, 11(2):147–148, February 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190132B.

Wise:1993:MIO

[WH93] Gary L. Wise and Eric B. Hall. On mutual independence. Statistics & Probability Letters, 17(5):395–398, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390260P.

Whitmore:1990:RSS

[Whi90] G. A. Whitmore. On the reliability of stochastic systems: a comment. Statistics & Probability Letters, 10(1):65–67, June 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090113L.

White:1999:FPR

[Whi99] Damien G. White. On fractal percolation in R^{2}. Statistics & Probability Letters, 45(2):187–190, November 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000589.

Wichura:1990:NDA

[Wic90] Michael J. Wichura. A note on Das’s approximation to Mills’s ratio. Statistics & Probability Letters, 10(4):297–299, September 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715290900459.

Wiens:1991:DAL

[Wie91] Douglas P. Wiens. Designs for approximately linear regression: two optimality properties of uniform designs. Statistics & Probability Letters, 12(3):217–221, September 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715291900812.

Wiens:1996:AGE

[Wie96] Douglas P. Wiens. Asymptotics of generalized m-estimation of regression and scale with fixed carriers, in an approximately linear model. Statistics & Probability Letters, 30(3):271–285, October 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295002308.

Wilfling:1996:LOP

[Wil96] Bernd Wilfling. Lorenz ordering of power-function order statistics. Statistics & Probability Letters, 30(4):313–319, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715295002340.

Wilcox:1998:STS

[Wil98] Rand R. Wilcox. Simulations on the Theil–Sen regression estimator with right-censored data. Statistics & Probability Letters, 39(1):43–47, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298000224.

Wirjanto:1996:LDU

[Wir96] T. S. Wirjanto. The limiting distributions of unit-root tests for data with cross-sectional and time-series dimensions. Statistics & Probability Letters, 30(1):73–77, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295002030.

Wise:1992:CMC

[Wis92] Gary L. Wise. A counterexample to a martingale characterization of a Wiener process. Statistics & Probability Letters, 15(5):337–338, December 8, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529290149Y.

Wisniewski:1996:EOS

[Wiś96] Mateusz Wiśniewski. On extreme-order statistics and point processes of exceedances in multivariate stationary Gaussian sequences. Statistics & Probability Letters, 29(1):55–59, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001557.

Wang:1993:DCP

[WJ93] Y. H. Wang and Shuixin Ji. Derivations of the compound Poisson distribution and process. Statistics & Probability Letters, 18(1):1–7, August 30, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529390091V.

Wang:1999:ELP

[WJ99] Qi-Hua Wang and Bing-Yi Jing. Empirical likelihood for partial linear models with fixed designs. Statistics & Probability Letters, 41(4):425–433, February 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002302.

Wells:1992:TFU

[WJT92] Martin T. Wells, Sreenivasa R. Jammalamadaka, and Ram C. Tiwari. Tests of fit using spacings statistics with estimated parameters. Statistics & Probability Letters, 13(5):365–372, April 7, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290109I.

Wan:1999:IFM

[WK99] Alan T. K. Wan and Hiroko Kurumai. An iterative feasible minimum mean squared error estimator of the disturbance variance in linear regression under asymmetric loss. Statistics & Probability Letters, 45(3):253–259, November 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000656.

Wu:1993:RSS

[WL93] Xizhi Wu and Zhen Luo. Residual sum of squares and multiple potential, diagnostics by a second order local approach. Statistics & Probability Letters, 16(4):289–296, March 16, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293901323.

Wang:1995:GAG

[WL95] S. J. Wang and Sik-Yum Lee. A geometric approach of the generalized least-squares estimation in analysis of covariance structures. Statistics & Probability Letters, 24(1):39–47, July 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400146Y.

Walker:1997:CPT

[WM97] Stephen Walker and Pietro Muliere. A characterisation of Pólya tree distributions. Statistics & Probability Letters, 31(3):163–168, January 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000284.

Witt:1994:RMA

[WMN94] Lee D. Witt, Joseph W. McKean, and Joshua D. Naranjo. Robust measures of association in the correlation model. Statistics & Probability Letters, 20(4):295–306, July 14, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900175.

Wong:1993:NIM

[Won93a] Augustine C. M. Wong. A note on inference for the mean parameter of the gamma distribution. Statistics & Probability Letters, 17(1):61–66, May 3, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390196P.

Wong:1993:MNS

[Won93b] Weng Kee Wong. Minimal number of support points for minimax optimal designs. Statistics & Probability Letters, 17(5):405– 409, August 5, 1993. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529390262H.

Wong:1995:EOD

[Won95] Weng Kee Wong. On the equivalence of D and G-optimal designs in heteroscedastic models. Statistics & Probability Letters, 25(4):317–321, December 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294002361.

Wood:1992:BOS

[Woo92] Andrew T. A. Wood. On the bias of order statistics in noni.i.d. samples. Statistics & Probability Letters, 15(4):285–291, November 13, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290164Z.

Wood:1995:WTC

[Woo95] Andrew T. A. Wood. When is a truncated covariance function on the line a covariance function on the circle? Statistics & Probability Letters, 24(2):157–164, August 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294001622.

Wu:1996:AEP

[WPH96] Lilian Shiao-Yen Wu, Jeffrey S. Pai, and J. R. M. Hosking. An algorithm for estimating parameters of state-space models. Statistics & Probability Letters, 28(2):99–106, June 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295000984.

Wang:1990:NCB

[WR90] Xiang Chen Wang and M. Bhaskara Rao. A note on convergence in Banach spaces of cotype p. Statistics & Probability Letters, 10(5):391–396, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715290900194.

Wood:1998:BGI

[WR98] G. R. Wood and J. M. Robertson. Buffon got it straight. Statistics & Probability Letters, 37(4):415–421, March 30, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001454.

Wright:1997:SAT

[Wri97] Tommy Wright. A simple algorithm for tighter exact upper confidence bounds with rare attributes in finite universes. Statistics & Probability Letters, 36(1):59–67, November 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297000497.

Wang:1995:SRS

[WRL95] Xaingchen Wang, M. Bhaskara Rao, and Deli Li. Some results on strong limit theorems for (LB)-space-valued random variables. Statistics & Probability Letters, 23(3):247–251, May 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400121N.

Weissfeld:1990:IDW

[WS90] Lisa A. Weissfeld and Helmut Schneider. Influence diagnostics for the Weibull model fit to censored data. Statistics & Probability Letters, 9(1):67–73, January 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090097Q.

Wu:1998:WAI

[WS98] Tiee-Jian Wu and Alfred Sepulveda. The weighted average information criterion for order selection in time series and regression models. Statistics & Probability Letters, 39(1):1–10, July 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298000030.

Wang:1990:MML

[WTC90] Song-Gui Wang, Siu-Keung Tse, and Shein-Chung Chow. On the measures of multicollinearity in least squares regression. Statistics & Probability Letters, 9(4):347–355, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529090145W.

Wu:1990:NBA

[Wu90] Tiee-Jian Wu. L_{p} nonuniform bounds for asymptotic normality of linear rank statistics. Statistics & Probability Letters, 9(4):

361–365, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090147Y.

Wu:1992:DMF

[Wu92] Eden K. H. Wu. Distributions minimizing Fisher information for scale in ϵ-contamination neighbourhoods. Statistics & Probability Letters, 14(5):373–383, July 11, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290098P.

Wu:1997:NCM

[Wu97a] Chufang Wu. New characterization of Marshall–Olkin-type distributions via bivariate random summation scheme. Statistics & Probability Letters, 34(2):171–178, June 2, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297827903.

Wu:1997:EKC

[Wu97b] Colin O. Wu. The effects of kernel choices in density estimation with biased data. Statistics & Probability Letters, 34(4):373– 383, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296002052.

Wu:1999:SCW

[Wu99a] Wei Biao Wu. On the strong convergence of a weighted sum. Statistics & Probability Letters, 44(1):19–22, August 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002879.

Wu:1999:CTF

[Wu99b] Zhen Wu. The comparison theorem of FBSDE. Statistics & Probability Letters, 44(1):1–6, August 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298002399.

Wang:1990:LLT

[WW90a] Mei Wang and Michael Woodroofe. A local limit theorem for sums of dependent random variables. Statistics & Probability Letters, 9(3):207–213, March 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090057E.

Wiens:1990:AMP

[WW90b] Douglas P. Wiens and K. H. Eden Wu. Asymptotic minimax properties of M-estimators of scale. Statistics & Probability Letters, 10(5):363–368, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090015Y.

Wang:1992:ORE

[WW92a] Yu Wang and Douglas Wiens. Optimal, robust R-estimators and test statistics in the linear model. Statistics & Probability Letters, 14(3):179–188, June 24, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715292900192.

Weerahandi:1992:SST

[WW92b] Samaradasa Weerahandi and Robert G. White. Some survey techniques for sampling rare subjects. Statistics & Probability Letters, 15(5):361–368, December 8, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290154W.

Wu:1994:PSN

[WW94] Xizhi Wu and Fanghuan Wan. A perturbation scheme for nonlinear models. Statistics & Probability Letters, 20(3):197–202, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900426.

Wahba:1995:BNZ

[WW95a] Grace Wahba and Yuedong Wang. Behavior near zero of the distribution of GCV smoothing parameter estimates. Statistics & Probability Letters, 25(2):105–111, November 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400211P.

Wang:1995:IMC

[WW95b] C. Y. Wang and Suojin Wang. On information matrices in case-control studies. Statistics & Probability Letters, 22(4):269– 274, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400076K.

Wen:1995:LTE

[WW95c] Liu Wen and Yang Weiguo. A limit theorem for the entropy density of nonhomogeneous Markov information source. Statistics & Probability Letters, 22(4):295–301, March 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400080R.

Wan:1996:DAB

[WW96] Fanghuan Wan and Xizhi Wu. A direct approach to a Bayesian sequential test for a normal mean. Statistics & Probability Letters, 27(1):37–41, March 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295000410.

Wen:1995:LRC

[WZ95] Liu Wen and Liu Zikuan. Likelihood ratio and a class of strong laws for the sequence of integer-valued random variables. Statistics & Probability Letters, 22(3):249–256, February 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529400073H.

Wen:1996:ETG

[WZ96] Liu Wen and Wang Zhongzhi. An extension of a theorem on gambling systems to arbitrary binary random variables. Statistics & Probability Letters, 28(1):51–58, June 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500081X.

Woodroofe:1999:IEG

[WZ99] Michael Woodroofe and Rong Zhang. Isotonic estimation for grouped data. Statistics & Probability Letters, 45(1):41–47, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715299000395.

Xiang:1995:NBE

[Xia95] Xiaojing Xiang. A note on the bias of L-estimators and a bias reduction procedure. Statistics & Probability Letters, 23(2):123– 127, May 1, 1995. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529400103F.

Xiao:1997:PDI

[Xia97] Yimin Xiao. Packing dimension of the image of fractional Brownian motion. Statistics & Probability Letters, 33(4):379–387, May 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296001514.

Xiaojun:1991:MAR

[XM91] Li Xiaojun and Joel M. Morris. On measuring asymmetry and the reliability of the skewness measure. Statistics & Probability Letters, 12(3):267–271, September 1991. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529190089A.

Xu:1997:ECB

[XY97] Jian-Lun Xu and Grace L. Yang. An exponential characterization based on a type II censored sample. Statistics & Probability Letters, 31(4):295–298, February 1, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000429.

Yamato:1990:PPR

[Yam90] Hajime Yamato. Probabilistic proofs of relations with Stirling numbers of the first kind by Dirichlet process. Statistics & Probability Letters, 10(3):189–192, August 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090072F.

Yang:1990:ERE

[Yan90] Song Yang. Efficient robust estimation of parameter in the random censorship model. Statistics & Probability Letters, 10(5):419–426, October 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529090023Z.

Yannaros:1991:PAR

[Yan91] Nikos Yannaros. Poisson approximation for random sums of Bernoulli random variables. Statistics & Probability Letters, 11(2):161–165, February 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529190135E.

Yang:1994:CLT

[Yan94] Song Yang. A Central Limit Theorem for functionals of the Kaplan–Meier estimator. Statistics & Probability Letters, 21 (5):337–345, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715294000263.

Yang:1996:OSR

[Yan96a] Song Yang. One-step robust parametric estimation with application to random censoring model. Statistics & Probability Letters, 26(3):225–232, February 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715296859700.

Yang:1996:WEP

[Yan96b] Song Yang. Weighted empiricals and the product-limit estimator in the multiplicative hazard and time transfer regression model. Statistics & Probability Letters, 30(1):17–24, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001964.

Yao:1990:ABC

[Yao90] Yi-Ching Yao. On the asymptotic behavior of a class of nonparametric tests for a change-point problem. Statistics & Probability Letters, 9(2):173–177, February 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290013U.

Yatracos:1991:SRP

[Yat91] Yannis G. Yatracos. On the species and related problems. Statistics & Probability Letters, 12(3):209–212, September 1991. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715291900797.

Yatracos:1996:LLS

[Yat96] Yannis G. Yatracos. Linear least squares regression: a different view. Statistics & Probability Letters, 29(2):143–148, August 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715295001670.

Yuan:1999:ADN

[YB99] Ke-Hai Yuan and Peter M. Bentler. On asymptotic distributions of normal theory MLE in covariance structure analysis under some nonnormal distributions. Statistics & Probability Letters, 42(2):107–113, April 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001710.

Yin:1990:DHB

[YC90] Y. Yin and R. J. Carroll. A diagnostic for heteroscedasticity based on the Spearman rank correlation. Statistics & Probability Letters, 10(1):69–76, June 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090114M.

Ycart:1992:IVM

[Yca92] Bernard Ycart. Integer valued Markov processes and exponential families. Statistics & Probability Letters, 14(1):71–78, May 4, 1992. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529290213O.

Yip:1993:EPS

[YF93] Paul Yip and Daniel Y. T. Fong. Estimating population size from a removal experiment. Statistics & Probability Letters, 16(2):129–135, January 27, 1993. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390157E.

Yang:1996:CMN

[YFL96] Zhen-Hai Yang, Kai-Tai Fang, and Jia-Juan Liang. A characterization of multivariate normal distribution and its application. Statistics & Probability Letters, 30(4):347–352, November 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715295002383.

Yin:1999:JDH

[Yin99] Chuancun Yin. The joint distribution of the hitting time and place to a sphere or spherical shell for Brownian motion with drift. Statistics & Probability Letters, 42(4):367–373, May 1, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715298002314.

Yip:1992:IPH

[Yip92] Paul Yip. An inference procedure for host-vector and venereal disease models. Statistics & Probability Letters, 14(5):363–371, July 11, 1992. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529290097O.

Yu:1996:LRT

[YL96] Philip L. H. Yu and K. Lam. Likelihood ratio test for the spacing between two adjacent location parameters. Statistics & Probability Letters, 26(1):43–49, January 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294002509.

Yamada:1999:TLS

[YL99] Shu Yamada and Dennis K. J. Lin. Three-level supersaturated designs. Statistics & Probability Letters, 45(1):31–39, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000383.

Yoshida:1999:SMP

[YLB99] O. Yoshida, J. G. Leite, and H. Bolfarine. Stochastic monotonicity properties of Bayes estimation of the population size for capture-recapture data. Statistics & Probability Letters, 42(3):257–266, April 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715298002089.

Yeo:1991:CBG

[YM91] G. F. Yeo and R. K. Milne. On characterizations of beta and gamma distributions. Statistics & Probability Letters, 11(3):239– 242, March 1991. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529190149L.

Yue:1995:MNS

[YM95] Xinnian Yue and Chunsheng Ma. Multivariate L_{√}-norm

symmetric distributions. Statistics & Probability Letters, 24 (4):281–288, September 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016771529400185B.

Yue:1999:VQS

[YM99] Rong-Xian Yue and Shi-Song Mao. On the variance of quadrature over scrambled nets and sequences. Statistics & Probability Letters, 44(3):267–280, September 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000188.

Yao:1996:CSM

[YR96] Tzy-Jyun Yao and Gregory C. Reinsel. Covariance structure models for clustered proportions. Statistics & Probability Letters, 29(1):61–70, August 15, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001565.

Yu:1998:APG

[YSLW98] Qiqing Yu, Anton Schick, Linxiong Li, and George Y. C. Wong. Asymptotic properties of the GMLE with case 2 interval-censored data. Statistics & Probability Letters, 37(3):223–228, March 16, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771529700120X.

Young:1994:IDP

[YSM94] Dean M. Young, John W. Seaman, Jr., and Laurie M. Meaux. Independence-distribution-preserving dependency structures for the modified likelihood ratio test for detecting unequal covariance matrices. Statistics & Probability Letters, 21(5):395–403, December 7, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294000379.

Yu:1993:AED

[Yu93] Qiqing Yu. Admissibility of the Empirical Distribution Function in discrete nonparametric invariant problems. Statistics & Probability Letters, 18(5):337–343, December 2, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529390025E.

Yu:1994:TSW

[Yu94] Kai Fun Yu. Truncating sample weights reduces variance. Statistics & Probability Letters, 19(4):267–269, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901759.

Yu:1996:NSA

[Yu96] Hao Yu. A note on strong approximation for quantile processes of strong mixing sequences. Statistics & Probability Letters, 30(1):1–7, September 30, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295001948.

Yuan:1996:SPL

[Yua96] Ao Yuan. Some properties of the Lynden–Bell estimator with truncated data. Statistics & Probability Letters, 27(4):375–384, May 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715295001026.

Zajic:1996:LDC

[Zaj96] Tim Zajic. Large deviations for a class of recursive algorithms. Statistics & Probability Letters, 26(2):135–140, February 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529500002X.

Zamar:1990:RAU

[Zam90] Ruben H. Zamar. Robustness against unexpected dependence in the location model. Statistics & Probability Letters, 9(4):367– 374, April 1990. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/016771529090148Z.

Zarepour:1999:BCH

[Zar99] M. Zarepour. Bootstrapping convex hulls. Statistics & Probability Letters, 45(1):55–63, October 15, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715299000425.

Zezhi:1994:CTS

[ZB94] Li Zezhi and Zhang Buchen. A convergence theorem for sums of dependent Hilbert space valued triangular arrays. Statistics & Probability Letters, 19(3):177–179, February 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294901015.

Zhu:1995:ATB

[ZC95] Li-Xing Zhu and Ping Cheng. The asymptotic tail behaviours of projection pursuit-type Kolmogorov statistics. Statistics & Probability Letters, 25(3):253–263, November 15, 1995. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294002292.

Zelterman:1990:TQI

[Zel90] Daniel Zelterman. On tests for qualitative interactions. Statistics & Probability Letters, 10(1):59–63, June 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529090112K.

Zhang:1992:RTB

[Zha92a] Zhiyi Zhang. Recovery tests in BIBDs with very small degrees of freedom for interblock errors. Statistics & Probability Letters, 15(3):197–202, October 21, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529290189C.

Zhang:1992:SFD

[Zha92b] Zhiyi Zhang. A spectral form of the dispersion model in designs with arbitrarily unequal block sizes. Statistics & Probability Letters, 15(4):313–319, November 13, 1992. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715292901685.

Zhang:1994:HOA

[Zha94a] Guoqiang Zhang. Higher-order asymptotic theory for discriminant analysis of Gaussian ARMA processes. Statistics & Probability Letters, 20(4):259–268, July 14, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294900132.

Zhang:1994:STD

[Zha94b] Sixiang Zhang. Singularity of two diffusion on C_{∞}. Statistics & Probability Letters, 19(2):143–145, January 27, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715294901473.

Zhao:1994:CMQ

[Zha94c] Yi Zhao. Covariance matrices of quadratic forms in elliptical distributions. Statistics & Probability Letters, 21(2):131–140, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167715294902208.

Zhang:1996:CCM

[Zha96] Li-Xin Zhang. Complete convergence of moving average processes under dependence assumptions. Statistics & Probability Letters, 30(2):165–170, October 15, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295002154.

Zhang:1997:ELC

[Zha97a] Biao Zhang. Empirical likelihood confidence intervals for Mfunctionals in the presence of auxiliary information. Statistics & Probability Letters, 32(1):87–97, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296000600.

Zhang:1997:APD

[Zha97b] Dixin Zhang. Asymptotic properties for Dirichlet processes indexed by a class of functions. Statistics & Probability Letters, 32(1):25–33, February 14, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715296000533.

Zhou:1996:NTP

[Zho96] Yong Zhou. A note on the TJW product-limit estimator for truncated and censored data. Statistics & Probability Letters, 26(4):381–387, March 1, 1996. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0167715295000356.

Zhou:1999:RAC

[Zho99] Mai Zhou. Regression analysis with censored data: Extensions of Koul–Susarla–Van Ryzin approach. Statistics & Probability Letters, 41(3):229–236, February 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529800159X.

Zielinski:1990:ROS

[Zie90] Ryszard Zieliński. Robustness of the one-sided Mann–Whitney– Wilcoxon test to dependency between samples. Statistics & Probability Letters, 10(4):291–295, September 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715290900448.

Zielinski:1998:USC

[Zie98] Ryszard Zieliński. Uniform strong consistency of sample quantiles. Statistics & Probability Letters, 37(2):115–119, February 15, 1998. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715297001089.

Zielinski:1999:BEN

[Zie99] Ryszard Zieliński. Best equivariant nonparametric estimator of a quantile. Statistics & Probability Letters, 45(1):79–84, October 15, 1999. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715299000450.

Zimmerman:1994:LDC

[Zim94] Dale L. Zimmerman. On the limiting distribution of and critical values for an origin-invariant bivariate Cramér-von Mises-type statistic. Statistics & Probability Letters, 20(3):189–195, June 22, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294900418.

Zhang:1993:APF

[ZK93] Ping Zhang and Abba M. Krieger. Appropriate penalties in the final prediction error criterion: a decision theoretic approach. Statistics & Probability Letters, 18(3):169–177, October 15, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715293902122.

Zhang:1997:BEB

[ZK97a] Shunpu Zhang and Rohana J. Karunamuni. Bayes and empirical Bayes estimation with errors in variables. Statistics & Probability Letters, 33(1):23–34, April 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771529600106X.

Zhao:1997:LDM

[ZK97b] Yi Zhao and Sadanori Konishi. Limit distributions of multivariate kurtosis and moments under Watson rotational symmetric distributions. Statistics & Probability Letters, 32(3):291–299, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296000867.

Zou:1997:AUE

[ZL97] Guohua Zou and Hua Liang. Admissibility of the usual estimators under error-in-variables superpopulation model. Statistics & Probability Letters, 32(3):301–309, March 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715296000879.

Zhu:1995:MTE

[ZN95] Li-Xing Zhu and Kai W. Ng. M-type estimators of regression function with applications. Statistics & Probability Letters, 25(2):133–144, November 1, 1995. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/016771529400215T.

Zografos:1994:ADE

[Zog94]  K. Zografos.  Asymptotic distributions of estimated f-

dissimilarity between populations in stratified random sampling. Statistics & Probability Letters, 21(2):147–151, September 23, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167715294902224.

Zhu:1996:ELR

[ZR96] Yiliang Zhu and Nancy Reid. Efficient likelihood ratio tests under P -ancillarity and P -sufficiency. Statistics & Probability Letters, 29(3):213–221, September 2, 1996. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0167715295001751.

Zacks:1998:OBF

[ZRB98] S. Zacks, A. Rogatko, and J. Babb. Optimal Bayesianfeasible dose escalation for cancer phase i trials. Statistics & Probability Letters, 38(3):215–220, June 15, 1998. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298000169.

Zhu:1997:PPM

[ZW97a]  Hong-Tu Zhu and Bo-Cheng Wei.  Preferred point α-

manifold and Amari’s α-connections. Statistics & Probability Letters, 36(3):219–229, December 15, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715297000655.

Zhu:1997:SNP

[ZW97b] Hong-Tu Zhu and Bo-Cheng Wei. Some notes on preferred point α-geometry and α-divergence function. Statistics & Probability Letters, 33(4):427–437, May 16, 1997. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715296001575.

Zacks:1999:EVC

[ZW99] Shelemyahu Zacks and Xiaodong Wang. Estimation of variance components in dynamic linear models. Statistics & Probability Letters, 41(3):325–330, February 1, 1999. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715298001680.

Zhou:1993:WWS

[ZZ93] Xiao-Wen Zhou and Jian-Wei Zhou. When will the sample paths be step function for two-parameter stochastic processes. Statistics & Probability Letters, 18(2):147–151, September 22, 1993. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016771529390183J.